A Bibliography of Papers in Lecture Notes in Computer Science (2002) (Part 3 of 4)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.12

Title word cross-reference

(2 + f(n)) [599]. 1 [642, 1201, 524]. 2 [526, 1036, 1273, 1251, 1236, 81]. 2^{m} [612]. 3 [872, 1761, 1036, 842, 1039, 639, 493]. 5 [1422]. ^{(2,128)} [522]. ^{∗} [409]. ^{2} [314]. ^{♯} [117]. st [1039]. ^{T M} [1337]. _{P ASS} [785]. _{Z} [738]. β [290]. δ [91]. ℓ [1181]. k [1487, 1181, 440, 1235, 917, 426, 640, 648, 916, 624, 501, 846]. λ [290]. m [540]. F_{q} [526]. q [94]. K [765]. O(\ log \) [402]. μ [797]. n [540, 652]. n − 3 [652]. Ω [775, 558, 776]. s, z, C, t [964].

-Ary [440]. -BM [91]. -Calculus [290, 797]. -Channel [624]. -D [872, 1273, 842, 1201]. -Dimensional [642]. -Gram [94]. -Grammar [846]. -Interval [81].

-Matching [91]. -Means [1487]. -Median [917]. -NN [501]. -out-of- [1181]. -Path [1236]. -Reduction [290]. -SAT [599]. -Trees [640]. -Truck [648]. -Uniform [426]. -Vertex [652].

.NET [1119].

/s [968].

1.0} [777]. 11monkeys3 [218].

2-Compromise [642]. 2-SAT [651]. 2.0 [785]. 2001 [1318, 280, 1321, 1339, 284, 221]. 2SAT [569].

3-Pass [541]. 3-SAT [651]. 3-Secure [535]. 3T [219].

4-Slot [748].

5-Regular [620]. 5dpo [238]. 5P [703].

802.11e [1254].

= [1362, 396].

A-Prolog [1391]. AAFG1 [525]. ABA [1001]. ABBA [1484]. Abbreviation [984]. Abduction [1366, 1388, 1361, 1399, 1397]. Abelian [445]. Abstract [1419, 1102, 50, 621, 611, 1547, 1418, 617, 66, 638, 650, 626, 32, 1103, 601, 455, 44]. AbstractCollection [590]. Abstracting [46, 39]. Abstraction [47, 49, 757, 51, 28, 1096, 30, 21, 23, 738, 747, 45]. Abstraction-Based [1096]. Abstractions [117, 43, 54, 1446, 55]. Academic [1076, 1075]. ACAS [577]. Accelerating [988]. Acceleration [1413]. Accelerometers [1562]. Acceptability [1359]. Acceptable [1252]. Access [1077, 1066, 1171, 1061, 1553, 1216, 1026, 1078, 1227, 933, 1028, 1068, 1328, 1228, 1002, 1017, 164, 1060, 1022, 996, 936, 1055, 1009, 976]. Accesses [1468]. Accessibility [1030, 939, 943, 1021, 937, 1053, 944, 1008]. Accessible [957, 1006, 1012, 995, 1069, 1081, 991, 994]. Accessing [715, 980, 1064, 1299, 959]. Accommodating [605]. Accumulating [63]. Accumulation [887]. Accumulators [928]. Accuracy [1637]. Accurate [184, 1036, 1730]. Achieve [943]. Achievements [1363]. Achieving [248]. Acoustic [1742]. Acquisition [549, 722]. Across [1339, 809, 1546]. Act [1065]. Acting [1243]. Action [1094, 1613, 33, 250]. Actions [1700, 1536, 44]. Activation [178]. Activation/Termination [178]. Active [468, 9, 1307, 645, 1713, 536]. Activity [1608, 1576, 1750, 1575, 674]. Activity-Dependent [1576]. Actuator [1037]. ad [1241, 372, 1240]. Adaptability [1706]. Adaptable [325]. Adaptation [1598, 108]. Adapted [1058]. Adapting [13, 1257]. Adaptive [1783, 1733, 1757, 972, 1012, 356, 1642, 1762, 1669, 1293, 1641, 1777, 1255, 950, 1471, 981, 1501, 1289, 1736, 1211]. Addition [542]. Addition-Subtraction [542]. Additional [181]. Addizionario [958]. Address [1175]. Adjusting [1135, 1414]. Administration [533]. Admissibility [1426]. Admissible [25]. Adolescents [987]. Adopting [319]. Adpated [837]. Advances [669]. Advantages [703]. Adverbs [711]. Adversaries [547]. Adversary [177]. AES [529]. Affine [1145, 900, 503]. Affordance [944]. against [546, 536]. Agent [72, 1281, 1476, 1479, 1542, 1531, 1288, 1484, 1289, 224, 1278, 1538, 165, 1539, 1285, 1280, 186, 1282, 1283, 560, 1541, 282, 1537, 1287, 1477, 1481, 1490]. Agent-Assisted [1476]. Agent-Based [1281, 1542]. Agent-Mediated [1278]. Agents [1530, 1354, 544, 1276, 1284, 1372, 1286, 201, 1535, 1373, 1544, 1475, 823, 953, 171, 172, 1725, 281, 1543, 1540, 7, 1478, 1489, 1740, 1483]. Aggregates [1447, 1380]. Aggregation [352]. AGILE [1476]. Agility [248]. AGILO [250]. Agnostic [997, 127]. Aid [1019]. Aided [979, 1018]. Air [1735]. Alarm [1056]. ALDMEISTER [801]. Algebra [571, 550, 577, 401, 576, 593]. Algebraic [580, 525, 388, 32, 1436, 362]. Algebras [1100]. Algorithm [1232, 1212, 1514, 1155, 1679, 135, 1626, 1487, 1757, 1240, 1181, 1521, 663, 887, 1308, 1136, 917, 879, 1159, 1467, 528, 442, 871, 1245, 1525, 1655, 1257, 1250, 637, 1251, 685, 1734, 1433, 674, 916, 504, 1410, 348, 1405, 641, 1182, 1217, 1201, 1738, 1448, 1764, 1144, 1625, 1256]. Algorithm-Based [674]. Algorithmic [737, 783]. Algorithms [1130, 570, 357, 185, 853, 378, 91, 1406, 1198, 584, 1782, 847, 603, 1235, 1669, 567, 1213, 88, 604, 640, 615, 147, 148, 1631, 1180, 1156, 624, 1158, 919, 1413, 1746, 634, 623, 425, 1218, 1199]. Aligned [1325]. Aligning [76]. Alignment [606]. Alive [913]. All-Pairs [348]. All-to-All [1217]. Allocation [283, 1405, 1533]. Allowable [1252]. Allowing [95]. Aloud [992]. Alpha [1527, 1029, 524]. Alpha-Beta [1527]. Alternate [1298, 1082]. Alternating [411]. Alternative [1534, 1392, 1067]. Alzheimer [555]. Ambiguity [714, 397]. AMEEDA [1137]. Amenable [1626]. American [1066]. AMIS [1012]. Amit [468]. Among [217, 106, 823, 1076]. Analog [1777]. Analysing [509]. Analysis [890, 1573, 1091, 432, 803, 498, 1125, 1647, 1582, 1632, 835, 1744, 1749, 1651, 1102, 671, 1677, 697, 1307, 1093, 490, 1045, 857, 1597, 749, 885, 813, 930, 1627, 528, 1210, 290, 676, 1505, 1595, 1665, 1708, 1433, 674, 1120, 287, 1336, 747, 919, 731, 1596, 1746, 1326, 61, 335, 523, 806, 1760, 1163, 1532, 1787, 1113, 1579, 830, 334, 512]. Analytic [445, 446]. Analyze [192]. Analyzers [1716]. Anaphora [693, 700]. Anaphoric [705]. AND/OR [621]. Animated [953]. Anisotropic [477]. Annealing [1633, 1753]. Annotating [819]. Annotation [1566, 903, 725]. Annotations [85, 327, 311]. Anomaly [673]. Anonymous [613]. Answer [64]. Answering [459, 449]. Antenna [1249, 1261]. Anthropological [1282]. Antirandomizing [415]. Any [1643]. AOP [101]. Aperiodic [1575]. API [1189]. APL [1329]. Appearance [666, 690]. Appearance-Based [666, 690]. Applets [1187]. Applicable [896]. Application [1165, 1767, 989, 92, 713, 316, 1203, 871, 635, 883, 681, 1542, 866, 1471, 66, 1758, 1484, 1502, 108, 78, 1019, 165, 107]. Application-Engineering [316]. Application-Oriented [1542]. Applications [1124, 464, 303, 145, 675, 1126, 1042, 374, 1548, 89, 409, 1227, 1160, 1224, 948, 1229, 1667, 457, 1196, 1556, 1218, 762, 15, 428, 334, 1137, 534, 1123]. Applied [1743, 1616, 1778, 1027, 1773, 690]. Apply [1135]. Applying [1582, 1652]. Approach [1779, 876, 1390, 780, 737, 1051, 1282, 1102, 112, 1233, 680, 109, 553, 1677, 870, 1491, 1661, 923, 857, 1642, 1042, 1613, 952, 1360, 282, 545, 1698, 201, 1432, 1059, 204, 1618, 1665, 1526, 983, 210, 388, 1692, 1434, 920, 918, 1472, 914, 1178, 1315, 767, 1300, 2, 165, 568, 613, 926, 888, 321, 1456, 205]. Approaches [331, 1494]. approximability [1133, 1134]. Approximate [95, 1294, 93, 1104, 414, 34, 1533]. Approximated [917, 845]. Approximating [412, 639]. Approximation [392, 48, 423, 1659, 1658, 625, 31, 1691, 1643, 78, 37, 38]. Approximations [346, 43]. APT [1733]. APT-Based [1733]. Arab [1016]. Arabic [856]. ARAIBO [278]. Arbitrage [1493]. Arc [85]. Archeological [500]. Architecting [1023]. Architectural [114, 1787]. Architecture [1247, 1561, 1530, 733, 1220, 824, 1613, 1541, 1167, 615, 1173, 1531, 208, 3, 1000, 219, 1224, 224, 1769, 324, 1179, 220, 1256, 1489, 322]. Architecture-Level [3]. Architectures [1124, 1150, 1231, 612, 323, 17, 1218, 339]. Archival [491]. ArchJava [114]. Arcs [867]. Area [1099, 1058, 1244]. Areas [1586]. Argument [1389]. Argumentation [1387, 1278]. Argumentation-Based [1387]. Arising [1097]. Arithmetic [950, 1614]. Arity [574]. ARMA [1747]. armed [144]. Arrangements [430]. Array [1037]. Arrays [615, 1249, 637, 1142]. Artefacts [331]. Arthur [598]. Artificial [1751, 1499, 1696, 1279]. Artisti [251]. ARTMAP [1639]. Ary [440]. Aspects [565, 783]. ASR [718]. Assembler [749]. Assembling [628]. Assembly [8, 595, 19, 341]. Assertions [1117]. Assessing [960]. Assessment [1751, 1086, 1203, 1368, 727, 988, 976]. Asset [338]. Assignment [1475].

Assignments [651]. Assist [975]. Assistance [1014, 989, 971]. Assisted [987, 990, 1476]. Assistive [1087, 951, 1086, 978, 1082, 1083, 1084, 1079, 1057, 976]. Associated [476]. Association [1171, 836, 1769, 1304]. Associative [1492, 1577, 827, 831, 1614]. Assumptions [530]. ASURA [266]. Asymmetric [883, 530]. Asymptotic [1720]. Asymptotically [1136]. Asynchronous [1184, 1608, 748, 1195, 547, 623]. ATM [1257]. Atomic [307, 105, 1166]. Atoms [783]. AtomsNet [484]. Attack [525, 542]. Attacks [1370, 536]. Attention [875, 1569, 170, 873, 1774]. Attentional [1768]. Attitudes [1066]. Attractor [1600]. Attributes [1515, 1446]. ATTUnited [221]. ATTUnited-2001 [221]. Auction [1533]. Auctions [1278]. Audio [498, 499, 1039, 1763, 670]. Audio-Visual [499, 1763]. Auditing [939]. Auditory [1023, 1018, 1017]. Augmentative [1067]. Augmented [366, 1321]. Austrian [1076]. Austrian-Wide [1076]. Authenticated [537]. Authentication [1555, 677]. Authentification [869]. Author [20, 60, 1246, 1275, 1291, 1312, 1350, 1375, 1400, 1416, 1439, 1529, 1545, 1568, 1790, 75, 99, 126, 154, 285, 315, 340, 431, 456, 469, 510, 548, 579, 594, 658, 691, 728, 763, 804, 834, 841, 932, 1088, 1105]. Authorization [532]. Auto [1492]. Auto-Associative [1492]. AutoBayes [788]. AutoBayes/CC [788]. Automata [411, 723, 655, 1620, 628, 355, 367, 368, 830, 656]. Automated [436, 709, 1445, 1388, 1010, 730, 1502, 1114, 489, 557]. Automatic [1785, 1775, 1126, 911, 1566, 903, 1660, 1433, 693, 722, 816, 914, 1540, 1132, 107, 788]. Automation [28, 1260]. Automed [1297, 1295, 1296]. Automotive [1547]. AutoMouse [1030]. Autonomous [112, 166, 1488, 54, 174, 1540, 1703]. Autoregressive [1760]. Autowrite [312].

Auxiliary [734]. Availability [1185]. Average [383]. Averages [128]. Avoidance [175]. Aware [72, 15]. Awareness [1069, 1563]. Axiomatic [289]. Axiomatising [389]. Axioms [358]. Axis [146].

B [730, 396, 760]. B2B [534]. BabyTigers [267]. Back [1552, 1765, 650]. Back-Propagation [1765]. Back-Walk-Free [650]. Backward [477]. Bagging [689]. Balancing [1128, 1161]. Banach [422]. Banach-Mazur [422]. band [718]. Bandit [144]. Bandwidth [1210]. Barrel [674]. Barrier [1216]. Barriers [1072]. Base [1483]. Based [890, 1232, 1644, 1281, 1788, 780, 931, 288, 1761, 1214, 1704, 1140, 185, 669, 1514, 507, 1744, 696, 1325, 1241, 239, 156, 666, 717, 875, 436, 1733, 1493, 1330, 1555, 1677, 694, 1181, 1126, 490, 1231, 192, 1387, 479, 482, 113, 1732, 887, 1216, 771, 1604, 1729, 904, 1294, 859, 1360, 1298, 454, 930, 905, 844, 1383, 1078, 555, 1116, 871, 1473, 684, 907, 1096, 151, 102, 118, 1594, 1525, 736, 1373, 360, 681, 499, 538, 337, 1542, 478, 1364, 1265, 208, 487, 683, 1244, 1656, 12, 678, 842, 86, 699, 1056, 674, 823, 170, 116, 965, 1384, 1606, 873]. Based [765, 918, 504, 1484, 1076, 1715, 1221, 62, 1472, 1299, 19, 496, 1667, 486, 1182, 78, 1326, 840, 250, 174, 175, 1540, 1774, 922, 492, 123, 2, 1452, 695, 1506, 165, 623, 1532, 1533, 1742, 568, 1463, 44, 1722, 1211, 215, 473, 926, 321, 481, 687, 334, 1457, 1512, 1676, 119, 1456, 1489, 1740, 1668, 493, 1401, 910, 690, 329, 912]. Based-On [1514]. Bases [815]. Basic [570, 439, 252, 950, 800, 345, 401]. Basis [1624, 1621, 1701, 756, 1668]. Basketball [499]. Baum [1759]. Baum-Welch [1759]. Bayes [151, 1676]. Bayes-Mixtures [151]. bayesian [848, 915, 904, 1665, 505]. BCI [1752]. BDD [765]. BDD-Based [765]. BDI [1290]. Be [790, 997]. Beaver [1484].

Bed [165]. Behalf [1243]. Behavior [1704, 222, 208, 178, 215]. Behavior-Based [208]. Behavioral [124]. Behaviors [269]. Behaviour [192, 224, 108]. Behaviours [72, 13]. Belief [446, 825]. Beliefs [1606]. Bend [646]. Benefit [1563]. Best [1399]. Beta [1527]. Better [92, 122, 521]. Between [1458, 507, 143, 952, 773, 1550, 542, 1538, 1539]. Beyond [10, 1749, 933]. Bi [1325, 771]. Bi-directional [1325]. Bialgebraic [384]. Bias [1787]. BiBa [521]. BIKMAS [1504]. Bilinear [538]. Bilingual [1319, 1333, 1317, 1332, 1323, 1327]. Bill [1742]. Bimodal [446]. Bin [366]. binary [52]. Binding [1638]. Bins [644]. Bio [1056]. Bio-Sensor [1056]. BioGRID [1229]. Bioinformatics [839, 1504]. Biologically [1573, 923]. Biomedical [720, 886]. Biomoleculago [1229]. Biophysical [1594]. Bipartite [1274, 425]. Biped [1694, 161]. BISECTION [392]. Bisimilarity [401]. Bisimulation [746]. BIST [1266]. Bit [1447, 122, 93]. Bit-Parallel [93]. Bit-Sliced [1447]. Bit-Vector [122]. Biter [188]. Black [1402]. Black-Box [1402]. Blackmailing [545]. Blind [1031, 1044, 1045, 1042, 1026, 1024, 1756, 1004, 1010, 998, 1080, 1028, 1040, 1041, 1034, 1062, 1022, 1035, 1001, 1009]. Blobs [200]. Block [97, 79]. Blocking [475, 1201]. Blocks [1656]. BlueGene [1131]. BlueGene/L [1131]. Blues [1615]. Blur [1767]. BM [91]. Bob [1352]. Böhm [297]. Boltzmann [1626, 382, 1614, 1166, 1625]. Bonkers [1059]. Books [933]. Boolean [40, 570, 780, 443, 789]. Boosting [147, 149]. Bound [362, 1159, 1467, 855]. Boundaries [1178]. Boundary [730]. Bounded [369, 1133, 392, 372, 359, 1286, 370, 358, 653, 601, 623, 798]. Bounded-Depth [358]. Bounds [131, 145, 393, 366, 144, 74, 1663, 1098, 391, 305, 142]. Box [1402]. Braid [525]. Braille [1050, 1033, 1049, 1051, 1048, 1002]. Brain [1750, 1505, 1586]. Brains [230]. Brainstormers [210]. Branch [370, 1159, 1467]. Branching [369]. Brazilian [719]. Breadth [1403]. Breadth-First [1403]. Bridge [1339]. Bridging [952, 448]. Brief [539]. BrlGraphEditor [1033]. Broadcast [516, 1216, 624, 672, 613]. Broadcasting [617, 1271, 631]. Broadcasts [490]. Browser [405, 972, 1469, 1023, 1003, 1565]. Browsing [471]. Bruijn [617]. Buffering [1258]. Builder [1469]. Building [1006, 819, 827, 1548, 207, 1, 821, 707]. Built [1263]. Built-in [1263]. Bulk [1134]. Bundling [26]. Burst [1596]. Bus [576, 774]. Business [935]. Busy [1484]. Button [985]. Buyers [1538]. Byte [745]. Bytecode [1421].

c [1153, 409, 968, 117, 1321]. C-E [1321]. c-Means [1153]. Cache [1269, 357, 1140, 376, 685, 31, 1178]. Cache-Oblivious [357]. Caches [1174]. Caching [1406, 443]. CADP [751]. Calculating [1447, 1328]. Calculations [848]. Calculi [752, 440, 445, 292]. Calculus [571, 436, 1425, 291, 298, 290, 797, 444, 1392, 447, 1723, 453]. Calibration [184]. Call [407, 1126, 1617]. Call-Graph [1126]. Calls [1252]. Camera [184, 1555]. Camera-Based [1555]. Can [790, 973]. Candidates [891]. Canonical [306]. Capabilities [1629, 1068]. Capability [1230]. Capacitated [1159]. Capacities [1620]. Capacity [1577]. Captioning [993]. Capture [1126]. Card [745, 1495]. Carrying [766, 767]. Cartographic [881]. Case [1370, 1549, 739, 732, 167, 551, 590, 317, 101, 116, 938, 1399, 1082, 840, 1480, 1463]. Case-Based [840, 1463]. Case-Study [739]. Cases [1412]. Cat [923]. Catalogues [1536]. CATCH [1126]. Categorical [1712, 1687]. Categorization [709]. Causes [1671].

CBIR [491]. CBR [1219]. CC [788]. CC4J [11]. CDMA [1273, 1249]. cdma2000 [1262]. Cell [1582, 1579]. Cells [1603, 1597, 1574, 111]. Cellular [1620, 367, 1643]. Centered [940]. Central [1595, 1596]. Centralised [1485]. Centric [1189, 457]. Century [1039]. Cerberus [268]. Certificates [1070]. Certification [788]. Certified [753]. CESR [246]. CFA [262, 243]. CFD [1202]. CG [820, 811, 817]. CGIF [811]. CGLF [811]. CGWorld [824]. Chain [1253]. Chains [542]. Challenge [1165, 839, 999, 463]. Challenged [1011]. Challenges [1110, 470, 1090, 492, 941, 1085]. ChaMeleons [223]. ChaMeleons-01 [223]. Champion [160]. chan [193]. Chance [954]. Change [1008]. Changing [1078]. Channel [624, 542]. Channels [1249, 762]. Character [898, 563, 1274, 1518, 922]. Characterisation [929]. Characteristic [875]. Characterization [30, 1595, 1596, 913]. Characterize [630]. Characterizing [380, 1152]. Characters [856, 985, 1669]. Cheating [519]. Checker [740]. Checkers [1421]. Checking [744, 300, 312, 395, 1096, 758, 21, 1098, 1099, 66, 32, 795, 798]. Checkpointing [1244]. Chemical [1502]. Children [970, 972, 974, 973, 971, 969]. Chinese [1316, 987, 1321, 1518]. Chip [1223]. Choice [997, 1289]. Chordal [425]. Chromosome [1660]. Church [353]. Cipher [522, 524]. Circuits [1585, 214]. Circular [1765, 430]. Circus} [754]. City [688]. Class [921, 120, 784, 590, 102, 912, 683, 894, 626, 123, 530, 662, 1690]. Class-Dependent [921]. Class-Discriminative [912]. Class-Independent [102]. Classes [598, 755, 526, 127]. Classification [899, 1441, 1427, 1521, 673, 748, 684, 561, 1507, 1755, 866, 1656, 1500, 148, 1660, 1662, 596, 891, 1703, 1518, 886, 662, 1448, 670, 926, 1456, 1742]. Classifier [890, 1508, 697, 885, 1495, 683, 924].

Classifier-Independent [890].

Classifier/Proper [697]. Classifiers [1634, 1783, 710, 1689, 925, 1509, 922]. Classify [1633]. Clausal [438, 1367, 784]. CLDS [1381]. CLEF [1319, 1349, 1314, 1320, 1333, 1322, 1317, 1318, 1321, 1334, 1328, 1339, 1329, 1332, 1342, 1340, 1327, 1315, 1337, 1348]. CLEF-2001 [1318, 1334, 1315]. Clickstream [835]. CLIOH [500]. Cliques [652]. CLIR [1331, 1325, 1330, 1347]. Clock [1118]. Clockwork [253]. Closed [373]. Closing [758]. Closure [294]. Cluster [1150, 616, 1487, 1185, 1264, 1255, 1118, 1709]. Clustered [1173]. Clustering [1630, 1487, 150, 887, 1460, 1658, 485, 1525, 854, 1153, 1303, 849, 918, 504, 1413, 1667, 1580, 1465, 1722, 1512, 1653]. Clusters [1133, 1692, 1201, 1137, 1152]. CM [239, 269]. CM-Dragons’01 [239]. CM-Pack’01 [269]. CMOS [182]. CMU [254]. CNF [778]. Co [1419, 611, 1321, 1443]. Co-evolutionary [1443]. Co-occurrence [1321]. Co-orthogonal [611]. Coach [233, 173]. Coaching [173]. Coarse [1143, 1224]. Coarse-Grain [1143]. CODACS [1179]. Code [739, 1127, 766, 745, 106, 749, 1210, 1273, 11, 80, 767, 788, 753]. Codebooks [56]. Codes [412, 611, 360, 1261, 535]. Codification [1620]. Coding [495, 1597, 709, 1519, 1772]. Coefficient [561]. Coevolution [926]. Cognitive [190, 1611, 1052, 1740]. Coherence [1574, 1178]. Collaboration [1083]. Collaborations [103]. Collaborative [1300, 1481]. Colleges [1082]. Collision [1245, 625]. Collocation [696]. Color [881, 182, 666, 681, 1520, 480, 477, 200]. Coloring [603, 1235, 567]. Colour [1422, 495, 860, 905]. Columnsort [1411]. Combination [222, 792, 897, 863, 1766, 924]. Combinatorial [1237, 53, 128, 1402, 613]. Combined [1315]. Combiners [885].

Combining [1419, 310, 1508, 666, 751, 710, 770, 575, 286, 1583, 788, 793, 453]. Commands [1011]. Commerce [1469, 1479, 1471, 713, 1472]. Commercial [741]. Committee [1636]. COMMON [85, 853]. Communication [1130, 632, 1133, 1589, 1037, 1483, 1553, 1043, 1206, 282, 748, 940, 1242, 352, 1466, 986, 1564, 624, 414, 1194, 1217, 1067, 946, 1285]. Communications [1186]. Communicative [1536]. Community [1082]. commutative [563]. Comon [1749]. Compact [602, 869, 527]. Compactly [1685]. Comparable [1324]. Comparative [1305, 1494, 1174, 1268, 690]. Comparing [381, 1080, 1318, 1586, 861, 1345]. Comparison [183, 1014, 853, 643, 1429, 1782, 885, 465, 1496, 1518, 886, 670]. Comparisons [1569]. Compatibility [124, 848]. Compensate [968]. Competencies [981]. Competitions [167]. Competitive [1724, 1136, 1713, 1142]. Compilation [125, 1138, 1173, 756]. Compiler [768]. Compilers [1138]. Complementarity [597]. Complete [419, 420, 878, 396]. Completeness [799, 441, 1385]. Completion [1681]. Complex [1582, 723, 1603, 807]. Complexities [130, 134]. Complexity [90, 131, 550, 359, 351, 1498, 386, 629, 797, 1718, 342, 1156, 414, 892, 133, 38]. Component [468, 10, 669, 1744, 1749, 1651, 6, 1116, 684, 1505, 1364, 16, 3, 14, 5, 1230, 1769, 1716, 334, 119, 18]. Component-Based [669, 684, 1364, 334]. Components [13, 325, 1197, 1708, 1, 19, 335, 110, 1742, 902]. Composable [103]. Composition [746, 1546, 716, 123, 301, 336]. Compound [711, 1322]. Comprehension [465]. Compressed [637, 1403]. Compression [56, 78, 79]. Compromise [642]. Compromises [1371]. Computable [422, 601, 141]. Computation [872, 381, 551, 520, 1431, 374, 555, 361, 547, 341, 1672, 557, 1157]. Computation-Based [555]. Computational [1351, 553, 386, 1250, 70, 1356, 1610, 342, 1462, 845, 1389, 18]. Computations [1143, 1106, 391]. Computer [1016, 571, 185, 1030, 987, 1555, 673, 1167, 1398, 973, 979, 981, 945, 1019, 941, 806, 582, 976, 576]. Computer-Aided [979]. Computers [1200, 1160, 969]. Computing [72, 1549, 1108, 1226, 614, 537, 778, 1559, 1547, 1548, 1551, 1557, 1244, 845, 597, 1560, 1556, 1172, 1562, 1238]. Concept [829, 819, 813, 809, 806, 833, 830]. Concepts [146, 819, 827, 136, 831, 1111, 39]. Conceptual [805, 1301, 823, 814, 826, 812]. Concerns [322]. Concluding [718]. Concurrency [117, 1094, 101]. Concurrent [1119, 1357]. Condition [1623]. Conditional [647]. Conditioned [437]. Conditioning [1705]. Conditions [301]. Conductances [1607]. Conducting [1538]. Confidence [1168]. Configurable [226]. Configuration [1560]. Conflicts [532]. Confluent [447, 305]. Congestion [1214]. Conjectures [572]. Conjugate [1786]. Conjunctions [359]. Conjunctive [66]. Connected [853, 618]. Connection [771, 1777, 447, 1385]. Connection-Based [771]. Connectionist [1619, 1612]. Connections [583]. Connector [12]. Considering [196]. Consistency [148]. Consistent [147]. Consonant [672]. Consonant-Vowel [672]. Consortium [933]. Conspiracy [1162]. Constant [361]. Constant-Round [361]. Constantly [1078]. Constrained [1130, 1488, 3, 2]. Constraint [1440, 375, 614, 549, 441, 374, 53, 1395, 566, 552, 1357, 1369, 59]. Constraints [307, 550, 1308, 52, 567, 42, 344, 1759, 178, 1380]. Construct [509]. Constructing [588, 98, 637, 724]. Construction [802, 701, 741, 28, 1151, 770, 355, 1725]. Constructions [1425]. Constructors [310]. Constructs [62]. Content [288, 507, 1567, 957, 490, 479, 904, 871, 1566, 1255, 504, 496, 492, 493]. Content-Adaptive [1255]. Content-Based [288, 507, 479, 904, 871, 504, 496, 492, 493]. Contents [1469, 716]. Context [1768, 790, 82, 827, 1675, 1559, 1523, 1544, 306, 938, 397, 833, 1288]. Context-Free [397]. Context-Rich [1544]. Context-Sensitive [790, 306, 938]. Contexts [559, 828, 17, 575]. Contextual [990, 915, 1460, 860, 808]. Continuous [1788, 706, 1626, 1732, 1515, 1264, 1745, 566, 925, 1543, 707]. Continuous-State [1788]. Continuum [37]. Contract [2]. Contract-Based [2]. Contracting [391]. Contractions [1049]. Contrast [1689]. Contrastive [1637]. Control [407, 1768, 1214, 1171, 749, 194, 757, 1216, 1485, 1604, 585, 1698, 1695, 202, 1270, 1257, 352, 1542, 1265, 43, 170, 978, 1585, 1178, 962, 1771, 214, 161]. Control-Flow [749]. Controlled [1162]. Controller [553]. Controllers [197, 1706]. Convergence [1212, 1734]. Conversation [467, 717]. Conversion [692]. Convex [131, 1157]. Convey [1025]. Convolutional [1754]. Convolutions [1727]. Cooperation [215]. Cooperation-Based [215]. Cooperative [459, 168, 206, 821, 174]. Coordinating [1276, 1475, 1716]. Coordination [156, 1534, 50, 1277, 1570, 1243]. CoPS [255]. CoPS-Team [255]. Copyright [999, 535]. Coq [1420, 1430, 1438]. CORBA [16]. Core [1411, 1097]. Corel [474]. Corpora [721, 725]. Corpus [1325, 1324, 672]. Correct [1155, 70]. Correcting [1256]. Correction [1502]. Correctness [182, 741, 1433]. Correlated [1608, 1605]. Correlation [1731, 674, 539, 924]. Correlations [931]. Correspondence [387]. Correspondences [900]. Cortex [1613, 674, 1579]. Cortical [1603, 1591, 1609]. Corticospinal [1604]. COS [522]. Coscheduling [1135]. Cost [189, 196, 1467, 678, 845, 1118, 889, 1747, 1123]. Cost-Sensitive [678]. Costs [1408, 1207]. COTS [1118]. Counter [770, 656]. Counter-Model [770]. Countermeasure [542]. Counting [651, 358]. Coupled [477, 1581]. Coupling [1210, 683]. Courses [980]. Covariance [893]. Cover [426]. Coverability [731]. Coverage [11]. Covering [414]. Coverings [652]. cp [1664]. CPG [1585]. Crawler [462]. Crawling [1470]. CREAM [1299]. Create [974]. Creating [1047, 1221]. Credit [1495]. Credulous [1387]. Criteria [954]. Criterion [170]. Critical [1751, 621, 1389, 321]. Critics [1429]. Critique [1382]. CrocaRoos [234]. Cross [1684, 1321, 1343, 1324, 1326, 1707]. Cross-Language [1321, 1343]. Cross-Lingual [1324, 1326]. Cross-Point [1707]. Cross-Validation [1684]. Crossing [620]. Crossover [1631]. Cryptanalysis [522, 524]. Cryptographic [360]. Cryptographically [746]. Cryptosystem [525]. Cryptosystems [612, 527]. CS [168, 203, 158]. CSL [1098]. CSP [392, 738]. Cue [903]. Cues [670]. Culture [319, 1078]. cum [1267]. Cumulant [1744]. Cuneate [1601]. Current [1592]. Current-to-Rate [1592]. Curricula [582]. Currying [308]. Curvature [863]. Curve [528, 610, 527]. Curves [494, 526, 634]. Customizable [1]. Customization [1566]. Customized [71]. Cut [434, 1426, 914]. Cut-Elimination [434]. Cuts [619]. CV [672]. Cyberoos’2001 [224]. Cycle [1227, 397]. Cycle-Free [397]. Cycle-Stealing [1227]. Cyclic [985]. Cylindrical [639]. CYNTHIA [1003].

D [872, 1761, 1036, 1273, 1251, 842, 1039, 1201, 639, 493]. D-Meshed [1251]. DAISY [933]. Data [781, 1633, 1748, 371, 1647, 518, 1193, 1731, 1687, 755, 399, 1677, 1569, 843, 383, 1305, 1297, 726, 839, 398, 1628, 1461, 589, 836, 1473, 740, 1151, 1517, 1444, 896, 1262, 1745, 676, 1474, 1147, 1563, 1153, 1253, 1302, 1303, 894, 564, 345, 1148, 1446, 1154, 1709, 918, 892, 919, 1395, 1667, 596, 1450, 1178, 1443, 1442, 138, 1310, 1448, 1179, 1638, 1653]. Data-Parallel [1193]. Database [1311, 712, 483, 1296, 1451, 1149, 1177, 463]. Databases [715, 459, 696, 642, 810, 836, 1147, 682, 1300, 1380, 493]. Datalog [1378, 62]. Datasets [1507]. DAWG [89]. Day [1496]. DBMS [1302]. DCGs [1396]. DCT [482]. DCTP [455]. DDDLIB [772]. De-noising [1688]. De-warping [877]. Dead [694]. Dead-End [694]. Deaf [992]. Deafblind [1043]. Debugger [1117]. Debugging [65, 346]. Decidability [294, 396]. Decidable [1381, 349, 309, 784]. Deciding [770, 410]. Decision [856, 921, 310, 1240, 1511, 192, 144, 792, 904, 1650, 27, 1151, 337, 443, 1513, 1578, 236, 873, 765, 1442, 1532, 1636, 1512, 1152, 45]. Decision-Making [337]. Decision-Support [1152]. Decisions [29, 864]. Declarative [69, 1379]. Decoding [680, 360, 1606]. Decomposition [1507, 635, 1707]. Decompositions [1294]. Decorrelation [1757]. Decoupled [1739]. Dedicated [1137, 1135]. Deduction [436, 551, 1431, 66, 1396]. Deductive [781, 1380]. Deep [224, 1114]. Defeating [545]. Deficient [939]. Deficits [968]. Definite [725]. Deflection [1212]. Deforesting [63]. Degenerate [1412]. Degradable [615]. Degree [392, 1079]. Delay [1252]. Delayed [1155, 1197, 1571]. Delayed-Initiation [1155]. Delays [1587]. Delegation [1287, 103]. Delegations [532]. Delivering [1004]. Delivery [1087, 1543]. Demand [1220, 585, 1161, 1179]. Demand-Data [1179]. Demand-Driven [1161]. Demonstrations [341]. Demonstrative [705]. Demonstrator [494]. Dendrites [1713]. Density [1720, 1495, 1649]. Depend [1631]. Dependable [632, 747]. Dependence [1109, 1486]. Dependencies [330]. Dependency [1294]. Dependent [921, 757, 1576, 1523, 1713, 1436]. Deployment [3, 2, 7, 15, 18]. Deposition [1166]. Depth [358]. Derivation [584, 1362, 557]. Derivational [305]. Derived [1435]. Deriving [90, 746, 1100, 292]. Descendants [295]. Descent [212, 1786]. Describing [762]. Description [1779, 268, 856, 278, 230, 276, 227, 262, 243, 233, 1377, 223, 238, 231, 712, 242, 764, 270, 225, 271, 776, 157, 255, 226, 259, 236, 232, 284, 786, 247, 254, 218, 788, 264, 228, 774, 235, 902, 244]. Descriptions [1031, 725]. Descriptive [584]. Descriptor [487]. Design [1232, 737, 1212, 1204, 188, 190, 553, 972, 423, 1012, 193, 940, 1270, 1455, 1531, 1058, 981, 966, 116, 171, 1011, 1236, 281, 457, 1536, 324, 214, 215, 462]. Design-Centric [457]. Designing [1038, 1030, 940, 1544, 163, 1196]. Designs [1146]. Despite [1571]. Destructive [61]. Detail [814]. Detect [1697]. Detecting [875, 688, 1692]. Detection [179, 1143, 1492, 669, 1752, 1782, 673, 908, 900, 625, 679, 681, 1495, 1754, 678, 209, 1770, 1601, 914, 1089, 506, 687, 1581]. Detector [1736]. Detectors [903, 1268]. Determined [507]. Determining [1521, 1308]. Deterministic [1219, 656, 139, 375, 62, 1182]. Developing [189, 1534, 71, 986, 1076]. Development [876, 1281, 745, 719, 1398, 741, 1190, 1473, 979, 338, 1272, 1364, 1260, 1609, 969, 775, 760, 216]. Developments [1016, 560]. Device [1014, 1032, 967, 1052]. Devices [537, 981, 14, 1057, 988, 1562]. Diagnosis [1483, 555, 30]. Diagram [1300]. Diagrams [1038, 1590, 27]. Dialog [714]. Dialogue [715]. Dialogues [1538]. Diatoms [850]. Dictionaries [1009]. Dictionary [609, 1330]. Dictionary-Based [1330].

Difference [102, 772]. Difference-Based [102]. Differences [1631]. Different [155, 906, 1586]. Differential [1757]. Differentially [1662]. Differentiation [1433]. Difficulties [972]. Diffie [538]. Diffusion [1593, 477]. Digit [897]. Digital [968, 1783, 995, 500, 998, 933, 991, 993, 492, 534]. Digraphs [617]. Dimension [152, 894]. Dimensional [1748, 495, 1731, 642, 1522, 95, 1465, 596, 1654, 1142, 1691]. Dimensionality [677]. Dimensions [633, 387, 128]. Direct [777]. Direction [1594]. directional [1325, 209, 175]. directional/Pin [179]. Directions [1353, 1786]. Dirty [233]. Disabilities [987, 1081, 981, 1008, 934, 1070, 1075, 936, 959, 962, 984, 976]. Disabled [1061, 942, 999, 973, 940, 1005, 1068, 986, 971, 1076, 938, 948, 969, 996, 941, 977]. Disaffected [961]. Disagreement [145]. Disambiguation [1325, 1538, 695]. Disappearing [941]. Disaster [281, 165]. Disconnection [444]. Discontinuity [704]. Discourse [1616]. Discourses [838]. Discover [1443]. Discovering [1277, 1444]. Discovery [77, 1561, 1282, 1461, 836, 1147, 54, 1454, 1452]. Discrete [1522, 1234, 1745, 610, 344, 55, 1759]. Discretization [1515, 33]. Discriminant [890]. Discriminate [879]. Discrimination [923, 1601]. Discriminative [1245, 912]. Disease [555]. Disenrollment [516]. disjoint [453]. Disjunctive [1368]. Disorder [975]. DISPARA [721]. Disparity [1771, 874]. Display [1426, 1032, 967]. Displayed [1765]. Displays [1002]. Dissimilarity [892]. Distance [609, 1140, 1241, 94, 1697, 1235, 604, 893, 1210, 1584, 851, 84, 1071, 1722]. Distance- [1235]. Distance-Based [1140]. Distance-Hereditary [604]. Distinguishability [542]. Distinguishable [847]. Distinguished [1203, 1239]. Distinguishing [403]. Distortion [912].

Distributed [1589, 464, 1108, 511, 6, 1220, 4, 1483, 1485, 1239, 1183, 1147, 1265, 1557, 1180, 100, 1000, 1299, 1089, 1113, 462, 774]. Distributing [721, 1187]. Distribution [862, 92, 376, 1567, 1655, 1255, 1635, 867]. Distributions [495]. Distributive [448]. Disturbances [1129]. Divergence [1637, 389]. Diversity [1261]. DMSumm [727]. DNA [614, 1662, 597]. DNF [657]. Do [735, 577]. Do-It-Yourself [577]. Doacross [1141]. Document [899, 1311, 857, 1460, 726, 1729, 1343, 1462, 1000, 1465, 1693, 877]. Documents [458, 819, 1045, 1047, 1025, 1004, 933, 1656, 1013, 1463, 1457]. Does [101, 1631, 1599, 68]. Dog [923]. Domain [1651, 717, 482, 1360, 1547, 204, 283, 575, 320, 527, 1456, 1209]. Domains [1441, 533, 1488, 798]. Domination [618]. Door [1078]. Doors [1065, 1006]. Double [1193, 1220]. Double-Scan [1193]. Doubling [542]. Down [865]. Dozen [233]. DPDA [410]. Dragons’01 [239]. Drawings [646, 1034]. drilling [1455]. Driven [732, 708, 1608, 323, 1161, 1179, 1638, 858, 45]. Driver [338]. DS [1249]. DS-CDMA [1249]. DSP [1256]. DTDs [1301]. DTSpin [758]. Dual [1215]. Dual-LAN [1215]. Dual-Path [1215]. Dummett [770]. Duration [436, 698]. during [1195, 564]. Dutch [1331, 1336, 253]. Dynamic [120, 1726, 1619, 1409, 6, 1567, 26, 9, 393, 1216, 1629, 1185, 580, 927, 1576, 1515, 1469, 581, 1303, 1148, 1537, 211, 24, 504, 1405, 108, 1753, 762, 1136]. Dynamical [1695]. Dynamically [103]. Dynamics [1788, 1768, 1748, 1587, 1591, 1571]. DZB [998].

E-commerce [713, 1472]. E-learning [957]. E-mail [1466, 1489]. E-marketing [837]. eAccessibility [934]. Each [1080]. Eagle [280]. Easily [1002]. Easy [955]. eBusiness [457]. ECDL [1074, 1071]. ECDL(R)

[1070, 1072]. Eclectic [1059]. eCommerce [937]. Economical [88]. Edge [1772, 1601]. Edges [200]. Edinburgh [795]. Edit [609, 851, 84]. Editor [1024]. Education [951, 1080, 950, 947, 945, 1071]. Educational [954]. Edutainment [961]. EEG [1752, 1732, 1750, 1755]. eEurope [934]. Effect [1169]. Effective [120, 387, 1495, 977, 1715, 1315]. Effectively [1051]. Effects [1322, 1593, 891, 734]. Efficiency [301]. Efficient [1779, 77, 872, 104, 1266, 182, 1420, 357, 518, 495, 1567, 622, 439, 1239, 604, 625, 1520, 612, 427, 637, 1244, 1414, 1673, 1554, 15, 789, 493]. Efficiently [1741]. Effort [26]. Eidetica [1338]. Eigen [264]. Eigenspace [854, 910]. Eigenspace-Based [910]. Eigenvectors [893]. ELAN} [303]. Elastic [1720]. Elderly [942, 940, 1053, 941]. Electric [204, 667]. Electricity [1496]. Electronic [744, 1059, 1479, 996]. Elementary [420]. Elimination [434, 552]. Elliptic [528, 610]. EM-Based [1668]. Email [1454, 1456]. Embeddable [71]. Embedded [745, 664, 1557]. Embedding [829, 769]. Emergent [1482]. Emerging [942, 941]. Emotional [1017]. Emotive [993]. Empirical [686, 724]. Employing [573, 992, 1163]. Employment [959]. Enabled [1233, 1758, 1230]. Enabler [1007]. Enabling [457, 15]. Encode [1618]. Encoded [1447, 1586]. Encoder [1222]. Encoding [122, 80, 1602]. Encodings [439]. Encrypted [518]. Encryption [530]. End [694]. Energy [371, 622, 1642, 625, 624, 1771]. Energy-Efficient [622, 625]. Enforcing [1178]. Engineering [288, 325, 113, 316, 1504, 339]. Engines [729]. English [1331, 811, 1316, 1321, 1339]. English-Dutch [1331]. Enhance [328, 1576]. Enhanced [673]. Enhancement [838, 865, 1267, 477]. Enhancing [1061]. Ensemble [1661, 689].

Ensembles [1630]. Ensuring [1182]. Enterprise [17, 5]. Entropic [1621]. Entropy [710, 128, 695]. Entropy-Based [695]. Entry [990, 234]. Enumeration [654]. Environment [72, 1552, 1305, 1015, 454, 1116, 1258, 1296, 1183, 1041, 54, 1, 1154, 1725, 577, 1299, 1267]. Environment-Aware [72]. Environments [1189, 1112, 1476, 1548, 151, 211, 3, 1560, 962, 177, 15]. Epilepsy [1609]. Episodes [1735]. Episodic [1612]. Epsilon [452]. Equal [649]. Equal-Flows [649]. Equality [444]. Equation [1166]. Equational [1101, 294]. Equations [1200, 1145, 347, 1203, 783, 1097]. Equilibria [351]. Equipped [1032]. Equivalence [1145, 409, 64, 1437, 410]. Equivariant [343]. Erlang [739]. Error [686, 1752, 1372, 1735, 1663]. Error-Tolerant [1372]. Errors [781, 1045, 1256]. ESAPS [339]. ESC [744]. ESC/Java [744]. Essence [418]. Essex [270, 225]. Establishing [1306, 320]. Establishment [762]. Estimate [893, 1635, 1734, 1494]. Estimated [250]. Estimates [1663]. Estimating [1563, 867]. Estimation [1684, 859, 198, 1228, 1648, 1691, 1652, 1714, 174, 1771, 471, 1562, 1123]. Estimations [1513]. Estimator [1743]. ESTs [76]. Eternity [587]. Ethernet [1215]. EU [1056]. Euclidean [423]. EUN [947]. Europe [1080]. European [1086, 952, 999]. Eurospider [1320]. Evaluate [950]. Evaluated [1337]. Evaluating [507, 828, 1405]. Evaluation [1031, 989, 1305, 1109, 751, 1294, 1213, 203, 1347, 1531, 981, 1197, 1120, 474, 1410, 816, 508, 1533, 1346]. Evaluational [953]. Evasion [417, 183]. Event [113, 1392]. Event-Based [113]. Events [1189]. Evidence [887, 198, 863]. Evolution [331, 4, 1620, 119]. Evolutionary [1491, 1659, 1696, 674, 332, 178, 1443]. Evolving [197, 1078, 1450]. Exact [381, 143, 74, 95]. Example [1431, 1069, 864, 858]. Example-Driven [858]. Examples [90, 608, 968]. Exception [1195]. Exceptions [291, 451]. Exclusion [1181, 623]. Executable [1420]. Executing [166, 1541, 1169]. Execution [1543]. Exemplified [954]. Exercise [588]. Exeter [1317]. Existence [134]. Existential [829, 833]. Expansion [898, 1327, 1326, 984]. Expeditious [1464]. Experience [838, 1066, 1028]. Experiences [940, 1027, 971]. Experiment [1523, 1309]. Experimental [1409, 643, 1330, 1406, 1604, 465, 1410, 341]. Experiments [895, 1320, 1317, 1329, 1332, 1315, 1348]. Expert [555, 1453, 914]. Expertise [1466]. Explicit [1425, 298, 848]. Exploitation [58]. Exploiting [1698, 1169]. Exploration [737]. Exploratory [1603, 1516, 1674, 919]. Exploring [143, 1034]. Explosion [839]. Exponential [357, 362]. Exponentially [1605]. Exponents [520]. Expressed [1662]. Expression [1633, 561, 1667, 1653]. Expressions [1024, 98, 1023, 699, 817]. Expressive [764]. Expressively [420]. Expressiveness [550]. Extended [1125, 621, 611, 1418, 617, 638, 650, 1739, 1103, 601, 1701]. Extending [605, 761, 916, 825]. Extensible [164]. Extension [574]. Extensional [297]. Extensions [1225]. Extracting [73, 74, 499]. Extraction [1632, 1459, 708, 1710, 1520, 1646, 1505, 1671, 1524, 1457]. Eye [980]. Eyeshot [156].

F [327]. F180 [184, 243]. Face [562, 1781, 669, 1782, 908, 906, 900, 684, 907, 1520, 909, 43, 678, 486, 677, 506, 687, 910]. FACERET [486]. Facilities [667]. Factor [1109, 1734, 1494, 1217]. Factorial [1759]. Factorization [78, 1199]. Factors [321]. Facts [1109]. Fading [1249]. Failure [1361, 1089, 1248]. Failures [101]. Fairness [1095, 1267, 1182]. False [1533]. False-Name-Proof [1533]. Families [349, 833, 339]. Family [316, 321]. Fast [416, 632, 187, 616, 663, 520, 1778, 95, 1216, 917, 398, 162, 916, 1158, 501, 521, 269]. Fastap [1551]. Faster [93, 348, 795]. FASTY [983]. Fatigue [1742]. Fatique [930]. Fault [1184, 1231, 1251, 641, 1089]. Fault-Tolerant [1184, 1251, 641, 1089]. Fax [479]. FC [226]. Feasibility [1550]. Feature [890, 921, 1508, 1582, 1720, 694, 1521, 857, 1572, 330, 328, 1646, 1505, 849, 889, 502, 1657, 891, 1670, 695, 874, 926, 1638, 329]. Feature-Based [694, 329]. Features [1510, 824, 482, 330, 1021, 906, 1525, 499, 889, 677, 1532, 1528, 1766]. Feautrier [1144]. Federated [459]. Feedback [1727, 977, 1324, 1571]. Feedforward [1784]. Fei [181]. FFT [1201]. FG [1267]. FG-LIW [1267]. Fictitious [1470]. Field [1768, 204, 1587, 14, 1625]. Fields [511, 666, 1574]. Fighters [240, 241]. File [1173, 484]. Filter [1701]. Filtering [1306, 1501, 462, 1489]. Filters [1675, 1739]. Filtersfor [94]. Financial [1733, 1498, 1760]. Find [916]. Finding [424, 92, 399, 88, 1745, 1662]. Finds [200]. Fine [585]. Fingerprint [1775]. Fingerprinting [535]. Fingertip [1018]. Finite [782, 411, 706, 347, 440, 1623, 409, 628, 1588]. Finite-State [706]. FIPA [1283]. FIPA-OS [1283]. Fire [1165, 1592, 1605, 1588, 1713, 1580, 284]. Firing [1598, 1605]. First [434, 349, 550, 1383, 940, 773, 326, 823, 160, 1403, 1384, 566]. First-Order [434, 550, 1383, 773, 326, 1384, 566]. Físchlár [492]. Fisher [1693]. Fishes [913]. Fitness [1631]. Fitting [882]. Fixpoints [1371]. Flat [1357]. Flexible [714, 574, 226, 15, 984]. Flipping [636]. Floor [175]. Flow [1590, 749, 1467, 1542, 391, 1172]. Flows [621, 1219, 649]. Fluctuations [1598]. Fly [1773]. fMRI [1569]. fMRI- [1569]. Focus [578]. Focused [1470, 292]. Focuses [1776]. Focusing [1741]. Focussing [802]. Folds [592]. FOOM [465]. Football [260]. Forecasting [1493, 1496]. Forests [1650]. Forgotten [1109]. Form [1522]. Formal [125, 1424, 732, 1417, 745, 744, 792, 813, 748, 28, 1432, 831, 396, 768, 806, 756]. Formalised [1426]. Formalism [820]. Formalisms [761]. Formality [741]. Formalized [811, 1438]. Formalized-English [811]. Formalizing [1427, 34]. Formally [746]. Format [1047]. Formation [1603, 573, 236]. Forms [1576]. Formulae [300, 1098]. Formulas [40, 780, 778, 443, 724, 828, 38]. Formulations [1392]. Forward [755, 680, 693, 477]. Foster [1070]. Foundations [570]. Foundedness [799]. Four [633, 266]. Fourier [1522, 1158]. Fourth [1756]. FPGA [185]. FPGA/VHDL [185]. FPGA/VHDL-Based [185]. Fractal [1787]. Fractional [1745]. Fractographs [930]. Fragment [442]. Frame [491, 811, 1683]. Frame-CG [811]. Frames [503, 502]. Framework [125, 989, 186, 6, 71, 709, 1270, 1517, 337, 1265, 32, 1011, 164, 1230, 795, 1506, 1490, 910]. Frameworks [1263]. France [243]. Fraud [1495]. Free [781, 1155, 1428, 650, 397]. Free-Style [1428]. Freeway [1737]. Frege [358]. Freiburg [168, 203, 158]. French [271, 1339]. Frequent [1440, 399]. Friedberg [1027]. FTB [1175]. FU [240, 241]. FU-Fighters [240, 241]. Fuel [1649]. Full [550, 607]. Fully [1183, 1303]. Fun2Mas [256]. Function [782, 605, 1684, 558, 847, 1592, 1659, 1658, 1036, 422, 1623, 127, 1691, 1643, 1052, 1701, 1668, 1707]. Functional [381, 1423, 1677, 751, 1294, 580, 554, 1586, 61]. Functions [1624, 494, 572, 1735, 742, 592, 172, 601, 539, 397]. Funeral [483]. Fusing [1776, 1453]. Fusion [885, 903, 257, 922, 1649]. Future [167, 1085]. Futures [1492]. Fuzzy [1632, 1639, 1511, 1521, 197, 1153, 1513, 1443, 1442, 1477, 1452, 1506, 1659]. Fuzzy-Neural [1511]. FX [1491].

Gait [1627, 979, 975]. Galois [583]. Game [46, 415, 351, 236, 250, 1279]. Games [183, 970, 400, 156, 25, 974, 380, 296, 1481]. Ganglion [1597]. Gap [952, 1550]. Gapped [94]. GARCH [1501]. Gardening [963]. Gas [1723]. Gateway [1260]. Gaussian [882, 1704, 1733, 1493, 929, 883, 884]. Gear [1455]. Gemplus [744]. Gender [670]. Gene [1633, 92, 1667, 1653]. Gene-Expression [1633]. General [73, 151, 1151, 1517, 1359, 450, 1137, 1438]. General-Purpose [1137]. Generalised [556, 201, 759, 391, 1157]. Generalization [1101, 131, 1629, 1663, 661, 44]. Generalized [1487, 855, 617, 547, 910]. Generated [494]. Generating [1584]. Generation [781, 782, 787, 441, 1383, 1566, 1175, 816, 743, 1638, 1348]. Generative [1620, 1755]. Generator [1204, 513, 1540]. Generators [1595, 1596]. Generic [10, 1605, 575]. Genes [1662]. Genetic [1514, 191, 1521, 561, 1251, 1631, 919, 1657, 1442, 568]. Genetic-Based [568]. Genome [76]. Genomes [595]. Genomic [839]. Genus [526]. Geometric [371, 129, 1738]. German [1336, 998]. GermanTeam [272]. Getting [1411]. GF [612]. GHC [1357]. Giessen [1027]. Giessen-Friedberg [1027]. Gigabit [1208]. Gilgul [9]. GIRT [1333]. Glass [666]. Glass-Markov [666]. Glaucomatous [866]. Global [156, 168, 1296, 1183, 345, 240, 1178, 512, 1581]. Global-Locally [1581]. Gloss [1345]. GMD [258]. GMD-Robots [258]. Goal [1725]. Goalkeeper [163]. Gödel [434, 770]. Gödel-Dummett [770]. Goes [192, 998].

Going [1552, 1040]. Gone [1059]. Good [416]. Gossiping [372, 631]. Governance [333]. GPS [1015]. Graded [797]. Gradient [1757, 1678, 212, 1786]. Gradual [767]. Grain [1143, 1224]. Gram [94, 1685, 1322]. Grammar [697, 699, 78, 846]. Grammar-Based [699, 78]. Grand [1165]. GrAnDe [786]. Graph [619, 1307, 1126, 850, 860, 844, 871, 1702, 854, 849, 823, 368, 869, 851, 1385, 652, 812, 848, 868]. Graph-Based [844]. Graphical [1510, 314, 924]. Graphics [1031, 1033, 858, 877]. GraphPlan [49]. Graphs [417, 1408, 702, 853, 400, 390, 620, 349, 697, 829, 819, 621, 618, 603, 813, 567, 604, 832, 1129, 809, 814, 653, 348, 808, 816, 826, 1404, 867, 425, 833, 1401]. GRASP [1236]. Gray [929]. Gray-Level [929]. Greedy [1679, 584, 148, 1533]. Grid [1232, 1226, 850, 645, 1154, 1230, 1250, 1111]. Grid-Computing [1226]. Grid-Packing [645]. Grids [18]. Ground [783]. Groundwork [719]. Group [525, 352, 517, 1089]. Grouped [889]. Groups [616, 347]. Growing [1723]. Guaranteed [1267]. Guaranteeing [1264]. Guarantees [150]. Guarded [442]. Guards [120]. Guide [573, 976]. Guided [1390]. Gurmukhi [878].

H [1502]. H.26L [1222]. Habitat [1058]. HAL [975]. HAL-3 [975]. Halfspaces [608]. Hamiltonian [604]. Hammerheads [254]. Hamming [1631]. Hand [563, 1669]. Hand-Written [563, 1669]. Handicapées [1058]. Handicapped [1003]. Handling [1195, 1301, 1300]. Handoff [1257]. Handwriting [1639]. Handwritten [898, 897, 869, 922]. Haptic [1036, 1052]. Hard [660, 401]. Hardness [392, 360]. Hardware [185, 1626, 1126, 1190, 1178]. Hardware-Amenable [1626]. Hardware-Software [1178]. Harmonisation [999]. Hashing [1168].

Haskell [585]. HAVAL [541]. Having [349]. Head [980, 967]. Hearing [968, 967]. Hebbian [1674]. Heilbronn [633]. Helli [227]. Helli-Respina [227]. Hellman [538]. Help [1015, 1054, 716]. Helpdesk [1617]. Helping [716]. Herbrand [783]. Hereditary [604]. Heteregenous [239]. Heterogeneous [1650, 1296, 251, 1374, 221, 1310]. Heteroscedastic [894]. Heuristic [47, 732, 27, 1236, 58]. Heuristic-Driven [732]. Heuristically [350]. Heuristics [696, 91, 1401, 779]. Heuristics-Based [696]. Hidden [1459, 1671, 1523, 1759, 918, 1089]. Hierarchical [1133, 931, 1781, 150, 1460, 726, 1449, 1461, 1517, 865, 208, 487, 1770, 35, 1217, 1703, 1766, 1676]. Hierarchically [1526, 1456]. Hierarchies [122, 1453, 123]. Hierarchy [354, 413]. High [1748, 1731, 1146, 1637, 1185, 395, 1138, 1210, 1160, 1154, 175, 922, 1649, 1172, 1108]. High-Accuracy [1637]. High-Availability [1185]. High-Dimensional [1748]. High-Level [1146, 395]. High-Speed [175]. Higher [1419, 796, 598, 309, 1418, 1356]. Higher-Order [796, 1356]. Highest [1079]. Highlighted [901]. Hilbert [452]. Hint [1140, 1500]. Histogram [929]. Histogramming [398]. Histograms [1709]. Historical [496]. HMM [680]. HMM/SVM [680]. HMMs [92]. HMPH [1058]. Hoare [734]. Hoc [1241, 1240, 372]. HOL [1424, 773]. Hole [179]. Holmes [544]. Home [1260, 1055]. Hop [622]. Hopfield [1788, 1452]. Host [1164]. Host-Parasite [1164]. Hosted [1292]. Hough [928]. HR [787]. HTML [1003, 1463]. Huffman [412]. Hull [1157]. Hulls [131]. Human [875, 906, 1627, 28, 595, 76, 1609, 1538, 1539]. Human-Automation [28]. Humanoid [193, 161]. Hummingbird [1337]. Hungarian [1050]. Huskies [276]. Huskies-01 [276]. Hybrid [179, 777, 435, 1639, 680, 553, 1373, 1164, 1118, 978, 55, 1114]. HyLoRes [777]. Hyper [632]. Hyper-rings [632]. Hyperbalanced [290]. Hyperbolic [368]. Hypercolumns [1600]. Hypercubes [1211]. Hyperelliptic [526, 527]. Hypergraph [1310]. Hypergraphs [426, 427]. Hypermedia [950]. Hyperparameter [1785]. Hyperresolution [784]. Hypothesis [476, 30, 1774, 1654].

I-tracings [496]. I/O [1189]. ICA [1748, 1738]. iCAMS [1564]. iCLEF [1344, 1345]. ICT [951]. Id [181]. Identifiability [1749]. Identification [1767, 1624, 850, 538, 912, 1747, 96, 1701, 488, 1554]. Identify [616]. Identifying [87, 1474, 569]. IEEE [1099, 1254]. II [1596, 513, 157]. Illumination [909]. ILP [1173]. Image [1767, 182, 665, 495, 1582, 507, 494, 875, 476, 857, 1572, 479, 482, 1778, 1765, 1762, 904, 905, 1777, 871, 483, 909, 478, 1274, 487, 470, 474, 1018, 503, 501, 56, 504, 1580, 477, 1769, 471, 1764, 473, 481, 877]. Image-Pair [476]. Imagery [1150, 491]. Images [881, 507, 475, 1697, 1776, 873, 869, 496, 493]. Imagination [960]. Imaging [1586, 175]. Immature [320]. Immediate [164]. Immune [421, 539]. Impact [1335, 1069]. Impaired [1016, 970, 974, 1047, 1015, 1010, 1020, 1080, 1027, 1000, 967, 1007, 1001, 1071, 1057, 1009]. Impairments [1030]. Impediments [693]. Imperative [580]. Imperfect [46]. Implementation [104, 565, 156, 190, 1305, 1222, 1516, 1360, 442, 1270, 313, 1259, 1531, 116, 1771, 214, 462, 1401]. Implementations [1549, 746]. Implementing [185, 1193, 1087, 580, 1727, 1302, 1119, 1230]. Implicit [1788, 1394]. Importance [1736]. Important [875]. Improve [198, 1083]. Improved [1744, 393, 426, 406, 602].

Improvement [1415, 1629, 1205, 1035, 301, 512]. Improvements [731]. Improves [1706]. Improving [1189, 1307, 300, 1294, 981, 1141, 1739, 391, 695, 956, 935]. In-memory [1177]. Inalienable [701]. Inapproximability [347, 426]. Including [959]. Inclusion [963, 992]. Inclusive [1011, 945, 956, 956]. Incomplete [895, 421, 437, 828]. Incompleteness [451]. Incompressible [1202]. Inconsistent [1453]. Incorporation [915, 198]. Increase [1607]. Increased [182]. Increasing [1752, 1059, 1263, 981, 1170]. Incremental [827, 736, 1682]. Indefinite [560, 452]. Independence [1623, 653]. Independent [890, 1744, 1749, 1651, 1168, 102, 517, 1505, 1764, 902]. Index [20, 60, 1246, 1275, 1291, 1312, 1350, 1375, 1400, 1416, 1439, 1529, 1545, 1568, 1790, 75, 99, 126, 154, 285, 315, 340, 431, 456, 469, 510, 548, 579, 594, 658, 691, 728, 763, 804, 834, 841, 932, 1088, 1105, 1447, 1493, 1414]. Indexed [589, 1175]. Indexer [1338]. Indexing [726, 500, 1451, 1462, 480, 1326]. indiGo [838]. Individual [1532, 1071]. Individually [990]. Individuals [992]. Indoors [1562]. Induces [1575]. Induction [1419, 551, 192, 1513, 45]. Inductive [1367, 554, 724, 138, 140, 576]. Inefficiency [1202]. Inequalities [772]. Inexact [854]. Inference [1511, 561, 138, 1676]. Inferring [608, 139]. InfiniBand [1213]. Infinite [67, 395, 1654, 798]. Infinite-State [395]. Infinity [571]. Influence [316, 1599]. Influenza [840]. Information [288, 46, 1306, 1377, 666, 805, 1065, 1316, 915, 1006, 1012, 955, 708, 942, 1597, 1043, 1460, 1576, 952, 1559, 1069, 1183, 1479, 499, 1272, 1068, 30, 1081, 1513, 1321, 1544, 809, 1328, 1662, 170, 1336, 1454, 700, 1564, 1586, 967, 1011, 722, 938, 1299, 1596, 58, 1672, 1404, 1017, 934, 936, 1673, 1346, 133, 531, 946, 463, 1457, 1668]. Information-Theoretic [30]. Informing [984]. Infrastructure [1281, 957, 1476, 16, 1111]. Infrastructures [515]. Infty [1029]. Inherence [397]. Initial [34, 1267, 234]. Initiation [1155]. Initiatives [1005]. Inlining [121]. Innovation [955]. Input [1598, 987, 985, 713, 1605, 1670, 1583, 1528]. Inputs [1165]. INS [1561]. INS/Twine [1561]. Insecurity [542]. Insider [536]. Inspections [1652]. Inspired [1573]. Instant [1227]. Instant-Access [1227]. Instantiation [329]. Institute [1027]. Institutionalizing [319]. Instruction [302, 1225, 1167, 1170]. Instruction-Level [1167, 1170]. Instructions [1171, 105, 1169]. Instruments [1086]. Integer [646]. Integers [793]. Integral [908]. Integrate [1592, 1605, 1588, 1713, 1580]. Integrate-and-Fire [1592, 1605, 1588, 1713, 1580]. Integrated [1102, 740, 561, 1300]. Integrating [570, 741, 1117]. Integration [1295, 1377, 560, 413, 1702, 444, 1018, 869, 1585, 1070, 552, 1310, 324]. Intellectual [962]. Intelligence [1499, 1279]. Intelligent [1276, 1469, 1328, 837, 982]. Intensional [299]. Intentional [1561]. Inter [464, 1607, 1209]. Inter-domain [1209]. Inter-organizational [464]. Inter-spike [1607]. Interaction [1642, 330, 859, 314, 54, 1278]. Interactions [1182, 1717, 1638]. Interactive [1038, 1227, 1342, 944, 486, 471]. Interchangeable [1341]. Interconnection [1206]. Interdependence [1127]. Interface [1014, 1012, 940, 773, 966, 55, 1035]. Interfaces [876, 815, 28, 981, 837, 982]. Interference [747, 1268, 1784]. Intermediaries [1243]. Internal [333, 1118, 138]. International [1083]. Internet [1016, 1207, 1209, 1355, 1468, 350, 980, 1004, 1445, 1255, 1250, 1271, 153, 111]. Internetworked [1265]. Interoperability [464, 823]. Interpolation [1745, 610].

Interpretation

[880, 211, 966, 1103, 1747, 1612, 830].

Interpreter [314, 116, 1357]. Interpretive [1124]. Intersection [354]. Interspike [1595]. Interval [1607, 567, 1595, 81]. Interventions [976]. Intraday [1491]. Introducing [1193]. Introduction [1313, 1186, 959, 332, 1285]. Intrusion [673]. Intuitionistic [769]. Invariance [1584]. Inverse [606, 1707]. Inverting [592]. Investigation [1325]. Investment [416]. Invited [1090]. Invocation [1195]. Involving [448]. IP [1208]. IPC [1172]. IPSIS [1044]. IR [1334, 1323, 1340]. IR-n [1334]. Irreducible [652]. Irregular [1143]. irst [1319]. Isabelle [768, 1422, 1424]. Isabelle/HOL [1424]. Isabelle/Isar [1422]. Isar [1422]. ISocRob [259]. Isomorphism [526]. Isotop [1719]. Issues [1108, 1393, 823, 1011]. IT-Certificates [1070]. Italian [1336]. ITC [1319]. ITC-irst [1319]. Item [398]. Items [399, 644]. Iteration [1745]. Iterative [1203, 1734, 33, 1606, 894]. Iterative-Refinement [33].

J [107]. J-Orchestra [107]. Japan [1080, 1046]. Japanese [965]. Java [1232, 125, 104, 745, 744, 4, 1407, 105, 590, 11, 768, 1187, 107]. Java-Based [1232]. JavaBean [5]. JHU [1329]. JHU/APL [1329]. JITty [311]. JML [109]. Job [1131]. Joint [1371, 166, 867]. JPEG [482]. JUnit [109].

K-Means [887, 1413]. k-NN [1508].

Kalman [1675, 1739, 1701]. Kansei [1449]. Karlsruhe [210]. KDD [1490]. Keep [1073]. Keepaway [176]. KENSEI [193]. KENSEI-chan [193]. Kernel [1761, 1577, 680, 1681, 676, 129, 674, 661, 1682, 1688, 662, 1686, 888, 690]. Kernel-Based [1761]. Kernels [686, 665, 1685, 1683, 1693, 132].

Kerneltron [668]. Key [515, 338, 612, 514, 529, 502, 543, 512, 1380]. Keyboarding [987]. keying [514]. Keypad [1551]. Keys [511]. Keystream [513]. KIF [811]. Kinematics [1707]. KISS [1073]. Kit [876]. Kleene [593]. Knapsack [365]. Knowledge [1619, 815, 1391, 836, 1455, 1147, 1458, 1504, 811, 437, 823, 1454, 446, 1453, 807, 1452, 1562, 1457]. Knowledge-Based [1457]. Known [1072]. KnowWare [1040]. Koblenz [229]. Kohonen [1778]. Kolmogorov [1645, 373]. Korean [701]. KOSPI [1492]. Kowalski [1385, 1352]. Kyushu [266].

L [1131, 396]. Labeled [1457]. Labeling [1638]. Labelled [385]. Labour [1072]. Labscape [1549]. LAESA [916]. Lambda [299]. LAN [1215]. Lana [112]. Landmark [1702]. Landmarks [1697]. Langley [1417]. Language [715, 876, 1376, 466, 1311, 9, 1393, 974, 1466, 1741, 1321, 1343, 742, 966, 171, 965, 1221, 173, 61, 707, 1357, 762, 846]. Languages [69, 354, 1355, 1042, 847, 137, 820, 1675, 1360, 1188, 353, 313, 355, 294, 1011, 1337, 139]. LANs [1254, 1268]. Large [1200, 706, 660, 1220, 896, 1741, 1097, 1153, 1142]. Large-Scale [1220]. Larger [1267]. Latency [650]. Later [1418]. Lateral [1717]. Lattice [1593, 56, 292, 569]. Lattices [819, 448]. Law [999]. Laws [350]. Lax [769]. Layer [1721, 1262, 1166, 1677]. Layered [635, 480, 293]. Layers [103]. Layout [1045, 1142]. Lazarus [228]. Lazy [798]. LCF [773]. LCF-Style [773]. Leading [1069]. League [184, 189, 159, 160, 266, 251, 234]. Leagues [155]. Leaky [1588]. Learn [987, 965, 39, 776]. Learnability [143, 657]. Learned [900]. Learners [961, 140]. Learning [1743, 838, 856, 1077, 1589, 183, 1781, 48, 145, 146, 1626, 987, 1733, 1493, 660, 1393, 819, 1491, 1476, 843, 1615, 1572, 1642, 847, 137, 1640, 1367, 1637, 1608, 1629, 1617, 1675, 1669, 1641, 1628, 1507, 1519, 1080, 866, 1475, 1674, 127, 210, 977, 56, 1759, 1713, 153, 1730, 502, 948, 1652, 1683, 35, 1789, 1694, 1700, 1763, 959, 1753, 1503, 36, 176, 1583, 177, 1071, 852, 670, 1625, 1766, 1481, 1486, 1539, 142, 1707, 1784, 957]. Learnt [1453]. Least [1679]. Least-Squares [1679]. Leave [1663, 1686]. Leave-One-Out [1663, 1686]. Leg [978]. Legacy [330]. Legal [1332, 1389]. Legged [159, 279, 197, 280, 267, 275, 266, 269]. Legislating [1066]. Legislation [1069]. Lemma [443]. Lempel [1519, 78]. Length [424, 92, 978, 305]. Lengths [304]. Less [26]. Level [1143, 1720, 1146, 1430, 1167, 395, 911, 952, 929, 546, 1662, 3, 814, 1532, 1170, 329, 1541, 1404]. Levels [1742, 1764, 45]. Levenshtein [94]. Levy [380]. Lexical [723, 1458]. Lexicon [697]. Lexicon-Grammar [697]. Libbraille [1002]. Libraries [1005, 996, 471]. Library [1237, 500, 998, 1191, 772, 1002, 1008, 492, 1001]. License [681]. Life [1539]. Life-Like [1539]. Lifetime [1240]. Lift [1072]. Light [791]. Like [1539]. Likelihood [1674, 873]. LILI [513]. LILI-II [513]. Limb [1052]. Limitations [1068]. Limited [161]. Limits [693]. Line [1266, 633, 319, 1639, 325, 644, 330, 645, 338, 337, 759, 326, 323, 327, 335, 324, 320, 79, 329, 322, 217]. Lineage [1297]. Linear [419, 862, 85, 780, 1783, 476, 438, 1640, 885, 525, 1623, 517, 1097, 894, 437, 136, 433, 1759, 892, 519, 1643, 1666, 425, 139, 1684, 561, 920, 37]. Linear-Threshold [136]. Linear-Time [438]. Linearly [807]. Lines [331, 317, 333, 318, 332, 336]. Lingual [1324, 1326, 983]. Linguistic [720, 1462]. Link [1249, 454]. Linkage [1665]. Linking [761, 505]. Linux [1044]. List [624]. Lists [1454]. Literacy [961, 1059]. literal [817]. Literature [1610]. Little [729]. Live [1271]. Live-Broadcasting [1271]. Liveness [1182]. Living [1058]. LIW [1267]. Load [1128, 1161, 1169]. Loading [120]. Local [419, 464, 131, 156, 189, 1724, 1240, 1721, 1669, 1451, 897, 1727, 1099, 86, 345, 503, 1405, 241]. Localisation [180, 212]. Locality [1205, 1141]. Localization [1781, 681, 209, 175, 269, 481, 205]. Localized [130]. Locally [1581]. Location [1563, 1564, 1562, 647]. Location-Awareness [1563]. Locative [697]. Locker [739]. Locomotion [1014, 199, 1696, 1585, 1694]. Logarithm [610]. Logarithmic [375, 363]. Logfiles [164]. Logic [1370, 72, 1390, 434, 766, 435, 390, 563, 1393, 1355, 1426, 438, 300, 65, 1366, 1146, 769, 1367, 771, 420, 197, 1360, 1383, 374, 773, 1373, 1353, 770, 70, 1364, 1356, 326, 437, 433, 1368, 559, 828, 808, 1389, 1395, 557, 1477, 552, 1369, 734]. Logic-Based [1373]. Logical [583, 1354, 74, 827, 795]. Logicality [556]. Logician [1351]. Logicism [1398]. Logics [419, 777, 288, 1381, 1377, 439, 440, 71, 764, 445, 446, 450]. LogP [1129]. LogP-Machines [1129]. Long [1031, 1615, 1210, 971, 1739]. Long-Term [1615, 1739]. LONGEST [85]. Longo [380]. Look [609]. Look-Up [609]. Lookahead [43]. Loop [1143, 801, 1270]. Loop-Level [1143]. Loops [1139, 304, 1141]. Losing [1598]. Lost [1758]. Loveland [373]. Low [1031, 189, 1021, 1118, 1019]. Low-Cost [1118]. Lower [1663, 362, 855, 892, 142]. Lower-Complexity [892]. LRP [271]. LS [1685]. LS-SVM [1685]. LSTM [1675]. LuckyStar [157]. Łukasiewicz [445]. LVQ [1639]. Lyapunov [1200].

MAC [1245, 1262]. Machine [183, 1626, 843, 668, 1317, 689, 1455, 866, 1692, 1652, 1682, 176, 1067, 1497]. Machines [1785, 675, 1679, 680, 671, 106, 664, 679, 1455, 1776, 1129, 678, 129, 659, 661, 1690, 682, 672, 1132, 1614, 1625, 687]. Macrocolumns [1578]. Macrostep [1437]. Magic [1379]. Magnetic [1505, 1562].

Mahalanobis [893]. mail [1466, 1489]. Mailing [1454]. Mainstream [1026]. Maintaining [1409, 534]. Maintenance [1305, 1239, 667]. Mainz [230]. Make [736, 101]. Make-Based [736]. Making [572, 337, 1558, 1532, 1055, 994, 1636]. MALLBA [1237]. Management [838, 714, 1499, 1733, 1185, 318, 1563, 1454, 1443, 531, 463, 1489]. Manager [715, 468, 1477]. Managing [1298]. Manifesto [844]. Manipulation [985]. Manufacturing [635]. Many [1074, 891]. Map [1726, 1720, 1711, 1728, 485, 1495, 1712, 1718, 1708, 1715, 1465, 1722, 1562, 1494]. Maple [572]. Mapping [709, 1723, 1137]. Mappings [1623, 892]. Maps [1731, 1697, 1721, 1778, 1460, 859, 1445, 63, 1727, 1040, 501, 1709, 1730, 1502, 486, 1717, 1581]. Margin [149]. Marine [1164]. Markers [181]. Market [1072, 1497, 1740]. Marketing [338, 837]. Markov [666, 144, 422, 1253, 1523, 1759, 918, 385]. Markovian [1602]. Markup [997]. Martingales [386]. Maryland [1345]. MAS [1284]. MASIVE [1480]. MASSAT [1488]. massive [1108]. Massively [1163]. Match [378]. Matches [88]. Matching [82, 476, 91, 309, 850, 298, 95, 871, 93, 854, 1274, 86, 80, 851, 596, 489, 472, 982, 81, 812, 848, 868]. Material [1025]. Materials [1071]. Math [1026]. Mathematical [570, 780, 565, 1025, 1024, 1023, 774]. Mathematics [1028, 1022, 806, 582]. MathWeb [576, 774]. matic [776]. Matrices [90]. Matrix [1681, 1685, 600, 893, 1722]. Matter [68]. Maturity [331]. Maximal [88, 87]. Maximize [1574]. Maximizing [644, 149]. Maximum [1092, 853, 1240, 621, 710, 1674, 1689, 873, 391, 695, 652]. Mazur [422]. Mazurkiewicz [420]. McGill [273]. MDL [1644]. MDPs [1290]. Mean [1092, 1587, 1625, 1248]. Mean-Field [1587].

Meaning [571]. Meanings [1394]. Means [1487, 887, 1153, 1413, 867, 1562]. Measure [1521, 912, 1672, 141]. Measurement [1087, 889, 1649, 1085]. Measures [131, 1621]. Measuring [1269, 1765, 1458, 379]. Mechanical [1037, 738]. Mechanism [1168, 748, 907, 102, 1118, 1384]. Mechanisms [1534, 1115, 1563]. Media [995, 1264, 1219, 1082, 991]. Median [917, 855, 845, 647]. Mediated [1278]. Mediation [1298]. Mediator [1299]. Medical [1506]. Meditation [40]. Meet [742]. MEGA [775]. Membership [1089, 613]. Memories [1577]. Memory [1139, 1171, 413, 1151, 623, 1612, 1177]. Mental [1036, 1041]. Merchant [1479]. Merge [641]. Merging [79, 140]. Merlin [598]. Mesh [1204]. Meshed [1251]. Message [352, 1221]. Messages [372]. Messaging [1289]. Meta [1680, 1226, 1386, 1430, 328, 1541, 794, 1357]. Meta- [1226]. Meta-interpreter [1357]. Meta-level [1541]. Meta-Model [328]. Meta-parameters [1680]. Meta-reasoning [1386, 1430]. Meta-theoretic [794]. Metacomputers [1165]. Metadata [1566]. Metamodel [1536]. Meteo [717]. Method [179, 1633, 562, 507, 1744, 92, 1510, 987, 1307, 1640, 121, 198, 1515, 893, 1195, 337, 759, 42, 1646, 1719, 1274, 724, 236, 978, 977, 1747, 335, 161, 1309]. Methodologies [465, 1518, 886]. Methodology [1349]. Methods [1751, 1788, 1417, 745, 1203, 1188, 844, 483, 676, 1205, 842, 345, 992, 1622, 734]. Metric [1702, 855, 1303, 888]. Metrics [1349, 1126, 1730]. MGTP [1383]. Microarray [1662]. Microcircuit [1610]. Microcircuits [1609]. Microelectrical [1037]. Microprocessors [1168]. Microsoft [1020]. Microstep [1437]. Middle [252, 162, 251]. Middle-Size [251]. Middlet [1471]. Middleware [468, 802, 1151, 1542, 206]. Migratory [1711]. MIKE [195]. Milan [256]. Milling [1699]. MIMIZU [1032]. MIN [392]. MIN-BISECTION [392]. MIN-CSP [392]. Mind [1539]. MINHO [260]. Minimal [600]. Minimalistic [1338]. Minimax [1634]. Minimization [686, 646, 35, 1784]. Minimum [1207, 145, 1409, 619, 1159, 1467, 89, 650, 1674]. Mining [77, 1440, 813, 1449, 836, 1451, 1151, 676, 1147, 1302, 345, 814, 1446, 1154, 837, 1395, 1450, 1443, 1442, 335, 1673, 1448, 1304, 1486, 1489]. Miro [206]. Mismatching [13]. Mispredicted [1169]. Missing [1648]. Mixability [134]. Mixed [1212, 1661, 1637]. Mixed-Signal [1637]. Mixture [882, 1655, 1635, 1648, 884, 1653]. Mixtures [862, 660, 151, 1693, 1760]. ML [920]. Mobile [1705, 185, 1241, 167, 544, 1483, 1469, 1563, 1471, 1058, 1244, 1564, 175, 1703, 7, 762, 217, 481, 1243, 1238]. Mobility [1075]. Modal [439, 828, 450]. Modalities [1014]. Mode [522]. Model [1165, 781, 1644, 782, 1624, 104, 1630, 1421, 1783, 476, 300, 461, 328, 726, 395, 441, 859, 1592, 1298, 930, 905, 1383, 1449, 374, 740, 1096, 758, 1594, 1525, 1479, 1520, 1745, 1755, 581, 1584, 1741, 21, 1098, 1099, 770, 443, 66, 1197, 1734, 724, 326, 1570, 32, 323, 624, 14, 1287, 918, 1773, 869, 1546, 19, 35, 821, 1667, 5, 1664, 1609, 807, 1560, 297, 1774, 623, 1310, 1676, 798]. Model-Based [859, 1298, 930, 905, 1525, 918]. Model-Checking [395]. Model-Driven [323]. Modeling [931, 1124, 1726, 565, 466, 92, 563, 330, 328, 1721, 1559, 674, 1299, 1585, 1693, 471, 1740]. Modelling [870, 515, 1698, 384, 1466, 1097, 920, 17, 884, 846]. Models [862, 882, 411, 1371, 1094, 908, 927, 783, 1298, 1685, 1188, 684, 854, 1041, 207, 1523, 1501, 1610, 437, 1301, 100, 1323, 1759, 873, 1389, 213, 1290, 707, 1760, 846, 1766, 1653, 1268]. Moderated [838]. Modern [117, 805, 954]. Modification [1637]. Modified [884, 543].

Modular [1204, 208, 1434, 111]. Modularity [290]. Modularized [214]. Module [968, 102, 1215]. Modules [102]. Molecular [341]. Moments [872]. Momentum [1641]. Monitoring [1266, 1483]. Monoids [404]. Monolingual [1319, 1335, 1339, 1332, 1336]. Monotonic [1380, 1387, 1592]. Monotonicity [1500]. Monuments [483]. Morphable [684]. Morphological [1336, 873, 722]. Morphologically [924]. Morphology [1706, 1599]. MorphoSys [1224]. Morton [1142]. Mosaics [859]. Most [596]. Most-Specific-Rule [596]. Motion [239, 187, 202, 977, 1601]. Motor [971, 1052, 984]. Motor- [971]. Mount [967]. Mouse [595]. Move [84]. Movement [980, 213]. Movements [1604]. Moves [25]. Movies [914, 489, 488]. Moving [491, 1414, 1774, 217, 463]. MP3D [1121]. MPEG [1526, 914]. MPI [1126, 1192, 1164]. MPI/OpenMP [1164]. Mpro [1340]. Mpro-IR [1340]. MR [1502]. MS [1021]. MSCs [395]. MST [1401]. MST-Based [1401]. MT [1345]. Multi [69, 1280, 186, 1282, 1687, 1283, 560, 1603, 1354, 1122, 1677, 144, 1483, 1541, 282, 718, 1455, 683, 169, 1537, 983, 1287, 920, 1288, 1691, 1690, 1289, 914, 1700, 1404, 557, 1477, 1533, 874, 762, 216, 1481, 1490]. Multi-agent [1280, 186, 1282, 1283, 560, 1541, 282, 1537, 1287, 1477, 1481, 1490]. Multi-agents [1483]. Multi-armed [144]. Multi-band [718]. Multi-categorical [1687]. Multi-Class [683, 1690]. Multi-context [1288]. Multi-dimensional [1691]. Multi-drilling [1455]. Multi-expert [914]. Multi-layer [1677]. Multi-level [1404]. Multi-linear [920]. Multi-lingual [983]. Multi-paradigm [69]. Multi-periodic [1122]. Multi-platform [216]. Multi-sensor [169]. Multi-stage [1481]. Multi-step [1700]. Multi-stream [1603, 718]. Multi-threaded [1354].

Multi-unit [1533]. Multi-valued [557]. Multi-way [762]. Multiagent [48, 1371, 188, 1534, 1488, 165, 1532, 1480, 1277]. Multicast [1211]. Multicasting [1207]. Multiclass [1678, 710]. Multicolorings [640]. Multicontext [559]. Multidimensional [383, 1301, 888]. Multifrontal [1204]. Multilayer [1618]. Multilevel [1500]. Multilingual [1316, 1333, 712, 725, 1348]. Multimedia [970, 987, 1225, 1012, 965, 1224, 1013, 1221, 1218, 484]. Multimodal [1628, 1583, 488]. Multiparty [1182]. Multipath [1249]. Multiple [1376, 1572, 1576, 87, 903, 478, 1776, 1772, 1261, 971, 1546, 1543, 1478, 45]. Multiplication [528]. Multiplicative [132]. Multiprocessor [1136]. Multiresolution [1762]. Multiscale [931]. Multiset [287]. Multisets [793]. Multisignature [536]. Multispectral [1150]. Multisubgraph [854]. Multithreaded [1189]. Multivariate [1517]. Multiversion [1155]. Multiwavelet [926]. Music [1061, 880, 1063, 1524, 1062, 1060]. MuTACLP [1376]. Mutated [96]. Mutation [800]. Mutual [623, 1539].

n [922, 1322, 1334]. N-Gram [1322].

n-Tuple [922]. Name [1533]. Names [712]. NanoJava [734]. Narrative [698]. NASA [1417]. Nash [351]. National [946]. Native [1011]. Natural [715, 466, 1757, 1678, 1572, 1574, 1466, 846]. Navier [1203]. Navigating [405, 1023]. Navigation [1647, 166, 196, 1015, 169, 492, 217]. Navigational [1272, 1773, 1149]. Nazar [1229]. Near [141]. Near-Optimal [141]. Nearest [916, 97]. Necessity [530]. Needs [1277, 954, 947]. Negation [815, 1361]. Negations [829]. Negotiating [1535]. Negotiation [1278]. Neighbor [1253]. Neighborhoods [1727]. Neighbors [97]. Neighbour [916]. Neighbours [916].

Nested [85, 1139]. Net [802, 1720, 428]. Nets [1095, 454, 314, 626, 864]. Network [1779, 1767, 1247, 1492, 1630, 1632, 1233, 1567, 1724, 423, 1590, 584, 1604, 1239, 1467, 1777, 1292, 1594, 1210, 1208, 1520, 1745, 1542, 710, 1645, 1584, 1099, 1251, 1706, 1591, 1588, 1076, 1643, 391, 1236, 1267, 1774, 1452, 1701]. Network-Hosted [1292]. Networking [1611, 1053]. Networks [1573, 1743, 1751, 1589, 1624, 1704, 1781, 371, 565, 1241, 1616, 1775, 372, 1240, 1122, 915, 622, 1678, 1206, 1721, 1216, 1629, 1765, 1762, 1213, 1658, 1699, 1710, 1627, 1623, 1600, 625, 1242, 1257, 1110, 1558, 1458, 1557, 1244, 631, 1754, 602, 1575, 1660, 1696, 1780, 1770, 1753, 649, 1787, 1636, 1602, 59, 1238, 493, 1784, 1737]. Neural [1779, 1767, 1634, 1743, 1751, 1589, 1624, 1768, 1492, 1630, 1632, 1616, 1775, 1724, 1511, 1590, 1597, 1721, 1613, 1516, 1629, 1617, 1762, 1658, 1699, 1777, 1627, 1695, 1623, 1600, 1520, 1495, 1745, 710, 1584, 1458, 1706, 1754, 1656, 1610, 1575, 626, 1660, 1696, 1570, 1780, 1770, 1773, 1643, 1789, 1585, 1672, 1736, 1774, 1753, 1571, 1787, 1636, 1723, 493, 1737]. Neural-Based [1656]. Neuro [1639, 1659, 1742]. Neuro-classification [1742]. Neuro-Fuzzy [1639, 1659]. Neurodynamical [1569]. Neuron [1569, 1592, 1605]. Neuronal [1602]. Neurone [1594]. Neurons [1603, 1587, 1588, 1713, 1580, 1763, 1771]. News [672]. Next [752, 801, 1496]. Next-Day [1496]. NFA [1734, 1494]. NFAs [98]. NGnet [1676]. NITStones2001 [231]. NL [815]. NLDA [1678]. NMDA [1607]. NN [1508, 1620, 501]. No [625, 1186]. Noble [1250]. Node [1006]. Noise [718, 1575, 1753]. Noise-Robust [718]. noising [1688]. Noisy [1441]. Non [890, 1091, 1133, 1108, 475, 1684, 563, 375, 1387, 1134, 1135, 1592, 52, 561, 894, 1339, 965, 62, 1746, 1182, 817, 37, 453, 848]. Non-approximability [1133, 1134].

Non-bayesian [848]. Non-binary [52]. Non-commutative [563]. Non-dedicated [1135]. Non-deterministic [375, 62, 1182]. Non-disjoint [453]. Non-English [1339]. Non-high [1108]. Non-iterative [894]. Non-Japanese [965]. Non-linear [1684, 561, 37]. Non-literal [817]. Non-massive [1108]. Non-monotonic [1387, 1592]. Non-parametric [890, 1746]. Non-repudiation [1091]. Non-retrieval [475]. Nonconvex [127]. Nondeclarative [1358]. Nondeterministic [374]. Nonholonomic [202]. Nonlinear [1726, 1641, 1719, 892, 1701, 1579]. Nonmonotonic [439]. Nonstationary [671]. Normal [634]. Normalization [1329]. Normalized [909]. Normative [1530]. North [1066]. Notation [1063, 19]. Note [1165, 1237, 1200, 1124, 1125, 1145, 1155, 1150, 1202, 1233, 1209, 1171, 1121, 1146, 1222, 520, 1216, 1185, 1135, 1213, 1159, 1190, 1245, 1210, 1183, 1208, 1195, 1131, 1244, 1141, 1175, 283, 1119, 1158, 1118, 1224, 1215, 1221, 1229, 1186, 1236, 1217, 1178, 1196, 1132, 1230, 1187, 1163, 1174, 1179, 1166, 1170, 1137, 1243, 779]. Noun [697, 1343]. Novel [562, 1592, 1698, 1223, 1019]. Novelty [1770]. NTCIR [1347]. Nuclear [1649]. Nucleus [1601]. Nullstellensatz [358]. Number [1133, 381, 619, 1487, 618, 644, 370, 1162, 1692, 653, 1614, 652]. Numbers [702, 421, 620, 422, 601]. Numeracy [961]. Numeric [553]. Numerical [1788, 1198, 1235, 1507, 564, 1433, 1446, 1166]. Nuprl [73].

O [1189]. OAR [335]. Object

[872, 104, 181, 1761, 666, 1270, 1263, 681, 70, 487, 842, 100, 667, 1780, 1714, 1703, 1149, 1196, 497, 1554, 493, 690]. Object-Based [487]. Object-Oriented [1270, 1263, 70]. Objectively [110]. Objects [1354, 9, 1036, 1119, 34, 1414, 463]. Oblivious [357, 376, 540]. Observable [1290]. Observation [156]. Observation-Based [156]. Observational [196]. Observations [925]. Obstacle [179, 175]. Obstacles [209, 217]. Obtaining [1183, 1662]. OCAWA [939]. Occam [638]. Occurrence [698, 1321]. Occurrences [87]. OCDMA [1268]. OCL [1424, 750]. OCL/RT [750]. OCR [878]. ODL [956]. Off [79]. Off-Line [79]. Offering [967]. Office [1021]. Offs [393, 350, 1089]. Older [934]. Omni [179, 209, 241, 175, 247]. Omni-directional [179, 209, 175]. Omni-directional/Pin-Hole [179]. Omnidirectional [199, 163, 200]. On-Line [1266, 633, 1639, 644, 645, 217]. One [94, 1663, 304, 404, 377, 521, 596, 1763, 662, 1686]. One-Class [662]. One-Dimensional [596]. One-Gapped [94]. One-Probe [377]. One-Rule [304]. One-Step [404]. One-Time [521]. Online [963, 1006, 366, 1641, 365, 505, 1739, 1008, 956]. Ontolingua [820]. Ontologies [464, 1282]. Ontology [818, 821, 45]. Ontology-Driven [45]. Open [537, 758, 345, 1385, 1067]. Opening [1065]. OpenMP [1125, 1126, 1164]. Operating [1269, 1697]. Operation [537]. Operational [1535]. Operations [97, 84]. Operator [873]. Operators [298, 448]. OPM [465]. Opponent [192, 213]. Opportunities [936, 941]. Optical [880, 1273]. Optimal [1130, 371, 122, 95, 1136, 1258, 151, 98, 855, 1273, 1745, 1671, 1706, 1662, 624, 141, 428]. Optimality [1144]. Optimally [85]. Optimisation [1237, 442, 561]. Optimising [1192, 1197]. Optimization [1165, 686, 1234, 1235, 1378, 29, 53, 12, 1120, 1402, 1657]. Optimizations [1208]. Optimized [77, 350]. Optimizing [151, 988]. Optimum [1259]. Option [1751, 664]. Options [335, 36]. Oracular [42]. Orange [253]. Orchestra [107]. Orchestrating [1106]. Order [1419, 434, 1705, 796, 135, 349, 550, 309, 1756, 1095, 1383, 1418, 773, 308, 1356, 326, 1384, 566, 1667, 867, 536]. Order-Specified [536]. Ordered [799, 1444, 138]. Ordering [782]. Orderings [790, 1359]. Orderly [602]. Organisational [1485, 1454]. Organising [1460, 1445, 1502, 1465]. Organization [331, 1729, 724]. Organizational [1536, 464]. Organizations [1537]. Organizing [1726, 1720, 1721, 1710, 1711, 1728, 485, 1727, 1718, 1708, 501, 1709, 1730, 1715, 486, 1722, 1717]. Orientation [181, 1772]. Oriented [820, 1270, 1263, 1542, 70, 1000, 938, 1196]. Orthogonal [378, 383, 646, 611]. Osaka [267, 261]. Other [1080]. Out-of-Core [1411, 1097]. Outcome [1086, 1085]. Outcomes [1087, 1083, 1084]. Outline [539]. Output [627, 713, 1643, 1670]. Overcomplete [1738]. Overfitting [1441]. Overhead [1125]. Overloaded [408]. Overview [968, 1314, 1190, 1557, 339]. Owaribito [242].

P [1220, 546, 626]. P-K [546]. P-Tree [1220]. PAC [40, 143, 144]. Pack’01 [269]. Packages [961, 329]. Packaging [8]. Packed [644]. Packet [1262, 596]. Packets [1611, 1758]. Packing [366, 645, 414]. PACOSY [940]. PageRank [630]. Pages [1459, 460]. Paillier [543]. Pair [476]. Pairing [697, 356]. Pairs [393, 87, 348]. Pairwise [683]. PALIO [938]. Paper [227, 1203, 1239, 157]. Papers [708]. Paradigm [802, 350, 580, 69]. Paradigms [381, 597]. Parallel [1130, 1200, 412, 1193, 1204, 146, 1202, 660, 1222, 1167, 1235, 1203, 1188, 1190, 93, 1227, 1147, 1117, 1153, 1162, 1148, 1219, 1158, 1161, 1236, 721, 1149, 1196, 1132, 1163, 1201, 1157, 1113, 1166, 1137, 1199, 1123]. Parallelism [1143, 413, 1167, 1177, 1170]. Parallelization [1138, 1151, 1141]. Parameterized [654, 1254]. Parameters [1767, 10, 63, 1635, 129, 1654, 527, 1680]. Parametric [187, 118, 606, 890, 1746].

Parametrized [1714]. Paramodulation [799]. Parasite [1164, 1187]. Pareto [151, 1728, 29]. Pareto-Optimal [151]. ParList [1158]. Parser [116]. Parsing [1396, 846]. Part [823]. Partial [378, 1095, 346, 640, 66, 388, 1668]. Partially [1063, 1453, 1290]. Participation [934]. Particle [1642]. Partition [872]. Partitioning [1408, 107]. Partitions [449]. Pass [541, 1763]. Passage [1334]. Passing [1221]. Passive [1554]. Patches [1659]. Path [424, 790, 591, 1215, 1410, 348, 1236, 132, 1211, 647]. Path-Based [1211]. Paths [393, 749, 363]. Pathways [1297]. Patient [940]. Pattern [77, 895, 675, 1521, 298, 83, 95, 1488, 883, 574, 1595, 86, 116, 1596, 1774, 1509, 81]. Pattern-Based [116]. Pattern-Terms [574]. Patterns [1440, 1608, 1750, 1451, 115, 1365, 138]. Pause [698]. PC [1118, 1035]. PCA [865, 1688, 1653]. PCAs [1648, 1693]. PCs [1201]. PD [1072, 1070]. PDF [920, 924]. PDMs [861]. Pedagogical [953]. Pedagogically [954]. Pedestrian [679]. Peeling [643]. Peer [1561, 1183]. Peer-to-Peer [1561, 1183]. Peer2Peer [484]. Peircean [818]. PELCR [1194]. Penalty [120]. Pennsylvania [279]. People [1016, 1030, 1045, 942, 1042, 1047, 1043, 1015, 999, 940, 1010, 1005, 1040, 986, 981, 1053, 977, 1034, 954, 1003, 1008, 934, 996, 1070, 936, 959, 962, 941, 1001, 1009]. Peptides [659]. Per-Item [398]. Perception [1036, 1018]. Perceptions [509]. Perceptron [135]. Perceptrons [1633, 1677, 1618, 1662]. Percepts [39]. Perceptually [875]. Perfect [361, 1161]. Perfectly [547]. Performance [1232, 1124, 1212, 1108, 1127, 507, 1241, 1225, 150, 1126, 1121, 383, 1109, 885, 751, 1138, 203, 1210, 1249, 1160, 43, 1120, 1154, 1261, 1186, 1267, 1114, 922, 1113, 334, 1268, 1309]. periodic [1122]. Periodicity [344].

Permissive [825]. Permutation [622].

Person [975, 488]. Personal

[509, 1043, 1271, 1056, 1055, 1489].

Personalisation [1445]. Personalization [1292, 837]. Personnes [1058]. Persons [939, 1063, 1020, 1007, 982, 1071, 1057, 976]. Perspective [1283, 1367, 1486]. Perspectives [1010]. Pervasive [1552, 1553, 1559, 1551]. Petri [1095]. Philosophy [1346]. Phoneme [964]. Phonetic [911]. Photographs [509]. Phrase [1343]. Phrases [899, 697]. Phylogeny [607]. Physical [1030, 981, 976]. Physically [986]. Physiological [1571]. Physiologically [1604]. Pictures [1036]. Piecewise [1640, 920, 1643]. Piecewise-Linear [1643]. PIM [1124]. Pin-Hole [179]. Pipelining [1205]. Piracy [546]. Place [1702]. Places [1040]. Plan [156, 338, 250]. Plan-Based [250]. Planar [1408, 602]. Planning [1370, 47, 166, 1382, 156, 50, 22, 437, 573]. Plasmas [1649]. Plastic [1591]. Plasticity [1608, 1576, 1599]. Plate [681]. Platform [188, 1604, 1557, 1119, 1229, 1556, 321, 216]. Platform-Based [321]. Platforms [339]. Plausible [923]. Play [544, 973, 1470]. Player [25]. Players [192, 221]. Playing [202]. PList [1158]. Pocket [958]. Point [862, 86, 508, 1707]. Point-Pattern [86]. Pointers [1610]. Pointing [1032]. Points [778, 569, 874, 639]. Pointwise [588]. Polarities [333]. Policies [1305, 151, 43, 946]. Policy [533]. Polish [702]. Polychotomic [122]. Polygon [635]. Polynomial [423, 600, 610, 138, 690]. Polynomial-Time [423]. Polynomially [358]. Polynomials [361]. Polytope [569]. Poor [1735]. Population [1772, 1587, 1606]. Pornographic [475]. Portable [1002, 1055]. Portal [1044]. Portfolio [1499, 1733]. Portrait [1351, 1352]. Portugal [226]. Portuguese [719, 711, 717, 705]. Pose [1714]. Position [90, 201, 1018]. Positions [98]. Positive [310, 138]. Possession [701]. Possibilistic [1096]. Possibilities [1010]. Post [1078, 1080]. Post-secondary [1078, 1080]. Potential [1642]. Potentials [1752]. pour [1058]. Power [1429, 350, 528, 975, 612, 1610, 667, 1562, 833, 248]. Power-Sum [612]. Powerdomain [379]. PowerList [1158]. PowerPoint [1020]. Practical [109, 952, 545, 1558, 97, 76, 165]. Practice [764]. Pragmatic [822]. PRAM [1121, 1176]. Pre [1441]. Pre-pruning [1441]. Precision [1649]. Precomputation [1170]. Precondition [1438]. Predication [1223]. Predictable [8]. Predicting [659]. Prediction [1232, 990, 152, 1735, 1699, 1498, 983, 1739, 1497, 1737]. Predictions [141]. Predictive [1711, 986, 342, 884, 133]. Predictors [1168]. Preemptive [408]. Preferences [1055]. Preferred [1047]. Prefetch [1171]. Prefetching [1406, 1253, 1169]. Prefrontal [1613]. Prehension [1570]. Preparing [1033]. Presentation [578]. Presentations [1020, 1221]. Presenting [1034]. Preserving [293]. Prevention [519]. Price [1493, 1496]. Pricing [1751, 1493]. Primality [511]. Primitive [543, 410]. Principal [1708, 1716, 418]. Principled [1517, 58]. Principles [382, 387, 1537]. Print [1061, 999, 1005, 1076, 1008, 996]. Printed [857, 1524]. Printers [1033]. Printing [1048]. Prior [1606, 141]. Priority [356]. Privacy [461, 546]. Proactive [1784]. Probabilistic [1091, 303, 1093, 1094, 927, 709, 655, 1095, 1432, 1096, 1099, 1104, 379, 864, 1103, 213, 174, 1693, 1673]. Probabilistically [250]. Probabilities [1328]. Probability [1514, 1631, 1090, 867]. Probe [377]. Problem [85, 1133, 633, 402, 145, 82, 1555, 423, 1661, 1134, 1159, 1467, 1116, 604, 409, 538, 650, 607, 648, 1500, 1715, 1652, 1236, 1114, 1385, 886, 1401]. Problems [41, 378, 660, 375, 614, 1658, 1378, 365, 363, 855, 1655, 308, 345, 448, 568, 81, 37, 428]. Procedure [466, 438, 792, 1385]. Procedures [310, 364, 1387, 443, 765]. Process [1743, 838, 1122, 316, 741, 1449, 836, 1270, 1100, 938, 401, 760]. Process-Oriented [938]. Processed [1045]. Processes [144, 1094, 67, 384, 23, 738, 385, 339]. Processing [1130, 182, 968, 815, 1122, 704, 1293, 398, 720, 1750, 1296, 1010, 1466, 1455, 1672, 1187, 725, 1464]. Processor [1163, 462]. Processors [1247, 1173]. Produced [1045]. Producing [1051, 1042]. Product [319, 331, 325, 330, 316, 317, 333, 338, 337, 326, 323, 1472, 327, 332, 335, 324, 320, 321, 329, 336, 322]. Production [318, 864]. Products [1469]. Professionals [1026]. Profiles [1649]. Profiling [69]. Prognosis [840]. Program [72, 73, 787, 67, 1363, 346, 981, 1362, 1442, 788, 756]. Programmable [1146]. Programmes [994]. Programming [1370, 1390, 222, 112, 191, 9, 1355, 65, 1366, 1367, 580, 927, 1431, 1681, 1360, 1188, 588, 1353, 1164, 1197, 24, 1368, 646, 64, 1176, 1395, 111, 1090, 1365, 552, 1369]. Programming-Based [1360]. Programs [369, 616, 4, 1192, 1407, 554, 62, 1113, 1438]. Progression [866]. Progressive [1186]. Project [838, 458, 713, 1056, 1179, 949, 1044]. Projection [882, 307, 1603, 1677, 1516, 908, 1719, 1674, 224, 868]. Projects [935]. Prolog [1391, 449, 1365]. Promise [1110]. Promote [963, 167, 1030]. Proof [802, 432, 803, 434, 766, 1101, 1382, 65, 1429, 1387, 771, 441, 454, 770, 1433, 573, 1384, 578, 767, 729, 1385, 775, 795, 1533, 756]. Proof-Carrying [766, 767]. Proof-Net [802]. Proof-Search [770]. Proof-Verifier [1384]. Proofs [74, 362]. Propagation [1165, 1765, 1252]. Proper [146, 697, 712]. Properties [1582, 1358, 599, 312, 759, 294]. Property [757, 514, 293]. Proposal [989, 1424, 977]. Propositional [1381, 438, 771, 446, 38]. Propositions [780]. Prosody [1025]. Prospects [1363]. Protecting [546]. Protection [535]. Protocol [1091, 622, 1216, 1245, 1099, 1244, 287, 1289, 19, 992, 1533]. Protocol-Based [19]. Protocols [1370, 1241, 65, 352]. Prototype [1121, 896, 1054, 119]. Prototype-Based [119]. Prototypes [467]. Prototyping [1122]. Provably [416]. Prover [1383, 450]. Provided [1169]. Providers [1546]. Providing [1015, 1078, 1082]. Proving [1419, 796, 1428, 1433, 1437, 576, 798]. Provision [1080, 1054, 1076]. Proxy [1269]. Pruning [1704, 1441]. Pseudo [1274, 1324]. Pseudo-bipartite [1274]. Pseudo-Relevance [1324]. PSPACE [401]. PSPACE-Hard [401]. PTAS [403]. Public [943, 1553, 515, 612, 543]. Public-Key [612, 543]. Pulsed [1777]. Puncturing [1690]. Pure [1357]. Purpose [1137]. Purse [744]. Pursuit [417, 882, 183, 1603, 1516, 1498, 1674]. Pursuit-Evasion [417, 183]. Pushdown [400]. Pushpin [1557]. Putting [1552]. PuzzleTool [1431].

QoS [1254, 15]. QoS-Aware [15]. QR} [1199, 1664]. QR-cp [1664]. Quadratic [511, 1590]. Quadruped [196, 199, 1696, 170, 212]. Quadrupedal [187]. Quality [1306, 1735, 1765, 1252]. Quantified [435, 443, 772, 569, 450]. Quantifier [552]. Quantifying [1528]. Quantization [1778, 1628, 1711]. Quantum [369, 1514, 655, 657, 656]. Quasi [122]. Quasi-Optimal [122]. Queries [459, 1316, 458, 1333, 591, 1450, 1456]. Query [1331, 1311, 378, 479, 904, 1294, 1293, 1296, 1148, 31, 1327, 1326, 1315, 142]. Query-Translation [1315]. Querying [390, 1310]. Quest [367, 789]. Question [449]. Queues [356]. Quick [1009]. Quickly [1051]. Quotient [1434].

R [1385]. Radar [1754]. Rademacher [130]. Radial [1624, 1621, 1701, 1668]. Radio [371, 372, 622, 625, 1242, 631]. Railway [740, 1404]. Random [421, 666, 382, 618, 1098, 1587, 128, 851, 1326, 867]. Randomized [417, 631]. Randomizing [361]. Randomly [494, 853]. Range [1447, 378, 383]. Range-Encoded [1447]. Rank [1199]. Rank-Revealing [1199]. Ranking [624, 504]. Rapid [1641]. Rat [674]. Rate [1598, 1669, 1592, 1605]. Rates [1752]. Ratio [1136, 1405]. Rational [172]. Ray [1161]. Razor [638]. RBF [1685, 1664]. RDF [819, 116]. Re [514]. Re-keying [514]. Reachability [349, 653]. Reactive [1530]. Reactivity [1189]. Read [735]. Reader [1024, 1046]. Reading [972, 1007, 1013, 1539]. Real [381, 1552, 1102, 1107, 422, 679, 1531, 211, 1176, 348, 601, 886, 1304]. Real-Number [381]. Real-Time [1102, 1107, 679, 1531, 211]. Real-Weighted [348]. Real-World [886]. Realistic [1107]. Reality [960, 1040, 959]. Realizability [395]. Realization [964]. Realizers [429]. Realizing [1172]. Reason [433]. Reasoning [114, 782, 1376, 1420, 50, 1387, 439, 567, 586, 764, 444, 1361, 794, 575, 840, 774, 1386, 1430]. Receptive [1574]. Receptor [1607]. Recogniser [1062]. Recognising [856, 1669]. Recognition [862, 562, 895, 1639, 675, 898, 666, 706, 680, 915, 870, 906, 1750, 1627, 684, 907, 883, 897, 909, 1523, 564, 842, 864, 667, 368, 869, 1714, 1763, 672, 1774, 1518, 922, 1622, 707, 1463, 497, 858, 1766, 877, 910, 690]. Recognizability [293]. Recognizing [213]. Recommendation [1472]. Recommending [938]. Recompilation [736]. Reconfigurable [1190, 1224, 1179]. Reconfiguration [6, 5]. Reconstruction [344, 1758]. Record [1506]. Records [1436]. Recovery [514, 512]. Recurrence [1145]. Recurrent [1781, 1617, 1660, 1789, 1753, 1787, 1701, 1602]. Recursive [459, 790, 421, 1616, 73, 1522, 905, 896, 410, 305, 1438]. Reddogs [273]. Reduce [1441]. Reduced [541]. Reducing [120, 1589, 845]. Reduction [1093, 1192, 290, 896, 894, 1664]. Reductions [374]. Redundancy [1174]. Reed [360]. Reengineering [330, 335]. References [705]. Refined [1307]. Refinement [755, 1093, 1094, 33, 747, 754, 593]. Reflection [100, 794]. Reflex [1585]. Reformulating [53]. Reformulation [41, 49, 52, 22, 59]. Refutations [358]. Region [175, 1774, 473, 1016]. Regions [875, 478]. Register [1173]. Regression [1644, 1680, 1640, 1685, 1655, 1691, 1497]. Regular [354, 620, 618, 847, 859, 98, 66, 591, 295, 1157]. Regularise [927]. Regularity [1468, 401]. Regulatory [1332]. Rehabilitation [1065, 979, 1052]. Reinforcement [1491, 1519, 210, 35, 1694, 1700, 36]. Reject [664]. Related [378, 1308, 1413]. Relation [808]. Relational [726, 810, 588, 742, 682]. Relations [583, 1702, 211, 808, 38]. Relationship [478]. Relationship-Based [478]. Relative [416]. Relaxation [24, 1638]. Relevance [1644, 1727, 1324, 1528]. Relevant [708, 954]. Remarks [718]. REMathEx [1024]. Remote [1266, 1068, 966, 1054, 1186, 164]. Remote-Write [1186]. Removable [365]. Rendering [1107, 1000]. Repeats [96]. Repertoire [1582]. Repetition [629]. Replacement [9, 685]. Replenishment [696]. Replication [16]. Report [1315]. Reports [714]. Repository [1295, 1565]. Representation [288, 1761, 1514, 1619, 1687, 1616, 820, 1036, 299, 811, 892, 807, 1764, 527].

Representations

[1572, 588, 897, 1100, 817, 1614]. Represented [783]. Representing [1391, 317, 808]. repudiation [1091]. Request [1255]. Requests [1149]. Requirements [570, 467, 1043, 252, 517]. Requires [386]. Requiring [1227]. Rerouting [1257]. Resampling [1633]. Rescue [207, 160, 283, 171, 284, 281, 165]. Research [1165, 1237, 1200, 1124, 1125, 1145, 1155, 1150, 188, 1202, 509, 1233, 167, 1209, 1171, 1121, 1146, 1222, 1216, 1185, 51, 1135, 1213, 718, 1159, 1190, 1245, 1210, 1183, 1208, 1195, 1131, 1244, 1141, 1175, 1119, 1158, 1118, 1224, 1215, 57, 1221, 1229, 1186, 1236, 539, 1217, 1178, 1008, 1385, 1196, 1132, 1230, 1187, 1163, 1556, 1174, 1179, 1166, 1170, 1137, 1243, 935]. Residential [1260]. Residual [289]. Resist [528]. Resolution [777, 432, 803, 182, 438, 359, 1245, 693, 1588, 700, 705, 1764]. Resolving [532]. Resonance [1505, 1588]. RESORT [1054]. Resource [739, 1561, 1381, 366, 1185, 1239, 3, 1443, 1477, 2, 1172]. Resource-Constrained [2]. Resource-Flow [1172]. Resources [1286, 1318, 1081, 1329, 1008]. Respina [227]. Response [1592, 1227]. Restoration [1767, 1762, 905]. Restricted [1626]. Restrictions [1321]. Results [1133, 155, 1314, 347, 426, 1005, 581, 406, 648, 34, 396, 935]. Retinal [1597]. Retraining [1780]. Retrieval [498, 495, 507, 1320, 1150, 494, 805, 1316, 491, 479, 482, 1335, 1729, 904, 860, 871, 483, 500, 1317, 478, 1274, 487, 470, 1321, 1334, 1339, 1336, 474, 503, 504, 502, 496, 486, 1673, 497, 1346, 473, 481, 493, 1309, 475]. Retrieve [1526]. Retrieving [1462]. Retroactive [1784]. Returns [416]. Reuse [1140, 1139]. Revealing [1199]. Reverse [1249]. Reverse-Link [1249]. Reversible [655, 355]. Review [843, 727, 954]. Revision [825]. Revisited [180, 439, 657, 289, 1527, 734]. Revocable [546]. Rewrite [303, 295, 305, 296]. Rewriter [311]. Rewriting [302, 300, 312, 291, 304, 404, 306, 289, 287, 293]. RFID [1554]. Rhetorical [708]. Rhythms [1595, 1596]. Rich [1582, 997, 1544]. Ridge [1654]. Rights [1280]. Rings [407, 632]. Risk [1155, 1760]. Risk-Free [1155]. RK [1205]. RMIT [284]. RMIT-on-Fire [284]. Robo [173]. RoboBase [164]. RoboCup [155, 159, 207, 281, 262, 184, 243, 188, 156, 279, 252, 162, 204, 283, 256, 246, 284]. RoboCup-2001 [155]. RoboCup-Rescue [207, 281]. RoboCuppers [250]. RobocupRescue [282]. RoboLog [229]. RoboMutts [274]. RoboRoos [248]. ROBOSIX [262, 243]. RoboSoccer [253]. Robot [185, 166, 167, 159, 191, 194, 196, 1695, 202, 280, 163, 267, 170, 266, 251, 260, 214, 217, 269, 216, 481, 248]. Robotic [210]. Robotics [206]. Robots [1704, 1705, 258, 239, 190, 197, 199, 162, 1558, 169, 212, 174, 175, 1703]. Robust [1743, 186, 1684, 1724, 1756, 718, 929, 865, 1274, 1747, 1746, 472, 1688, 269]. Robustness [1429, 1253]. Rogi [244]. Role [698, 1389, 677, 1614]. Rolling [230]. Roobots [245]. Rosser [353]. Rotation [1374]. Rotations [95]. Rough [1521]. Round [361]. Routed [1211]. Routing [1212, 41, 371, 1241, 1209, 622, 1206, 351, 1617, 1213, 1250, 1251, 1259, 602]. Rovers [270]. RSA [511, 520]. RSTIndex} [1462]. RT [750]. Rule [304, 1710, 596, 1452, 1506, 1304]. Rule-Based [1506]. Rules [1632, 835, 1307, 836, 1444, 1435, 1362, 1443]. Running [193]. Runs [1338]. Runtime [1171]. Russian [1333].

s [968, 785]. S.P.Q.R [263, 275]. Sacrifice [1186]. Safe [121, 61, 119]. Safeness [736]. Safety [1272, 759]. SAHARA [1546]. Salesman [1715]. Saliency [1697, 473].

Salient [508]. Sample [893]. Samples [895, 1662]. Sampling [382, 1635, 1736, 1157]. SAT [780, 651, 599, 1488]. Satellite [1150]. Satisfaction [375, 614, 53, 566]. Satisfiability [778, 1403, 568, 789]. Satisfying [651]. SB [1121]. SB-PRAM [1121]. SBCe [219]. Scalability [1283, 1225, 1115, 1152]. Scalable [1561, 1233, 1567, 767, 1196, 1673]. Scalar [528]. Scale [1220, 874]. SCALEA [1113]. Scaling [660, 810, 1525, 1789, 888]. Scaling-Up [1525]. Scan [1193, 1567]. Scenarii [564]. Scenario [12]. Scenario-Based [12]. Scenarios [467, 1541]. Scene [489, 692]. Scenes [688]. Sceptical [1387]. Schedule [529, 1564]. Scheduler [1189, 1132]. Scheduler-Centric [1189]. Scheduling [1130, 41, 408, 364, 1128, 1613, 1134, 1136, 567, 1131, 406, 1129, 1148, 1219, 1154, 1194, 1543, 1374, 1144]. Schema [1295, 1297, 813, 1461, 1435]. Schemas [708]. Scheme [686, 1613, 1173, 538, 514, 1223, 1267, 477, 523, 536]. Schemes [516, 73, 423, 896, 517, 530]. Schmitt [1695]. School [1552, 947]. Schütte [434]. Schütte-Tait [434]. Science [1026, 1398, 806, 582]. Sciences [1027]. Scientific [708, 1010, 1230]. Scientist [1351]. Scope [336]. Scopes [113]. Scoping [337]. Score [90, 1684, 1329, 1062]. Scores [1524]. Screen [1046]. Script [878]. SDL [758]. SDL-Systems [758]. Seagoing [1272]. Seamless [413]. Search [781, 47, 364, 383, 771, 1378, 860, 27, 1488, 770, 1162, 1593, 24, 1403, 501, 377, 1527, 812]. Searches [1114]. Searching [378]. SearchServer [1337]. Second [1705, 135, 1756, 1347, 308, 867, 451]. Second-Order [1705, 135, 308, 867]. secondary [1078, 1080]. Secret [520, 517, 519]. Section [1065]. Secure [1004, 361, 547, 514, 1090, 535, 536]. Security [1370, 1091, 65, 287, 541, 1186, 543, 530, 531, 512]. See [967]. See-through [967]. Segmentation [881, 911, 1777, 170, 1780, 1580]. Segments [96, 639]. Select [458]. Select-Project [458]. Selected [1108]. Selecting [373, 1654, 1636]. Selection [890, 1644, 1508, 1140, 1783, 302, 732, 1510, 1415, 1680, 403, 201, 1273, 30, 1253, 1718, 1343, 889, 1667, 1657, 250, 891, 1509, 1670, 695, 178, 1676]. Selective [964]. Self [124, 180, 1726, 1704, 1720, 1181, 1721, 1460, 1710, 1445, 1711, 1728, 485, 151, 628, 1727, 1718, 1197, 1708, 724, 1175, 501, 1709, 1730, 1715, 1414, 1502, 486, 1465, 175, 988, 1722, 1717, 205]. Self-Adjusting [1414]. Self-Assembling [628]. Self-indexed [1175]. Self-Localisation [180]. Self-Localization [175, 205]. Self-optimising [1197]. Self-Optimizing [151, 988]. Self-Organising [1460, 1445, 1502, 1465]. Self-Organization [724]. Self-Organizing [1726, 1720, 1721, 1710, 1711, 1728, 485, 1727, 1718, 1708, 501, 1709, 1730, 1715, 486, 1722, 1717]. Self-stabilizing [1181]. Self-Typed [124]. Selfish [351]. Sellers [1538]. SEM [779]. Semantic [498, 464, 346, 499, 1458, 100, 700, 1326, 1310, 1463, 451, 812, 761]. Semantically [997]. Semantics [1423, 1424, 1459, 703, 818, 1535, 343, 1437, 123, 473, 452, 38, 1285]. Semantics-Based [123, 473]. SemCor [1328]. Semi [1661, 1136, 362, 1461, 56, 601, 1457]. Semi-algebraic [362]. Semi-computable [601]. Semi-dynamic [1136]. Semi-Lattice [56]. Semi-structured [1461, 1457]. Semi-supervised [1661]. Semiconcept [832]. Semidefinite [1681]. Semigroups [559]. SEN [952]. Sensation [1018]. Sense [101, 1018, 695]. Sensibility [1632]. Sensing [186, 168]. Sensitive [790, 82, 1675, 1594, 306, 678, 1402, 938]. Sensitivity [1598]. Sensor [1110, 1558, 1706, 1557, 1056, 170, 169].

Sensors [182, 217, 1562]. Sensory [968, 674, 977, 1583]. Sentence [816]. Separability [1666]. Separate [125]. Separation [1756, 322]. Sequence [835, 680, 1451, 574, 97, 606]. Sequencer [1184]. Sequences [85, 1582, 1139, 1572, 373, 1780, 913, 1753, 1666, 1602]. Sequencing [1725]. Sequent [445, 433, 1435, 557]. Sequential [1440, 843, 1655, 23, 918, 1784]. Series [1743, 671, 192]. Server [460, 1255, 1260]. Servers [1233, 1264, 1219]. Service [1266, 1087, 1004, 1258, 1264, 1080, 1262, 1272, 1054, 1260, 1076, 1546, 1565, 18]. Services [1231, 942, 1292, 1000, 1758, 110, 1001]. Session [959]. Set [421, 796, 898, 307, 1225, 627, 778, 414, 1384, 64]. Sets [67, 896, 1153, 1303, 295, 81]. Setting [531]. Several [183]. Severe [968, 1030, 981]. Shading [901, 863]. Shallow [1336]. Shape [862, 666, 487, 901, 863, 472, 861]. Shape-from-Shading [901, 863]. Shared [511, 1192, 1151, 623]. Sharif [246]. Sharing [310, 518, 106, 517, 519, 1178, 484]. Sharpening [638]. Sheffield [1344]. Sherlock [544]. Shift [1502]. Shifting [904]. Shifts [1623]. Ship [1259]. Shock [852]. Short [1102, 283, 1073, 521, 535, 1540]. Shortest [393, 1410, 348]. Shostak [791, 286]. Should [997]. Side [1379, 542, 734]. Sighted [1063]. Sign [966, 965]. Signal [968, 1637, 1732, 927, 659]. Signaling [698, 1254]. Signalling [740]. Signals [1724, 1522, 1711]. Signatures [521, 534]. Signing [521]. Signs [115]. SILC [1056]. Silhouettes [842]. Silicon [668]. SIMD [1223]. Similar [616, 1603, 1574]. Similarities [1503]. Similarity [507, 1440, 1521, 1458, 724, 86]. Simple [120, 184, 109, 528, 1574, 1755, 809, 97, 1073, 569, 623, 1579]. Simplicity [141]. Simplified [438]. Simply [1399]. Simpson [748]. Simulate [587, 358]. Simulated [1633, 201, 1725]. Simulation [1124, 968, 156, 755, 1455, 1164, 160, 283, 284, 165, 1532, 234]. Simulations [1202]. Simulator [219, 281]. SINAI [1328]. Single [1752, 622, 1569, 985, 1732, 1721, 645, 1755]. Single-Hop [622]. Single-Neuron [1569]. Single-Trial [1755]. SIP [1127]. Site [939, 1622, 1029]. Sites [943]. Situation [468, 1723]. Situational [990]. Situations [250]. Size [516, 494, 189, 372, 252, 162, 625, 1635, 246, 251, 68, 248]. Sized [644, 366]. Skeleton [1193, 1191]. Skeletons [1237, 1232]. Skills [1059, 950, 1076]. Skin [1520]. Skin-Color [1520]. Sleep [1732]. Sliced [1447]. Slices [1135]. Slot [748]. Slovakia [1028]. Slow [1582]. Small [1573, 609, 189, 246, 1593, 170, 367, 1562, 248]. Small-World [1573, 1593]. SMO [685]. Smooth [187, 1643]. Smoothed [1709]. SMP [1152]. SNePSwD [825]. Snooker [490]. SOAP [1265]. SOAP-Based [1265]. SoapBox [1556]. Soccer [166, 190, 279, 191, 201, 202, 169, 210, 173, 174, 176, 214, 216]. Social [963, 955]. Societies [1282]. Society [955, 942, 934, 936]. Software [467, 331, 325, 1171, 757, 317, 318, 1364, 1058, 1197, 1, 3, 823, 1064, 76, 965, 954, 1652, 1178, 1560, 332, 2, 1538, 7, 320, 576, 774, 336, 339, 18]. Solomon [360]. Solution [1097, 1163, 1166]. Solutions [26, 37]. Solve [699, 1652]. Solver [1715]. Solvers [789]. Solving [85, 570, 780, 1200, 402, 796, 614, 1116, 591, 772]. SOM [1729]. Some [895, 131, 1381, 1341, 581, 1392, 971, 34, 295, 451]. Somesthetic [1018]. SOMs [1714]. Sony [159, 197, 280]. Sort [1415]. Sorting [356, 641]. Sound [746]. Source [1127, 1241, 1756, 329]. Source-Level [329]. Sources [1202, 1474, 1310]. Space [104, 375, 88, 363, 855, 637, 24, 170, 926, 1737]. Space- [104]. Space-Economical [88]. Spaces [1654]. Spanish [711, 1335, 699]. Spanning [1409, 370, 1159, 602]. Sparrows [265]. Sparse [1200, 1679, 674, 348, 1682]. Sparsifying [1685]. Spatial [390, 478, 892, 1583]. Spatio [211, 1570]. Spatio-temporal [211]. Spatiotemporal [23]. Speaker [912]. Speaking [1025]. Special [950, 954, 947, 959]. Specialization [1482]. Species [913]. Specific [90, 1360, 575, 596]. Specification [1102, 744, 1146, 1360, 388, 1288]. Specifications [781, 1420, 732, 584, 346, 587]. Specified [536]. Specifying [464, 172]. Spectra [1505, 1502]. Spectral [849, 842, 1746]. Spectrally [480]. Speech [968, 706, 713, 911, 1763, 672, 707]. Speed [978, 141, 175]. Speeding [1175, 1700, 1149]. Speeding-up [1700]. Spike [1608, 1713, 1607]. Spike-Driven [1608]. Spike-Timing [1713]. Spiking [1587, 1763]. Spin [666, 493]. Splice [1622]. Splitting [1322, 80]. Spoken [876, 1061, 994]. Spontaneous [1575]. Sports [903]. Spotting [964]. SQL [1450]. Squares [1679]. Squid [1269]. SSA [1664]. Stability [1203, 1646, 1667]. Stability-Based [1667]. Stabilized [1571]. Stabilizing [1209, 1181]. Stable [1371, 778, 1641]. Stackelberg [406]. Stage [1639, 1481]. Staging [1732]. STaMs [1284]. Standard [1072, 948, 173]. Star [354, 1414]. STAR-Tree [1414]. Start [1114]. State [1788, 706, 67, 395, 198, 24, 174, 1565, 44, 1737]. Statecharts [172]. States [1084]. Static [744, 362, 290, 61]. Station [1483]. Statistical [1031, 642, 870, 906, 1323, 722, 1103, 96, 1397, 1486, 142]. Statistics [402, 1756, 1602]. Status [1036]. Stave [1524]. Stealing [1227]. Steiner [607, 1401]. Stemming [1335, 1337]. Step [404, 823, 978, 954, 1700]. Stepwise [1208]. Stereo [179, 1697, 927, 209]. Stimuli [1602]. Stochastic [369, 1102, 1669, 1698, 1100, 1097, 1164, 23, 373, 1593, 1588, 212, 1786]. Stock [1493, 1497]. Stokes [1203]. Stooges [237]. Stopper [80]. Storage [1577, 1228, 1405]. Story [1540]. Strategies [416, 1325, 303, 302, 1093, 1277, 406, 24, 1362, 295, 1008, 1481]. Strategy [400, 943, 1258, 1255, 226, 147, 1114, 177, 311]. Stream [1122, 1264, 522, 524, 1603, 718]. Stream-Processing [1122]. Streaming [1258]. Streams [399, 398]. Strengthening [529]. Stricter [386]. String [402, 82, 304, 93, 89, 80, 851, 403]. Strings [83, 87, 845]. Strong [1385, 401, 530]. Strongly [460, 1571, 425]. Structural [856, 870, 852]. Structurally [879]. Structure [1615, 757, 351, 698, 630, 808, 1442]. Structured [933, 138, 1456, 1461, 1457]. Structures [371, 357, 805, 383, 515, 1011, 1436, 596, 1666, 1463, 1486]. Students [1026, 1024, 1078, 1080, 1028, 1081, 1000, 967, 1076, 1022, 1075, 1027]. Studies [1370]. Studio [1064]. Study [1549, 739, 1241, 732, 167, 1409, 643, 1406, 1468, 590, 317, 1262, 1494, 953, 116, 794, 1082, 1609, 1084, 1480, 1174, 1268, 690]. Studying [1075]. Style [434, 1428, 441, 773]. Sub [1016, 403]. Sub-region [1016]. Subcubic [603]. Subgraph [853]. Subject [1456]. Subjectively [507]. SUBSEQUENCE [85]. Subsequences [373]. Subset [378]. Subsets [921]. Substitution [298]. Substitutions [409]. Subtitled [994]. Subtitles [994]. Subtraction [542]. Subtyping [118]. Successful [321]. Successive [868]. Sufficient [799]. Suffix [637]. Suffixes [89]. Suite [733]. Sum [612]. Summaries [506]. Summarization [505]. Summarize [708]. Summary [51, 57]. Superexponential [304]. Superlogarithmic [424]. Superposition [441]. Supertrees [636]. Supervised [895, 1669, 1662, 1503, 1661]. Supervision [194]. Supplementing [760]. Support [960, 1785, 1549, 1281, 464, 1687, 1112, 675, 1679, 671, 1680, 1476, 664, 668, 1762, 1449, 679, 689, 1776, 678, 3, 1692, 659, 1690, 1747, 1746, 682, 672, 1075, 1509, 677, 670, 687, 1497, 1152]. Supported [1685].

Supporting [108, 164]. Supports [1269]. Surface [428]. Surfaces [1427, 901, 634]. Surfing [461]. Surveillance [1770]. Survey [1016, 675, 1330, 1386, 709, 1293, 1368, 1389, 1046]. Survivable [423]. Survival [1247]. SVD [1664]. SVM [686, 680, 663, 1685, 684, 683, 667, 1758, 690]. Swarm [222, 1482, 1484]. Sweden [277]. Sweep [759, 1412]. Sweep-Line [759]. Sweeping [376]. Symbol [870]. Symbolic [69, 400, 553, 300, 27, 555, 561, 1100, 564, 626, 577, 557, 1123]. Symbolic/Numeric [553]. Symbols [574]. Symmetric [363]. Symmetry [782, 1581, 779]. Synapse [1704, 1576]. Synaptic [1608, 1587]. Synchronization [1216, 1118, 762]. Synchronized [313]. Synchronous [1134, 1249, 1575]. Syntactic [800]. Syntactical [722]. Syntax [1419, 703, 343]. Synthesis [752, 400, 1146, 1363, 394, 554, 788]. Synthesizing [753]. System [179, 1232, 1266, 1269, 185, 1639, 1320, 188, 717, 987, 1087, 1012, 1298, 1048, 555, 940, 1473, 979, 776, 162, 1469, 1347, 975, 1249, 1032, 1272, 1131, 1164, 986, 1251, 1271, 1471, 878, 1557, 1334, 1504, 699, 1056, 163, 966, 1223, 1054, 1034, 1402, 1773, 1747, 486, 721, 1543, 786, 293, 557, 1149, 692, 1673, 455, 165, 1052, 1701, 1067, 216, 788, 15, 481, 1457, 512, 1490, 774, 200, 339]. System-on-Chip [1223]. Systematic [1190, 324]. Systems [1280, 1549, 1184, 1726, 1145, 571, 1619, 507, 989, 186, 1358, 1102, 112, 1283, 560, 6, 408, 1122, 915, 312, 1277, 1037, 1642, 113, 937, 1641, 1190, 1559, 394, 1698, 358, 758, 1273, 1097, 1104, 70, 1265, 1482, 1244, 626, 1328, 1180, 433, 446, 1228, 1287, 55, 747, 1288, 948, 17, 1289, 944, 1182, 1089, 1404, 396, 707, 305, 1571, 1480, 531, 1152]. Systolic [612].

t [968]. t/ [968]. T3 [1448]. Tableau [436, 440, 442, 446, 453]. Tableau-Based [436]. Tableaux [435, 441, 451]. Tables [602]. Tabulation [1397]. Tactical [1419]. Tactile [1033, 1038, 1037, 1032]. Tagged [1172]. Tags [1554]. Tait [434]. Take [955, 935]. Taking [703, 1079]. Talk [1090]. Talkie [195]. Talking [1042, 933]. Tangent [897]. Target [915, 1175]. Task [1134, 709, 1129, 1339, 283]. Task-Graphs [1129]. Tasks [166, 439, 1333, 1544, 1154, 1773]. Taxonomy [1306]. TCM [1261]. TCP [1267]. Teach [948]. Teachers [951]. Teaching [188, 987, 1059, 826, 956]. Team [268, 278, 230, 276, 227, 262, 256, 243, 233, 223, 189, 279, 238, 231, 242, 270, 225, 271, 263, 157, 255, 226, 280, 259, 246, 267, 236, 232, 275, 219, 266, 284, 251, 173, 260, 277, 247, 253, 254, 218, 177, 269, 264, 228, 235, 244, 159]. Teamwork [252, 226]. Technical [1025, 971, 1034]. Technique [120, 1249, 578, 62, 1736]. Techniques [1331, 856, 183, 48, 51, 1135, 1658, 860, 1138, 1322, 866, 1496, 345, 1462, 1586, 743, 1327, 508, 37]. Technologies [942, 1043, 954, 1111, 936]. Technology [468, 1016, 1065, 1087, 951, 1086, 1078, 1053, 944, 1082, 1083, 1084, 1079, 946, 976]. Technology-Based [1078]. TeDUB [1034]. Tele [1054]. Tele-Help [1054]. Telephone [717]. Telephone-Based [717]. Temporal [419, 1376, 711, 766, 438, 300, 1286, 420, 1750, 1574, 1444, 29, 1117, 699, 1570, 1753, 1599, 861, 1602, 830, 59, 211]. Temporally [1155, 1444, 44]. Term [312, 1615, 971, 722, 1739, 293, 1503, 305, 1345]. Term-for-Term [1345]. Termination [1307, 1432, 306, 1361, 178]. Terms [872, 1341, 574, 299]. Terrain [635]. Tertiary [1228]. Test [511, 733, 732, 476, 1338, 1729, 743, 884, 165, 1029]. Test-Bed [165]. Testability [1263]. Testbed [560, 176]. Testing [109, 778, 427, 730, 724, 1103, 385]. Tests [1263]. Texels [859]. Texprint} [879]. Text [1050, 77, 990, 137, 813, 709, 710, 983, 814, 692, 1673, 877, 1489]. Text-to-Scene [692]. Text/Graphics [877]. Texts [720, 1010, 699, 705]. Textual [1025, 1544, 1011]. Texture [930, 681, 1163, 926, 902]. Texture-Based [681]. Textures [879]. TFA [1733, 1493]. Thai [1330]. Their [1697, 51, 478]. Thematic [959]. Theorem [556, 1419, 1422, 429, 1749, 796, 787, 1427, 1428, 1383, 1433, 794, 450, 451, 576, 798]. Theoretic [1423, 30, 794]. Theories [124, 310, 1376, 1158, 34, 286, 453]. Theory [1101, 145, 1616, 509, 73, 1493, 349, 1569, 65, 1597, 904, 764, 581, 404, 342, 1156, 573, 289, 236, 1384, 447, 341, 649, 1771, 1279]. Theory-Based [1493]. Thin [120]. Think [992]. Thomson [1332]. threaded [1354]. Threatening [840]. Three [633, 495, 91, 1494]. Threshold [516, 517, 136]. throughout [999]. Tickets [623]. Tiered [205]. Tiled [1253]. Tiling [1139]. Time [419, 85, 1743, 104, 1102, 402, 1107, 671, 423, 438, 192, 74, 1135, 1258, 679, 1098, 637, 1531, 1244, 437, 211, 624, 1228, 1402, 391, 521, 1789, 138, 425, 1172, 1248, 1737]. Time-Based [1244]. Time-Efficient [104]. Time-Scaling [1789]. Time-Sensitive [1402]. Time-Tagged [1172]. Timed [1184, 384]. Timeouts [586]. Times [1214]. Timetable [1404]. Timing [749, 33, 1713, 1090]. Tissue [626]. TL [974]. TNO [1318]. Token [1181]. Token-Based [1181]. Tolerance [1231, 567]. Tolerant [1184, 1372, 1251, 641, 1089]. Tolerating [547]. Tomography [344]. Tool [876, 960, 1269, 1752, 509, 616, 939, 1126, 312, 1699, 1564, 1113, 756, 1137]. Tools [1112, 1047, 1021, 19]. Top [865]. Topic [1348, 1456]. Topographic [1731]. Topological [643, 1712, 1412]. Topologies [1239]. Topology [1207, 1215]. Torque [161]. Total [388]. Tourist [938]. Tournament [155, 588]. Trace [404, 1174].

Traces [419, 420]. Tracing [1297, 1115, 546, 1161]. tracings [496]. Track [1335, 1342, 1327]. Trackies [261]. Tracking [239, 136]. Tracks [1319, 200]. Tractable [419, 31]. Trade [393, 350, 1089]. Trade-Offs [393, 350, 1089]. Trademarks [485]. Tradeoffs [299]. Trading [1427, 1491]. Traditional [1033]. Traffic [1254]. Trained [1675, 1701]. Training [1633, 895, 898, 1679, 663, 951, 1663, 1659, 979, 1645, 981, 953, 667, 1736, 959, 1071, 956, 1668]. Traitor [546]. Transactional [17, 1478]. Transducer [301]. Transducers [706, 627]. Transducing [293]. Transfer [1576, 809, 540]. Transform [1522, 928]. Transformation [1358, 1297, 67, 1566, 907, 1158, 62]. Transformation-Based [907]. Transformational [584]. Transformations [65]. Transitions [187]. Transitive [701]. Translation [1331, 1050, 1325, 1317, 1318, 1328, 1329, 1323, 1315, 1067, 15]. Translation-Based [1325]. Translations [1343]. Translator [966]. Transmission [1214, 1752, 1249]. Transmit [1261]. Transparent [1231]. Transposer [968, 964]. Travel [959, 1737]. Traveling [1715]. Treatment [702]. Tree [856, 921, 1415, 1220, 192, 137, 122, 1460, 1159, 1151, 313, 607, 294, 1442, 138, 44, 846, 301, 1512, 1414]. Tree-Based [44]. Tree-Tuple [313]. Trees [46, 1441, 1409, 550, 1511, 440, 370, 1650, 640, 1513, 602, 380, 852, 45]. Treewidth [1407]. Trend [1492]. Trends [1016]. Trial [1752, 1755]. Triangle [633]. Trigger [1695]. Triggering [1307]. Trilearn [235]. Triphasic [1595, 1596]. Truck [648, 1543]. True [1094, 893]. Trust [515, 1479]. Trustworthy [767]. Truth [474]. TsinghuAeolus [220]. TTree [44]. TTSLI [313]. Tumours [1505]. Tuning [1785, 1127, 1579]. Tuple [313, 922]. Tutorial [948]. TV [994]. Twine [1561]. Two [1549, 1145, 411, 1639, 25, 1429, 1522, 1430, 83, 95, 546, 894, 1496, 886, 1142, 205]. Two-Class [894]. Two-Dimensional [1522, 1142]. Two-Level [1430, 546]. Two-Pattern [83]. Two-Player [25]. Two-Stage [1639]. Two-Tiered [205]. Two-Way [411]. Type [781, 1767, 1423, 73, 701, 121, 589, 119]. Type-Indexed [589]. Type-Safe [121, 119]. Type-Theoretic [1423]. Typed [124, 460]. Types [589, 118, 66, 1434]. Typhoon [676]. Typing [982]. Typings [418].

Ubiquitous

[1549, 1547, 1548, 1557, 1560, 1556, 15].

UCFL [353]. UDP [1208]. UDP/IP [1208]. Ulm [265]. Ultrametric [666]. Ultrasonic [217]. UML [327, 760]. UML-F [327]. Unanticipated [108]. Unbiased [924]. Uncertain [1453]. Uncertainty [1698]. Understanding [1769]. Unfoldings [1095]. Unification [1418, 574, 308]. Uniform [1189, 426, 1229, 448, 140]. Uniformly [96]. UNILANG [173]. Uninitialized [394]. Union [608, 999]. Unique [88]. Unit [109, 1533]. United [266, 1084]. UniTesK [733]. Units [1621, 723, 1578, 672]. Universal [367, 1011, 936, 397]. Universalization [416]. Universes [1425]. University [261, 279, 1027]. Unknown [1487, 1635]. Unlocking [1006]. Unsafe [1248]. Unscaled [1623]. Unsupervised [1779, 1724, 1732, 1505, 1780, 1766]. UNSW [159]. until [1098]. Untrusted [1555]. Up-Trend [1492]. Updates [72, 61, 132]. Updating [1390]. UPMC [262, 243]. UPMC-CFA [262, 243]. Upper [1052]. UQ [234, 248]. Urban [1015]. Usability [642, 328, 937, 1011]. Usable [1005]. Usage [1513, 954]. Use [1306, 1042, 827, 1043, 98]. Used [282]. Useful [1057]. Usefulness [1016, 328]. User [876, 509, 1468, 1043, 940, 465, 207, 1471, 966, 837, 934, 1035, 471, 1046]. User-Adpated [837]. User-Centered [940]. User-Modeling [471]. Users [1031, 1445, 1041, 1011, 1035, 1019, 984, 1243]. Using [1331, 781, 1767, 856, 1751, 921, 686, 1624, 1768, 746, 1492, 498, 899, 1781, 371, 1632, 511, 1731, 47, 1189, 188, 666, 706, 745, 744, 1775, 671, 1316, 987, 870, 1521, 192, 1297, 1483, 1597, 67, 1025, 1721, 1460, 751, 908, 813, 1762, 1021, 859, 860, 1015, 1699, 27, 1777, 346, 1627, 1445, 684, 485, 1151, 204, 1020, 1655, 679, 338, 1520, 1470, 854, 950, 1505, 1164, 1274, 1523, 487, 66, 1776, 1458, 1321, 842, 326, 1462, 1660, 1662, 209, 1433, 1301, 433, 1692, 814, 236, 1175, 646, 667, 1564, 1158, 1323, 920, 1648, 977, 542, 1770, 1007]. Using [747, 1709, 630, 1054, 504, 1288, 722, 1072, 575, 1324, 1267, 1465, 743, 1326, 596, 1404, 672, 557, 335, 221, 7, 1187, 1532, 1310, 1172, 874, 217, 670, 481, 1256, 1581, 329]. Utaclir [1322]. Utility [1475, 172, 250]. Utility- [250]. Utilizing [1756, 1533]. UvA [235]. UW [276].

Validation [1684, 346, 326, 884, 1506]. Validity [534]. VAlloy [742]. Value [1092, 1168, 51, 1648]. valued [557]. Values [1533]. Variabilities [316]. Variability [1607, 317, 324]. Variable [644, 366]. Variable-Sized [644, 366]. Variables [796, 1192, 587, 832, 574, 138, 734]. Variance [118]. Variance-Based [118]. Variant [309, 1729]. Variation [318, 601, 336]. Variational [1676]. Variations [852]. Various [1014]. Varying [1258]. VBR [1219]. Vector [1785, 1644, 1687, 675, 1679, 671, 1680, 122, 1778, 664, 668, 1628, 1711, 679, 689, 1776, 678, 1692, 659, 1690, 1472, 1747, 1746, 682, 1326, 672, 1509, 677, 670, 687, 1497]. Vector-Based [1326]. Vectors [897, 849]. Vehicle [41, 864]. Veneti [251]. Vergence [1771]. Verification [1775, 67, 757, 792, 751, 1360, 748, 590, 1104, 55, 19, 743, 1774, 768]. Verified [746, 1421]. Verifier [745, 1384]. Verifying [739, 1122, 521]. Version [785, 451]. Versions [541]. versus [551, 110, 656]. Vertex [426, 652]. Vertical [209]. Vertices [621, 370, 867]. Vessels [1272]. VHDL-Based [185]. VI [289]. via [1553, 361, 602, 1475, 892, 1657, 1707]. Vickie [949]. Video [1779, 498, 1220, 490, 1222, 500, 1574, 499, 903, 470, 1526, 1780, 505, 1770, 502, 913, 492, 506, 497]. Video-on-Demand [1220]. Videoconference [1780]. Videophones [1057]. View [168, 1460, 842]. View-Based [842]. Viewpoint [821]. ViperRoos [249, 189]. Virtual [960, 106, 1005, 1040, 742, 947, 996, 959, 962, 1001, 734]. Virtualization [1146]. Viseme [1523]. Vision [179, 1031, 185, 239, 189, 1021, 844, 162, 1550, 209, 163, 240, 241, 174, 1019, 481, 200]. Vision-Based [239, 174, 481]. Visual [1603, 875, 1569, 1036, 485, 499, 1584, 170, 873, 1773, 1017, 1763, 670, 1579, 39]. Visualisation [1230]. Visualization [1647, 1731, 1063, 1517, 207, 1709, 471]. Visually [1016, 970, 939, 974, 1047, 1015, 1010, 1020, 1080, 1027, 1000, 1007, 1011, 1003, 1001, 1071, 1057, 1009]. Visually-Impaired [1080]. Viterbi [1256]. VITIPI [989]. Vivo [1502, 1505]. VLSI [1637, 1777, 615]. VLSI/WSI [615]. Vocabulary [706, 997, 1741]. Vocational [1070]. Voice [1758, 944]. Voice-Enabled [1758]. VoiceXML [944]. VoIP [1252]. Volatile [1497]. Voluntary [1604]. Voting [710]. Vowel [672]. VQ [912]. VQ-based [912]. vs [1429, 353, 1194, 1089].

W [801]. Wagner [429]. Walk

[643, 650, 977]. Walk-disabled [977]. Walkie [195]. Walkie-Talkie [195]. Walking [978, 269, 161]. Walks [851]. Walls [1553]. Warehouse [1305, 1473]. Warehouses [1474, 1148]. warping [877]. Watermark [496]. Watermarking [1651]. Wavelet [931, 1742]. Wavelet-Based [931].

Waves [840]. Way [411, 109, 1292, 806, 762]. Ways [1074, 693]. Weak [298, 1623, 1242, 134]. Weakest [1438]. Wear [1699]. Weather [717]. Web} [823, 1031, 459, 1647, 835, 1044, 1459, 939, 1106, 972, 1406, 819, 1231, 761, 943, 461, 1553, 820, 1081, 1259, 1253, 1462, 630, 837, 1299, 721, 110, 822]. Web-Oriented [820]. Webspace [1309]. Weight [1637, 1257, 1273, 1472]. Weight-Vector [1472]. Weighted [706, 912, 348, 863]. Weighting [1508]. Weights [1408, 1521, 1777]. Welch [1759]. Well [799, 1072]. Well-Foundedness [799]. Wheeled [263]. Whitening [662]. Who [992]. Wide [1406, 1244, 1076, 1106]. Widening [336]. Width [369]. Wiener [1624]. Wilcoxon [1729]. WinBraille [1049, 1051]. Window [1267]. Windows [688, 578, 1046, 1041]. Winnow [136]. Wired [1267]. Wired-cum-Wireless [1267]. Wireless [1241, 1239, 1258, 1245, 1257, 1262, 1099, 1267, 1254, 762]. within [609, 674, 1454, 1580, 96, 1230, 1348]. without [181, 1598, 928, 1178, 848]. Wizards [225]. Word [1045, 724, 816, 1693, 448, 695]. WordNet [1321, 719]. Words [629, 1458, 958]. WordsEye [692]. Work [981]. Working [1333, 1525, 981]. Workloads [1109, 1177, 1152]. Workplace [1069]. Workshop [1347]. World [1573, 1552, 544, 198, 1593, 993, 886, 1304, 1106, 1406]. World-Wide [1106]. Worldwide [1111]. Wormhole [1211]. Wormhole-Routed [1211]. Wright [558, 280]. Write [1186]. Writing [989]. Written [563, 1669, 705]. Wrong [415]. WSD [694]. WSI [615]. WWW [1300].

XML [458, 870, 1473, 1301, 116, 1007, 1221, 1464]. XML-Based [1473, 1221]. XTR [610].

YabAI [160]. Years [1418]. Yielding [141]. Yields [1582]. Yorkist [844]. Young [961, 972, 973, 969]. Yourself [577, 1055]. YowAI [232]. YowAI-2001 [232]. Yue [181]. Yue-Fei [181].

z [968, 730]. Zeng01 [236]. Zheng [523]. Zheng-Seberry [523]. Ziv [1519, 78].

References

Meur:2002:EBC

[1] Anne-Françoise Le Meur, Charles Consel, and Benoît Escrig. An environment for building customizable software components. Lecture Notes in Computer Science, 2370:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2370/23700001.htm;   http:/
                                              /link.springer-ny.com/link/service/
                                              series/0558/papers/2370/23700001.
                                              pdf.

Sommer:2002:CBA

[2] Nicolas Le Sommer and Frédéric Guidec. A contract-based approach of resource-constrained software deployment. Lecture Notes in Computer Science, 2370:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2370/23700015.htm;   http:/
                                              /link.springer-ny.com/link/service/
                                              series/0558/papers/2370/23700015.
                                              pdf.

Mikic-Rakic:2002:ALS

[3] Marija Mikic-Rakic and Nenad Medvidovic. Architecture-level support for software component deployment in resource constrained environments. Lecture Notes in Computer Science, 2370: 31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700031.
pdf.

Eisenbach:2002:EDJ

[4] Susan Eisenbach, Chris Sadler, and Shakil Shaikh. Evolution of distributed Java programs. Lecture Notes in Computer Science, 2370:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700051.
pdf.

Rutherford:2002:REJ

[5] Matthew J. Rutherford, Kenneth Anderson, Antonio Carzaniga, Dennis Heimbigner, and Alexander L. Wolf. Reconfiguration in the enterprise JavaBean component model. Lecture Notes in Computer Science, 2370:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700067.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2370/23700067.
                                        pdf.

Chen:2002:CFD

[6] Xuejun Chen and Martin Simons. A component framework for dynamic reconfiguration of distributed systems. Lecture Notes in Computer Science, 2370:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2370/23700082.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2370/23700082.
                                        pdf.

Sudmann:2002:SDU

[7] Nils P. Sudmann and Dag Johansen. Software deployment using mobile agents. Lecture Notes in Computer Science, 2370:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2370/23700097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2370/23700097.
                                        pdf.

Hissam:2002:PPA

[8] Scott A. Hissam, Gabriel A. Moreno, Judith A. Stafford, and Kurt C. Wallnau. Packaging predictable assembly. Lecture Notes in Computer Science, 2370:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2370/23700108.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2370/23700108.
pdf.

Costanza:2002:DRA

[9] Pascal Costanza. Dynamic replacement of active objects in the gilgul programming language. Lecture Notes in Computer Science, 2370:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700125.
pdf.

Assmann:2002:BGC

[10] Uwe Aßmann. Beyond generic component parameters. Lecture Notes in Computer Science, 2370:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700141.
pdf.

Kniesel:2002:CCC

[11] Günter Kniesel and Michael Austermann. CC4J — code coverage for Java. Lecture Notes in Computer Science, 2370:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700155.
pdf.

Lowe:2002:SBC

[12] Welf Löwe and Markus Noga. Scenariobased connector optimization. Lecture Notes in Computer Science, 2370:170– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700170.
pdf.

Bracciali:2002:ACM

[13] Andrea Bracciali, Antonio Brogi, and Carlos Canal. Adapting components with mismatching behaviours. Lecture Notes in Computer Science, 2370:185– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700185.
pdf.

Nierstrasz:2002:CMF

[14] Oscar Nierstrasz, Gabriela Arévalo, Stéphane Ducasse, Roel Wuyts, Andrew P. Black, Peter O. Müller, Christian Zeidler, Thomas Genssler, and Reinier van den Born. A component model for field devices. Lecture Notes in Computer Science, 2370:200– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700200.
pdf.

Wichadakul:2002:TSE

[15] Duangdao Wichadakul and Klara Nahrstedt. A translation system for enabling flexible and efficient deployment of QoS-aware applications in ubiquitous environments. Lecture Notes in Computer Science, 2370:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700210.
pdf.

Marangozova:2002:ICC

[16] Vania Marangozova and Daniel Hagimont. An infrastructure for CORBA component replication. Lecture Notes in Computer Science, 2370:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2370/23700222.
pdf.

Poernomo:2002:AES

[17] Iman Poernomo, Ralf Reussner, and Heinz Schmidt. Architectures of enterprise systems: Modelling transactional contexts. Lecture Notes in Computer Science, 2370:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2370/23700233.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2370/23700233.
                                        pdf.

vonLaszewski:2002:SCS

[18] Gregor von Laszewski, Eric Blau, Michael Bletzinger, Jarek Gawor, Peter Lane, Stuart Martin, and Michael Russell. Software, component, and service deployment in computational grids. Lecture Notes in Computer Science, 2370:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2370/23700244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2370/23700244.
                                        pdf.

Rapicault:2002:MNT

[19] Pascal Rapicault, Jean-Paul Rigault, and Luc Bourlier. Model, notation, and tools for verification of protocolbased components assembly. Lecture Notes in Computer Science, 2370:257– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2370/23700257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2370/23700257.
                                        pdf.

Anonymous:2002:AIa

[20] Anonymous. Author index. Lecture Notes in Computer Science, 2370:269– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2370/2370auth.pdf.

Kurshan:2002:MCA

[21] Robert P. Kurshan. Model checking and abstraction. Lecture Notes in Computer Science, 2371:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710001.
pdf.

Long:2002:RP

[22] Derek Long, Maria Fox, and Muna Hamdi. Reformulation in planning. Lecture Notes in Computer Science, 2371:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2371/23710018.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2371/23710018.
pdf.

Mahadevan:2002:SAS

[23] Sridhar Mahadevan. Spatiotemporal abstraction of stochastic sequential processes. Lecture Notes in Computer Science, 2371:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710033.
pdf.

Mingozzi:2002:SSR

[24] Aristide Mingozzi. State space relaxation and search strategies in dynamic programming. Lecture Notes in Computer Science, 2371:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710051.
pdf.

Cazenave:2002:AMT

[25] Tristan Cazenave. Admissible moves in two-player games. Lecture Notes in Computer Science, 2371:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710052.
pdf.

Choueiry:2002:DBL

[26] Berthe Y. Choueiry and Amy M. Davis. Dynamic bundling: Less effort for more solutions. Lecture Notes in Computer Science, 2371:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710064.
pdf.

Hansen:2002:SHS

[27] Eric Hansen, Rong Zhou, and Zhengzhu Feng. Symbolic heuristic search using decision diagrams. Lecture Notes in Computer Science, 2371:

83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710083.
pdf.

Heymann:2002:CHA

[28] Michael Heymann and Asaf Degani. On the construction of human-automation interfaces by formal abstraction. Lecture Notes in Computer Science, 2371: 99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710099.
pdf.

Khatib:2002:POT

[29] Lina Khatib, Paul Morris, and Robert Morris. Pareto optimization of temporal decisions. Lecture Notes in Computer Science, 2371:116–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710116.
pdf.

Kumar:2002:ITC

[30] T. K. Satish Kumar. An informationtheoretic characterization of abstraction in diagnosis and hypothesis selection. Lecture Notes in Computer Science, 2371:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2371/23710126.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2371/23710126.pdf.

Miranker:2002:TQC

[31] Daniel Miranker, Malcolm C. Taylor, and Anand Padmanaban. A tractable query cache by approximation. Lecture Notes in Computer Science, 2371: 140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710140.
                                        pdf.

Mukhopadhyay:2002:AFA

[32] Supratik Mukhopadhyay and Andreas Podelski. An algebraic framework for abstract model checking. Lecture Notes in Computer Science, 2371:152– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710152.
                                        pdf.

Neller:2002:ATD

[33] Todd W. Neller. Action timing discretization with iterative-refinement. Lecture Notes in Computer Science, 2371:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2371/23710170.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2371/23710170.
pdf.

Parmar:2002:FAO

[34] Aarati Parmar. Formalizing approximate objects and theories: Some initial results. Lecture Notes in Computer Science, 2371:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2371/23710178.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2371/23710178.pdf.

Ravindran:2002:MMH

[35] Balaraman Ravindran and Andrew G. Barto. Model minimization in hierarchical reinforcement learning. Lecture Notes in Computer Science, 2371:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710196.
pdf.

Stolle:2002:LOR

[36] Martin Stolle and Doina Precup. Learning options in reinforcement learning. Lecture Notes in Computer Science, 2371:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2371/23710212.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2371/23710212.pdf.

Vu:2002:ATN

[37] Xuan-Ha Vu, Djamila Sam-Haroud, and Marius-Calin Silaghi. Approximation techniques for non-linear problems with continuum of solutions. Lecture Notes in Computer Science, 2371:224– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710224.
                                        pdf.

Zanuttini:2002:ARP

[38] Bruno Zanuttini. Approximation of relations by propositional formulas: Complexity and semantics. Lecture Notes in Computer Science, 2371:242– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710242.
                                        pdf.

Zucker:2002:AVP

[39] Jean-Daniel Zucker, Nicolas Bredeche, and Lorenza Saitta. Abstracting visual percepts to learn concepts. Lecture Notes in Computer Science, 2371: 256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710256.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2371/23710256.
pdf.

Apolloni:2002:PMB

[40] Bruno Apolloni, Fabio Baraghini, and Giorgio Palmas. PAC meditation on Boolean formulas. Lecture Notes in Computer Science, 2371:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710274.
pdf.

Beck:2002:RVR

[41] J. Christopher Beck, Patrick Prosser, and Evgeny Selensky. On the reformulation of vehicle routing problems and scheduling problems. Lecture Notes in Computer Science, 2371:282– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710282.
pdf.

Kumar:2002:OCM

[42] T. K. Satish Kumar and Richard Dearden. The oracular constraints method. Lecture Notes in Computer Science, 2371:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2371/23710290.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2371/23710290.
pdf.

Levner:2002:PLC

[43] Ilya Levner, Vadim Bulitko, Omid Madani, and Russell Greiner. Performance of lookahead control policies in the face of abstractions and approximations. Lecture Notes in Computer Science, 2371:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710299.
pdf.

Uther:2002:TTB

[44] William T. B. Uther and Manuela M. Veloso. TTree: Tree-based state generalization with temporally abstract actions. Lecture Notes in Computer Science, 2371:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710308.
pdf.

Zhang:2002:ODI

[45] Jun Zhang, Adrian Silvescu, and Vasant Honavar. Ontology-driven induction of decision trees at multiple levels of abstraction. Lecture Notes in Computer Science, 2371:316– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2371/23710316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710316.
pdf.

Billings:2002:AII

[46] Darse Billings. Abstracting imperfect information game trees. Lecture Notes in Computer Science, 2371:324– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710324.
pdf.

Botea:2002:UAH

[47] Adi Botea. Using abstraction for heuristic search and planning. Lecture Notes in Computer Science, 2371:326– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710326.
pdf.

Bowling:2002:ATM

[48] Michael Bowling. Approximation techniques in multiagent learning. Lecture Notes in Computer Science, 2371:328– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710328.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2371/23710328.
                                        pdf.

Buettner:2002:ARG

[49] Daniel Buettner. Abstraction and reformulation in GraphPlan. Lecture Notes in Computer Science, 2371:330– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710330.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710330.
                                        pdf.

Clement:2002:ARP

[50] Bradley J. Clement. Abstract reasoning for planning and coordination. Lecture Notes in Computer Science, 2371: 331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710331.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710331.
                                        pdf.

Gelman:2002:RSA

[51] Irit Askira Gelman. Research summary: Abstraction techniques, and their value. Lecture Notes in Computer Science, 2371:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710333.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710333.
                                        pdf.

Glaubius:2002:RNB

[52] Robert Glaubius. Reformulation of non-binary constraints. Lecture Notes in Computer Science, 2371:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710335.
pdf.

Kumar:2002:RCO

[53] T. K. Satish Kumar. Reformulating combinatorial optimization as constraint satisfaction. Lecture Notes in Computer Science, 2371:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710336.
pdf.

McGovern:2002:ADA

[54] Amy McGovern. Autonomous discovery of abstractions through interaction with an environment. Lecture Notes in Computer Science, 2371:338– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710338.
pdf.

Oishi:2002:IVD

[55] Meeko Oishi. Interface verification: Discrete abstractions of hybrid systems. Lecture Notes in Computer Science, 2371:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710340.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710340.
                                        pdf.

Okubo:2002:LSL

[56] Yoshiaki Okubo and Xiaobo Li. Learning semi-lattice codebooks for image compression. Lecture Notes in Computer Science, 2371:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710342.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710342.
                                        pdf.

Pickett:2002:RS

[57] Marc Pickett. Research summary. Lecture Notes in Computer Science, 2371: 344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2371/23710344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2371/23710344.
                                        pdf.

Ruml:2002:PEH

[58] Wheeler Ruml. Principled exploitation of heuristic information. Lecture Notes in Computer Science, 2371:345– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710345.
pdf.

Xu:2002:RTC

[59] Lin Xu. Reformulation of temporal constraint networks. Lecture Notes in Computer Science, 2371:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2371/23710347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2371/23710347.
pdf.

Anonymous:2002:AIb

[60] Anonymous. Author index. Lecture Notes in Computer Science, 2371:349– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2371/2371auth.pdf.

Shankar:2002:SAS

[61] Natarajan Shankar. Static analysis for safe destructive updates in a functional language. Lecture Notes in Computer Science, 2372:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2372/23720001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2372/23720001.
                                        pdf.

Potikas:2002:TTD

[62] Petros Potikas, Panos Rondogiannis, and Manolis Gergatsoulis. A transformation technique for datalog programs based on non-deterministic constructs. Lecture Notes in Computer Science, 2372:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2372/23720025.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2372/23720025.
                                        pdf.

Kakehi:2002:DPA

[63] Kazuhiko Kakehi, Robert Glück, and Yoshihiko Futamura. On deforesting parameters of accumulating maps. Lecture Notes in Computer Science, 2372: 46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2372/23720046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2372/23720046.
                                        pdf.

Osorio:2002:EAS

[64] Mauricio Osorio, Juan A. Navarro, and José Arrazola. Equivalence in answer set programming. Lecture Notes in Computer Science, 2372:57– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2372/23720057.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2372/23720057.
pdf.

Delzanno:2002:PTT

[65] Giorgio Delzanno and Sandro Etalle. Proof theory, transformations, and logic programming for debugging security protocols. Lecture Notes in Computer Science, 2372:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2372/23720076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2372/23720076.
pdf.

Leuschel:2002:ACP

[66] Michael Leuschel and Stefan Gruner. Abstract conjunctive partial deduction using regular types and its application to model checking. Lecture Notes in Computer Science, 2372:91– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2372/23720091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2372/23720091.
pdf.

Fioravanti:2002:VSI

[67] Fabio Fioravanti, Alberto Pettorossi, and Maurizio Proietti. Verification of sets of infinite state processes using program transformation. Lecture Notes in Computer Science, 2372:111– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2372/23720111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2372/23720111.
                                        pdf.

Vanhoof:2002:WSD

[68] Wim Vanhoof and Maurice Bruynooghe. When size does matter. Lecture Notes in Computer Science, 2372:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2372/23720129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2372/23720129.
                                        pdf.

Albert:2002:SPM

[69] Elvira Albert and Germán Vidal. Symbolic profiling for multi-paradigm declarative languages. Lecture Notes in Computer Science, 2372:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2372/23720148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2372/23720148.
                                        pdf.

Lau:2002:COO

[70] Kung-Kiu Lau and Mario Ornaghi. Correct object-oriented systems in computational logic. Lecture Notes in Computer Science, 2372:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2372/23720168.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2372/23720168.
pdf.

Ferre:2002:FDE

[71] Sébastien Ferré and Olivier Ridoux. A framework for developing embeddable customized logics. Lecture Notes in Computer Science, 2372:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2372/23720191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2372/23720191.
pdf.

Alferes:2002:CEA

[72] José J. Alferes, Antonio Brogi, João A. Leite, and Luís M. Pereira. Computing environment-aware agent behaviours with logic program updates. Lecture Notes in Computer Science, 2372:216– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2372/23720216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2372/23720216.
pdf.

Caldwell:2002:EGR

[73] James L. Caldwell. Extracting general recursive program schemes in nuprl’s type theory. Lecture Notes in Computer Science, 2372:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2372/23720233.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2372/23720233.
                                        pdf.

Ferrari:2002:EET

[74] Mauro Ferrari, Camillo Fiorentini, and Mario Ornaghi. Extracting exact time bounds from logical proofs. Lecture Notes in Computer Science, 2372:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2372/23720245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2372/23720245.
                                        pdf.

Anonymous:2002:AIc

[75] Anonymous. Author index. Lecture Notes in Computer Science, 2372:267– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2372/2372auth.pdf.

Ogasawara:2002:PSA

[76] Jun Ogasawara and Shinichi Morishita. Practical software for aligning ESTs to human genome. Lecture Notes in Computer Science, 2373:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730001.
                                        pdf.

Arimura:2002:ETM

[77] Hiroki Arimura. Efficient text mining with optimized pattern discovery. Lecture Notes in Computer Science, 2373: 17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730017.
pdf.

Rytter:2002:ALZ

[78] Wojciech Rytter. Application of Lempel–Ziv factorization to the approximation of grammar-based compression. Lecture Notes in Computer Science, 2373:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2373/23730020.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2373/23730020.pdf.

Wan:2002:BML

[79] Raymond Wan and Alistair Moffat. Block merging for off-line compression. Lecture Notes in Computer Science, 2373:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2373/23730032.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2373/23730032.
pdf.

Rautio:2002:SMS

[80] Jussi Rautio, Jani Tanninen, and Jorma Tarhio. String matching with stopper encoding and code splitting. Lecture Notes in Computer Science, 2373:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730042.
                                        pdf.

Vialette:2002:PMP

[81] Stéphane Vialette. Pattern matching problems over 2-interval sets. Lecture Notes in Computer Science, 2373:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730053.
                                        pdf.

Chakaravarthy:2002:PCS

[82] Venkatesan T. Chakaravarthy and Rajasekar Krishnamurthy. The problem of context sensitive string matching. Lecture Notes in Computer Science, 2373: 64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730064.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730064.
                                        pdf.

Franek:2002:TPS

[83] František Franek, Jiandong Jiang, Weilin Lu, and William F. Smyth. Two-pattern strings. Lecture Notes in Computer Science, 2373:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730076.
pdf.

Shapira:2002:EDM

[84] Dana Shapira and James A. Storer. Edit distance with move operations. Lecture Notes in Computer Science, 2373:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2373/23730085.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2373/23730085.
pdf.

Alber:2002:TOS

[85] Jochen Alber, Jens Gramm, Jiong Guo, and Rolf Niedermeier. Towards optimally solving the LONGEST COMMON SUBSEQUENCE problem for sequences with nested arc annotations in linear time. Lecture Notes in Computer Science, 2373:99– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730099.
pdf.

Makinen:2002:LSB

[86] Veli Mäkinen and Esko Ukkonen. Local similarity based point-pattern matching. Lecture Notes in Computer Science, 2373:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2373/23730115.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2373/23730115.pdf.

Iliopoulos:2002:IOM

[87] Costas S. Iliopoulos, Christos Makris, Spiros Sioutas, Athanasios Tsakalidis, and Kostas Tsichlas. Identifying occurrences of maximal pairs in multiple strings. Lecture Notes in Computer Science, 2373:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730133.
                                        pdf.

Hon:2002:SEA

[88] Wing-Kai Hon and Kunihiko Sadakane. Space-economical algorithms for finding maximal unique matches. Lecture Notes in Computer Science, 2373:144– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730144.
                                        pdf.

Inenaga:2002:MDA

[89] Shunsuke Inenaga, Masayuki Takeda, Ayumi Shinohara, Hiromasa Hoshino, and Setsuo Arikawa. The minimum DAWG for all suffixes of a string and its applications. Lecture Notes in Computer Science, 2373:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730153.
pdf.

Akutsu:2002:CDP

[90] Tatsuya Akutsu, Hideo Bannai, Satoru Miyano, and Sascha Ott. On the complexity of deriving position specific score matrices from examples. Lecture Notes in Computer Science, 2373:168– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730168.
pdf.

Crochemore:2002:THM

[91] Maxime Crochemore, Costas S. Iliopoulos, Thierry Lecroq, Wojciech Plandowski, and Wojciech Rytter. Three heuristics for δ-matching:δ-BM algorithms. Lecture Notes in Computer Science, 2373:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730178.
pdf.

Brejova:2002:BML

[92] Brona Brejová and Tomáš Vinar. A better method for length distribution modeling in HMMs and its application to gene finding. Lecture Notes in Computer Science, 2373:190– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730190.
                                        pdf.

Hyyro:2002:FBP

[93] Heikki Hyyrö and Gonzalo Navarro. Faster bit-parallel approximate string matching. Lecture Notes in Computer Science, 2373:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2373/23730203.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2373/23730203.pdf.

Burkhardt:2002:OGG

[94] Stefan Burkhardt and Juha Kärkkäinen. One-gapped q-gram filtersfor Levenshtein distance. Lecture Notes in Computer Science, 2373:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730225.
                                        pdf.

Fredriksson:2002:OEF

[95] Kimmo Fredriksson, Gonzalo Navarro, and Esko Ukkonen. Optimal exact and fast approximate two dimensional pattern matching allowing rotations. Lecture Notes in Computer Science, 2373: 235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730235.
pdf.

Sahinalp:2002:SIU

[96] S. Cenk Sahinalp, Evan Eichler, Paul Goldberg, Petra Berenbrink, Tom Friedetzky, and Funda Ergun. Statistical identification of uniformly mutated segments within repeats. Lecture Notes in Computer Science, 2373:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730249.
pdf.

Muthukrishnan:2002:SPS

[97] S. Muthu Muthukrishnan and S. Cenk Sahinalp. Simple and practical sequence nearest neighbors with block operations. Lecture Notes in Computer Science, 2373:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2373/23730262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2373/23730262.
pdf.

Ilie:2002:CNO

[98] Lucian Ilie and Sheng Yu. Constructing NFAs by optimal use of positions in regular expressions. Lecture Notes in Computer Science, 2373:279– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2373/23730279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2373/23730279.
                                        pdf.

Anonymous:2002:AId

[99] Anonymous. Author index. Lecture Notes in Computer Science, 2373:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2373/2373ind.pdf.

Meseguer:2002:SMD

[100] José Meseguer and Carolyn Talcott. Semantic models for distributed object reflection. Lecture Notes in Computer Science, 2374:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740001.
                                        pdf.

Kienzle:2002:ADI

[101] Jörg Kienzle and Rachid Guerraoui. AOP: Does it make sense? the case of concurrency and failures. Lecture Notes in Computer Science, 2374:37– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740037.
pdf.

Ichisugi:2002:DBM

[102] Yuuji Ichisugi and Akira Tanaka. Difference-based modules: A classindependent module mechanism. Lecture Notes in Computer Science, 2374: 62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740062.
pdf.

Ostermann:2002:DCC

[103] Klaus Ostermann. Dynamically composable collaborations with delegation layers. Lecture Notes in Computer Science, 2374:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740089.
pdf.

Bacon:2002:STE

[104] David F. Bacon, Stephen J. Fink, and David Grove. Space- and time-efficient implementation of the Java object model. Lecture Notes in Computer Science, 2374:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740111.
                                        pdf.

Hovemeyer:2002:AIJ

[105] David Hovemeyer, William Pugh, and Jaime Spacco. Atomic instructions in Java. Lecture Notes in Computer Science, 2374:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740133.
                                        pdf.

Czajkowski:2002:CSA

[106] Grzegorz Czajkowski, Laurent Daynès, and Nathaniel Nystrom. Code sharing among virtual machines. Lecture Notes in Computer Science, 2374:155– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740155.
                                        pdf.

Tilevich:2002:JOA

[107] Eli Tilevich and Yannis Smaragdakis. J-orchestra: Automatic Java application partitioning. Lecture Notes in Computer Science, 2374:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740178.
pdf.

Redmond:2002:SUD

[108] Barry Redmond and Vinny Cahill. Supporting unanticipated dynamic adaptation of application behaviour. Lecture Notes in Computer Science, 2374:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2374/23740205.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2374/23740205.
pdf.

Cheon:2002:SPA

[109] Yoonsik Cheon and Gary T. Leavens. A simple and practical approach to unit testing: The JML and JUnit way. Lecture Notes in Computer Science, 2374: 231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740231.
pdf.

Szyperski:2002:OCV

[110] Clemens Szyperski. Objectively: Components versus Web services. Lecture Notes in Computer Science, 2374:256– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2374/23740256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740256.
                                        pdf.

Rinat:2002:MIP

[111] Ran Rinat and Scott Smith. Modular Internet programming with cells. Lecture Notes in Computer Science, 2374: 257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740257.
                                        pdf.

Bryce:2002:LAP

[112] Ciarán Bryce, Chrislain Razafimahefa, and Michel Pawlak. Lana: An approach to programming autonomous systems. Lecture Notes in Computer Science, 2374:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740281.
                                        pdf.

Fiege:2002:EEB

[113] Ludger Fiege, Mira Mezini, Gero Mühl, and Alejandro P. Buchmann. Engineering event-based systems with scopes. Lecture Notes in Computer Science, 2374:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2374/23740309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740309.
pdf.

Aldrich:2002:ARA

[114] Jonathan Aldrich, Craig Chambers, and David Notkin. Architectural reasoning in ArchJava. Lecture Notes in Computer Science, 2374:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740334.
pdf.

Noble:2002:PS

[115] James Noble and Robert Biddle. Patterns as signs. Lecture Notes in Computer Science, 2374:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740368.
pdf.

Neumann:2002:PBD

[116] Gustaf Neumann and Uwe Zdun. Pattern-based design and implementation of an XML and RDF parser and interpreter: A case study. Lecture Notes in Computer Science, 2374: 392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740392.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740392.
                                        pdf.

Benton:2002:MCA

[117] Nick Benton, Luca Cardelli, and Cédric Fournet. Modern concurrency abstractions for C^{♯}. Lecture Notes in Computer Science, 2374:415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740415.
                                        pdf.

Igarashi:2002:VBS

[118] Atsushi Igarashi and Mirko Viroli. On variance-based subtyping for parametric types. Lecture Notes in Computer Science, 2374:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740441.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2374/23740441.
                                        pdf.

Zenger:2002:TSP

[119] Matthias Zenger. Type-safe prototypebased component evolution. Lecture Notes in Computer Science, 2374:470– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2374/23740470.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2374/23740470.
pdf.

Arnold:2002:TGS

[120] Matthew Arnold and Barbara G. Ryder. Thin guards: A simple and effective technique for reducing the penalty of dynamic class loading. Lecture Notes in Computer Science, 2374:498– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740498.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740498.
pdf.

Glew:2002:TSM

[121] Neal Glew and Jens Palsberg. Typesafe method inlining. Lecture Notes in Computer Science, 2374:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740525.
pdf.

Filman:2002:PEB

[122] Robert E. Filman. Polychotomic encoding: A better quasi-optimal bitvector encoding of tree hierarchies. Lecture Notes in Computer Science, 2374: 545–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740545.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2374/23740545.
pdf.

Snelting:2002:SBC

[123] Gregor Snelting and Frank Tip. Semantics-based composition of class hierarchies. Lecture Notes in Computer Science, 2374:562–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2374/23740562.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2374/23740562.pdf.

Alagic:2002:BCS

[124] Suad Alagić and Svetlana Kouznetsova. Behavioral compatibility of selftyped theories. Lecture Notes in Computer Science, 2374:585–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740585.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740585.
pdf.

Ancona:2002:FFJ

[125] Davide Ancona, Giovanni Lagorio, and Elena Zucca. A formal framework for Java separate compilation. Lecture Notes in Computer Science, 2374:609– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2374/23740609.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2374/23740609.
pdf.

Anonymous:2002:AIe

[126] Anonymous. Author index. Lecture Notes in Computer Science, 2374:637– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2374/2374auth.pdf.

Mendelson:2002:ALN

[127] Shahar Mendelson and Robert C. Williamson. Agnostic learning nonconvex function classes. Lecture Notes in Computer Science, 2375:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750001.
pdf.

Mendelson:2002:ECD

[128] Shahar Mendelson and Roman Vershynin. Entropy, combinatorial dimensions and random averages. Lecture Notes in Computer Science, 2375:14– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750014.
pdf.

Mendelson:2002:GPK

[129] Shahar Mendelson. Geometric parameters of kernel machines. Lecture Notes in Computer Science, 2375:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750029.
pdf.

Bartlett:2002:LRC

[130] Peter L. Bartlett, Olivier Bousquet, and Shahar Mendelson. Localized Rademacher complexities. Lecture Notes in Computer Science, 2375:44– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750044.
pdf.

Bousquet:2002:SLM

[131] Olivier Bousquet, Vladimir Koltchinskii, and Dmitriy Panchenko. Some local measures of complexity of convex hulls and generalization bounds. Lecture Notes in Computer Science, 2375: 59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750059.
pdf.

Takimoto:2002:PKM

[132] Eiji Takimoto and Manfred K. Warmuth. Path kernels and multiplicative updates. Lecture Notes in Computer Science, 2375:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750074.
pdf.

Vyugin:2002:PCI

[133] Michael V. Vyugin and Vladimir V. V’yugin. Predictive complexity and information. Lecture Notes in Computer Science, 2375:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750090.
pdf.

Kalnishkan:2002:MEW

[134] Yuri Kalnishkan and Michael V. Vyugin. Mixability and the existence of weak complexities. Lecture Notes in Computer Science, 2375:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750105.
pdf.

Cesa-Bianchi:2002:SOP

[135] Nicolò Cesa-Bianchi, Alex Conconi, and Claudio Gentile. A secondorder perceptron algorithm. Lecture Notes in Computer Science, 2375:121– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750121.
                                        pdf.

Mesterharm:2002:TLT

[136] Chris Mesterharm. Tracking linearthreshold concepts with winnow. Lecture Notes in Computer Science, 2375: 138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750138.
                                        pdf.

Fernau:2002:LTL

[137] Henning Fernau. Learning tree languages from text. Lecture Notes in Computer Science, 2375:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750153.
                                        pdf.

Suzuki:2002:PTI

[138] Yusuke Suzuki, Ryuta Akanuma, Takayoshi Shoudai, Tetsuhiro Miyahara, and Tomoyuki Uchida. Polynomial time inductive inference of ordered tree patterns with internal structured variables from positive data. Lecture Notes in Computer Science, 2375:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750169.
pdf.

delaHiguera:2002:IDL

[139] Colin de la Higuera and Jose Oncina. Inferring deterministic linear languages. Lecture Notes in Computer Science, 2375:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750185.
pdf.

Zilles:2002:MUI

[140] Sandra Zilles. Merging uniform inductive learners. Lecture Notes in Computer Science, 2375:201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750201.
pdf.

Schmidhuber:2002:SPN

[141] Jürgen Schmidhuber. The speed prior: A new simplicity measure yielding near-optimal computable predictions. Lecture Notes in Computer Science, 2375:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750216.
                                        pdf.

Yang:2002:NLB

[142] Ke Yang. New lower bounds for statistical query learning. Lecture Notes in Computer Science, 2375:229– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750229.
                                        pdf.

Bshouty:2002:ELB

[143] Nader H. Bshouty, Jeffrey C. Jackson, and Christino Tamon. Exploring learnability between exact and PAC. Lecture Notes in Computer Science, 2375: 244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750244.
                                        pdf.

Even-Dar:2002:PBM

[144] Eyal Even-Dar, Shie Mannor, and Yishay Mansour. PAC bounds for multi-armed bandit and Markov decision processes. Lecture Notes in Computer Science, 2375:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2375/23750255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750255.
pdf.

Bshouty:2002:BMD

[145] Nader H. Bshouty and Lynn Burroughs. Bounds for the minimum disagreement problem with applications to learning theory. Lecture Notes in Computer Science, 2375:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750271.
pdf.

Bshouty:2002:PLA

[146] Nader H. Bshouty and Lynn Burroughs. On the proper learning of axis parallel concepts. Lecture Notes in Computer Science, 2375:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750287.
pdf.

Lugosi:2002:CSB

[147] Gábor Lugosi and Nicolas Vayatis. A consistent strategy for boosting algorithms. Lecture Notes in Computer Science, 2375:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2375/23750303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750303.
                                        pdf.

Mannor:2002:CGA

[148] Shie Mannor, Ron Meir, and Tong Zhang. The consistency of greedy algorithms for classification. Lecture Notes in Computer Science, 2375:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750319.
                                        pdf.

Ratsch:2002:MMB

[149] Gunnar Rätsch and Manfred K. Warmuth. Maximizing the margin with boosting. Lecture Notes in Computer Science, 2375:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750334.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2375/23750334.
                                        pdf.

Dasgupta:2002:PGH

[150] Sanjoy Dasgupta. Performance guarantees for hierarchical clustering. Lecture Notes in Computer Science, 2375: 351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2375/23750351.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2375/23750351.
pdf.

Hutter:2002:SOP

[151] Marcus Hutter. Self-optimizing and pareto-optimal policies in general environments based on Bayes-mixtures. Lecture Notes in Computer Science, 2375:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2375/23750364.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2375/23750364.
pdf.

Fortnow:2002:PD

[152] Lance Fortnow and Jack H. Lutz. Prediction and dimension. Lecture Notes in Computer Science, 2375:380– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750380.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750380.
pdf.

Papadimitriou:2002:LI

[153] Christos Papadimitriou. Learning the Internet. Lecture Notes in Computer Science, 2375:396–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2375/23750396.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2375/23750396.
pdf.

Anonymous:2002:AIf

[154] Anonymous. Author index. Lecture Notes in Computer Science, 2375:397– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2375/2375ind.pdf.

Birk:2002:TRD

[155] Andreas Birk, Silvia Coradeschi, and Satoshi Tadokoro. The tournament results of the different leagues of robocup-2001. Lecture Notes in Computer Science, 2377:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770001.
                                        pdf.

Cai:2002:GPL

[156] Yunpeng Cai, Jiang Chen, Jinyi Yao, and Shi Li. Global planning from local eyeshot: An implementation of observation-based plan coordination in RoboCup simulation games. Lecture Notes in Computer Science, 2377:12– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770012.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770012.
                                        pdf.

Kiat:2002:LIT

[157] Ng Beng Kiat, Quek Yee Ming, Tay Boon Hock, Yuen Suen Yee, and Simon Koh. LuckyStar II — team description paper. Lecture Notes in Computer Science, 2377:22– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770022.
pdf.

Weigel:2002:CF

[158] Thilo Weigel, Alexander Kleiner, Florian Diesch, Markus Dietl, Jens-Steffen Gutmann, Bernhard Nebel, Patrick Stiegeler, and Boris Szerbakowski. CS freiburg 2001. Lecture Notes in Computer Science, 2377:26– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770026.
pdf.

Chen:2002:URS

[159] Spencer Chen, Martin Siu, Thomas Vogelgesang, Tak Fai Yik, Bernhard Hengst, Son Bao Pham, and Claude Sammut. The UNSW RoboCup 2001 Sony Legged Robot League Team. Lecture Notes in Computer Science, 2377: 39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770039.
pdf.

Morimoto:2002:YFR

[160] Takeshi Morimoto, Kenji Kono, and Ikuo Takeuchi. YabAI: The first rescue simulation league champion. Lecture Notes in Computer Science, 2377: 49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770049.
                                        pdf.

Yamasaki:2002:CMH

[161] Fuminori Yamasaki, Ken Endo, Minoru Asada, and Hiroaki Kitano. A control method for humanoid biped walking with limited torque. Lecture Notes in Computer Science, 2377:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770060.
                                        pdf.

Jamzad:2002:FVS

[162] M. Jamzad, B. S. Sadjad, V. S. Mirrokni, M. Kazemi, H. Chitsaz, A. Heydarnoori, M. T. Hajiaghai, and E. Chiniforooshan. A fast vision system for middle size robots in RoboCup. Lecture Notes in Computer Science, 2377: 71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770071.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2377/23770071.
pdf.

Menegatti:2002:DOV

[163] Emanuele Menegatti, Francesco Nori, Enrico Pagello, Carlo Pellizzari, and Davide Spagnoli. Designing an omnidirectional vision system for a goalkeeper robot. Lecture Notes in Computer Science, 2377:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770081.
pdf.

Sear:2002:REF

[164] John A. Sear and Rupert W. Ford. RoboBase: An extensible framework supporting immediate remote access to logfiles. Lecture Notes in Computer Science, 2377:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770092.
pdf.

Takahashi:2002:ABA

[165] Tomoichi Takahashi, Satoshi Tadokoro, Masayuki Ohta, and Nobuhiro Ito. Agent based approach in disaster rescue simulation — from testbed of multiagent system to practical application. Lecture Notes in Computer Science, 2377:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770102.
                                        pdf.

Buck:2002:PEJ

[166] Sebastian Buck, Michael Beetz, and Thorsten Schmitt. Planning and executing joint navigation tasks in autonomous robot soccer. Lecture Notes in Computer Science, 2377:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770112.
                                        pdf.

Casper:2002:CSH

[167] Jennifer Casper, Mark Micire, Jeff Hyams, and Robin Murphy. A case study of how mobile robot competitions promote future research. Lecture Notes in Computer Science, 2377:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770123.
                                        pdf.

Dietl:2002:CFG

[168] Markus Dietl, Jens-Steffen Gutmann, and Bernhard Nebel. CS freiburg: Global view by cooperative sensing. Lecture Notes in Computer Science, 2377:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770133.
pdf.

Marques:2002:MSN

[169] Carlos F. Marques and Pedro U. Lima. Multi-sensor navigation for soccer robots. Lecture Notes in Computer Science, 2377:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770144.
pdf.

Mitsunaga:2002:VAC

[170] Noriaki Mitsunaga and Minoru Asada. Visual attention control by sensor space segmentation for a small quadruped robot based on information criterion. Lecture Notes in Computer Science, 2377:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770154.
pdf.

Noda:2002:LDR

[171] Itsuki Noda, Tomoichi Takahashi, Shuji Morita, Tetsuhiko Koto, and Satoshi Tadokoro. Language design for rescue agents. Lecture Notes in Computer Science, 2377:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770164.
                                        pdf.

Obst:2002:SRA

[172] Oliver Obst. Specifying rational agents with statecharts and utility functions. Lecture Notes in Computer Science, 2377:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770173.
                                        pdf.

Reis:2002:CUS

[173] Luis Paulo Reis and Nuno Lau. COACH UNILANG — A standard language for coaching a (robo)soccer team. Lecture Notes in Computer Science, 2377:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2377/23770183.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770183.
                                        pdf.

Schmitt:2002:CPS

[174] Thorsten Schmitt, Robert Hanek, Sebastian Buck, and Michael Beetz. Cooperative probabilistic state estimation for vision-based autonomous soccer robots. Lecture Notes in Computer Science, 2377:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2377/23770193.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2377/23770193.pdf.

Sekimori:2002:HSO

[175] Daisuke Sekimori, Tomoya Usui, Yasuhiro Masutani, and Fumio Miyazaki. High-speed obstacle avoidance and self-localization for mobile robots based on omni-directional imaging of floor region. Lecture Notes in Computer Science, 2377:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770204.
pdf.

Stone:2002:KSM

[176] Peter Stone and Richard S. Sutton. Keepaway soccer: A machine learning testbed. Lecture Notes in Computer Science, 2377:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770214.
pdf.

Takahashi:2002:SLT

[177] Yasutake Takahashi, Takashi Tamura, and Minoru Asada. Strategy learning for a team in adversary environments. Lecture Notes in Computer Science, 2377:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770224.
                                        pdf.

Uchibe:2002:EBS

[178] Eiji Uchibe, Masakazu Yanase, and Minoru Asada. Evolutionary behavior selection with activation/ termination constraints. Lecture Notes in Computer Science, 2377:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770234.
                                        pdf.

Adorni:2002:SOD

[179] Giovanni Adorni, Luca Bolognini, Stefano Cagnoni, and Monica Mordonini. Stereo obstacle detection method for a hybrid omni-directional/ pin-hole vision system. Lecture Notes in Computer Science, 2377:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770244.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2377/23770244.
pdf.

Bach:2002:SLR

[180] Joscha Bach and Michael Gollin. Selflocalisation revisited. Lecture Notes in Computer Science, 2377:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770251.
pdf.

Baltes:2002:YFO

[181] Jacky Baltes. Yue-fei: Object orientation and id without additional markers. Lecture Notes in Computer Science, 2377:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770257.
pdf.

Baltes:2002:EIP

[182] Jacky Baltes. Efficient image processing for increased resolution and color correctness of CMOS image sensors. Lecture Notes in Computer Science, 2377:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770263.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2377/23770263.
pdf.

Baltes:2002:CSM

[183] Jacky Baltes and Yongjoo Park. Comparison of several machine learning techniques in pursuit-evasion games. Lecture Notes in Computer Science, 2377:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770269.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770269.
pdf.

Benosman:2002:SAC

[184] Ryad Benosman, Jerome Douret, and Jean Devars. A simple and accurate camera calibration for the F180 RoboCup league. Lecture Notes in Computer Science, 2377:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770275.
pdf.

Bianchi:2002:ICV

[185] Reinaldo A. C. Bianchi and Anna H. Reali-Costa. Implementing computer vision algorithms in hardware: An FPGA/VHDL-based vision system for a mobile robot. Lecture Notes in Computer Science, 2377:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2377/23770281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770281.
pdf.

Bonarini:2002:FRS

[186] Andrea Bonarini, Matteo Matteucci, and Marcello Restelli. A framework for robust sensing in multiagent systems. Lecture Notes in Computer Science, 2377:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770287.
pdf.

Bruce:2002:FPT

[187] James Bruce, Scott Lenser, and Manuela Veloso. Fast parametric transitions for smooth quadrupedal motion. Lecture Notes in Computer Science, 2377:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770293.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770293.
pdf.

Buhler:2002:BPT

[188] Paul Buhler and José M. Vidal. Biter: A platform for the teaching and research of multiagent systems’ design using RoboCup. Lecture Notes in Computer Science, 2377:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770299.
                                        pdf.

Chang:2002:VDL

[189] Mark M. Chang, Brett Browning, and Gordon F. Wyeth. ViperRoos: Developing a low cost local vision team for the small size league. Lecture Notes in Computer Science, 2377:305– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770305.
                                        pdf.

Castelpietra:2002:DIC

[190] C. Castelpietra, A. Guidotti, L. Iocchi, D. Nardi, and R. Rosati. Design and implementation of cognitive soccer robots. Lecture Notes in Computer Science, 2377:312–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770312.
                                        pdf.

Ciesielski:2002:GPR

[191] Vic Ciesielski, Dylan Mawhinney, and Peter Wilson. Genetic programming for robot soccer. Lecture Notes in Computer Science, 2377:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770319.
pdf.

Drucker:2002:TGU

[192] Christian Drücker, Sebastian Hübner, Ubbo Visser, and Hans-Georg Weland. ”as time goes by” — using time series based decision tree induction to analyze the behaviour of opponent players. Lecture Notes in Computer Science, 2377: 325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770325.
pdf.

Demura:2002:KCD

[193] Kosei Demura, Nobuhiro Tachi, Tetsuya Maekawa, and Tamaki Ueno. KENSEI-chan: Design of a humanoid for running. Lecture Notes in Computer Science, 2377:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770331.
pdf.

Figueras:2002:SRC

[194] Albert Figueras, Joan Colomer, Thor I. Fossen, and J. Lluis de la Rosa. Supervision of robot control. Lecture Notes in Computer Science, 2377:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770337.
                                        pdf.

Frank:2002:WTM

[195] Ian Frank, Kumiko Tanaka-Ishii, Hitoshi Matsubara, and Eiichi Osawa. Walkie-talkie MIKE. Lecture Notes in Computer Science, 2377:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770343.
                                        pdf.

Fukase:2002:QRN

[196] Takeshi Fukase, Masahiro Yokoi, Yuichi Kobayashi, Ryuichi Ueda, Hideo Yuasa, and Tamio Arai. Quadruped robot navigation considering the observational cost. Lecture Notes in Computer Science, 2377:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770350.
                                        pdf.

Gu:2002:EFL

[197] Dongbing Gu and Huosheng Hu. Evolving fuzzy logic controllers for Sony legged robots. Lecture Notes in Computer Science, 2377:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770356.
pdf.

Haker:2002:MIN

[198] Martin Haker, André Meyer, Daniel Polani, and Thomas Martinetz. A method for incorporation of new evidence to improve world state estimation. Lecture Notes in Computer Science, 2377:362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770362.
pdf.

Hengst:2002:OLQ

[199] Bernhard Hengst, Darren Ibbotson, Son Bao Pham, and Claude Sammut. Omnidirectional locomotion for quadruped robots. Lecture Notes in Computer Science, 2377:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770368.
pdf.

vHundelshausen:2002:OVS

[200] Felix v. Hundelshausen, Sven Behnke, and Raúl Rojas. An omnidirectional vision system that finds and tracks color edges and blobs. Lecture Notes in Computer Science, 2377:374– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770374.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770374.
                                        pdf.

Hunter:2002:GAP

[201] Matthew Hunter and Huosheng Hu. A generalised approach to position selection for simulated soccer agents. Lecture Notes in Computer Science, 2377: 380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770380.
                                        pdf.

Indiveri:2002:MCN

[202] Giovanni Indiveri. On the motion control of a nonholonomic soccer playing robot. Lecture Notes in Computer Science, 2377:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770387.
                                        pdf.

Isekenmeier:2002:EPC

[203] Guido Isekenmeier, Bernhard Nebel, and Thilo Weigel. Evaluation of the performance of CS freiburg 1999 and CS freiburg 2000. Lecture Notes in Computer Science, 2377:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770393.
pdf.

Johansson:2002:UEF

[204] Stefan J. Johansson and Alessandro Saffiotti. Using the electric field approach in the RoboCup domain. Lecture Notes in Computer Science, 2377: 399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770399.
pdf.

deJong:2002:TTA

[205] Frank de Jong, Jurjen Caarls, Robert Bartelds, and Pieter P. Jonker. A two-tiered approach to self-localization. Lecture Notes in Computer Science, 2377:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770405.
pdf.

Kraetzschmar:2002:MMC

[206] Gerhard Kraetzschmar, Hans Utz, Stefan Sablatnög, Stefan Enderle, and Günther Palm. Miro — middleware for cooperative robotics. Lecture Notes in Computer Science, 2377:411– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770411.
                                        pdf.

Kuwata:2002:BUM

[207] Yoshitaka Kuwata and Atsushi Shinjoh. Building user models for robocuprescue visualization. Lecture Notes in Computer Science, 2377:417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770417.
                                        pdf.

Lenser:2002:MHB

[208] Scott Lenser, James Bruce, and Manuela Veloso. A modular hierarchical behavior-based architecture. Lecture Notes in Computer Science, 2377: 423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770423.
                                        pdf.

Matsuoka:2002:LOD

[209] Takeshi Matsuoka, Manabu Araoka, Tsutomu Hasegawa, Akira Mohri, Motoji Yamamoto, Toshihiro Kiriki, Nobuhiro Ushimi, Takuya Sugimoto, Jyun’ichi Inoue, and Yuuki Yamaguchi. Localization and obstacles detection using omni-directional vertical stereo vision. Lecture Notes in Computer Science, 2377:429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770429.
pdf.

Merke:2002:KBR

[210] A. Merke and M. Riedmiller. Karlsruhe brainstormers — A reinforcement learning approach to robotic soccer. Lecture Notes in Computer Science, 2377:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770435.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770435.
pdf.

Miene:2002:IST

[211] Andrea Miene and Ubbo Visser. Interpretation of spatio-temporal relations in real-time and dynamic environments. Lecture Notes in Computer Science, 2377:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2377/23770441.

                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2377/23770441.pdf.

Pham:2002:SGD

[212] Son Bao Pham, Bernhard Hengst, Darren Ibbotson, and Claude Sammut. Stochastic gradient descent localisation in quadruped robots. Lecture Notes in Computer Science, 2377:447– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770447.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770447.
                                        pdf.

Riley:2002:RPO

[213] Patrick Riley and Manuela Veloso. Recognizing probabilistic opponent movement models. Lecture Notes in Computer Science, 2377:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770453.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770453.
                                        pdf.

Tu:2002:DIS

[214] Kuo-Yang Tu. Design and implementation of a soccer robot with modularized control circuits. Lecture Notes in Computer Science, 2377:459– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770459.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2377/23770459.
pdf.

Wang:2002:CBB

[215] Hui Wang, Han Wang, Chunmiao Wang, and William Y. C. Soh. Cooperation-based behavior design. Lecture Notes in Computer Science, 2377:465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770465.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770465.
pdf.

Wang:2002:MPS

[216] Hui Wang, Han Wang, Chunmiao Wang, and William Y. C. Soh. Multi-platform soccer robot development system. Lecture Notes in Computer Science, 2377:471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770471.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770471.
pdf.

Ushimi:2002:LNM

[217] Nobuhiro Ushimi, Motoji Yamamoto, Jyun’ichi Inoue, Takuya Sugimoto, Manabu Araoka, Takeshi Matsuoka, Toshihiro Kiriki, Yuuki Yamaguchi, Tsutomu Hasegawa, and Akira Mohri. On-line navigation of mobile robot among moving obstacles using ultrasonic sensors. Lecture Notes in Computer Science, 2377:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770477.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770477.
                                        pdf.

Suzuki:2002:TD

[218] Keisuke Suzuki, Naotaka Tanaka, and Mio Yamamoto. 11monkeys3 team description. Lecture Notes in Computer Science, 2377:484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770484.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770484.
                                        pdf.

Nazemi:2002:ASS

[219] Eslam Nazemi, Mahmood Rahmani, and Bahman Radjabalipour. 3T architecture for the SBCe simulator team. Lecture Notes in Computer Science, 2377:487–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2377/23770487.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770487.
                                        pdf.

Yao:2002:AT

[220] Jinyi Yao, Jiang Chen, Yunpeng Cai, and Shi Li. Architecture of TsinghuAeolus. Lecture Notes in Computer Science, 2377:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770491.
pdf.

Stone:2002:AUH

[221] Peter Stone. ATTUnited-2001: Using heterogeneous players. Lecture Notes in Computer Science, 2377:495– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770495.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770495.
pdf.

Browne:2002:BCS

[222] Keen Browne, Jon McCune, Adam Trost, David Evans, and David Brogan. Behavior combination and swarm programming. Lecture Notes in Computer Science, 2377:499–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770499.
pdf.

Carpenter:2002:CTD

[223] Paul Carpenter, Patrick Riley, Gal Kaminka, Manuela Veloso, Ignacio Thayer, and Robert Wang. Chameleons-01 team description. Lecture Notes in Computer Science, 2377: 503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770503.
                                        pdf.

Prokopenko:2002:CDB

[224] Mikhail Prokopenko, Peter Wang, and Thomas Howard. Cyberoos’2001: “Deep Behaviour Projection” agent architecture. Lecture Notes in Computer Science, 2377:507–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2377/23770507.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2377/23770507.pdf.

Hu:2002:EWT

[225] Huosheng Hu, Kostas Kostiadis, Matthew Hunter, and Nikolaos Kalyviotis. Essex wizards 2001 team description. Lecture Notes in Computer Science, 2377:511–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2377/23770511.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2377/23770511.pdf.

Lau:2002:FPT

[226] Nuno Lau and Luis Paulo Reis. FC portugal 2001 team description: Flexible teamwork and configurable strategy. Lecture Notes in Computer Science, 2377:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770515.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770515.
pdf.

B:2002:HRT

[227] Bahador Nooraei B., Siavash Rahbar N., and Omid Aladini. Helli-respina 2001 team description paper. Lecture Notes in Computer Science, 2377:519– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770519.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770519.
pdf.

Yuen:2002:LTD

[228] Anthony Yuen. Lazarus team description. Lecture Notes in Computer Science, 2377:522–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2377/23770522.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2377/23770522.pdf.

Murray:2002:RK

[229] Jan Murray, Oliver Obst, and Frieder Stolzenburg. RoboLog koblenz 2001. Lecture Notes in Computer Science, 2377:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770526.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770526.
                                        pdf.

Arnold:2002:TDM

[230] A. Arnold, F. Flentge, Ch. Schneider, G. Schwandtner, Th. Uthmann, and M. Wache. Team description mainz rolling brains 2001. Lecture Notes in Computer Science, 2377:531– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770531.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770531.
                                        pdf.

Esaki:2002:TDN

[231] Tetsuya Esaki, Taku Sakushima, Yoshiki Asai, and Nobuhiro Ito. Team description of NITStones2001. Lecture Notes in Computer Science, 2377:535– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770535.
                                        pdf.

Nakayama:2002:TYD

[232] Koji Nakayama and Ikuo Takeuchi. Team YowAI-2001 description. Lecture Notes in Computer Science, 2377:539– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770539.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2377/23770539.
pdf.

Buttinger:2002:DDT

[233] Sean Buttinger, Marco Diedrich, Leo Hennig, Angelika Hoenemann, Philipp Huegelmeyer, Andreas Nie, Andres Pegam, Collin Rogowski, Claus Rollinger, Timo Steffens, and Wilfried Teiken. The dirty dozen team and coach description. Lecture Notes in Computer Science, 2377:543–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770543.
pdf.

Wyeth:2002:UCI

[234] Gordon Wyeth, Mark Venz, Helen Mayfield, Jun Akiyama, and Rex Heathwood. UQ CrocaRoos: An initial entry to the simulation league. Lecture Notes in Computer Science, 2377: 547–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770547.
pdf.

deBoer:2002:UTT

[235] Remco de Boer, Jelle Kok, and Frans Groen. UvA trilearn 2001 team description. Lecture Notes in Computer Science, 2377:551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770551.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770551.
                                        pdf.

Morishita:2002:ZTD

[236] Takuya Morishita, Hiroki Shimora, Kouichirou Hiratsuka, Takenori Kubo, Kyoichi Hiroshima, Raiko Funakami, Junji Nishino, Tomohiro Odaka, and Hisakazu Ogura. Zeng01 team description: Formation decision method using game theory. Lecture Notes in Computer Science, 2377:555–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770555.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770555.
                                        pdf.

Baltes:2002:S

[237] Jacky Baltes. 4 stooges. Lecture Notes in Computer Science, 2377:559– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770559.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770559.
                                        pdf.

Costa:2002:TD

[238] Paulo Costa, Armando Sousa, Paulo Marques, Pedro Costa, Susana Gaio, and António Moreira. 5dpo team description. Lecture Notes in Computer Science, 2377:563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2377/23770563.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2377/23770563.pdf.

Browning:2002:CDV

[239] Brett Browning, Michael Bowling, James Bruce, Ravi Balasubramanian, and Manuela Veloso. CM-Dragons’01 — vision-based motion tracking and heteregenous robots. Lecture Notes in Computer Science, 2377:567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770567.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770567.
pdf.

Rojas:2002:FFG

[240] Raúl Rojas, Sven Behnke, Achim Liers, and Lars Knipping. FU-fighters 2001 (global vision). Lecture Notes in Computer Science, 2377:571–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770571.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770571.
pdf.

Rojas:2002:FFO

[241] Raúl Rojas, Felix von Hundelshausen, Sven Behnke, and Bernhard Frötschl. FU-fighters omni 2001 (local vision). Lecture Notes in Computer Science, 2377:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770575.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770575.
pdf.

Hibino:2002:OTD

[242] Shinya Hibino, Yukiharu Kodama, Yasunori Nagasaka, Tomoichi Takahashi, Kazuhito Murakami, and Tadashi Naruse. Owaribito — A team description. Lecture Notes in Computer Science, 2377:579–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770579.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770579.
pdf.

Bras:2002:RFT

[243] Francis Bras, Ryad Benosman, André Anglade, Seko Latidine, Olivier Martino, Aude Lagardere, Alain Boun, Alain Testa, and Gilles Cordurié. RoboCup 2001 (F180) team description: RoboSix UPMC-CFA (france). Lecture Notes in Computer Science, 2377:583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2377/23770583.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770583.
pdf.

delaRosa:2002:RTD

[244] J. Ll. de la Rosa, B. Innocenti, M. Montaner, A. Figueras, I. Muñoz, and J. A. Ramon. Rogi team description. Lecture Notes in Computer Science, 2377: 587–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770587.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770587.
pdf.

Thomas:2002:R

[245] Jason Thomas, Kenichi Yoshimura, and Andrew Peel. Roobots. Lecture Notes in Computer Science, 2377:591– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770591.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770591.
pdf.

Manzuri:2002:SCS

[246] Mohammad Taghi Manzuri, Hamid Reza Chitsaz, Reza Ghorbani, Pooya Karimian, Alireza Mirazi, Mehran Motamed, Roozbeh Mottaghi, and Payam Sabzmeydani. Sharif CESR small size Robocup team. Lecture Notes in Computer Science, 2377:595–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770595.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770595.
pdf.

Sekimori:2002:TDT

[247] Daisuke Sekimori, Nobuhito Mori, Junichi Ieda, Wataru Matsui, Osamu Miyake, Tomoya Usui, Yukihisa Tanaka, Dong Pyo Kim, Tetsuhiro Maeda, Hirokazu Sugimoto, Ryouhei Fujimoto, Masaya Enomoto, Yasuhiro Masutani, and Fumio Miyazaki. The team description of the team OMNI. Lecture Notes in Computer Science, 2377:599–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2377/23770599.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770599.
                                        pdf.

Wyeth:2002:URA

[248] Gordon Wyeth, David Ball, David Cusack, and Adrian Ratnapala. UQ RoboRoos: Achieving power and agility in a small size robot. Lecture Notes in Computer Science, 2377:603– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770603.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770603.
                                        pdf.

Chang:2002:V

[249] Mark M. Chang and Gordon F. Wyeth. ViperRoos 2001. Lecture Notes in Computer Science, 2377:607– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770607.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2377/23770607.
pdf.

Schmitt:2002:ARU

[250] Thorsten Schmitt, Sebastian Buck, and Michael Beetz. AGILO RoboCuppers 2001: Utility- and plan-based action selection based on probabilistically estimated game situations. Lecture Notes in Computer Science, 2377: 611–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770611.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770611.
pdf.

Pagello:2002:AVH

[251] E. Pagello, M. Bert, M. Barbon, E. Menegatti, C. Moroni, C. Pellizzari, D. Spagnoli, and S. Zaffalon. Artisti veneti: An heterogeneous robot team for the 2001 middle-size league. Lecture Notes in Computer Science, 2377: 616–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770616.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770616.
pdf.

Jamzad:2002:BRT

[252] M. Jamzad, H. Chitsaz, A. Foroughnassirai, R. Ghorbani, M. Kazemi, V. S. Mirrokni, and B. S. Sadjad. Basic requirements for a teamwork in middle size RoboCup. Lecture Notes in Computer Science, 2377:621– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770621.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770621.
                                        pdf.

Spaan:2002:COD

[253] Matthijs Spaan, Marco Wiering, Robert Bartelds, Raymond Donkervoort, Pieter Jonker, and Frans Groen. Clockwork Orange: The Dutch RoboSoccer team. Lecture Notes in Computer Science, 2377:627– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770627.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770627.
                                        pdf.

Stancliff:2002:CHT

[254] Steve Stancliff, Ravi Balasubramanian, Tucker Balch, Rosemary Emery, Kevin Sikorski, and Ashley Stroupe. CMU hammerheads 2001 team description. Lecture Notes in Computer Science, 2377:631–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770631.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770631.
                                        pdf.

Lafrenz:2002:CTD

[255] R. Lafrenz, M. Becht, T. Buchheim, P. Burger, G. Hetzel, G. Kindermann, M. Schanz, M. Schulé, and P. Levi. CoPS-team description. Lecture Notes in Computer Science, 2377: 635–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770635.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770635.
pdf.

Bonarini:2002:FMR

[256] Andrea Bonarini, Giovanni Invernizzi, Fabio Marchese, Matteo Matteucci, Marcello Restelli, and Domenico Sorrenti. Fun2Mas: The Milan Robocup team. Lecture Notes in Computer Science, 2377:639–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770639.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770639.
pdf.

Matsuoka:2002:F

[257] Takeshi Matsuoka, Motoji Yamamoto, Nobuhiro Ushimi, Jyun’ichi Inoue, Takuya Sugimoto, Manabu Araoka, Toshihiro Kiriki, Yuuki Yamaguchi, Tsutomu Hasegawa, and Akira Mohri. Fusion. Lecture Notes in Computer Science, 2377:643–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770643.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2377/23770643.
                                        pdf.

Bredenfeld:2002:GR

[258] Ansgar Bredenfeld, Vlatko Becanovic, Thomas Christaller, Horst Günther, Giovanni Indiveri, Hans-Ulrich Kobialka, Paul-Gerhard Plöger, and Peter Schöll. GMD-robots. Lecture Notes in Computer Science, 2377:648– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770648.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770648.
                                        pdf.

Lima:2002:ITD

[259] Pedro Lima, Luis Custódio, Bruno Damas, Manuel Lopes, Carlos Marques, Luis Toscano, and Rodrigo Ventura. ISocRob 2001 team description. Lecture Notes in Computer Science, 2377:653–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2377/23770653.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770653.
                                        pdf.

Ribeiro:2002:MRF

[260] Fernando Ribeiro, Carlos Machado, Sérgio Sampaio, and Bruno Martins. MINHO robot football team for 2001. Lecture Notes in Computer Science, 2377:657–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2377/23770657.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2377/23770657.
pdf.

Takahashi:2002:OUT

[261] Yasutake Takahashi, Shoichi Ikenoue, Shujiro Inui, Kouichi Hikita, Yutaka Katoh, and Minoru Asada. Osaka University “Trackies 2001”. Lecture Notes in Computer Science, 2377:661– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770661.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770661.
pdf.

Benosman:2002:RUC

[262] Ryad Benosman, Francis Bras, Frederic Bach, Simon Boulay, Emmanuelle Cahn, Sylvain Come, Gilles Cordurié, Jarlegan Marie Annick, Lapied Loic, Cyrille Potereau, Franck Richard, Samedi Sath, Xavier Vasseur, and Pascal Vincent. ROBOSIX UPMC-CFA: RoboCup team description. Lecture Notes in Computer Science, 2377:665– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770665.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770665.
pdf.

Iocchi:2002:PQR

[263] Luca Iocchi, Daniele Baldassari, Flavio Cappelli, Alessandro Farinelli, Giorgio Grisetti, Floris Maathuis, and Daniele Nardi. S.P.Q.R. wheeled team. Lecture Notes in Computer Science, 2377: 669–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770669.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770669.
pdf.

Yoshida:2002:TDE

[264] Kazuo Yoshida, Ryoichi Tsuzaki, Junichi Kougo, Takaaki Okabe, Nobuyuki Kurihara, Daiki Sakai, Ryotaku Hayashi, and Hikari Fujii. Team description eigen. Lecture Notes in Computer Science, 2377:673–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770673.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770673.
pdf.

Utz:2002:US

[265] Hans Utz, Gerd Mayer, Dominik Maschke, Alexander Neubeck, Peter Schaeffer, Philipp Baer, Ingmar Baetge, Jan Fischer, Roland Holzer, Markus Lauer, Alexander Reisser, Florian Sterk, Günther Palm, and Gerhard Kraetzschmar. The ulm sparrows 2001. Lecture Notes in Computer Science, 2377:677–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770677.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2377/23770677.
pdf.

Oda:2002:AKU

[266] Kentaro Oda, Takeshi Ohashi, Shuichi Kouno, Kunio Gohara, Toyohiro Hayashi, Takeshi Kato, Yuki Katsumi, and Toshiyuki Ishimura. ASURA: Kyushu united team in the four legged robot league. Lecture Notes in Computer Science, 2377:681–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770681.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770681.
pdf.

Mitsunaga:2002:BOL

[267] Noriaki Mitsunaga, Yukie Nagai, Tomohiro Ishida, Taku Izumi, and Minoru Asada. BabyTigers 2001: Osaka legged robot team. Lecture Notes in Computer Science, 2377:685–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770685.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770685.
pdf.

Akin:2002:CTD

[268] H. Levent Akin, Andon Topalov, and Okyay Kaynak. Cerberus 2001 team description. Lecture Notes in Computer Science, 2377:689–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770689.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770689.
                                        pdf.

Uther:2002:CPF

[269] William Uther, Scott Lenser, James Bruce, Martin Hock, and Manuela Veloso. CM-Pack’01: Fast legged robot walking, robust localization, and team behaviors. Lecture Notes in Computer Science, 2377:693–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770693.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770693.
                                        pdf.

Hu:2002:ERT

[270] Huosheng Hu, Dongbing Gu, Dragos Golubovic, Bo Li, and Zhengyu Liu. Essex rovers 2001 team description. Lecture Notes in Computer Science, 2377:697–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2377/23770697.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770697.
                                        pdf.

Hugel:2002:FLT

[271] Vincent Hugel, Olivier Stasse, Patrick Bonnin, and Pierre Blazevic. French LRP team’s description. Lecture Notes in Computer Science, 2377:701– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770701.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770701.
pdf.

Brunn:2002:G

[272] Ronnie Brunn, Uwe Düffert, Matthias Jüngel, Tim Laue, Martin Lötzsch, Sebastian Petters, Max Risler, Thomas Röfer, Kai Spiess, and Andreas Sztybryc. GermanTeam 2001. Lecture Notes in Computer Science, 2377:705– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770705.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770705.
pdf.

Sud:2002:MR

[273] Daniel Sud, Francois Cayouette, Gu Jin Hua, and Jeremy Cooperstock. McGill reddogs. Lecture Notes in Computer Science, 2377:709– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770709.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770709.
pdf.

Clarke:2002:R

[274] Kate Clarke, Stephen Dempster, Ian Falcao, Bronwen Jones, Daniel Rudolph, Alan Blair, Chris Mc-Carthy, Dariusz Walter, and Nick Barnes. RoboMutts ++. Lecture Notes in Computer Science, 2377:713– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770713.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770713.
                                        pdf.

Nardi:2002:PQR

[275] D. Nardi, V. Bonifaci, C. Castelpietra, U. Di Iorio, A. Guidotti, L. Iocchi, M. Salerno, and F. Zonfrilli. S.P.Q.R. legged team. Lecture Notes in Computer Science, 2377:717–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770717.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770717.
                                        pdf.

Azari:2002:TDU

[276] D. Azari, J. K. Burns, K. Deshmukh, D. Fox, D. Grimes, C. T. Kwok, R. Pitkanen, A. P. Shon, and P. Tressel. Team description: UW huskies-01. Lecture Notes in Computer Science, 2377: 721–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770721.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770721.
                                        pdf.

Saffiotti:2002:TS

[277] A. Saffiotti, A. Björklund, S. Johansson, and Z. Wasik. Team sweden. Lecture Notes in Computer Science, 2377: 725–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770725.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770725.
pdf.

Arai:2002:TDA

[278] Tamio Arai, Takeshi Fukase, Ryuichi Ueda, Yuichi Kobayashi, and Takanobu Kawabe. The team description of ARAIBO. Lecture Notes in Computer Science, 2377:730–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2377/23770730.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2377/23770730.pdf.

Chitta:2002:UPR

[279] Sachin Chitta, William Sacks, Jim Ostrowski, Aveek Das, and P. K. Mishra. The University of Pennsylvania RoboCup legged soccer team. Lecture Notes in Computer Science, 2377: 734–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770734.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770734.
pdf.

Li:2002:WES

[280] Xiang Li, Zefeng Zhang, Lei Jiang, and Xiaoping Chen. Wright Eagle 2001 — Sony legged robot team. Lecture Notes in Computer Science, 2377:739– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770739.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770739.
                                        pdf.

Sakushima:2002:DAD

[281] Taku Sakushima, Tetsuya Esaki, Yoshiki Asai, Nobuhiro Ito, and Koichi Wada. A design of agents for the disaster simulator on robocup-rescue. Lecture Notes in Computer Science, 2377: 743–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2377/23770743.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2377/23770743.
                                        pdf.

Habibi:2002:AMA

[282] Jafar Habibi, Ali Nouri, and Mazda Ahmadi. An approach to multi-agent communication used in RobocupRescue. Lecture Notes in Computer Science, 2377:747–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2377/23770747.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2377/23770747.pdf.

Nair:2002:TAR

[283] Ranjit Nair, Takayuki Ito, Milind Tambe, and Stacy Marsella. Task allocation in the RoboCup rescue simulation domain: A short note. Lecture Notes in Computer Science, 2377:751– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770751.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770751.
pdf.

Padgham:2002:TDR

[284] Lin Padgham, John Thangarajah, David Poutakidis, and Chandaka Fernando. Team description for RMITon-Fire: Robocup Rescue Simulation Team 2001. Lecture Notes in Computer Science, 2377:755–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2377/23770755.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2377/23770755.
pdf.

Anonymous:2002:AIg

[285] Anonymous. Author index. Lecture Notes in Computer Science, 2377:759– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2377/2377indx.pdf.

Shankar:2002:CST

[286] Natarajan Shankar and Harald Rueß. Combining shostak theories. Lecture Notes in Computer Science, 2378:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780001.
                                        pdf.

Mitchell:2002:MRS

[287] John C. Mitchell. Multiset rewriting and security protocol analysis. Lecture Notes in Computer Science, 2378:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780019.
                                        pdf.

Baader:2002:ELC

[288] Franz Baader. Engineering of logics for the content-based representation of information. Lecture Notes in Computer Science, 2378:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780023.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780023.
                                        pdf.

Mellies:2002:ART

[289] Paul-André Melliès. Axiomatic rewriting theory VI residual theory revisited. Lecture Notes in Computer Science, 2378:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2378/23780024.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2378/23780024.
pdf.

Kennaway:2002:SAM

[290] Richard Kennaway, Zurab Khasidashvili, and Adolfo Piperno. Static analysis of modularity of β-reduction in the hyperbalanced λ-calculus. Lecture Notes in Computer Science, 2378: 51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780051.
pdf.

Faure:2002:ERC

[291] Germain Faure and Claude Kirchner. Exceptions in the rewriting calculus. Lecture Notes in Computer Science, 2378:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2378/23780066.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2378/23780066.
pdf.

Struth:2002:DFL

[292] Georg Struth. Deriving focused lattice calculi. Lecture Notes in Computer Science, 2378:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2378/23780083.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2378/23780083.pdf.

Seki:2002:LTT

[293] Hiroyuki Seki, Toshinori Takai, Youhei Fujinaka, and Yuichi Kaji. Layered transducing term rewriting system and its recognizability preserving property. Lecture Notes in Computer Science, 2378:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780098.
pdf.

Ohsaki:2002:DCP

[294] Hitoshi Ohsaki and Toshinori Takai. Decidability and closure properties of equational tree languages. Lecture Notes in Computer Science, 2378:114– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780114.
pdf.

Rety:2002:RSD

[295] Pierre Réty and Julie Vuotto. Regular sets of descendants by some rewrite strategies. Lecture Notes in Computer Science, 2378:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780129.
pdf.

Waldmann:2002:RG

[296] Johannes Waldmann. Rewrite games. Lecture Notes in Computer Science, 2378:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2378/23780144.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2378/23780144.
pdf.

Severi:2002:EBM

[297] Paula Severi and Fer-Jan de Vries. An extensional Böhm model. Lecture Notes in Computer Science, 2378:159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780159.
pdf.

Forest:2002:WCE

[298] Julien Forest. A weak calculus with explicit operators for pattern matching and substitution. Lecture Notes in Computer Science, 2378:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780174.
pdf.

Liang:2002:TIR

[299] Chuck Liang and Gopalan Nadathur. Tradeoffs in the intensional representation of lambda terms. Lecture Notes in Computer Science, 2378:192– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780192.
pdf.

Deharbe:2002:ISM

[300] David Déharbe, Anamaria Martins Moreira, and Christophe Ringeissen. Improving symbolic model checking by rewriting temporal logic formulae. Lecture Notes in Computer Science, 2378: 207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780207.
pdf.

Voigtlander:2002:CEI

[301] Janis Voigtländer. Conditions for efficiency improvement by tree transducer composition. Lecture Notes in Computer Science, 2378:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780222.
pdf.

Bravenboer:2002:RSI

[302] Martin Bravenboer and Eelco Visser. Rewriting strategies for instruction selection. Lecture Notes in Computer Science, 2378:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2378/23780237.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2378/23780237.pdf.

Bournez:2002:PRS

[303] Olivier Bournez and Claude Kirchner. Probabilistic rewrite strategies. applications to ELAN . Lecture Notes in Computer Science, 2378:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780252.
pdf.

Geser:2002:LSL

[304] Alfons Geser. Loops of superexponential lengths in one-rule string rewriting. Lecture Notes in Computer Science, 2378:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2378/23780267.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2378/23780267.
pdf.

Tahhan-Bittar:2002:RDL

[305] Elias Tahhan-Bittar. Recursive derivational length bounds for confluent term rewrite systems. Lecture Notes in Computer Science, 2378:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780281.
                                        pdf.

Lucas:2002:TCC

[306] Salvador Lucas. Termination of (canonical) context-sensitive rewriting. Lecture Notes in Computer Science, 2378: 296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780296.
                                        pdf.

Charatonik:2002:ASC

[307] Witold Charatonik and Jean-Marc Talbot. Atomic set constraints with projection. Lecture Notes in Computer Science, 2378:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780311.
                                        pdf.

Levy:2002:CSO

[308] Jordi Levy and Mateu Villaret. Currying second-order unification problems. Lecture Notes in Computer Science, 2378:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2378/23780326.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2378/23780326.
pdf.

Dougherty:2002:DVH

[309] Dan Dougherty and Tomasz Wierzbicki. A decidable variant of higher order matching. Lecture Notes in Computer Science, 2378:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2378/23780340.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2378/23780340.pdf.

Baader:2002:CDP

[310] Franz Baader and Cesare Tinelli. Combining decision procedures for positive theories sharing constructors. Lecture Notes in Computer Science, 2378:352– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780352.
pdf.

vandePol:2002:JRS

[311] Jaco van de Pol. JITty: A rewriter with strategy annotations. Lecture Notes in Computer Science, 2378:367– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2378/23780367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2378/23780367.
pdf.

Durand:2002:ATC

[312] Irène Durand. Autowrite: A tool for checking properties of term rewriting systems. Lecture Notes in Computer Science, 2378:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780371.
                                        pdf.

Lecland:2002:TIT

[313] Benoit Lecland and Pierre Réty. TTSLI: An implementation of treetuple synchronized languages. Lecture Notes in Computer Science, 2378:376– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780376.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780376.
                                        pdf.

Lippi:2002:GII

[314] Sylvain Lippi. in^{2} : A graphical interpreter for interaction nets. Lecture Notes in Computer Science, 2378:380– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2378/23780380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2378/23780380.
                                        pdf.

Anonymous:2002:AIh

[315] Anonymous. Author index. Lecture Notes in Computer Science, 2378:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2378/2378auth.pdf.

Geyer:2002:IVA

[316] Lars Geyer and Martin Becker. On the influence of variabilities on the application-engineering process of a product family. Lecture Notes in Computer Science, 2379:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790001.
pdf.

Jaring:2002:RVS

[317] Michel Jaring and Jan Bosch. Representing variability in software product lines: A case study. Lecture Notes in Computer Science, 2379:15– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790015.
pdf.

Krueger:2002:VMS

[318] Charles W. Krueger. Variation management for software production lines. Lecture Notes in Computer Science, 2379:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2379/23790037.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2379/23790037.
pdf.

Bockle:2002:AIP

[319] Günter Böckle, Jesús Bermejo Muñoz, Peter Knauber, Charles W. Krueger, Julio Cesar Sampaio do Prado Leite, Frank van der Linden, Linda Northrop, Michael Stark, and David M. Weiss. Adopting and institutionalizing a product line culture. Lecture Notes in Computer Science, 2379:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790049.
pdf.

Voget:2002:ESP

[320] Stefan Voget and Martin Becker. Establishing a software product line in an immature domain. Lecture Notes in Computer Science, 2379:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790060.
pdf.

Wijnstra:2002:CFS

[321] Jan Gerben Wijnstra. Critical factors for a successful platform-based product family approach. Lecture Notes in Computer Science, 2379:68– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790068.
pdf.

vanZyl:2002:PLA

[322] Jay van Zyl. Product line architecture and the separation of concerns. Lecture Notes in Computer Science, 2379: 90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790090.
pdf.

Muthig:2002:MDP

[323] Dirk Muthig and Colin Atkinson. Model-driven product line architectures. Lecture Notes in Computer Science, 2379:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2379/23790110.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2379/23790110.pdf.

Thiel:2002:SIV

[324] Steffen Thiel and Andreas Hein. Systematic integration of variability into product line architecture design. Lecture Notes in Computer Science, 2379:

130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790130.
                                        pdf.

Brown:2002:ACS

[325] T. John Brown, Ivor Spence, Peter Kilpatrick, and Danny Crookes. Adaptable components for software product line engineering. Lecture Notes in Computer Science, 2379:154– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790154.
                                        pdf.

Mannion:2002:UFO

[326] Mike Mannion. Using first-order logic for product line model validation. Lecture Notes in Computer Science, 2379: 176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790176.
                                        pdf.

Pree:2002:PLA

[327] Wolfgang Pree, Marcus Fontoura, and Bernhard Rumpe. Product line annotations with UML-F. Lecture Notes in Computer Science, 2379:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790188.
pdf.

Fey:2002:FMM

[328] Dániel Fey, Róbert Fajta, and András Boros. Feature modeling: A meta-model to enhance usability and usefulness. Lecture Notes in Computer Science, 2379:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790198.
pdf.

vanDeursen:2002:FBP

[329] Arie van Deursen, Merijn de Jonge, and Tobias Kuipers. Featurebased product line instantiation using source-level packages. Lecture Notes in Computer Science, 2379:217– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790217.
pdf.

Ferber:2002:FID

[330] Stefan Ferber, Jürgen Haag, and Juha Savolainen. Feature interaction and dependencies: Modeling features for reengineering a legacy product line. Lecture Notes in Computer Science, 2379:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790235.
                                        pdf.

Bosch:2002:MES

[331] Jan Bosch. Maturity and evolution in software product lines: Approaches, artefacts and organization. Lecture Notes in Computer Science, 2379:257– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790257.
                                        pdf.

Simon:2002:EIS

[332] Daniel Simon and Thomas Eisenbarth. Evolutionary introduction of software product lines. Lecture Notes in Computer Science, 2379:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790272.
                                        pdf.

Jolley:2002:GPI

[333] Truman M. Jolley, David J. Kasik, and Conrad E. Kimball. Governance polarities of internal product lines. Lecture Notes in Computer Science, 2379: 284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790284.
pdf.

Yacoub:2002:PAC

[334] Sherif Yacoub. Performance analysis of component-based applications. Lecture Notes in Computer Science, 2379: 299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790299.
pdf.

Smith:2002:UOA

[335] Dennis Smith, Liam O’ Brien, and John Bergey. Using the options analysis for reengineering (OAR) method for mining components for a product line. Lecture Notes in Computer Science, 2379: 316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790316.
pdf.

vanOmmering:2002:WSS

[336] Rob van Ommering and Jan Bosch. Widening the scope of software product lines — from variation to composition. Lecture Notes in Computer Science, 2379:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790328.
                                        pdf.

Kishi:2002:MPL

[337] Tomoji Kishi, Natsuko Noda, and Takuya Katayama. A method for product line scoping based on a decisionmaking framework. Lecture Notes in Computer Science, 2379:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790348.
                                        pdf.

Kang:2002:UMP

[338] Kyo C. Kang, Patrick Donohoe, Eunman Koh, Jaejoon Lee, and Kwanwoo Lee. Using a marketing and product plan as a key driver for product line asset development. Lecture Notes in Computer Science, 2379:366– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2379/23790366.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2379/23790366.
                                        pdf.

vanderLinden:2002:ESA

[339] Frank van der Linden. Engineering software architectures, processes and platforms for system families — ESAPS overview. Lecture Notes in Computer Science, 2379:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2379/23790383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2379/23790383.
pdf.

Anonymous:2002:AIi

[340] Anonymous. Author index. Lecture Notes in Computer Science, 2379:399– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2379/2379auth.pdf.

Reif:2002:MAC

[341] John H. Reif. Molecular assembly and computation: From theory to experimental demonstrations. Lecture Notes in Computer Science, 2380:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800001.
pdf.

Marathe:2002:TPC

[342] Madhav V. Marathe. Towards a predictive computational complexity theory. Lecture Notes in Computer Science, 2380:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800022.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800022.
                                        pdf.

Pitts:2002:ESS

[343] Andrew M. Pitts. Equivariant syntax and semantics. Lecture Notes in Computer Science, 2380:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800032.
                                        pdf.

Lungo:2002:DTR

[344] Alberto Del Lungo, Andrea Frosini, Maurice Nivat, and Laurent Vuillon. Discrete tomography: Reconstruction under periodicity constraints. Lecture Notes in Computer Science, 2380:38– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800038.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800038.
                                        pdf.

Mannila:2002:LGM

[345] Heikki Mannila. Local and global methods in data mining: Basic techniques and open problems. Lecture Notes in Computer Science, 2380:57– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2380/23800057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800057.
pdf.

Hermenegildo:2002:PDV

[346] M. Hermenegildo, G. Puebla, F. Bueno, and P. López-García. Program debugging and validation using semantic approximations and partial specifications. Lecture Notes in Computer Science, 2380:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800069.
pdf.

Engebretsen:2002:IRE

[347] Lars Engebretsen, Jonas Holmerin, and Alexander Russell. Inapproximability results for equations over finite groups. Lecture Notes in Computer Science, 2380:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800073.
pdf.

Pettie:2002:FAP

[348] Seth Pettie. A faster all-pairs shortest path algorithm for realweighted sparse graphs. Lecture Notes in Computer Science, 2380:85– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800085.
                                        pdf.

Colcombet:2002:FGH

[349] Thomas Colcombet. On families of graphs having a decidable first order theory with reachability. Lecture Notes in Computer Science, 2380:98– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800098.
                                        pdf.

Fabrikant:2002:HOT

[350] Alex Fabrikant, Elias Koutsoupias, and Christos H. Papadimitriou. Heuristically optimized trade-offs: A new paradigm for power laws in the Internet. Lecture Notes in Computer Science, 2380:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2380/23800110.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2380/23800110.pdf.

Fotakis:2002:SCN

[351] Dimitris Fotakis, Spyros Kontogiannis, Elias Koutsoupias, Marios Mavronicolas, and Paul Spirakis. The structure and complexity of Nash equilibria for a selfish routing game. Lecture Notes in Computer Science, 2380:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800123.
pdf.

Khanna:2002:CMA

[352] Sanjeev Khanna, Joseph (Seffi) Naor, and Dan Raz. Control message aggregation in group communication protocols. Lecture Notes in Computer Science, 2380:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800135.
pdf.

Jurdzinski:2002:CRL

[353] Tomasz Jurdziński and Krzysztof Lorys. Church–Rosser languages vs. UCFL. Lecture Notes in Computer Science, 2380:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2380/23800147.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2380/23800147.pdf.

Bala:2002:IRL

[354] Sebastian Bala. Intersection of regular languages and star hierarchy. Lecture Notes in Computer Science, 2380: 159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800159.
                                        pdf.

Lombardy:2002:CRA

[355] Sylvain Lombardy. On the construction of reversible automata for reversible languages. Lecture Notes in Computer Science, 2380:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800170.
                                        pdf.

Elmasry:2002:PQP

[356] Amr Elmasry. Priority queues, pairing, and adaptive sorting. Lecture Notes in Computer Science, 2380:183– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800183.
                                        pdf.

Bender:2002:ESE

[357] Michael A. Bender, Richard Cole, and Rajeev Raman. Exponential structures for efficient cacheoblivious algorithms. Lecture Notes in Computer Science, 2380:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800195.
pdf.

Impagliazzo:2002:BDF

[358] Russell Impagliazzo and Nathan Segerlind. Bounded-depth Frege systems with counting axioms polynomially simulate Nullstellensatz refutations. Lecture Notes in Computer Science, 2380:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2380/23800208.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2380/23800208.pdf.

Esteban:2002:CRB

[359] Juan Luis Esteban, Nicola Galesi, and Jochen Messner. On the complexity of resolution with bounded conjunctions. Lecture Notes in Computer Science, 2380:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2380/23800220.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2380/23800220.
pdf.

Kiayias:2002:CHB

[360] Aggelos Kiayias and Moti Yung. Cryptographic hardness based on the decoding of Reed–Solomon codes. Lecture Notes in Computer Science, 2380: 232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800232.
                                        pdf.

Ishai:2002:PCR

[361] Yuval Ishai and Eyal Kushilevitz. Perfect constant-round secure computation via perfect randomizing polynomials. Lecture Notes in Computer Science, 2380:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2380/23800244.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2380/23800244.pdf.

Grigoriev:2002:ELB

[362] Dima Grigoriev, Edward A. Hirsch, and Dmitrii V. Pasechnik. Exponential lower bound for static semialgebraic proofs. Lecture Notes in Computer Science, 2380:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800257.
                                        pdf.

Jakoby:2002:PPS

[363] Andreas Jakoby and Maciej Liskiewicz. Paths problems in symmetric logarithmic space. Lecture Notes in Computer Science, 2380:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800269.
pdf.

Damaschke:2002:SSP

[364] Peter Damaschke. Scheduling search procedures. Lecture Notes in Computer Science, 2380:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2380/23800281.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2380/23800281.pdf.

Iwama:2002:ROK

[365] Kazuo Iwama and Shiro Taketomi. Removable online knapsack problems. Lecture Notes in Computer Science, 2380:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800293.
pdf.

Epstein:2002:NBV

[366] Leah Epstein, Steve Seiden, and Rob van Stee. New bounds for variablesized and resource augmented online bin packing. Lecture Notes in Computer Science, 2380:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2380/23800306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800306.
                                        pdf.

Ollinger:2002:QSU

[367] Nicolas Ollinger. The quest for small universal cellular automata. Lecture Notes in Computer Science, 2380:318– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800318.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800318.
                                        pdf.

Papazian:2002:HRG

[368] Christophe Papazian and Eric Rémila. Hyperbolic recognition by graph automata. Lecture Notes in Computer Science, 2380:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2380/23800330.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2380/23800330.pdf.

Ablayev:2002:QSB

[369] Farid Ablayev, Cristopher Moore, and Christopher Pollett. Quantum and stochastic branching programs of bounded width. Lecture Notes in Computer Science, 2380:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800343.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2380/23800343.
pdf.

Gargano:2002:STB

[370] Luisa Gargano, Pavol Hell, Ladislav Stacho, and Ugo Vaccaro. Spanning trees with bounded number of branch vertices. Lecture Notes in Computer Science, 2380:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800355.
pdf.

Beier:2002:EOR

[371] René Beier, Peter Sanders, and Naveen Sivadasan. Energy optimal routing in radio networks using geometric data structures. Lecture Notes in Computer Science, 2380:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800366.
pdf.

Christersson:2002:GBS

[372] Malin Christersson, Leszek Gasieniec, and Andrzej Lingas. Gossiping with bounded size messages in ad hoc radio networks. Lecture Notes in Computer Science, 2380:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2380/23800377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800377.
                                        pdf.

Merkle:2002:KLS

[373] Wolfgang Merkle. The Kolmogorov– Loveland stochastic sequences are not closed under selecting subsequences. Lecture Notes in Computer Science, 2380:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2380/23800390.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800390.
                                        pdf.

Hearn:2002:NCL

[374] Robert A. Hearn and Erik D. Demaine. The nondeterministic constraint logic model of computation: Reductions and applications. Lecture Notes in Computer Science, 2380:401– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800401.
                                        pdf.

Dalmau:2002:CSP

[375] Víctor Dalmau. Constraint satisfaction problems in non-deterministic logarithmic space. Lecture Notes in Computer Science, 2380:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2380/23800414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800414.
pdf.

Brodal:2002:COD

[376] Gerth Stølting Brodal and Rolf Fagerberg. Cache oblivious distribution sweeping. Lecture Notes in Computer Science, 2380:426–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800426.
pdf.

Ostlin:2002:OPS

[377] Anna Östlin and Rasmus Pagh. One-probe search. Lecture Notes in Computer Science, 2380:439–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800439.
pdf.

Charikar:2002:NAS

[378] Moses Charikar, Piotr Indyk, and Rina Panigrahy. New algorithms for subset query, partial match, orthogonal range searching, and related problems. Lecture Notes in Computer Science, 2380: 451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800451.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800451.
                                        pdf.

Martin:2002:MPP

[379] Keye Martin, Michael Mislove, and James Worrell. Measuring the probabilistic powerdomain. Lecture Notes in Computer Science, 2380:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800463.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800463.
                                        pdf.

Ong:2002:GCL

[380] C.-H. L. Ong and P. Di Gianantonio. Games characterizing Levy–Longo trees. Lecture Notes in Computer Science, 2380:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800476.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800476.
                                        pdf.

Bauer:2002:CFP

[381] Andrej Bauer, Martín Hötzel Escardó, and Alex Simpson. Comparing functional paradigms for exact real-number computation. Lecture Notes in Computer Science, 2380:488– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800488.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2380/23800488.
pdf.

Duchon:2002:RSB

[382] Philippe Duchon, Philippe Flajolet, Guy Louchard, and Gilles Schaeffer. Random sampling from Boltzmann principles. Lecture Notes in Computer Science, 2380:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800501.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800501.
pdf.

Duch:2002:APO

[383] Amalia Duch and Conrado Martínez. On the average performance of orthogonal range search in multidimensional data structures. Lecture Notes in Computer Science, 2380:514–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800514.
pdf.

Kick:2002:BMT

[384] Marco Kick. Bialgebraic modelling of timed processes. Lecture Notes in Computer Science, 2380:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800525.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800525.
                                        pdf.

vanBreugel:2002:TLM

[385] Franck van Breugel, Steven Shalit, and James Worrell. Testing labelled Markov processes. Lecture Notes in Computer Science, 2380:537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800537.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800537.
                                        pdf.

Hitchcock:2002:WCC

[386] John M. Hitchcock and Jack H. Lutz. Why computational complexity requires stricter martingales. Lecture Notes in Computer Science, 2380:549– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800549.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800549.
                                        pdf.

Hitchcock:2002:CPE

[387] John M. Hitchcock. Correspondence principles for effective dimensions. Lecture Notes in Computer Science, 2380: 561–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800561.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2380/23800561.
pdf.

Meseguer:2002:TAP

[388] José Meseguer and Grigore Rosu. A total approach to partial algebraic specification. Lecture Notes in Computer Science, 2380:572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800572.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800572.
pdf.

Lohrey:2002:AD

[389] Markus Lohrey, Pedro R. D’Argenio, and Holger Hermanns. Axiomatising divergence. Lecture Notes in Computer Science, 2380:585–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800585.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800585.
pdf.

Cardelli:2002:SLQ

[390] Luca Cardelli, Philippa Gardner, and Giorgio Ghelli. A spatial logic for querying graphs. Lecture Notes in Computer Science, 2380:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800597.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2380/23800597.
pdf.

Radzik:2002:ITB

[391] Tomasz Radzik. Improving time bounds on maximum generalised flow computations by contracting the network. Lecture Notes in Computer Science, 2380:611–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2380/23800611.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2380/23800611.pdf.

Berman:2002:AHB

[392] Piotr Berman and Marek Karpinski. Approximation hardness of bounded degree MIN-CSP and MIN-BISECTION. Lecture Notes in Computer Science, 2380:623–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2380/23800623.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2380/23800623.pdf.

Demetrescu:2002:IBN

[393] Camil Demetrescu and Giuseppe F. Italiano. Improved bounds and new trade-offs for dynamic all pairs shortest paths. Lecture Notes in Computer Science, 2380:633–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800633.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2380/23800633.
pdf.

Henzinger:2002:SUS

[394] Thomas A. Henzinger, Sriram C. Krishnan, Orna Kupferman, and Freddy Y. C. Mang. Synthesis of uninitialized systems. Lecture Notes in Computer Science, 2380:644–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800644.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800644.
pdf.

Genest:2002:ISH

[395] Blaise Genest, Anca Muscholl, Helmut Seidl, and Marc Zeitoun. Infinitestate high-level MSCs: Model-checking and realizability. Lecture Notes in Computer Science, 2380:657–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800657.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800657.
pdf.

Senizergues:2002:BDR

[396] Géraud Sénizergues. L(A) = L(B)? decidability results from complete formal systems. Lecture Notes in Computer Science, 2380:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800037.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800037.

pdf. Gödel Prize 2002.

Wich:2002:UIC

[397] Klaus Wich. Universal inherence of cycle-free context-free ambiguity functions. Lecture Notes in Computer Science, 2380:669–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800669.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800669.
                                        pdf.

Guha:2002:HDS

[398] Sudipto Guha, Piotr Indyk, S. Muthukrishnan, and Martin J. Strauss. Histogramming data streams with fast per-item processing. Lecture Notes in Computer Science, 2380:681–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800681.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800681.
                                        pdf.

Charikar:2002:FFI

[399] Moses Charikar, Kevin Chen, and Martin Farach-Colton. Finding frequent items in data streams. Lecture Notes in Computer Science, 2380:693– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800693.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2380/23800693.
pdf.

Cachat:2002:SSS

[400] Thierry Cachat. Symbolic strategy synthesis for games on pushdown graphs. Lecture Notes in Computer Science, 2380:704–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2380/23800704.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2380/23800704.
pdf.

Srba:2002:SBR

[401] Jiří Srba. Strong bisimilarity and regularity of basic process algebra is PSPACE-hard. Lecture Notes in Computer Science, 2380:716–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800716.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800716.
pdf.

Brodal:2002:SSS

[402] Gerth Stølting Brodal, Rune B. Lyngsø, Anna Östlin, and Christian N. S. Pedersen. Solving the string statistics problem in time O(\ log \). Lecture Notes in Computer Science, 2380:728–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800728.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800728.
                                        pdf.

Deng:2002:PDS

[403] Xiaotie Deng, Guojun Li, Zimao Li, Bin Ma, and Lusheng Wang. A PTAS for distinguishing (sub)string selection. Lecture Notes in Computer Science, 2380:740–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2380/23800740.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800740.
                                        pdf.

Kuske:2002:TOS

[404] Dietrich Kuske and Markus Lohrey. On the theory of one-step rewriting in trace monoids. Lecture Notes in Computer Science, 2380:752–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800752.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800752.
                                        pdf.

Bielecki:2002:NB

[405] Michał Bielecki, Jan Hidders, Jan Paredaens, Jerzy Tyszkiewicz, and Jan Van den Bussche. Navigating with a browser. Lecture Notes in Computer Science, 2380:764–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800764.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2380/23800764.
pdf.

Kumar:2002:IRS

[406] V. S. Anil Kumar and Madhav V. Marathe. Improved results for Stackelberg scheduling strategies. Lecture Notes in Computer Science, 2380:776– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800776.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800776.
pdf.

Adamy:2002:CCR

[407] Udo Adamy, Christoph Ambuehl, R. Sai Anand, and Thomas Erlebach. Call control in rings. Lecture Notes in Computer Science, 2380:788–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800788.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800788.
pdf.

Chrobak:2002:PSO

[408] Marek Chrobak, Leah Epstein, John Noga, Jiří Sgall, Rob van Stee, Tomáš Tichý, and Nodari Vakhania. Preemptive scheduling in overloaded systems. Lecture Notes in Computer Science, 2380:800–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2380/23800800.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800800.
                                        pdf.

Karhumaki:2002:EPF

[409] J. Karhumäki and L. P. Lisovik. The equivalence problem of finite substitutions on ab^{∗} c, with applications. Lecture Notes in Computer Science, 2380: 812–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800812.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800812.
                                        pdf.

Stirling:2002:DDE

[410] Colin Stirling. Deciding DPDA equivalence is primitive recursive. Lecture Notes in Computer Science, 2380:821– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800821.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800821.
                                        pdf.

Bojanczyk:2002:TWA

[411] Mikołaj Bojańczyk. Two-way alternating automata and finite models. Lecture Notes in Computer Science, 2380: 833–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800833.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2380/23800833.
pdf.

Berman:2002:AHC

[412] Piotr Berman, Marek Karpinski, and Yakov Nekrich. Approximating Huffman codes in parallel. Lecture Notes in Computer Science, 2380:845–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800845.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800845.
pdf.

Fantozzi:2002:SIP

[413] Carlo Fantozzi, Andrea Pietracaprina, and Geppino Pucci. Seamless integration of parallelism and memory hierarchy. Lecture Notes in Computer Science, 2380:856–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800856.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800856.
pdf.

Nisan:2002:CCA

[414] Noam Nisan. The communication complexity of approximate set packing and covering. Lecture Notes in Computer Science, 2380:868–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800868.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2380/23800868.
                                        pdf.

Doerr:2002:AWG

[415] Benjamin Doerr. Antirandomizing the wrong game. Lecture Notes in Computer Science, 2380:876–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800876.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800876.
                                        pdf.

Akcoglu:2002:FUI

[416] Karhan Akcoglu, Petros Drineas, and Ming-Yang Kao. Fast universalization of investment strategies with provably good relative returns. Lecture Notes in Computer Science, 2380:888– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800888.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800888.
                                        pdf.

Adler:2002:RPE

[417] Micah Adler, Harald Räcke, Naveen Sivadasan, Christian Sohler, and Berthold Vöcking. Randomized pursuit-evasion in graphs. Lecture Notes in Computer Science, 2380:901– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800901.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2380/23800901.
pdf.

Wells:2002:EPT

[418] J. B. Wells. The essence of principal typings. Lecture Notes in Computer Science, 2380:913–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800913.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800913.
pdf.

Adsul:2002:CTL

[419] Bharat Adsul and Milind Sohoni. Complete and tractable local linear time temporal logics over traces. Lecture Notes in Computer Science, 2380:926– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800926.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800926.
pdf.

Gastin:2002:EEC

[420] Paul Gastin and Madhavan Mukund. An elementary expressively complete temporal logic for Mazurkiewicz traces. Lecture Notes in Computer Science, 2380:938–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2380/23800938.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2380/23800938.
pdf.

Brattka:2002:RNI

[421] Vasco Brattka. Random numbers and an incomplete immune recursive set. Lecture Notes in Computer Science, 2380:950–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2380/23800950.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800950.
                                        pdf.

Hertling:2002:BMC

[422] Peter Hertling. A Banach-Mazur computable but not Markov computable function on the computable real numbers. Lecture Notes in Computer Science, 2380:962–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2380/23800962.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2380/23800962.pdf.

Czumaj:2002:PTA

[423] Artur Czumaj, Andrzej Lingas, and Hairong Zhao. Polynomialtime approximation schemes for the Euclidean survivable network design problem. Lecture Notes in Computer Science, 2380:973–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23800973.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23800973.
                                        pdf.

Bjorklund:2002:FPS

[424] Andreas Björklund and Thore Husfeldt. Finding a path of superlogarithmic length. Lecture Notes in Computer Science, 2380:985–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800985.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800985.
pdf.

Uehara:2002:LTA

[425] Ryuhei Uehara. Linear time algorithms on chordal bipartite and strongly chordal graphs. Lecture Notes in Computer Science, 2380:993– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23800993.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23800993.
pdf.

Holmerin:2002:IIR

[426] Jonas Holmerin. Improved inapproximability results for vertex cover on kuniform hypergraphs. Lecture Notes in Computer Science, 2380:1005–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23801005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23801005.
pdf.

Kohayakawa:2002:ETH

[427] Yoshiharu Kohayakawa, Brendan Nagle, and Vojtěch Rödl. Efficient testing of hypergraphs. Lecture Notes in Computer Science, 2380:1017–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23801017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23801017.
                                        pdf.

Wu:2002:ONS

[428] Xiaodong Wu and Danny Z. Chen. Optimal net surface problems with applications. Lecture Notes in Computer Science, 2380:1029–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23801029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23801029.
                                        pdf.

Bonichon:2002:WTR

[429] Nicolas Bonichon, Bertrand Le Saëc, and Mohamed Mosbah. Wagner’s theorem on realizers. Lecture Notes in Computer Science, 2380:1043–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2380/23801043.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2380/23801043.
                                        pdf.

Liberatore:2002:CA

[430] Vincenzo Liberatore. Circular arrangements. Lecture Notes in Computer Science, 2380:1054–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2380/23801054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2380/23801054.
pdf.

Anonymous:2002:AIj

[431] Anonymous. Author index. Lecture Notes in Computer Science, 2380:1067– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2380/2380auth.pdf.

Baaz:2002:PARa

[432] Matthias Baaz. Proof analysis by resolution. Lecture Notes in Computer Science, 2381:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810001.
pdf.

Miller:2002:ULL

[433] Dale Miller and Elaine Pimentel. Using linear logic to reason about sequent systems. Lecture Notes in Computer Science, 2381:2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810002.
pdf.

Baaz:2002:STS

[434] Matthias Baaz and Agata Ciabattoni. A schütte-tait style cut-elimination proof for first-order Gödel logic. Lecture Notes in Computer Science, 2381: 24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810024.
pdf.

Blackburn:2002:TQH

[435] Patrick Blackburn and Maarten Marx. Tableaux for quantified hybrid logic. Lecture Notes in Computer Science, 2381:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2381/23810038.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2381/23810038.
pdf.

Chetcuti-Sperandio:2002:TBA

[436] Nathalie Chetcuti-Sperandio. Tableaubased automated deduction for duration calculus. Lecture Notes in Computer Science, 2381:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2381/23810053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810053.
pdf.

Mayer:2002:LTL

[437] Marta Cialdea Mayer and Carla Limongelli. Linear time logic, conditioned models, and planning with incomplete knowledge. Lecture Notes in Computer Science, 2381:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810070.
pdf.

Degtyarev:2002:SCR

[438] Anatoli Degtyarev, Michael Fisher, and Boris Konev. A simplified clausal resolution procedure for propositional linear-time temporal logic. Lecture Notes in Computer Science, 2381:85– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810085.
pdf.

Eiter:2002:MNL

[439] Thomas Eiter, Volker Klotz, Hans Tompits, and Stefan Woltran. Modal nonmonotonic logics revisited: Efficient encodings for the basic reasoning tasks. Lecture Notes in Computer Science, 2381:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2381/23810100.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2381/23810100.pdf.

Ferrari:2002:TCL

[440] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. Tableau calculi for the logics of finite k-ary trees. Lecture Notes in Computer Science, 2381:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810115.
                                        pdf.

Giese:2002:MGS

[441] Martin Giese. A model generation style completeness proof for constraint tableaux with superposition. Lecture Notes in Computer Science, 2381:130– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810130.
                                        pdf.

Hladik:2002:IOT

[442] Jan Hladik. Implementation and optimisation of a tableau algorithm for the guarded fragment. Lecture Notes in Computer Science, 2381:145– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2381/23810145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810145.
pdf.

Letz:2002:LMC

[443] Reinhold Letz. Lemma and model caching in decision procedures for quantified Boolean formulas. Lecture Notes in Computer Science, 2381:160– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810160.
pdf.

Letz:2002:IER

[444] Reinhold Letz and Gernot Stenz. Integration of equality reasoning into the disconnection calculus. Lecture Notes in Computer Science, 2381:176– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810176.
pdf.

Metcalfe:2002:ASC

[445] George Metcalfe, Nicola Olivetti, and Dov Gabbay. Analytic sequent calculi for Abelian and Łukasiewicz logics. Lecture Notes in Computer Science, 2381:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2381/23810191.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810191.
                                        pdf.

Nguyen:2002:ATS

[446] Linh Anh Nguyen. Analytic tableau systems for propositional bimodal logics of knowledge and belief. Lecture Notes in Computer Science, 2381:206– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810206.
                                        pdf.

Petermann:2002:CTC

[447] Uwe Petermann. A confluent theory connection calculus. Lecture Notes in Computer Science, 2381:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810221.
                                        pdf.

Sofronie-Stokkermans:2002:UWP

[448] Viorica Sofronie-Stokkermans. On uniform word problems involving bridging operators on distributive lattices. Lecture Notes in Computer Science, 2381: 235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810235.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2381/23810235.
pdf.

tenCate:2002:QAP

[449] Balder ten Cate and Chung chieh Shan. Question answering: From partitions to prolog. Lecture Notes in Computer Science, 2381:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810251.
pdf.

Thion:2002:GTP

[450] V. Thion, S. Cerrito, and Marta Cialdea Mayer. A general theorem prover for quantified modal logics. Lecture Notes in Computer Science, 2381:266– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810266.
pdf.

Willard:2002:SNE

[451] Dan E. Willard. Some new exceptions for the semantic tableaux version of the second incompleteness theorem. Lecture Notes in Computer Science, 2381: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810281.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2381/23810281.
                                        pdf.

Wirth:2002:NIS

[452] Claus-Peter Wirth. A new indefinite semantics for Hilbert’s epsilon. Lecture Notes in Computer Science, 2381: 298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810298.
                                        pdf.

Zarba:2002:TCC

[453] Calogero G. Zarba. A tableau calculus for combining non-disjoint theories. Lecture Notes in Computer Science, 2381:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2381/23810315.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810315.
                                        pdf.

Habert:2002:LPE

[454] L. Habert, J.-M. Notin, and D. Galmiche. LINK: A proof environment based on proof nets. Lecture Notes in Computer Science, 2381:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2381/23810330.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2381/23810330.
                                        pdf.

Stenz:2002:DSA

[455] Gernot Stenz. DCTP 1.2 — system abstract. Lecture Notes in Computer Science, 2381:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2381/23810335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2381/23810335.
pdf.

Anonymous:2002:AIk

[456] Anonymous. Author index. Lecture Notes in Computer Science, 2381:341– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2381/2381auth.pdf.

Segev:2002:EDC

[457] Arie Segev. Enabling design-centric eBusiness applications. Lecture Notes in Computer Science, 2382:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2382/23820001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820001.
pdf.

Cohen:2002:SPQ

[458] Sara Cohen, Yaron Kanza, and Yehoshua Sagiv. Select-project queries over XML documents. Lecture Notes in Computer Science, 2382:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2382/23820002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820002.
pdf.

Balaban:2002:ACR

[459] Mira Balaban, Nikolai Berezansky, and Ehud Gudes. Answering cooperative recursive queries in Web federated databases. Lecture Notes in Computer Science, 2382:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2382/23820014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820014.
pdf.

Draheim:2002:STS

[460] Dirk Draheim and Gerald Weber. Strongly typed server pages. Lecture Notes in Computer Science, 2382:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2382/23820029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820029.
pdf.

Elovici:2002:NPM

[461] Yuval Elovici, Bracha Shapira, and Adlai Maschiach. A new privacy model for Web surfing. Lecture Notes in Computer Science, 2382:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2382/23820045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820045.
pdf.

Zeinalipour-Yazti:2002:DID

[462] Demetrios Zeinalipour-Yazti and Marios Dikaiakos. Design and implementation of a distributed crawler and filtering processor. Lecture Notes in Computer Science, 2382:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2382/23820058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820058.
pdf.

Wolfson:2002:MOI

[463] Ouri Wolfson. Moving objects information management: The database challenge. Lecture Notes in Computer Science, 2382:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2382/23820075.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2382/23820075.pdf.

Benaroch:2002:SLO

[464] Michel Benaroch. Specifying local ontologies in support of semantic interoperability of distributed interorganizational applications. Lecture Notes in Computer Science, 2382:90– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2382/23820090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2382/23820090.
                                        pdf.

Kabeli:2002:FOM

[465] Judith Kabeli and Peretz Shoval. FOOM and OPM methodologies — experimental comparison of user comprehension. Lecture Notes in Computer Science, 2382:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2382/23820107.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2382/23820107.pdf.

Bollen:2002:NLM

[466] Peter Bollen. The natural language modeling procedure. Lecture Notes in Computer Science, 2382:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2382/23820123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2382/23820123.
                                        pdf.

Bahn:2002:CAS

[467] David Bahn, J. David Naumann, and Shawn Curley. Conversation about software requirements with prototypes and scenarios. Lecture Notes in Computer Science, 2382:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2382/23820147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820147.
pdf.

Adi:2002:SMC

[468] Asaf Adi, David Botzer, and Opher Etzion. The situation manager component of amit — active middleware technology. Lecture Notes in Computer Science, 2382:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2382/23820158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2382/23820158.
pdf.

Anonymous:2002:AIl

[469] Anonymous. Author index. Lecture Notes in Computer Science, 2382:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2382/2382auth.pdf.

Lew:2002:CIV

[470] Michael S. Lew, Nicu Sebe, and John P. Eakins. Challenges of image and video retrieval. Lecture Notes in Computer Science, 2383:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830001.
pdf.

Tian:2002:VEU

[471] Qi Tian, Baback Moghaddam, and Thomas S. Huang. Visualization, estimation and user-modeling for interactive browsing of image libraries. Lecture Notes in Computer Science, 2383: 7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830007.
                                        pdf.

Sebe:2002:RSM

[472] Nicu Sebe and Michael Lew. Robust shape matching. Lecture Notes in Computer Science, 2383:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830017.
                                        pdf.

Wang:2002:SBI

[473] Wei Wang, Yuqing Song, and Aidong Zhang. Semantics-based image retrieval by region saliency. Lecture Notes in Computer Science, 2383:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830029.
                                        pdf.

Muller:2002:TAC

[474] Henning Müller, Stephane Marchand-Maillet, and Thierry Pun. The truth about Corel — evaluation in image retrieval. Lecture Notes in Computer Science, 2383:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830038.
pdf.

Bosson:2002:NRB

[475] Alison Bosson, Gavin C. Cawley, Yi Chan, and Richard Harvey. Non-retrieval: Blocking pornographic images. Lecture Notes in Computer Science, 2383:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830050.
pdf.

Cox:2002:LIP

[476] Gregory Cox and Gerhard de Jager. A linear image-pair model and the associated hypothesis test for matching. Lecture Notes in Computer Science, 2383: 61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830061.
pdf.

Smolka:2002:CFB

[477] Bogdan Smolka and Konstantinos N. Plataniotis. On the coupled forward and backward anisotropic diffusion scheme for color image enhancement. Lecture Notes in Computer Science, 2383:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2383/23830070.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830070.
                                        pdf.

Ko:2002:MRT

[478] ByoungChul Ko and Hyeran Byun. Multiple regions and their spatial relationship-based image retrieval. Lecture Notes in Computer Science, 2383: 81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830081.
                                        pdf.

Fauzi:2002:QFC

[479] Mohammad F. A. Fauzi and Paul H. Lewis. Query by fax for contentbased image retrieval. Lecture Notes in Computer Science, 2383:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830091.
                                        pdf.

Qiu:2002:SLC

[480] Guoping Qiu and Kin-Man. Spectrally layered color indexing. Lecture Notes in Computer Science, 2383:100– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830100.
pdf.

Wolf:2002:UIR

[481] Jürgen Wolf, Wolfram Burgard, and Hans Burkhardt. Using an image retrieval system for vision-based mobile robot localization. Lecture Notes in Computer Science, 2383:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830108.
pdf.

Feng:2002:JIR

[482] Guocan Feng and Jianmin Jiang. JPEG image retrieval based on features from DCT domain. Lecture Notes in Computer Science, 2383:120– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830120.
pdf.

Howell:2002:IRM

[483] A. Jonathan Howell and David S. Young. Image retrieval methods for a database of funeral monuments. Lecture Notes in Computer Science, 2383: 129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830129.
pdf.

deBruijn:2002:AMP

[484] Willem de Bruijn and Michael S. Lew. AtomsNet: Multimedia peer2peer file sharing. Lecture Notes in Computer Science, 2383:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830138.
pdf.

Hussain:2002:VCT

[485] Mustaq Hussain, John Eakins, and Graham Sexton. Visual clustering of trademarks using the selforganizing map. Lecture Notes in Computer Science, 2383:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830147.
pdf.

Ruiz-del-Solar:2002:FIF

[486] Javier Ruiz del Solar and Pablo Navarrete. FACERET: An interactive face retrieval system based on self-organizing maps. Lecture Notes in Computer Science, 2383:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830157.
pdf.

Leung:2002:OBI

[487] Man-Wai Leung and Kwok-Leung Chan. Object-based image retrieval using hierarchical shape descriptor. Lecture Notes in Computer Science, 2383: 165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830165.
pdf.

Vendrig:2002:MPI

[488] Jeroen Vendrig and Marcel Worring. Multimodal person identification in movies. Lecture Notes in Computer Science, 2383:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830175.
pdf.

Schaffalitzky:2002:ASM

[489] F. Schaffalitzky and A. Zisserman. Automated scene matching in movies. Lecture Notes in Computer Science, 2383: 186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830186.
                                        pdf.

Denman:2002:CBA

[490] H. Denman, N. Rea, and A. Kokaram. Content based analysis for video from snooker broadcasts. Lecture Notes in Computer Science, 2383:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830198.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830198.
                                        pdf.

Enser:2002:RAM

[491] Peter G. B. Enser and Criss J. Sandom. Retrieval of archival moving imagery — CBIR outside the frame? Lecture Notes in Computer Science, 2383: 206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830206.
                                        pdf.

Smeaton:2002:CCB

[492] Alan F. Smeaton. Challenges for content-based navigation of digital video in the Físchlár digital library. Lecture Notes in Computer Science, 2383:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2383/23830215.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2383/23830215.
pdf.

deAlarcon:2002:SIN

[493] Pedro A. de Alarcón, Alberto D. Pascual-Montano, and José M. Carazo. Spin images and neural networks for efficient content-based retrieval in 3D object databases. Lecture Notes in Computer Science, 2383:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830225.
pdf.

Brucale:2002:SFI

[494] A. Brucale, M. d’Amico, M. Ferri, L. Gualandri, and A. Lovato. Size functions for image retrieval: A demonstrator on randomly generated curves. Lecture Notes in Computer Science, 2383: 235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830235.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2383/23830235.
                                        pdf.

Berens:2002:ECT

[495] Jeff Berens and Graham D. Finlayson. An efficient coding of three dimensional colour distributions for image retrieval. Lecture Notes in Computer Science, 2383:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830245.
                                        pdf.

Riley:2002:CBR

[496] K. Jonathan Riley and John P. Eakins. Content-based retrieval of historical watermark images: I-tracings. Lecture Notes in Computer Science, 2383:253– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830253.
                                        pdf.

Visser:2002:ORV

[497] Rene Visser, Nicu Sebe, and Erwin Bakker. Object recognition for video retrieval. Lecture Notes in Computer Science, 2383:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830262.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2383/23830262.
pdf.

Bakker:2002:SVR

[498] Erwin M. Bakker and Michael S. Lew. Semantic video retrieval using audio analysis. Lecture Notes in Computer Science, 2383:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830271.
pdf.

Kim:2002:ESI

[499] Kyungsu Kim, Junho Choi, Namjung Kim, and Pankoo Kim. Extracting semantic information from basketball video based on audiovisual features. Lecture Notes in Computer Science, 2383:278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830278.
pdf.

Huang:2002:VIR

[500] Jeffrey Huang, Deepa Umamaheswaran, and Mathew Palakal. Video indexing and retrieval for archeological digital library, CLIOH. Lecture Notes in Computer Science, 2383:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2383/23830289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830289.
                                        pdf.

Oh:2002:FNI

[501] Kun Seok Oh, Aghbari Zaher, and Pan Koo Kim. Fast k-NN image search with self-organizing maps. Lecture Notes in Computer Science, 2383:299– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830299.
                                        pdf.

Pickering:2002:VRF

[502] Marcus J. Pickering, Stefan M. Rüger, and David Sinclair. Video retrieval by feature learning in key frames. Lecture Notes in Computer Science, 2383: 309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2383/23830309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2383/23830309.
                                        pdf.

Obdrzalek:2002:LAF

[503] Štepán Obdrzálek and Jirí Matas. Local affine frames for image retrieval. Lecture Notes in Computer Science, 2383:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2383/23830318.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2383/23830318.
pdf.

Park:2002:RAU

[504] Gunhan Park, Yunju Baek, and Heung-Kyu Lee. A ranking algorithm using dynamic clustering for contentbased image retrieval. Lecture Notes in Computer Science, 2383:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830328.
pdf.

Orriols:2002:OBV

[505] Xavier Orriols and Xavier Binefa. Online Bayesian video summarization and linking. Lecture Notes in Computer Science, 2383:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830338.
pdf.

Viallet:2002:FDV

[506] Jean Emmanuel Viallet and Olivier Bernier. Face detection for video summaries. Lecture Notes in Computer Science, 2383:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2383/23830348.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2383/23830348.pdf.

Black:2002:MEP

[507] John A. Black, Jr., Gamal Fahmy, and Sethuraman Panchanathan. A method for evaluating the performance of content-based image retrieval systems based on subjectively determined similarity between images. Lecture Notes in Computer Science, 2383:356– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830356.
pdf.

Sebe:2002:ESP

[508] N. Sebe, Q. Tian, E. Loupias, M. Lew, and T. Huang. Evaluation of salient point techniques. Lecture Notes in Computer Science, 2383:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2383/23830367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830367.
pdf.

Burke:2002:PCT

[509] Mary A. Burke. Personal construct theory as a research tool for analysing user perceptions of photographs. Lecture Notes in Computer Science, 2383: 378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2383/23830378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2383/23830378.
pdf.

Anonymous:2002:AIm

[510] Anonymous. Author index. Lecture Notes in Computer Science, 2383:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2383/2383ind.pdf.

Biehl:2002:NDP

[511] Ingrid Biehl and Tsuyoshi Takagi. A new distributed primality test for shared RSA keys using quadratic fields. Lecture Notes in Computer Science, 2384:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840001.
pdf.

Yang:2002:SAI

[512] Yanjiang Yang, Feng Bao, and Robert H. Deng. Security analysis and improvement of the global key recovery system. Lecture Notes in Computer Science, 2384:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840017.
pdf.

Clark:2002:LIK

[513] A. Clark, Ed Dawson, J. Fuller, J. Golić, H-J. Lee, William Millan, S-J. Moon, and L. Simpson. The LILI-II keystream generator. Lecture Notes in Computer Science, 2384:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840025.
                                        pdf.

Kurnio:2002:SRK

[514] Hartono Kurnio, Rei Safavi-Naini, and Huaxiong Wang. A secure re-keying scheme with key recovery property. Lecture Notes in Computer Science, 2384:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840040.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840040.
                                        pdf.

Henderson:2002:MTS

[515] Marie Henderson, Robert Coulter, Ed Dawson, and Eiji Okamoto. Modelling trust structures for public key infrastructures. Lecture Notes in Computer Science, 2384:56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840056.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2384/23840056.
pdf.

Barwick:2002:SBT

[516] S. G. Barwick, W.-A. Jackson, Keith M. Martin, and Peter R. Wild. Size of broadcast in threshold schemes with disenrollment. Lecture Notes in Computer Science, 2384:71– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840071.
pdf.

King:2002:RGI

[517] Brian King. Requirements for group independent linear threshold secret sharing schemes. Lecture Notes in Computer Science, 2384:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840089.
pdf.

Bennett:2002:ESE

[518] Krista Bennett, Christian Grothoff, Tzvetan Horozov, and Ioana Patrascu. Efficient sharing of encrypted data. Lecture Notes in Computer Science, 2384:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2384/23840107.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2384/23840107.
                                        pdf.

Pieprzyk:2002:CPL

[519] Josef Pieprzyk and Xian-Mo Zhang. Cheating prevention in linear secret sharing. Lecture Notes in Computer Science, 2384:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2384/23840121.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2384/23840121.pdf.

Fischer:2002:NFC

[520] Wieland Fischer and Jean-Pierre Seifert. Note on fast computation of secret RSA exponents. Lecture Notes in Computer Science, 2384:136– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840136.
                                        pdf.

Reyzin:2002:BTB

[521] Leonid Reyzin and Natan Reyzin. Better than BiBa: Short onetime signatures with fast signing and verifying. Lecture Notes in Computer Science, 2384:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840144.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2384/23840144.
pdf.

Wu:2002:CSCa

[522] Hongjun Wu and Feng Bao. Cryptanalysis of stream cipher COS^{(2,128)} Mode I. Lecture Notes in Computer Science, 2384:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2384/23840154.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2384/23840154.
pdf.

Soldera:2002:AZS

[523] David Soldera, Jennifer Seberry, and Chengxin Qu. The analysis of zheng-seberry scheme. Lecture Notes in Computer Science, 2384:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840159.
pdf.

Wu:2002:CSCb

[524] Hongjun Wu. Cryptanalysis of stream cipher Alpha1. Lecture Notes in Computer Science, 2384:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840169.
pdf.

Hughes:2002:LAA

[525] James Hughes. A linear algebraic attack on the AAFG1 braid group cryptosystem. Lecture Notes in Computer Science, 2384:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840176.
pdf.

Choie:2002:ICH

[526] Y. Choie and D. Yun. Isomorphism classes of hyperelliptic curves of genus 2 over F_{q}. Lecture Notes in Computer Science, 2384:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840190.
pdf.

Zhang:2002:CRD

[527] Fangguo Zhang, Shengli Liu, and Kwangjo Kim. Compact representation of domain parameters of hyperelliptic curve cryptosystems. Lecture Notes in Computer Science, 2384:203– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840203.
pdf.

Hitchcock:2002:NEC

[528] Yvonne Hitchcock and Paul Montague. A new elliptic curve scalar multiplication algorithm to resist simple power analysis. Lecture Notes in Computer Science, 2384:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840214.
pdf.

May:2002:SKS

[529] Lauren May, Matt Henricksen, William Millan, Gary Carter, and Ed Dawson. Strengthening the key schedule of the AES. Lecture Notes in Computer Science, 2384:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840226.
pdf.

Steinfeld:2002:NSA

[530] Ron Steinfeld, Joonsang Baek, and Yuliang Zheng. On the necessity of strong assumptions for the security of a class of asymmetric encryption schemes. Lecture Notes in Computer Science, 2384:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840241.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2384/23840241.
                                        pdf.

Warren:2002:SMI

[531] M. J. Warren and L. M. Batten. Security management: An information systems setting. Lecture Notes in Computer Science, 2384:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840257.
                                        pdf.

Ruan:2002:RCA

[532] Chun Ruan and Vijay Varadharajan. Resolving conflicts in authorization delegations. Lecture Notes in Computer Science, 2384:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840271.
                                        pdf.

Hitchens:2002:PAD

[533] M. Hitchens, Vijay Varadharajan, and G. Saunders. Policy administration domains. Lecture Notes in Computer Science, 2384:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840286.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2384/23840286.
pdf.

Zhou:2002:MVD

[534] Jianying Zhou. Maintaining the validity of digital signatures in B2B applications. Lecture Notes in Computer Science, 2384:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840303.
pdf.

Sebe:2002:SSF

[535] Francesc Sebé and Josep Domingo-Ferrer. Short 3-secure fingerprinting codes for copyright protection. Lecture Notes in Computer Science, 2384:316– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840316.
pdf.

Tada:2002:OSM

[536] Mitsuru Tada. An order-specified multisignature scheme secure against active insider attacks. Lecture Notes in Computer Science, 2384:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840328.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2384/23840328.
pdf.

England:2002:AOO

[537] Paul England and Marcus Peinado. Authenticated operation of open computing devices. Lecture Notes in Computer Science, 2384:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840346.
pdf.

Kim:2002:NIS

[538] Myungsun Kim and Kwangjo Kim. A new identification scheme based on the bilinear Diffie–Hellman problem. Lecture Notes in Computer Science, 2384: 362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840362.
pdf.

Roy:2002:BOR

[539] Bimal Roy. A brief outline of research on correlation immune functions. Lecture Notes in Computer Science, 2384: 379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840379.
pdf.

Mu:2002:OT

[540] Yi Mu, Junqi Zhang, and Vijay Varadharajan. m out of n oblivious transfer. Lecture Notes in Computer Science, 2384:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840395.
pdf.

Park:2002:SRV

[541] Sangwoo Park, Soo Hak Sung, Seongtaek Chee, and Jongin Lim. On the security of reduced versions of 3-pass HAVAL. Lecture Notes in Computer Science, 2384:406–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840406.
pdf.

Okeya:2002:ISC

[542] Katsuyuki Okeya and Kouichi Sakurai. On insecurity of the side channel attack countermeasure using addition-subtraction chains under distinguishability between addition and doubling. Lecture Notes in Computer Science, 2384:420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840420.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2384/23840420.
                                        pdf.

Sakurai:2002:SMP

[543] Kouichi Sakurai and Tsuyoshi Takagi. On the security of a modified Paillier public-key primitive. Lecture Notes in Computer Science, 2384:436– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840436.
                                        pdf.

Cubaleska:2002:HPS

[544] Biljana Cubaleska, Weidong Qiu, and Markus Schneider. How to play Sherlock Holmes in the world of mobile agents. Lecture Notes in Computer Science, 2384:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2384/23840449.
                                        pdf.

Han:2002:PAD

[545] Dong-Guk Han, Hye-Young Park, Young-Ho Park, Sangjin Lee, Dong Hoon Lee, and Hyung-Jin Yang. A practical approach defeating blackmailing. Lecture Notes in Computer Science, 2384: 464–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2384/23840464.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2384/23840464.
pdf.

Kim:2002:PAP

[546] Hyun-Jeong Kim, Dong Hoon Lee, and Moti Yung. Privacy against piracy: Protecting two-level revocable P-K traitor tracing. Lecture Notes in Computer Science, 2384:482– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840482.
pdf.

Kumar:2002:APS

[547] M. V. N. Ashwin Kumar, K. Srinathan, and C. Pandu Rangan. Asynchronous perfectly secure computation tolerating generalized adversaries. Lecture Notes in Computer Science, 2384:497– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2384/23840497.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2384/23840497.
pdf.

Anonymous:2002:AIn

[548] Anonymous. Author index. Lecture Notes in Computer Science, 2384:513– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2384/2384auth.pdf.

Freuder:2002:CA

[549] Eugene C. Freuder. Constraint acquisition. Lecture Notes in Computer Science, 2385:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2385/23850001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2385/23850001.pdf.

Colmerauer:2002:ECF

[550] Alain Colmerauer. Expressiveness and complexity of full first-order constraints in the algebra of trees. Lecture Notes in Computer Science, 2385: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850002.
                                        pdf.

Deplagne:2002:DVC

[551] Eric Deplagne and Claude Kirchner. Deduction versus computation: The case of induction. Lecture Notes in Computer Science, 2385:4– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850004.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850004.
                                        pdf.

Sturm:2002:IQE

[552] Thomas Sturm. Integration of quantifier elimination with constraint logic programming. Lecture Notes in Computer Science, 2385:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850007.
pdf.

Chetty:2002:THS

[553] Madhu Chetty. Towards a hybrid symbolic/numeric computational approach in controller design. Lecture Notes in Computer Science, 2385:12– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850012.
pdf.

Kitzelmann:2002:ISF

[554] Emanuel Kitzelmann, Ute Schmid, Martin Mühlpfordt, and Fritz Wysotzki. Inductive synthesis of functional programs. Lecture Notes in Computer Science, 2385:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2385/23850026.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2385/23850026.pdf.

Herrero:2002:SCB

[555] Begoña Herrero, Luis M. Laita, Eugenio Roanes-Lozano, Víctor Maojo, Luis de Ledesma, JoséCrespo, and Laura Laita. A symbolic computationbased expert system for alzheimer’s disease diagnosis. Lecture Notes in Computer Science, 2385:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850038.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850038.
                                        pdf.

Aiguier:2002:GLT

[556] Marc Aiguier, Diane Bahrami, and Catherine Dubois. On a generalised logicality theorem. Lecture Notes in Computer Science, 2385:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850051.
                                        pdf.

Smirnova:2002:USC

[557] Elena Smirnova. Using symbolic computation in an automated sequent derivation system for multi-valued logic. Lecture Notes in Computer Science, 2385:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2385/23850064.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2385/23850064.pdf.

Corless:2002:WF

[558] Robert M. Corless and D. J. Jeffrey. The Wright ω function. Lecture Notes in Computer Science, 2385:76– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850076.
pdf.

Nossum:2002:MLS

[559] Rolf Nossum and Luciano Serafini. Multicontext logic for semigroups of contexts. Lecture Notes in Computer Science, 2385:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850090.
pdf.

Campbell:2002:IIT

[560] J. A. Campbell. Indefinite integration as a testbed for developments in multi-agent systems. Lecture Notes in Computer Science, 2385:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850102.
pdf.

Hunter:2002:EIG

[561] Andrew Hunter. Expression inference — genetic symbolic classification integrated with non-linear coefficient optimisation. Lecture Notes in Computer Science, 2385:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850117.
                                        pdf.

Bai:2002:NFR

[562] Li Bai and Yihui Liu. A novel face recognition method. Lecture Notes in Computer Science, 2385:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850128.
                                        pdf.

Castaing:2002:NCL

[563] Jacqueline Castaing. Non-commutative logic for hand-written character modeling. Lecture Notes in Computer Science, 2385:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2385/23850136.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2385/23850136.pdf.

Loriette-Rougegrez:2002:NSD

[564] S. Loriette-Rougegrez. From numerical to symbolic data during the recognition of scenarii. Lecture Notes in Computer Science, 2385:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850154.
pdf.

Bernhaupt:2002:MMN

[565] Regina Bernhaupt and Jochen Pfalzgraf. On mathematical modeling of networks and implementation aspects. Lecture Notes in Computer Science, 2385:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850168.
pdf.

Ratschan:2002:CFO

[566] Stefan Ratschan. Continuous firstorder constraint satisfaction. Lecture Notes in Computer Science, 2385:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850181.
pdf.

Golumbic:2002:CAT

[567] Martin Charles Golumbic and Assaf Siani. Coloring algorithms for tolerance graphs: Reasoning and scheduling with interval constraints. Lecture Notes in Computer Science, 2385:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850196.
                                        pdf.

Tounsi:2002:GBA

[568] Mohamed Tounsi. A genetic-based approach for satisfiability problems. Lecture Notes in Computer Science, 2385: 208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850208.
                                        pdf.

Subramani:2002:ISQ

[569] K. Subramani. On identifying simple and quantified lattice points in the 2SAT polytope. Lecture Notes in Computer Science, 2385:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850217.
                                        pdf.

Audemard:2002:IBM

[570] Gilles Audemard, Piergiorgio Bertoli, Alessandro Cimatti, Artur Korniłowicz, and Roberto Sebastiani. Integrating Boolean and mathematical solving: Foundations, basic algorithms, and requirements. Lecture Notes in Computer Science, 2385:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850231.
pdf.

Beeson:2002:MIC

[571] Michael Beeson and Freek Wiedijk. The meaning of infinity in calculus and computer algebra systems. Lecture Notes in Computer Science, 2385: 246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850246.
pdf.

Colton:2002:MCA

[572] Simon Colton. Making conjectures about Maple functions. Lecture Notes in Computer Science, 2385:259–274, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850259.
pdf.

Meier:2002:ETF

[573] Andreas Meier, Volker Sorge, and Simon Colton. Employing theory formation to guide proof planning. Lecture Notes in Computer Science, 2385:275– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850275.
                                        pdf.

Kutsia:2002:USV

[574] Temur Kutsia. Unification with sequence variables and flexible arity symbols and its extension with patternterms. Lecture Notes in Computer Science, 2385:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2385/23850290.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2385/23850290.pdf.

Ranise:2002:CGD

[575] Silvio Ranise. Combining generic and domain specific reasoning by using contexts. Lecture Notes in Computer Science, 2385:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2385/23850305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2385/23850305.
                                        pdf.

Zimmer:2002:ITP

[576] Jürgen Zimmer and Louise A. Dennis. Inductive theorem proving and computer algebra in the Math-Web software bus. Lecture Notes in Computer Science, 2385:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850319.
pdf.

Pinkus:2002:ADI

[577] Ayal Z. Pinkus and Serge Winitzki. Y ACAS: A do-it-yourself symbolic algebra environment. Lecture Notes in Computer Science, 2385:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850332.
pdf.

Piroi:2002:FWN

[578] Florina Piroi and Bruno Buchberger. Focus windows: A new technique for proof presentation. Lecture Notes in Computer Science, 2385:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2385/23850337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2385/23850337.
pdf.

Anonymous:2002:AIo

[579] Anonymous. Author index. Lecture Notes in Computer Science, 2385:343– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2385/2385auth.pdf.

Giegerich:2002:IAD

[580] Robert Giegerich and Peter Steffen. Implementing algebraic dynamic programming in the functional and the imperative programming paradigm. Lecture Notes in Computer Science, 2386: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860001.
                                        pdf.

Kozen:2002:SRD

[581] Dexter Kozen. Some results in dynamic model theory. Lecture Notes in Computer Science, 2386:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860021.
                                        pdf.

Wing:2002:MCS

[582] Jeannette M. Wing. Mathematics in computer science curricula. Lecture Notes in Computer Science, 2386:22– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860022.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860022.
                                        pdf.

Backhouse:2002:LRG

[583] Kevin Backhouse and Roland Backhouse. Logical relations and Galois connections. Lecture Notes in Computer Science, 2386:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2386/23860023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2386/23860023.
pdf.

Duran:2002:TDG

[584] Juan Eduardo Durán. Transformational derivation of greedy network algorithms from descriptive specifications. Lecture Notes in Computer Science, 2386:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2386/23860040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2386/23860040.
pdf.

Harrison:2002:FCD

[585] William Harrison, Tim Sheard, and James Hook. Fine control of demand in haskell. Lecture Notes in Computer Science, 2386:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2386/23860068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2386/23860068.
pdf.

Hayes:2002:RAT

[586] Ian J. Hayes. Reasoning about timeouts. Lecture Notes in Computer Science, 2386:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2386/23860094.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2386/23860094.pdf.

Hesselink:2002:EVS

[587] Wim H. Hesselink. Eternity variables to simulate specifications. Lecture Notes in Computer Science, 2386: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860117.
                                        pdf.

Hinze:2002:CTR

[588] Ralf Hinze. Constructing tournament representations: An exercise in pointwise relational programming. Lecture Notes in Computer Science, 2386:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860131.
                                        pdf.

Hinze:2002:TID

[589] Ralf Hinze, Johan Jeuring, and Andres Löh. Type-indexed data types. Lecture Notes in Computer Science, 2386:

148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2386/23860148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2386/23860148.
pdf.

Huisman:2002:VJA

[590] Marieke Huisman. Verification of Java’s AbstractCollection class: A case study. Lecture Notes in Computer Science, 2386:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2386/23860175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2386/23860175.
pdf.

Liu:2002:SRP

[591] Yanhong A. Liu and Fuxiang Yu. Solving regular path queries. Lecture Notes in Computer Science, 2386:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2386/23860195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2386/23860195.
pdf.

Mu:2002:IFF

[592] Shin-Cheng Mu and Richard Bird. Inverting functions as folds. Lecture Notes in Computer Science, 2386:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860209.
                                        pdf.

vonWright:2002:KAR

[593] Joakim von Wright. From Kleene algebra to refinement algebra. Lecture Notes in Computer Science, 2386:233– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2386/23860233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2386/23860233.
                                        pdf.

Anonymous:2002:AIp

[594] Anonymous. Author index. Lecture Notes in Computer Science, 2386:263– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2386/23860263.pdf.

Myers:2002:AHM

[595] Gene Myers. The assembly of the human and mouse genomes. Lecture Notes in Computer Science, 2387:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870001.
                                        pdf.

Sahni:2002:DSO

[596] Sartaj Sahni. Data structures for onedimensional packet classification using most-specific-rule matching. Lecture Notes in Computer Science, 2387:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870002.
pdf.

Salomaa:2002:DCP

[597] Arto Salomaa. DNA complementarity and paradigms of computing. Lecture Notes in Computer Science, 2387: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870003.
pdf.

Cai:2002:HAM

[598] Jin-Yi Cai, Denis Charles, A. Pavan, and Samik Sengupta. On higher Arthur–Merlin classes. Lecture Notes in Computer Science, 2387:18– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870018.
pdf.

Deng:2002:SP

[599] Xiaotie Deng, C. H. Lee, Yunlei Zhao, and Hong Zhu. (2 + f(n))- SAT and its properties. Lecture Notes in Computer Science, 2387:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870028.
pdf.

Hoang:2002:MPM

[600] Thanh Minh Hoang and Thomas Thierauf. On the minimal polynomial of a matrix. Lecture Notes in Computer Science, 2387:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870037.
pdf.

Rettinger:2002:CRF

[601] Robert Rettinger, Xizhong Zheng, and Burchard von Braunmühl. Computable real functions of bounded variation and semi-computable real numbers (extended abstract). Lecture Notes in Computer Science, 2387:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870047.
pdf.

Lu:2002:ICR

[602] Hsueh-I Lu. Improved compact routing tables for planar networks via orderly spanning trees. Lecture Notes in Computer Science, 2387:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870057.
pdf.

Gabow:2002:CAS

[603] Harold N. Gabow and San Skulrattanakulchai. Coloring algorithms on subcubic graphs. Lecture Notes in Computer Science, 2387:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870067.
pdf.

Hsieh:2002:EAH

[604] Sun yuan Hsieh, Chin wen Ho, Tsan sheng Hsu, and Ming tat Ko. Efficient algorithms for the Hamiltonian problem on distance-hereditary graphs. Lecture Notes in Computer Science, 2387: 77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870077.
pdf.

Boyar:2002:EAF

[605] Joan Boyar, Lene M. Favrholdt, Kim S. Larsen, and Morten N. Nielsen. Extending the accommodating function. Lecture Notes in Computer Science, 2387:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870087.
                                        pdf.

Sun:2002:IPS

[606] Fangting Sun, David Fernández-Baca, and Wei Yu. Inverse parametric sequence alignment. Lecture Notes in Computer Science, 2387:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870097.
                                        pdf.

Lu:2002:FST

[607] Chin Lung Lu, Chuan Yi Tang, and Richard Chia-Tung Lee. The full Steiner tree problem in phylogeny. Lecture Notes in Computer Science, 2387: 107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870107.
                                        pdf.

Akutsu:2002:IUH

[608] Tatsuya Akutsu and Sascha Ott. Inferring a union of halfspaces from examples. Lecture Notes in Computer Science, 2387:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870117.
pdf.

Arslan:2002:DLW

[609] Abdullah N. Arslan and Ömer Eğecioğlu. Dictionary look-up within small edit distance. Lecture Notes in Computer Science, 2387:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870127.
pdf.

Lange:2002:PIE

[610] Tanja Lange and Arne Winterhof. Polynomial interpolation of the elliptic curve and XTR discrete logarithm. Lecture Notes in Computer Science, 2387:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870137.
pdf.

Grolmusz:2002:COC

[611] Vince Grolmusz. Co-orthogonal codes (extended abstract). Lecture Notes in Computer Science, 2387:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870144.
                                        pdf.

Kim:2002:EPS

[612] Nam-Yeun Kim, Won-Ho Lee, and Kee-Young Yoo. Efficient power-sum systolic architectures for public-key cryptosystems in GF(2^{m} ). Lecture Notes in Computer Science, 2387:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870153.
                                        pdf.

Wang:2002:CAA

[613] Huaxiong Wang and Josef Pieprzyk. A combinatorial approach to anonymous membership broadcast. Lecture Notes in Computer Science, 2387:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870162.
                                        pdf.

Dantsin:2002:SCS

[614] Evgeny Dantsin and Alexander Wolpert. Solving constraint satisfaction problems with DNA computing. Lecture Notes in Computer Science, 2387:171– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870171.
pdf.

Jigang:2002:NAA

[615] Wu Jigang, Heiko Schröder, and Thambipillai Srikanthan. New architecture and algorithms for degradable VLSI/WSI arrays. Lecture Notes in Computer Science, 2387:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870181.
pdf.

Carter:2002:CFT

[616] Casey Carter and Nicholas Tran. Cluster: A fast tool to identify groups of similar programs. Lecture Notes in Computer Science, 2387:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870191.
pdf.

Kikuchi:2002:BGB

[617] Yosuke Kikuchi, Shingo Osawa, and Yukio Shibata. Broadcasting in generalized de Bruijn digraphs (extended abstract). Lecture Notes in Computer Science, 2387:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870200.
                                        pdf.

Duckworth:2002:CDN

[618] William Duckworth and Bernard Mans. On the connected domination number of random regular graphs. Lecture Notes in Computer Science, 2387: 210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870210.
                                        pdf.

Chandran:2002:NMC

[619] L. Sunil Chandran and L. Shankar Ram. On the number of minimum cuts in a graph. Lecture Notes in Computer Science, 2387:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870220.
                                        pdf.

Chia:2002:CNR

[620] G. L. Chia and C. S. Gan. On crossing numbers of 5-regular graphs. Lecture Notes in Computer Science, 2387: 230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870230.
pdf.

Desmedt:2002:MFC

[621] Yvo Desmedt and Yongge Wang. Maximum flows and critical vertices in AND/OR graphs (extended abstract). Lecture Notes in Computer Science, 2387:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2387/23870238.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2387/23870238.
pdf.

Datta:2002:NEE

[622] Amitava Datta and Albert Y. Zomaya. New energy-efficient permutation routing protocol for single-hop radio networks. Lecture Notes in Computer Science, 2387:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2387/23870249.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2387/23870249.pdf.

Takamura:2002:SME

[623] Masataka Takamura and Yoshihide Igarashi. Simple mutual exclusion algorithms based on bounded tickets on the asynchronous shared memory model. Lecture Notes in Computer Science, 2387:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870259.
pdf.

Nakano:2002:TEO

[624] Koji Nakano. Time and energy optimal list ranking algorithms on the k-channel broadcast communication model. Lecture Notes in Computer Science, 2387:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870269.
pdf.

Jurdzinski:2002:EES

[625] Tomasz Jurdziński, Mirosław Kutyłowski, and Jan Zatopiański. Energy-efficient size approximation of radio networks with no collision detection. Lecture Notes in Computer Science, 2387:279– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870279.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2387/23870279.
pdf.

Martin-Vide:2002:NCS

[626] C. Martín-Vide, J. Pazos, G. Păun, and A. Rodríguez-Patón. A new class of symbolic abstract neural nets: Tissue P systems. Lecture Notes in Computer Science, 2387:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870290.
pdf.

Czyzowicz:2002:TSO

[627] Jurek Czyzowicz, Wojciech Fraczak, and Andrzej Pelc. Transducers with set output. Lecture Notes in Computer Science, 2387:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870300.
pdf.

Klein:2002:SAF

[628] Andreas Klein and Martin Kutrib. Selfassembling finite automata. Lecture Notes in Computer Science, 2387:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870310.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2387/23870310.
pdf.

Ilie:2002:RCW

[629] Lucian Ilie, Sheng Yu, and Kaizhong Zhang. Repetition complexity of words. Lecture Notes in Computer Science, 2387:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2387/23870320.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2387/23870320.
pdf.

Pandurangan:2002:UPC

[630] Gopal Pandurangan, Prabhakar Raghavan, and Eli Upfal. Using PageRank to characterize Web structure. Lecture Notes in Computer Science, 2387:330– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870330.
pdf.

Liu:2002:RBG

[631] Ding Liu and Manoj Prabhakaran. On randomized broadcasting and gossiping in radio networks. Lecture Notes in Computer Science, 2387:340– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870340.
pdf.

Altman:2002:FDC

[632] Tom Altman, Yoshihide Igarashi, and Kazuhiro Motegi. Fast and dependable communication in hyper-rings. Lecture Notes in Computer Science, 2387:350– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870350.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870350.
pdf.

Barequet:2002:LHT

[633] Gill Barequet. The on-line Heilbronn’s triangle problem in three and four dimensions. Lecture Notes in Computer Science, 2387:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870360.
pdf.

Schaefer:2002:ANC

[634] Marcus Schaefer, Eric Sedgwick, and Daniel Štefankovič. Algorithms for normal curves and surfaces. Lecture Notes in Computer Science, 2387:370– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870370.
pdf.

Ilinkin:2002:TPD

[635] Ivaylo Ilinkin, Ravi Janardan, and Michiel Smid. Terrain polygon decomposition, with application to layered manufacturing. Lecture Notes in Computer Science, 2387:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870381.
pdf.

Chen:2002:SF

[636] D. Chen, O. Eulenstein, David Fernández-Baca, and M. Sanderson. Supertrees by flipping. Lecture Notes in Computer Science, 2387:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870391.
pdf.

Lam:2002:STE

[637] Tak-Wah Lam, Kunihiko Sadakane, Wing-Kin Sung, and Siu-Ming Yiu. A space and time efficient algorithm for constructing compressed suffix arrays. Lecture Notes in Computer Science, 2387:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2387/23870401.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2387/23870401.
pdf.

Li:2002:SOR

[638] Ming Li, John Tromp, and Paul Vitányi. Sharpening Occam’s razor (extended abstract). Lecture Notes in Computer Science, 2387:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870411.
pdf.

Zhu:2002:APC

[639] Binhai Zhu. Approximating 3D points with cylindrical segments. Lecture Notes in Computer Science, 2387:420– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870420.
pdf.

Ito:2002:AMP

[640] Takehiro Ito, Takao Nishizeki, and Xiao Zhou. Algorithms for the multicolorings of partial k-trees. Lecture Notes in Computer Science, 2387:430– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870430.
pdf.

Ravikumar:2002:FTM

[641] B. Ravikumar. A fault-tolerant merge sorting algorithm. Lecture Notes in Computer Science, 2387:440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870440.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870440.
pdf.

Brankovis:2002:CUD

[642] Ljiljana Brankoviś and Jozef Širáň. 2compromise usability in 1-dimensional statistical databases. Lecture Notes in Computer Science, 2387:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870448.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870448.
pdf.

Chen:2002:ESC

[643] Danny Z. Chen, Shuang Luan, and Jinhui Xu. An experimental study and comparison of topological peeling and topological walk. Lecture Notes in Computer Science, 2387:456– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870456.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870456.
pdf.

Epstein:2002:LMN

[644] Leah Epstein and Lene M. Favrholdt. On-line maximizing the number of items packed in variable-sized bins. Lecture Notes in Computer Science, 2387:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2387/23870467.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2387/23870467.
pdf.

Fujita:2002:LGP

[645] Satoshi Fujita. On-line grid-packing with a single active grid. Lecture Notes in Computer Science, 2387:476– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870476.
pdf.

Mutzel:2002:BMO

[646] Petra Mutzel and René Weiskircher. Bend minimization in orthogonal drawings using integer programming. Lecture Notes in Computer Science, 2387: 484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870484.
pdf.

Wang:2002:CLM

[647] Biing-Feng Wang, Shan-Chyun Ku, and Yong-Hsian Hsieh. The conditional location of a median path. Lecture Notes in Computer Science, 2387: 494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870494.
pdf.

Ma:2002:NRT

[648] Weimin Ma, Yinfeng Xu, Jane You, James Liu, and Kanliang Wang. New results on the k-truck problem. Lecture Notes in Computer Science, 2387: 504–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870504.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870504.
pdf.

Srinathan:2002:TEF

[649] K. Srinathan, Pranava R. Goundan, M. V. N. Ashwin Kumar, R. Nandakumar, and C. Pandu Rangan. Theory of equal-flows in networks. Lecture Notes in Computer Science, 2387:514– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870514.
pdf.

Lin:2002:MBW

[650] Yaw-Ling Lin. Minimum back-walk-free latency problem (extended abstract). Lecture Notes in Computer Science, 2387:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2387/23870525.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2387/23870525.pdf.

Dahllof:2002:CSA

[651] Vilhelm Dahllöf, Peter Jonsson, and Magnus Wahlström. Counting satisfying assignments in 2-SAT and 3-SAT. Lecture Notes in Computer Science, 2387:535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870535.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870535.
pdf.

Tomescu:2002:MNI

[652] Ioan Tomescu. On the maximum number of irreducible coverings of an nvertex graph by n − 3 cliques. Lecture Notes in Computer Science, 2387: 544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870544.
pdf.

Nickelsen:2002:RGB

[653] Arfst Nickelsen and Till Tantau. On reachability in graphs with bounded independence number. Lecture Notes in Computer Science, 2387:554–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870554.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870554.
                                        pdf.

Fernau:2002:PE

[654] Henning Fernau. On parameterized enumeration. Lecture Notes in Computer Science, 2387:564–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870564.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870564.
                                        pdf.

Golovkins:2002:PRA

[655] Marats Golovkins and Maksim Kravtsev. Probabilistic reversible automata and quantum automata. Lecture Notes in Computer Science, 2387:574– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2387/23870574.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2387/23870574.
                                        pdf.

Yamasaki:2002:QVD

[656] Tomohiro Yamasaki, Hirotada Kobayashi, and Hiroshi Imai. Quantum versus deterministic counter automata. Lecture Notes in Computer Science, 2387:584– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870584.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870584.
pdf.

Jackson:2002:QDL

[657] Jeffrey C. Jackson, Christino Tamon, and Tomoyuki Yamakami. Quantum DNF learnability revisited. Lecture Notes in Computer Science, 2387:595– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2387/23870595.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2387/23870595.
pdf.

Anonymous:2002:AIq

[658] Anonymous. Author index. Lecture Notes in Computer Science, 2387:605– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2387/2387auth.pdf.

Mukherjee:2002:PSP

[659] Neelanjan Mukherjee and Sayan Mukherjee. Predicting signal peptides with support vector machines. Lecture Notes in Computer Science, 2388:

1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880001.
pdf.

Collobert:2002:SLL

[660] Ronan Collobert, Yoshua Bengio, and Samy Bengio. Scaling large learning problems with hard parallel mixtures. Lecture Notes in Computer Science, 2388:8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880008.
pdf.

Navarrete:2002:GKM

[661] Pablo Navarrete and Javier Ruiz del Solar. On the generalization of kernel machines. Lecture Notes in Computer Science, 2388:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880024.
pdf.

Tax:2002:KWO

[662] David M. J. Tax and Piotr Juszczak. Kernel whitening for one-class classification. Lecture Notes in Computer Science, 2388:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2388/23880040.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2388/23880040.pdf.

Dong:2002:FST

[663] Jian xiong Dong, Adam Krzyzak, and Ching Y. Suen. A fast SVM training algorithm. Lecture Notes in Computer Science, 2388:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880053.
pdf.

Fumera:2002:SVM

[664] Giorgio Fumera and Fabio Roli. Support vector machines with embedded reject option. Lecture Notes in Computer Science, 2388:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880068.
pdf.

Barla:2002:IK

[665] Annalisa Barla, Emanuele Franceschi, Francesca Odone, and Alessandro Verri. Image kernels. Lecture Notes in Computer Science, 2388:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880083.
                                        pdf.

Caputo:2002:CCS

[666] B. Caputo, Gy. Dorkó, and H. Niemann. Combining color and shape information for appearance-based object recognition using ultrametric spin glass-markov random fields. Lecture Notes in Computer Science, 2388:97– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880097.
                                        pdf.

Nakajima:2002:MTE

[667] Chikahito Nakajima and Massimiliano Pontil. Maintenance training of electric power facilities using object recognition by SVM. Lecture Notes in Computer Science, 2388:112– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880112.
                                        pdf.

Genov:2002:KSV

[668] Roman Genov and Gert Cauwenberghs. Kerneltron: Support vector ‘machine’ in silicon. Lecture Notes in Computer Science, 2388:120– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880120.
pdf.

Bileschi:2002:ACB

[669] Stanley M. Bileschi and Bernd Heisele. Advances in componentbased face detection. Lecture Notes in Computer Science, 2388:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880135.
pdf.

Walawalkar:2002:SVL

[670] L. Walawalkar, Mohammad Yeasin, Anand M. Narasimhamurthy, and Rajeev Sharma. Support vector learning for gender classification using audio and visual cues: A comparison. Lecture Notes in Computer Science, 2388: 144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880144.
pdf.

Chang:2002:ANT

[671] Ming-Wei Chang, Chih-Jen Lin, and Ruby C. Weng. Analysis of nonstationary time series using support vector machines. Lecture Notes in Computer Science, 2388:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880160.
pdf.

Sekhar:2002:RCV

[672] C. Chandra Sekhar, Kazuya Takeda, and Fumitada Itakura. Recognition of consonant-vowel (CV) units of speech in a broadcast news corpus using support vector machines. Lecture Notes in Computer Science, 2388:171– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880171.
pdf.

Fugate:2002:ADE

[673] Mike Fugate and James R. Gattiker. Anomaly detection enhanced classification in computer intrusion detection. Lecture Notes in Computer Science, 2388:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880186.
pdf.

Milanova:2002:SCK

[674] Mariofanna Milanova, Tomasz G. Smolinski, Grzegorz M. Boratyn, Jacek M. Zurada, and Andrzej Wrobel. Sparse correlation kernel analysis and evolutionary algorithm-based modeling of the sensory activity within the rat’s barrel cortex. Lecture Notes in Computer Science, 2388:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880198.
pdf.

Byun:2002:ASV

[675] Hyeran Byun and Seong-Whan Lee. Applications of support vector machines for pattern recognition: A survey. Lecture Notes in Computer Science, 2388:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2388/23880213.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2388/23880213.pdf.

Kitamoto:2002:TAD

[676] Asanobu Kitamoto. Typhoon analysis and data mining with kernel methods. Lecture Notes in Computer Science, 2388:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2388/23880237.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2388/23880237.
                                        pdf.

Smeraldi:2002:SVF

[677] Fabrizio Smeraldi, Josef Bigun, and Wulfram Gerstner. Support vector features and the role of dimensionality in face authentication. Lecture Notes in Computer Science, 2388:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880249.
                                        pdf.

Ma:2002:FDB

[678] Yong Ma and Xiaoqing Ding. Face detection based on cost-sensitive support vector machines. Lecture Notes in Computer Science, 2388:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880260.
                                        pdf.

Kang:2002:RTP

[679] Seonghoon Kang, Hyeran Byun, and Seong-Whan Lee. Real-time pedestrian detection using support vector machines. Lecture Notes in Computer Science, 2388:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880268.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2388/23880268.
pdf.

Chakrabartty:2002:FDK

[680] Shantanu Chakrabartty and Gert Cauwenberghs. Forward decoding kernel machines: A hybrid HMM/SVM approach to sequence recognition. Lecture Notes in Computer Science, 2388: 278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880278.
pdf.

Kim:2002:CTB

[681] Kwang In Kim, Keechul Jung, and Jin Hyung Kim. Color texture-based object detection: An application to license plate localization. Lecture Notes in Computer Science, 2388:293– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880293.
pdf.

Ruping:2002:SVM

[682] Stefan Rüping. Support vector machines in relational databases. Lecture Notes in Computer Science, 2388:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880310.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880310.
                                        pdf.

Li:2002:MCS

[683] Zeyu Li, Shiwei Tang, and Shuicheng Yan. Multi-class SVM classifier based on pairwise coupling. Lecture Notes in Computer Science, 2388:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880321.
                                        pdf.

Huang:2002:FRU

[684] Jennifer Huang, Volker Blanz, and Bernd Heisele. Face recognition using component-based SVM classification and morphable models. Lecture Notes in Computer Science, 2388:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880334.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880334.
                                        pdf.

Li:2002:NCR

[685] Jianmin Li, Bo Zhang, and Fuzong Lin. A new cache replacement algorithm in SMO. Lecture Notes in Computer Science, 2388:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880342.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2388/23880342.
pdf.

Ayat:2002:OSK

[686] Nedjem-Eddine Ayat, Mohamed Cheriet, and Ching Y. Suen. Optimization of the SVM kernels using an empirical error minimization scheme. Lecture Notes in Computer Science, 2388:354– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880354.
pdf.

Xi:2002:FDB

[687] Dihua Xi and Seong-Whan Lee. Face detection based on support vector machines. Lecture Notes in Computer Science, 2388:370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2388/23880370.
pdf.

Johansson:2002:DWC

[688] Björn Johansson and Fredrik Kahl. Detecting windows in city scenes. Lecture Notes in Computer Science, 2388:388– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2388/23880388.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2388/23880388.
                                        pdf.

Kim:2002:SVM

[689] Hyun-Chul Kim, Shaoning Pang, Hong-Mo Je, Daijin Kim, and Sung-Yang Bang. Support vector machine ensemble with bagging. Lecture Notes in Computer Science, 2388:397– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880397.
                                        pdf.

dosSantos:2002:CSP

[690] Eulanda Miranda dos Santos and Herman Martins Gomes. A comparative study of polynomial kernel SVM applied to appearance-based object recognition. Lecture Notes in Computer Science, 2388:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2388/23880408.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2388/23880408.
                                        pdf.

Anonymous:2002:AIr

[691] Anonymous. Author index. Lecture Notes in Computer Science, 2388:419– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2388/2388ind.pdf.

Sproat:2002:WTS

[692] Richard Sproat. WordsEye: A textto-scene conversion system. Lecture Notes in Computer Science, 2389:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890001.
pdf.

Mitkov:2002:AAR

[693] Ruslan Mitkov. Automatic anaphora resolution: Limits, impediments, and ways forward. Lecture Notes in Computer Science, 2389:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890003.
pdf.

Cucchiarelli:2002:FBW

[694] Alessandro Cucchiarelli and Paola Velardi. Feature-based WSD: Why we are at a dead-end. Lecture Notes in Computer Science, 2389:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890005.
pdf.

Suarez:2002:IFS

[695] Armando Suárez and Manuel Palomar. Improving feature selection for maximum entropy-based word sense disambiguation. Lecture Notes in Computer Science, 2389:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890015.
pdf.

Bolshakov:2002:HBR

[696] Igor A. Bolshakov and Alexander Gelbukh. Heuristics-based replenishment of collocation databases. Lecture Notes in Computer Science, 2389:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890025.
pdf.

Constant:2002:ALP

[697] Matthieu Constant. On the analysis of locative phrases with graphs and lexicon-grammar: The classifier/proper noun pairing. Lecture Notes in Computer Science, 2389:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890033.
pdf.

Oliveira:2002:RPO

[698] Miguel Oliveira. The role of pause occurrence and pause duration in the signaling of narrative structure. Lecture Notes in Computer Science, 2389:43– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890043.
pdf.

Martinez-Barco:2002:GBS

[699] Patricio Martínez-Barco, Estela Saquete, and Rafael Muñoz. A grammarbased system to solve temporal expressions in Spanish texts. Lecture Notes in Computer Science, 2389:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890053.
pdf.

Munoz:2002:SIA

[700] R. Muñoz, M. Saiz-Noeda, and A. Montoyo. Semantic information in anaphora resolution. Lecture Notes in Computer Science, 2389:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890063.
pdf.

Chung:2002:TTI

[701] Gohsran Chung. A type of transitive inalienable possession construction in korean. Lecture Notes in Computer Science, 2389:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890071.
                                        pdf.

Bogacki:2002:TNP

[702] Krzysztof Bogacki. The treatment of numbers in polish by graphs. Lecture Notes in Computer Science, 2389:75– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890075.
                                        pdf.

Coheur:2002:SST

[703] Luísa Coheur and Nuno Mamede. From syntax to semantics: Taking advantages of 5P. Lecture Notes in Computer Science, 2389:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890079.
                                        pdf.

Cori:2002:PD

[704] Marcel Cori. Processing discontinuity. Lecture Notes in Computer Science, 2389:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2389/23890083.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2389/23890083.
pdf.

SantAnna:2002:RDA

[705] Victor Sant’Anna and Vera L. S. de Lima. Resolution of demonstrative anaphoric references in Portuguese written texts. Lecture Notes in Computer Science, 2389:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890087.
pdf.

Caseiro:2002:LVC

[706] Diamantino Caseiro and Isabel Trancoso. Large vocabulary continuous speech recognition using weighted finite-state transducers. Lecture Notes in Computer Science, 2389:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890091.
pdf.

Souto:2002:BLM

[707] Nuno Souto, Hugo Meinedo, and João P. Neto. Building language models for continuous speech recognition systems. Lecture Notes in Computer Science, 2389:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890101.
pdf.

Ellouze:2002:RIE

[708] Mariem Ellouze and Abdelmajid Ben Hamadou. Relevant information extraction driven with rhetorical schemas to summarize scientific papers. Lecture Notes in Computer Science, 2389: 111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890111.
pdf.

Giorgetti:2002:MAS

[709] Daniela Giorgetti, Irina Prodanof, and Fabrizio Sebastiani. Mapping an automated survey coding task into a probabilistic text categorization framework. Lecture Notes in Computer Science, 2389:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890115.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2389/23890115.
pdf.

Koehn:2002:CMM

[710] Philipp Koehn. Combining multiclass maximum entropy text classifiers with neural network voting. Lecture Notes in Computer Science, 2389:125– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890125.
pdf.

Baptista:2002:CTA

[711] Jorge Baptista and Dolors Català Guitart. Compound temporal adverbs in Portuguese and in Spanish. Lecture Notes in Computer Science, 2389:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890133.
pdf.

Grass:2002:DMD

[712] Thierry Grass, Denis Maurel, and Odile Piton. Description of a multilingual database of proper names. Lecture Notes in Computer Science, 2389: 137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890137.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2389/23890137.
pdf.

Freitas:2002:PSI

[713] Diamantino Freitas, António Moura, Daniela Braga, Helder Ferreira, João Paulo Teixeira, Maria João Barros, Paulo Gouveia, and Vagner Latsch. A project of speech input and output in an E-commerce application. Lecture Notes in Computer Science, 2389:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890141.
pdf.

Bucher:2002:ARF

[714] Kerstin Bücher, Michael Knorr, and Bernd Ludwig. Ambiguity reports for flexible dialog management. Lecture Notes in Computer Science, 2389:151– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890151.
pdf.

Abreu:2002:NLD

[715] Salvador Abreu, Paulo Quaresma, Luis Quintano, and Irene Rodrigues. A natural language dialogue manager for accessing databases. Lecture Notes in Computer Science, 2389:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2389/23890161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890161.
pdf.

Silveira:2002:HCH

[716] Milene Silveira, Maria Carmelita Dias, and Violeta Quental. Helping the composition of help contents. Lecture Notes in Computer Science, 2389:171– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890171.
pdf.

Cardoso:2002:MTB

[717] Pedro Cardoso, Luis Flores, Thibault Langlois, and João Neto. Meteo: A telephone-based Portuguese conversation system in weather domain. Lecture Notes in Computer Science, 2389: 175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890175.
pdf.

Hagen:2002:CRM

[718] Astrid Hagen. Concluding remarks on multi-band and multi-stream research for noise-robust ASR. Lecture Notes in Computer Science, 2389:179– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890179.
                                        pdf.

Dias-da-Silva:2002:GDB

[719] Bento C. Dias da Silva, Mirna F. de Oliveira, and Helio R. de Moraes. Groundwork for the development of the Brazilian Portuguese Wordnet. Lecture Notes in Computer Science, 2389:189– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890189.
                                        pdf.

Hagege:2002:LPB

[720] Caroline Hagège, Ágnes Sándor, and Anne Schiller. Linguistic processing of biomedical texts. Lecture Notes in Computer Science, 2389:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890197.
                                        pdf.

Santos:2002:DSD

[721] Diana Santos. DISPARA, a system for distributing parallel corpora on the Web. Lecture Notes in Computer Science, 2389:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2389/23890209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890209.
pdf.

Paulo:2002:UMS

[722] Joana Lúcio Paulo, Margarita Correia, Nuno J. Mamede, and Caroline Hagège. Using morphological, syntactical, and statistical information for automatic term acquisition. Lecture Notes in Computer Science, 2389:219– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890219.
pdf.

Carvalho:2002:CLU

[723] Paula Carvalho, Cristina Mota, and Elisabete Ranchhod. Complex lexical units and automata. Lecture Notes in Computer Science, 2389:229– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890229.
pdf.

Makagonov:2002:CEF

[724] Pavel Makagonov and Mikhail Alexandrov. Constructing empirical formulas for testing word similarity by the inductive method of model self-organization. Lecture Notes in Computer Science, 2389:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890239.
                                        pdf.

Vieira:2002:MCA

[725] Renata Vieira, Susanne Salmon-Alt, and Emmanuel Schang. Multilingual corpora annotation for processing definite descriptions. Lecture Notes in Computer Science, 2389:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890249.
                                        pdf.

Gelbukh:2002:RDM

[726] Alexander Gelbukh, Grigori Sidorov, and Adolfo Guzmán-Arenas. Relational data model in document hierarchical indexing. Lecture Notes in Computer Science, 2389:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2389/23890259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2389/23890259.
                                        pdf.

Pardo:2002:DRA

[727] Thiago Alexandre Salgueiro Pardo and Lucia Helena Machado Rino. DM-Summ: Review and assessment. Lecture Notes in Computer Science, 2389: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2389/23890263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2389/23890263.
pdf.

Anonymous:2002:AIs

[728] Anonymous. Author index. Lecture Notes in Computer Science, 2389:275– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2389/2389indx.pdf.

Shankar:2002:LEP

[729] Natarajan Shankar. Little engines of proof. Lecture Notes in Computer Science, 2391:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910001.
pdf.

Legeard:2002:ABT

[730] Bruno Legeard, Fabien Peureux, and Mark Utting. Automated boundary testing from Z and B. Lecture Notes in Computer Science, 2391:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910021.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2391/23910021.
                                        pdf.

Ratsaby:2002:ICA

[731] Gil Ratsaby, Baruch Sterin, and Shmuel Ur. Improvements in coverability analysis. Lecture Notes in Computer Science, 2391:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910041.
                                        pdf.

Burguillo-Rial:2002:HDT

[732] Juan C. Burguillo-Rial, Manuel J. Fernández-Iglesias, Francisco J. González-Castaño, and Martín Llamas-Nistal. Heuristic-driven test case selection from formal specifications. A case study. Lecture Notes in Computer Science, 2391:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2391/23910057.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2391/23910057.pdf.

Bourdonov:2002:UTS

[733] Igor B. Bourdonov, Alexander S. Kossatchev, Victor V. Kuliamin, and Alexander K. Petrenko. UniTesK test suite architecture. Lecture Notes in Computer Science, 2391:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2391/23910077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910077.
pdf.

vonOheimb:2002:HLN

[734] David von Oheimb and Tobias Nipkow. Hoare logic for NanoJava: Auxiliary variables, side effects, and virtual methods revisited. Lecture Notes in Computer Science, 2391:89– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910089.
pdf.

Bicarregui:2002:DR

[735] Juan C. Bicarregui. Do not read this. Lecture Notes in Computer Science, 2391:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2391/23910106.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2391/23910106.
pdf.

Jorgensen:2002:SMB

[736] Niels Jørgensen. Safeness of makebased incremental recompilation. Lecture Notes in Computer Science, 2391: 126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910126.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2391/23910126.
pdf.

Barner:2002:AAD

[737] Sharon Barner, Shoham Ben-David, Anna Gringauze, Baruch Sterin, and Yaron Wolfsthal. An algorithmic approach to design exploration. Lecture Notes in Computer Science, 2391:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910146.
pdf.

Mota:2002:MAC

[738] Alexandre Mota, Paulo Borba, and Augusto Sampaio. Mechanical abstraction of CSP_{Z} processes. Lecture Notes in Computer Science, 2391:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910163.
pdf.

Arts:2002:VEC

[739] Thomas Arts, Clara Benac Earle, and John Derrick. Verifying Erlang code: A resource locker case-study. Lecture Notes in Computer Science, 2391:184– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910184.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2391/23910184.
pdf.

Huber:2002:TIM

[740] Michael Huber and Steve King. Towards an integrated model checker for railway signalling data. Lecture Notes in Computer Science, 2391:204– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910204.
pdf.

Hall:2002:CCI

[741] Anthony Hall. Correctness by construction: Integrating formality into a commercial development process. Lecture Notes in Computer Science, 2391:224– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910224.
pdf.

Marinov:2002:VVF

[742] Darko Marinov and Sarfraz Khurshid. VAlloy — virtual functions meet a relational language. Lecture Notes in Computer Science, 2391:234– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910234.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2391/23910234.
                                        pdf.

Rusu:2002:VUT

[743] Vlad Rusu. Verification using test generation techniques. Lecture Notes in Computer Science, 2391:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910252.
                                        pdf.

Catano:2002:FSS

[744] Néstor Cataño and Marieke Huisman. Formal specification and static checking of Gemplus’ electronic purse using ESC/Java. Lecture Notes in Computer Science, 2391:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910272.
                                        pdf.

Casset:2002:DEV

[745] Ludovic Casset. Development of an embedded verifier for Java card byte code using formal methods. Lecture Notes in Computer Science, 2391:290– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910290.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2391/23910290.
pdf.

Backes:2002:DCS

[746] Michael Backes, Christian Jacobi, and Birgit Pfitzmann. Deriving cryptographically sound implementations using composition and formally verified bisimulation. Lecture Notes in Computer Science, 2391:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910310.
pdf.

Pahl:2002:IAD

[747] Claus Pahl. Interference analysis for dependable systems using refinement and abstraction. Lecture Notes in Computer Science, 2391:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910330.
pdf.

Henderson:2002:FCV

[748] N. Henderson and S. E. Paynter. The formal classification and verification of simpson’s 4-slot asynchronous communication mechanism. Lecture Notes in Computer Science, 2391:350– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2391/23910350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910350.
                                        pdf.

Fidge:2002:TAA

[749] C. J. Fidge. Timing analysis of assembler code control-flow paths. Lecture Notes in Computer Science, 2391:370– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910370.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910370.
                                        pdf.

Cengarle:2002:TOR

[750] María Victoria Cengarle and Alexander Knapp. Towards OCL/RT. Lecture Notes in Computer Science, 2391: 390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910390.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910390.
                                        pdf.

Garavel:2002:CFV

[751] Hubert Garavel and Holger Hermanns. On combining functional verification and performance evaluation using CADP. Lecture Notes in Computer Science, 2391:410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910410.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2391/23910410.
pdf.

Basin:2002:NSC

[752] David Basin. The next 700 synthesis calculi. Lecture Notes in Computer Science, 2391:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910430.
pdf.

Whalen:2002:SCC

[753] Michael Whalen, Johann Schumann, and Bernd Fischer. Synthesizing certified code. Lecture Notes in Computer Science, 2391:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910431.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910431.
pdf.

Sampaio:2002:RC

[754] Augusto Sampaio, Jim Woodcock, and Ana Cavalcanti. Refinement in Circus. Lecture Notes in Computer Science, 2391:451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2391/23910451.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2391/23910451.
pdf.

Cavalcanti:2002:FSD

[755] Ana Cavalcanti and David A. Naumann. Forward simulation for data refinement of classes. Lecture Notes in Computer Science, 2391:471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910471.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910471.
                                        pdf.

Wildman:2002:FBP

[756] Luke Wildman. A formal basis for a program compilation proof tool. Lecture Notes in Computer Science, 2391: 491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910491.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910491.
                                        pdf.

Firley:2002:PDA

[757] Thomas Firley and Ursula Goltz. Property dependent abstraction of control structure for software verification. Lecture Notes in Computer Science, 2391: 511–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910511.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910511.
                                        pdf.

Ioustinova:2002:COS

[758] Natalia Ioustinova, Natalia Sidorova, and Martin Steffen. Closing open SDL-systems for model checking with DTSpin. Lecture Notes in Computer Science, 2391:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2391/23910531.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2391/23910531.pdf.

Kristensen:2002:GSL

[759] Lars Michael Kristensen and Thomas Mailund. A generalised sweep-line method for safety properties. Lecture Notes in Computer Science, 2391:549– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910549.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910549.
pdf.

Treharne:2002:SUD

[760] Helen Treharne. Supplementing a UML development process with B. Lecture Notes in Computer Science, 2391:568– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2391/23910568.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2391/23910568.
pdf.

Dong:2002:SWE

[761] Jin Song Dong, Jing Sun, and Hai Wang. Semantic Web for extending and linking formalisms. Lecture Notes in Computer Science, 2391:587– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910587.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910587.
                                        pdf.

Umedu:2002:LDW

[762] Takaaki Umedu, Yoshiki Terashima, Keiichi Yasumoto, Akio Nakata, Teruo Higashino, and Kenichi Taniguchi. A language for describing wireless mobile applications with dynamic establishment of multi-way synchronization channels. Lecture Notes in Computer Science, 2391:607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2391/23910607.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2391/23910607.
                                        pdf.

Anonymous:2002:AIt

[763] Anonymous. Author index. Lecture Notes in Computer Science, 2391:625– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2391/2391auth.pdf.

Horrocks:2002:RED

[764] Ian Horrocks. Reasoning with expressive description logics: Theory and practice. Lecture Notes in Computer Science, 2392:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920001.
pdf.

Pan:2002:BBD

[765] Guoqiang Pan, Ulrike Sattler, and Moshe Y. Vardi. BDD-based decision procedures for K. Lecture Notes in Computer Science, 2392:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920016.
pdf.

Bernard:2002:TLP

[766] Andrew Bernard and Peter Lee. Temporal logic for proof-carrying code. Lecture Notes in Computer Science, 2392: 31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920031.
pdf.

Schneck:2002:GAM

[767] Robert R. Schneck and George C. Necula. A gradual approach to a more trustworthy, yet scalable, proof-carrying code. Lecture Notes in Computer Science, 2392:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920047.
                                        pdf.

Strecker:2002:FVJ

[768] Martin Strecker. Formal verification of a Java compiler in isabelle. Lecture Notes in Computer Science, 2392:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920063.
                                        pdf.

Egly:2002:ELL

[769] Uwe Egly. Embedding Lax logic into intuitionistic logic. Lecture Notes in Computer Science, 2392:78– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920078.
                                        pdf.

Larchey-Wendling:2002:CPS

[770] Dominique Larchey-Wendling. Combining proof-search and countermodel construction for deciding Gödeldummett logic. Lecture Notes in Computer Science, 2392:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920094.
pdf.

Galmiche:2002:CBP

[771] Didier Galmiche and Daniel Méry. Connection-based proof search in propositional BI logic. Lecture Notes in Computer Science, 2392:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920111.
pdf.

Moller:2002:DLS

[772] Jesper B. Møller. DDDLIB: A library for solving quantified difference inequalities. Lecture Notes in Computer Science, 2392:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920129.
pdf.

Hurd:2002:LSI

[773] Joe Hurd. An LCF-style interface between HOL and first-order logic. Lecture Notes in Computer Science, 2392: 134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920134.
pdf.

Zimmer:2002:SDM

[774] Jürgen Zimmer and Michael Kohlhase. System description: The MathWeb software bus for distributed mathematical reasoning. Lecture Notes in Computer Science, 2392:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920139.
pdf.

Siekmann:2002:PD

[775] Jörg Siekmann, Christoph Benzmüller, Vladimir Brezhnev, Lassaad Cheikhrouhou, Armin Fiedler, Andreas Franke, Helmut Horacek, Michael Kohlhase, Andreas Meier, Erica Melis, Markus Moschner, Immanuel Normann, Martin Pollet, Volker Sorge, Carsten Ullrich, Claus-Peter Wirth, and Jürgen Zimmer. Proof development with ΩMEGA. Lecture Notes in Computer Science, 2392:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2392/23920144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920144.
pdf.

Jamnik:2002:LMS

[776] Mateja Jamnik, Manfred Kerber, and Martin Pollet. LearnΩ matic: System description. Lecture Notes in Computer Science, 2392:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920150.
pdf.

Areces:2002:HDR

[777] Carlos Areces and Juan Heguiabehere. HyLoRes 1.0 : Direct resolution for hybrid logics. Lecture Notes in Computer Science, 2392:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920156.
pdf.

Goldberg:2002:TSC

[778] Eugene Goldberg. Testing satisfiability of CNF formulas by computing a stable set of points. Lecture Notes in Computer Science, 2392:161– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2392/23920161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920161.
                                        pdf.

delaTour:2002:NSH

[779] Thierry Boy de la Tour. A note on symmetry heuristics in SEM. Lecture Notes in Computer Science, 2392:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920181.
                                        pdf.

Audemard:2002:SBA

[780] Gilles Audemard, Piergiorgio Bertoli, Alessandro Cimatti, Artur Korniłowicz, and Roberto Sebastiani. A SAT based approach for solving formulas over Boolean and linear mathematical propositions. Lecture Notes in Computer Science, 2392:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920195.
                                        pdf.

Ahrendt:2002:DSE

[781] Wolfgang Ahrendt. Deductive search for errors in free data type specifications using model generation. Lecture Notes in Computer Science, 2392:211– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2392/23920211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920211.
pdf.

Audemard:2002:RSF

[782] Gilles Audemard and Belaid Benhamou. Reasoning by symmetry and function ordering in finite model generation. Lecture Notes in Computer Science, 2392:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920226.
pdf.

Gramlich:2002:AAH

[783] Bernhard Gramlich and Reinhard Pichler. Algorithmic aspects of Herbrand models represented by ground atoms with ground equations. Lecture Notes in Computer Science, 2392:241– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920241.
pdf.

Georgieva:2002:NCC

[784] Lilia Georgieva, Ullrich Hustadt, and Renate A. Schmidt. A new clausal class decidable by hyperresolution. Lecture Notes in Computer Science, 2392:260– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920260.
                                        pdf.

Weidenbach:2002:V

[785] Christoph Weidenbach, Uwe Brahm, Thomas Hillenbrand, Enno Keen, Christian Theobald, and Dalibor Topić. S_{P ASS} Version 2.0. Lecture Notes in Computer Science, 2392:275– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920275.
                                        pdf.

Schulz:2002:SDG

[786] Stephan Schulz and Geoff Sutcliffe. System description: GrAnDe 1.0. Lecture Notes in Computer Science, 2392: 280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920280.
                                        pdf.

Colton:2002:HPT

[787] Simon Colton. The HR program for theorem generation. Lecture Notes in Computer Science, 2392:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2392/23920285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920285.
pdf.

Whalen:2002:ACC

[788] Michael Whalen, Johann Schumann, and Bernd Fischer. AutoBayes/ CC — combining program synthesis with automatic code certification — system description. Lecture Notes in Computer Science, 2392:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920290.
pdf.

Zhang:2002:QEB

[789] Lintao Zhang and Sharad Malik. The quest for efficient Boolean satisfiability solvers. Lecture Notes in Computer Science, 2392:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920295.
pdf.

Borralleras:2002:RPO

[790] Cristina Borralleras, Salvador Lucas, and Albert Rubio. Recursive path orderings can be context-sensitive. Lecture Notes in Computer Science, 2392: 314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920314.
                                        pdf.

Ganzinger:2002:SL

[791] Harald Ganzinger. Shostak light. Lecture Notes in Computer Science, 2392: 332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920332.
                                        pdf.

Ford:2002:FVC

[792] Jonathan Ford and Natarajan Shankar. Formal verification of a combination decision procedure. Lecture Notes in Computer Science, 2392:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920347.
                                        pdf.

Zarba:2002:CMI

[793] Calogero G. Zarba. Combining multisets with integers. Lecture Notes in Computer Science, 2392:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920363.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2392/23920363.
pdf.

Paulson:2002:RTS

[794] Lawrence C. Paulson. The reflection theorem: A study in metatheoretic reasoning. Lecture Notes in Computer Science, 2392:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920377.
pdf.

Stump:2002:FPC

[795] Aaron Stump and David L. Dill. Faster proof checking in the edinburgh logical framework. Lecture Notes in Computer Science, 2392:392– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920392.
pdf.

Brown:2002:SSV

[796] Chad E. Brown. Solving for set variables in higher-order theorem proving. Lecture Notes in Computer Science, 2392:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2392/23920408.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2392/23920408.
pdf.

Kupferman:2002:CGC

[797] Orna Kupferman, Ulrike Sattler, and Moshe Y. Vardi. The complexity of the graded μ-calculus. Lecture Notes in Computer Science, 2392:423– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920423.
                                        pdf.

deMoura:2002:LTP

[798] Leonardo de Moura, Harald Rueß, and Maria Sorea. Lazy theorem proving for bounded model checking over infinite domains. Lecture Notes in Computer Science, 2392:438– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920438.
                                        pdf.

Bofill:2002:WFS

[799] Miquel Bofill and Albert Rubio. Wellfoundedness is sufficient for completeness of ordered paramodulation. Lecture Notes in Computer Science, 2392: 456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920456.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920456.
                                        pdf.

Lynch:2002:BSM

[800] Christopher Lynch and Barbara Morawska. Basic syntactic mutation. Lecture Notes in Computer Science, 2392:471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2392/23920471.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2392/23920471.
pdf.

Hillenbrand:2002:NWA

[801] Thomas Hillenbrand and Bernd Löchner. The next W ALD-MEISTER loop. Lecture Notes in Computer Science, 2392:486–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920486.
pdf.

Andreoli:2002:FPN

[802] Jean Marc Andreoli. Focussing proof-net construction as a middleware paradigm. Lecture Notes in Computer Science, 2392:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2392/23920501.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2392/23920501.
pdf.

Baaz:2002:PARb

[803] Matthias Baaz. Proof analysis by resolution. Lecture Notes in Computer Science, 2392:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2392/23920517.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2392/23920517.
                                        pdf.

Anonymous:2002:AIu

[804] Anonymous. Author index. Lecture Notes in Computer Science, 2392:533– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2392/2392auth.pdf.

Carpineto:2002:CSM

[805] Claudio Carpineto. Conceptual structures in modern information retrieval. Lecture Notes in Computer Science, 2393:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2393/23930001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930001.
                                        pdf.

Stumme:2002:FCA

[806] Gerd Stumme. Formal concept analysis on its way from mathematics to computer science. Lecture Notes in Computer Science, 2393:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930002.
pdf.

Sarbo:2002:LCM

[807] Janos Sarbo and József Farkas. A linearly complex model for knowledge representation. Lecture Notes in Computer Science, 2393:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930020.
pdf.

Pollandt:2002:RGS

[808] Silke Pollandt. Relation graphs: A structure for representing relations in contextual logic of relations. Lecture Notes in Computer Science, 2393:34– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930034.
pdf.

Malik:2002:ITA

[809] Grit Malik. Information transfer across simple concept graphs. Lecture Notes in Computer Science, 2393:48– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2393/23930048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930048.
                                        pdf.

Hereth:2002:RSD

[810] Joachim Hereth. Relational scaling and databases. Lecture Notes in Computer Science, 2393:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930062.
                                        pdf.

Martin:2002:KRC

[811] Philippe Martin. Knowledge representation in CGLF, CGIF, KIF, frame-CG and formalized-english. Lecture Notes in Computer Science, 2393:77– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930077.
                                        pdf.

Zhong:2002:CGM

[812] Jiwei Zhong, Haiping Zhu, Jianming Li, and Yong Yu. Conceptual graph matching for semantic search. Lecture Notes in Computer Science, 2393:92– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930092.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2393/23930092.
pdf.

Gatzemeier:2002:TSM

[813] Felix H. Gatzemeier and Oliver Meyer. Text schema mining using graphs and formal concept analysis. Lecture Notes in Computer Science, 2393:107– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930107.
pdf.

Montes-y-Gomez:2002:TMD

[814] Manuel Montes y Gómez, Alexander Gelbukh, and Aurelio López-López. Text mining at detail level using conceptual graphs. Lecture Notes in Computer Science, 2393:122– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930122.
pdf.

Boytcheva:2002:PNN

[815] Svetla Boytcheva, Albena Strupchanska, and Galia Angelova. Processing negation in NL interfaces to knowledge bases. Lecture Notes in Computer Science, 2393:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930137.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930137.
                                        pdf.

Reidsma:2002:AGE

[816] Dennis Reidsma. Automatic generation and evaluation of sentence graphs out of word graphs. Lecture Notes in Computer Science, 2393:151– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930151.
                                        pdf.

Scharfe:2002:CRN

[817] Henrik Schärfe. CG representations of non-literal expressions. Lecture Notes in Computer Science, 2393:166– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930166.
                                        pdf.

Farkas:2002:POS

[818] József Farkas and Janos Sarbo. A peircean ontology of semantics. Lecture Notes in Computer Science, 2393: 177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930177.
                                        pdf.

Delteil:2002:BCL

[819] Alexandre Delteil, Catherine Faron, and Rose Dieng. Building concept lattices by learning concepts from RDF graphs annotating Web documents. Lecture Notes in Computer Science, 2393:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2393/23930191.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2393/23930191.
pdf.

Gerbe:2002:CFO

[820] Olivier Gerbé and Guy W. Mineau. The CG formalism as an ontolingua for Web-oriented representation languages. Lecture Notes in Computer Science, 2393:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930205.
pdf.

Ribiere:2002:VMC

[821] Myriam Ribière and Rose Dieng-Kuntz. A viewpoint model for cooperative building of an ontology. Lecture Notes in Computer Science, 2393: 220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930220.
pdf.

deMoor:2002:TPW

[822] Aldo de Moor, Mary Keeler, and Gary Richmond. Towards a pragmatic Web. Lecture Notes in Computer Science, 2393:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2393/23930235.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930235.
                                        pdf.

Mineau:2002:FST

[823] Guy W. Mineau. A first step toward the Knowledge Web: Interoperability issues among conceptual graph based software agents Part I. Lecture Notes in Computer Science, 2393:250– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930250.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930250.
                                        pdf.

Dobrev:2002:CAF

[824] Pavlin Dobrev and Kristina Toutanova. CGWorld — architecture and features. Lecture Notes in Computer Science, 2393:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2393/23930261.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930261.
                                        pdf.

Pimentel:2002:ESP

[825] César F. Pimentel and Maria R. Cravo. Extending SNePSwD with permissive belief revision. Lecture Notes in Computer Science, 2393:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930271.
pdf.

Scharfe:2002:TCG

[826] Henrik Schärfe, Ulrik Petersen, and Peter hrstrøm. On teaching conceptual graphs. Lecture Notes in Computer Science, 2393:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930285.
pdf.

Ferre:2002:UAC

[827] Sébastien Ferré and Olivier Ridoux. The use of associative concepts in the incremental building of a logical context. Lecture Notes in Computer Science, 2393:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930299.
pdf.

Obiedkov:2002:MLE

[828] Sergei Obiedkov. Modal logic for evaluating formulas in incomplete contexts. Lecture Notes in Computer Science, 2393:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930314.
                                        pdf.

Dau:2002:EEG

[829] Frithjof Dau. An embedding of existential graphs into concept graphs with negations. Lecture Notes in Computer Science, 2393:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2393/23930326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930326.
                                        pdf.

Wolff:2002:IAT

[830] Karl Erich Wolff. Interpretation of automata in temporal concept analysis. Lecture Notes in Computer Science, 2393:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2393/23930341.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2393/23930341.
                                        pdf.

Priss:2002:AFC

[831] Uta Priss. Associative and formal concepts. Lecture Notes in Computer Science, 2393:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2393/23930354.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2393/23930354.pdf.

Klinger:2002:SGV

[832] Julia Klinger. Semiconcept graphs with variables. Lecture Notes in Computer Science, 2393:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930369.
pdf.

Wille:2002:ECG

[833] Rudolf Wille. Existential concept graphs of power context families. Lecture Notes in Computer Science, 2393: 382–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2393/23930382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2393/23930382.
pdf.

Anonymous:2002:AIv

[834] Anonymous. Author index. Lecture Notes in Computer Science, 2393:397– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2393/2393ind.pdf.

Blanc:2002:SRW

[835] Erika Blanc and Paolo Giudici. Sequence rules for Web clickstream analysis. Lecture Notes in Computer Science, 2394:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2394/23940001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2394/23940001.
pdf.

Hipp:2002:DMA

[836] Jochen Hipp, Ulrich Güntzer, and Gholamreza Nakhaeizadeh. Data mining of association rules and the process of knowledge discovery in databases. Lecture Notes in Computer Science, 2394: 15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2394/23940015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2394/23940015.
pdf.

Perner:2002:IMW

[837] P. Perner and G. Fiss. Intelligent E-marketing with Web mining, personalization, and useradpated interfaces. Lecture Notes in Computer Science, 2394:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2394/23940037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2394/23940037.
pdf.

Althoff:2002:IPE

[838] Klaus-Dieter Althoff, Ulrike Becker-Kornstaedt, Björn Decker, Andreas Klotz, Edda Leopold, Jörg Rech, and Angi Voss. The indiGo project: Enhancement of experience management and process learning with moderated discourses. Lecture Notes in Computer Science, 2394:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2394/23940053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2394/23940053.
pdf.

Glass:2002:GDE

[839] Änne Glass and Thomas Karopka. Genomic data explosion — the challenge for bioinformatics? Lecture Notes in Computer Science, 2394:80– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2394/23940080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2394/23940080.
pdf.

Schmidt:2002:CBR

[840] Rainer Schmidt and Lothar Gierl. Case-based reasoning for prognosis of threatening influenza waves. Lecture Notes in Computer Science, 2394:99– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2394/23940099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2394/23940099.
                                        pdf.

Anonymous:2002:AIw

[841] Anonymous. Author index. Lecture Notes in Computer Science, 2394:109– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2394/2394auth.pdf.

Macrini:2002:SMV

[842] Diego Macrini, Ali Shokoufandeh, Sven Dickinson, Kaleem Siddiqi, and Steven Zucker. Spectral methods for view-based 3-D object recognition using silhouettes. Lecture Notes in Computer Science, 2396:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960001.
                                        pdf.

Dietterich:2002:MLS

[843] Thomas G. Dietterich. Machine learning for sequential data: A review. Lecture Notes in Computer Science, 2396: 15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960015.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2396/23960015.
pdf.

Hancock:2002:GBM

[844] Edwin Hancock and Richard C. Wilson. Graph-based methods for vision: A yorkist manifesto. Lecture Notes in Computer Science, 2396:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960031.
pdf.

Martinez-Hinarejos:2002:RCC

[845] Carlos D. Martínez-Hinarejos, Alfonso Juan, Francisco Casacuberta, and Ramón Mollineda. Reducing the computational cost of computing approximated median strings. Lecture Notes in Computer Science, 2396:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960047.
pdf.

Verdu-Mas:2002:TGM

[846] Jose L. Verdú-Mas, Mikel L. Forcada, Rafael C. Carrasco, and Jorge Calera-Rubio. Tree k-grammar models for natural language modelling and parsing. Lecture Notes in Computer Science, 2396:56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960056.
                                        pdf.

Fernau:2002:ALF

[847] Henning Fernau and Agnes Radl. Algorithms for learning function distinguishable regular languages. Lecture Notes in Computer Science, 2396:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960064.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960064.
                                        pdf.

vanWyk:2002:NBG

[848] B. J. van Wyk and M. A. van Wyk. Non-bayesian graph matching without explicit compatibility calculations. Lecture Notes in Computer Science, 2396: 74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960074.
                                        pdf.

Luo:2002:SFV

[849] Bin Luo, Richard C. Wilson, and Edwin R. Hancock. Spectral feature vectors for graph clustering. Lecture Notes in Computer Science, 2396:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2396/23960083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960083.
pdf.

Fischer:2002:IDG

[850] Stefan Fischer, Kaspar Gilomen, and Horst Bunke. Identification of diatoms by grid graph matching. Lecture Notes in Computer Science, 2396: 94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960094.
pdf.

Robles-Kelly:2002:SED

[851] Antonio Robles-Kelly and Edwin R. Hancock. String edit distance, random walks and graph matching. Lecture Notes in Computer Science, 2396: 104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960104.
pdf.

Torsello:2002:LSV

[852] Andrea Torsello and Edwin R. Hancock. Learning structural variations in shock trees. Lecture Notes in Computer Science, 2396:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960113.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960113.
                                        pdf.

Bunke:2002:CAM

[853] H. Bunke, P. Foggia, C. Guidobaldi, C. Sansone, and M. Vento. A comparison of algorithms for maximum common subgraph on randomly connected graphs. Lecture Notes in Computer Science, 2396:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960123.
                                        pdf.

Kosinov:2002:IMM

[854] Serhiy Kosinov and Terry Caelli. Inexact multisubgraph matching using graph eigenspace and clustering models. Lecture Notes in Computer Science, 2396:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960133.
                                        pdf.

Jiang:2002:OLB

[855] Xiaoyi Jiang and Horst Bunke. Optimal lower bound for generalized median problems in metric space. Lecture Notes in Computer Science, 2396: 143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2396/23960143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960143.
pdf.

Amin:2002:SDR

[856] Adnan Amin. Structural description to recognising Arabic characters using decision tree learning techniques. Lecture Notes in Computer Science, 2396: 152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960152.
pdf.

Duong:2002:FAP

[857] Jean Duong, Myrian Côté, and Hubert Emptoz. Feature approach for printed document image analysis. Lecture Notes in Computer Science, 2396: 159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960159.
pdf.

Wenyin:2002:EDG

[858] Liu Wenyin. Example-driven graphics recognition. Lecture Notes in Computer Science, 2396:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960168.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960168.
                                        pdf.

Gimelfarb:2002:ETR

[859] Georgy Gimel’farb. Estimation of texels for regular mosaics using modelbased interaction maps. Lecture Notes in Computer Science, 2396:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960177.
                                        pdf.

Gregory:2002:UGS

[860] Lee Gregory and Josef Kittler. Using graph search techniques for contextual colour retrieval. Lecture Notes in Computer Science, 2396:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960186.
                                        pdf.

Tassone:2002:CST

[861] Ezra Tassone, Geoff West, and Svetha Venkatesh. Comparing shape and temporal PDMs. Lecture Notes in Computer Science, 2396:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960195.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2396/23960195.
pdf.

Al-Shaher:2002:LSR

[862] Abdullah A. Al-Shaher and Edwin R. Hancock. Linear shape recognition with mixtures of point distribution models. Lecture Notes in Computer Science, 2396:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960205.
pdf.

Sartori:2002:CWE

[863] Fabio Sartori and Edwin R. Hancock. Curvature weighted evidence combination for shape-from-shading. Lecture Notes in Computer Science, 2396:216– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960216.
pdf.

Michaelsen:2002:PDP

[864] E. Michaelsen and U. Stilla. Probabilistic decisions in production nets: An example from vehicle recognition. Lecture Notes in Computer Science, 2396: 225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960225.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2396/23960225.
                                        pdf.

Langs:2002:HTE

[865] Georg Langs, Horst Bischof, and Walter G. Kropatsch. Hierarchical top down enhancement of robust PCA. Lecture Notes in Computer Science, 2396:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2396/23960234.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960234.
                                        pdf.

Lazarescu:2002:AML

[866] Mihai Lazarescu, Andrew Turpin, and Svetha Venkatesh. An application of machine learning techniques for the classification of glaucomatous progression. Lecture Notes in Computer Science, 2396:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960243.
                                        pdf.

Serratosa:2002:EJP

[867] Francesc Serratosa, René Alquézar, and Alberto Sanfeliu. Estimating the joint probability distribution of random vertices and arcs by means of second-order random graphs. Lecture Notes in Computer Science, 2396:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2396/23960252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960252.
pdf.

vanWyk:2002:SPG

[868] B. J. van Wyk, M. A. van Wyk, and H. E. Hanrahan. Successive projection graph matching. Lecture Notes in Computer Science, 2396:263– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960263.
pdf.

Popel:2002:CGM

[869] Denis V. Popel. Compact graph model of handwritten images: Integration into authentification and recognition. Lecture Notes in Computer Science, 2396:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2396/23960272.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2396/23960272.
pdf.

Delalandre:2002:SSA

[870] Mathieu Delalandre, Pierre Héroux, Sébastien Adam, Eric Trupin, and Jean-Marc Ogier. A statistical and structural approach for symbol recognition, using XML modelling. Lecture Notes in Computer Science, 2396:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960281.
                                        pdf.

Hlaoui:2002:NAG

[871] Adel Hlaoui and Shengrui Wang. A new algorithm for graph matching with application to contentbased image retrieval. Lecture Notes in Computer Science, 2396:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960291.
                                        pdf.

Azuela:2002:ECD

[872] Juan Humberto Sossa Azuela, Francisco Cuevas de la Rosa, and Héctor Benitez. Efficient computation of 3-D moments in terms of an object’s partition. Lecture Notes in Computer Science, 2396:301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960301.
                                        pdf.

Palenichka:2002:VAO

[873] Roman M. Palenichka. A visual attention operator based on morphological models of images and maximum likelihood decision. Lecture Notes in Computer Science, 2396:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960310.
pdf.

Ulusoy:2002:DUF

[874] Ilkay Ulusoy, Edwin R. Hancock, and Ugur Halici. Disparity using feature points in multi scale. Lecture Notes in Computer Science, 2396:320– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960320.
pdf.

Cheoi:2002:DPI

[875] Kyungjoo Cheoi and Yillbyung Lee. Detecting perceptually important regions in an image based on human visual attention characteristic. Lecture Notes in Computer Science, 2396:329– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960329.
pdf.

Alam:2002:DSL

[876] Hassan Alam, Ahmad Fuad Rezaur Rahman, Timotius Tjahjadi, Hua Cheng, Paul Llid Aman Kumar, Rachmat Hartono, Yulia Tarnikova, and Che Wilcox. Development of spoken language user interfaces: A tool kit approach. Lecture Notes in Computer Science, 2396:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960339.
                                        pdf.

Wu:2002:DIW

[877] Changhua Wu and Gady Agam. Document image de-warping for text/ graphics recognition. Lecture Notes in Computer Science, 2396:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960348.
                                        pdf.

Lehal:2002:COS

[878] G. S. Lehal and Chandan Singh. A complete OCR system for Gurmukhi script. Lecture Notes in Computer Science, 2396:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2396/23960358.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2396/23960358.pdf.

Grau:2002:TNA

[879] Antoni Grau, Joan Climent, Francesc Serratosa, and Alberto Sanfeliu. Texprint: A new algorithm to discriminate textures structurally. Lecture Notes in Computer Science, 2396:368– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960368.
pdf.

Droettboom:2002:OMI

[880] Michael Droettboom, Ichiro Fujinaga, and Karl MacMillan. Optical music interpretation. Lecture Notes in Computer Science, 2396:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960378.
pdf.

Azuela:2002:SCC

[881] Juan Humberto Sossa Azuela, Aurelio Velázquez, and Serguei Levachkine. On the segmentation of color cartographic images. Lecture Notes in Computer Science, 2396:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960387.
pdf.

Aladjem:2002:PPF

[882] Mayer Aladjem. Projection pursuit fitting Gaussian mixture models. Lecture Notes in Computer Science, 2396:396– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960396.
                                        pdf.

Kato:2002:AGA

[883] Tsuyoshi Kato, Shinichiro Omachi, and Hirotomo Aso. Asymmetric Gaussian and its application to pattern recognition. Lecture Notes in Computer Science, 2396:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960405.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960405.
                                        pdf.

Sadeghi:2002:MPV

[884] Mohammad Sadeghi and Josef Kittler. Modified predictive validation test for Gaussian mixture modelling. Lecture Notes in Computer Science, 2396:414– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960414.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960414.
                                        pdf.

Fumera:2002:PAC

[885] Giorgio Fumera and Fabio Roli. Performance analysis and comparison of linear combiners for classifier fusion. Lecture Notes in Computer Science, 2396: 424–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960424.
pdf.

Somorjai:2002:CTC

[886] Ray Somorjai, Arunas Janeliunas, Richard Baumgartner, and Sarunas Raudys. Comparison of two classification methodologies on a realworld biomedical problem. Lecture Notes in Computer Science, 2396:433– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960433.
pdf.

Fred:2002:EAC

[887] Ana Fred and Anil K. Jain. Evidence accumulation clustering based on the K-means algorithm. Lecture Notes in Computer Science, 2396:442– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960442.
pdf.

Webb:2002:KAM

[888] Andrew Webb. A kernel approach to metric multidimensional scaling. Lecture Notes in Computer Science, 2396: 452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960452.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960452.
                                        pdf.

Paclik:2002:FSM

[889] Pavel Paclík, Robert P. W. Duin, Geert M. P. van Kempen, and Reinhard Kohlus. On feature selection with measurement cost and grouped features. Lecture Notes in Computer Science, 2396:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960461.
                                        pdf.

Abe:2002:CIF

[890] Naoto Abe, Mineichi Kudo, and Masaru Shimbo. Classifier-independent feature selection based on nonparametric discriminant analysis. Lecture Notes in Computer Science, 2396: 470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960470.
                                        pdf.

Schulerud:2002:EMF

[891] Helene Schulerud and Fritz Albregtsen. Effects of many feature candidates in feature selection and classification. Lecture Notes in Computer Science, 2396: 480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960480.
pdf.

Pekalska:2002:SRD

[892] Elzbieta Pekalska and Robert P. W. Duin. Spatial representation of dissimilarity data via lower-complexity linear and nonlinear mappings. Lecture Notes in Computer Science, 2396:488– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960488.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960488.
pdf.

Iwamura:2002:MET

[893] Masakazu Iwamura, Shinichiro Omachi, and Hirotomo Aso. A method to estimate the true Mahalanobis distance from eigenvectors of sample covariance matrix. Lecture Notes in Computer Science, 2396:498–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960498.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2396/23960498.
                                        pdf.

Loog:2002:NIH

[894] M. Loog and R. P. W. Duin. Noniterative heteroscedastic linear dimension reduction for two-class data. Lecture Notes in Computer Science, 2396: 508–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960508.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960508.
                                        pdf.

Barandela:2002:SES

[895] R. Barandela, F. J. Ferri, and T. Nájera. Some experiments in supervised pattern recognition with incomplete training samples. Lecture Notes in Computer Science, 2396:518– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960518.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960518.
                                        pdf.

Kim:2002:RPR

[896] Sang-Woon Kim and B. J. Oommen. Recursive prototype reduction schemes applicable for large data sets. Lecture Notes in Computer Science, 2396:528– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960528.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2396/23960528.
pdf.

Keysers:2002:CTV

[897] Daniel Keysers, Roberto Paredes, Hermann Ney, and Enrique Vidal. Combination of tangent vectors and local representations for handwritten digit recognition. Lecture Notes in Computer Science, 2396:538–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960538.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960538.
pdf.

Cano:2002:TSE

[898] Javier Cano, Juan-Carlos Perez-Cortes, Joaquim Arlandis, and Rafael Llobet. Training set expansion in handwritten character recognition. Lecture Notes in Computer Science, 2396:548– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960548.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960548.
pdf.

Bakus:2002:DCU

[899] Jan Bakus and Mohamed Kamel. Document classification using phrases. Lecture Notes in Computer Science, 2396: 557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960557.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960557.
                                        pdf.

Hamouz:2002:FDL

[900] Miroslav Hamouz, Josef Kittler, Jiri Matas, and Petr Bílek. Face detection by learned affine correspondences. Lecture Notes in Computer Science, 2396: 566–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960566.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960566.
                                        pdf.

Ragheb:2002:SSH

[901] Hossein Ragheb and Edwin R. Hancock. Shape-from-shading for highlighted surfaces. Lecture Notes in Computer Science, 2396:576–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960576.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960576.
                                        pdf.

deRidder:2002:TDI

[902] Dick de Ridder, Robert P. W. Duin, and Josef Kittler. Texture description by independent components. Lecture Notes in Computer Science, 2396:587– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960587.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2396/23960587.
pdf.

Kittler:2002:FMC

[903] Josef Kittler, Marco Ballette, W. J. Christmas, Edward Jaser, and Kieron Messer. Fusion of multiple cue detectors for automatic sports video annotation. Lecture Notes in Computer Science, 2396:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2396/23960597.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2396/23960597.pdf.

Giacinto:2002:QSB

[904] Giorgio Giacinto and Fabio Roli. Query shifting based on Bayesian decision theory for content-based image retrieval. Lecture Notes in Computer Science, 2396:607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960607.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960607.
pdf.

Haindl:2002:RMB

[905] Michal Haindl. Recursive model-based colour image restoration. Lecture Notes in Computer Science, 2396:617– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960617.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2396/23960617.
                                        pdf.

Haddadnia:2002:HFR

[906] Javad Haddadnia, Majid Ahmadi, and Karim Faez. Human face recognition with different statistical features. Lecture Notes in Computer Science, 2396: 627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960627.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960627.
                                        pdf.

Huang:2002:TBM

[907] Yea-Shuan Huang and Yao-Hong Tsai. A transformation-based mechanism for face recognition. Lecture Notes in Computer Science, 2396:636–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960636.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960636.
                                        pdf.

Garcia-Mateos:2002:FDU

[908] Ginés García-Mateos, Alberto Ruiz, and Pedro E. Lopez de Teruel. Face detection using integral projection models. Lecture Notes in Computer Science, 2396:644–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960644.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2396/23960644.
pdf.

Ko:2002:INF

[909] Jaepil Ko, Eunju Kim, and Heyran Byun. Illumination normalized face image for face recognition. Lecture Notes in Computer Science, 2396:654– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960654.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960654.
pdf.

delSolar:2002:TGE

[910] Javier Ruiz del Solar and Pablo Navarrete. Towards a generalized eigenspacebased face recognition framework. Lecture Notes in Computer Science, 2396: 662–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960662.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960662.
pdf.

Gomez:2002:ASS

[911] Jon Ander Gómez and María José Castro. Automatic segmentation of speech at the phonetic level. Lecture Notes in Computer Science, 2396:672– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960672.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2396/23960672.
pdf.

Kinnunen:2002:CDW

[912] Tomi Kinnunen and Ismo Kärkkäinen. Class-discriminative weighted distortion measure for VQ-based speaker identification. Lecture Notes in Computer Science, 2396:681–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2396/23960681.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2396/23960681.pdf.

Semani:2002:AFS

[913] Dahbia Semani, Christophe Saint-Jean, Carl Frélicot, Thierry Bouwmans, and Pierre Courtellemont. Alive fishes species characterization from video sequences. Lecture Notes in Computer Science, 2396:689–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960689.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960689.
pdf.

Santo:2002:ACD

[914] Massimo De Santo, Gennaro Percannella, Carlo Sansone, Roberto Santoro, and Mario Vento. Automatic cut detection in MPEG movies: A multi-expert approach. Lecture Notes in Computer Science, 2396:699–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2396/23960699.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960699.
pdf.

Copsey:2002:BNI

[915] Keith Copsey and Andrew Webb. Bayesian networks for incorporation of contextual information in target recognition systems. Lecture Notes in Computer Science, 2396:709–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960709.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960709.
pdf.

Moreno-Seco:2002:ELF

[916] Francisco Moreno-Seco, Luisa Micó, and Jose Oncina. Extending LAESA fast nearest neighbour algorithm to find the k nearest neighbours. Lecture Notes in Computer Science, 2396:718– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960718.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960718.
pdf.

Gomez-Ballester:2002:FAM

[917] Eva Gómez-Ballester, Luisa Micó, and Jose Oncina. A fast approximated k-median algorithm. Lecture Notes in Computer Science, 2396:725– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960725.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960725.
                                        pdf.

Panuccio:2002:HMM

[918] Antonello Panuccio, Manuele Bicego, and Vittorio Murino. A hidden Markov model-based approach to sequential data clustering. Lecture Notes in Computer Science, 2396:734– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960734.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960734.
                                        pdf.

Perez-Jimenez:2002:GAE

[919] Alberto Perez-Jimenez and Juan-Carlos Perez-Cortes. Genetic algorithms for exploratory data analysis. Lecture Notes in Computer Science, 2396:743–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2396/23960743.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960743.
                                        pdf.

Nyssen:2002:PML

[920] Edgard Nyssen, Naren Naik, and Bart Truyen. Piecewise multi-linear PDF modelling, using an ML approach. Lecture Notes in Computer Science, 2396: 752–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960752.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960752.
pdf.

Aoki:2002:DTU

[921] Kazuaki Aoki and Mineichi Kudo. Decision tree using class-dependent feature subsets. Lecture Notes in Computer Science, 2396:761–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960761.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960761.
pdf.

Sirlantzis:2002:FTB

[922] Konstantinos Sirlantzis, Sanaul Hoque, Michael C. Fairhurst, and Ahmad Fuad Rezaur Rahman. Fusion of n-tuple based classifiers for high performance handwritten character recognition. Lecture Notes in Computer Science, 2396:770–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960770.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960770.
pdf.

Draper:2002:BPA

[923] Bruce A. Draper, Kyungim Baek, and Jeff Boody. A biologically plausible approach to cat and dog discrimination. Lecture Notes in Computer Science, 2396:779–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2396/23960779.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960779.
                                        pdf.

Windridge:2002:MUC

[924] David Windridge and Josef Kittler. Morphologically unbiased classifier combination through graphical PDF correlation. Lecture Notes in Computer Science, 2396:789–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960789.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960789.
                                        pdf.

Sakano:2002:CUC

[925] Hitoshi Sakano and Takashi Suenaga. Classifiers under continuous observations. Lecture Notes in Computer Science, 2396:798–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960798.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960798.
                                        pdf.

Wang:2002:TCB

[926] Jing-Wein Wang. Texture classification based on coevolution approach in multiwavelet feature space. Lecture Notes in Computer Science, 2396:806– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960806.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960806.
pdf.

Gimelfarb:2002:PSM

[927] Georgy Gimel’farb and Uri Lipowezky. Probabilistic signal models to regularise dynamic programming stereo. Lecture Notes in Computer Science, 2396:814–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2396/23960814.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2396/23960814.
pdf.

Imiya:2002:HTA

[928] Atsushi Imiya, Tetsu Hada, and Ken Tatara. The Hough transform without the accumulators. Lecture Notes in Computer Science, 2396:823–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2396/23960823.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2396/23960823.
pdf.

Inesta:2002:RGL

[929] José Manuel Iñesta and Jorge Calera-Rubio. Robust gray-level histogram Gaussian characterisation. Lecture Notes in Computer Science, 2396:833– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960833.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960833.
                                        pdf.

Haindl:2002:MBF

[930] Michal Haindl and Hynek Lauschmann. Model-based fatique fractographs texture analysis. Lecture Notes in Computer Science, 2396:842–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2396/23960842.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960842.
                                        pdf.

Azimifar:2002:HMM

[931] Zohreh Azimifar, Paul Fieguth, and Ed Jernigan. Hierarchical multiscale modeling of wavelet-based correlations. Lecture Notes in Computer Science, 2396:850–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2396/23960850.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2396/23960850.
                                        pdf.

Anonymous:2002:AIx

[932] Anonymous. Author index. Lecture Notes in Computer Science, 2396:861– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2396/2396ind.pdf.

Kerscher:2002:SAD

[933] George Kerscher. Structured access to documents, digital talking books, and beyond: The DAISY consortium. Lecture Notes in Computer Science, 2398: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980001.
pdf.

Session:2002:EEU

[934] Introduction to the Special Thematic Session. eEurope — eAccessibility — user participation participation of people with disabilities and older people in the information society. Lecture Notes in Computer Science, 2398: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980003.
pdf.

totheWorkshop:2002:RBI

[935] Introduction to the Workshop. From research to business improving the take up of results of research projects. Lecture Notes in Computer Science, 2398: 6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980006.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980006.
                                        pdf.

Session:2002:UAI

[936] Introduction to the Special Thematic Session. Universal access to information society technologies: Opportunities for people with disabilities. Lecture Notes in Computer Science, 2398:8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980008.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980008.
                                        pdf.

Gladstone:2002:AUE

[937] Keith Gladstone, Cathy Rundle, and Tara Alexander. Accessibility and usability of eCommerce systems. Lecture Notes in Computer Science, 2398:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980011.
                                        pdf.

Pieper:2002:RCS

[938] M. Pieper. Recommending contextsensitive and process-oriented tourist information to the disabled — the PALIO case. Lecture Notes in Computer Science, 2398:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2398/23980019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980019.
pdf.

Chene:2002:WSA

[939] Denis Chêne and Michel Hoël. Web site accessibility auditing tool for visually deficient persons OCAWA. Lecture Notes in Computer Science, 2398:27– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980027.
pdf.

Holzinger:2002:UCI

[940] Andreas Holzinger. User-centered interface design for disabled and elderly people: First experiences with designing a patient communication system (PACOSY). Lecture Notes in Computer Science, 2398:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980033.
pdf.

Stephanidis:2002:DCE

[941] Constantine Stephanidis. The disappearing computer: Emerging opportunities and challenges for disabled and elderly people. Lecture Notes in Computer Science, 2398:41– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980041.
                                        pdf.

Emiliani:2002:NTS

[942] Pier Luigi Emiliani. New technologies and services for disabled and elderly people in the emerging information society. Lecture Notes in Computer Science, 2398:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980049.
                                        pdf.

Duchateau:2002:SAA

[943] Sylvie Duchateau, Denis Boulay, Claudine Tchang-Ayo, and Dominique Burger. A strategy to achieve the accessibility of public Web sites. Lecture Notes in Computer Science, 2398:58– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980058.
                                        pdf.

Privat:2002:AAV

[944] Régis Privat, Nadine Vigouroux, Philippe Truillet, and Bernard Oriola. Accessibility and affordance for voice interactive systems with the VoiceXML technology. Lecture Notes in Computer Science, 2398:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980061.
pdf.

Session:2002:CIE

[945] Introduction to the Special Thematic Session. Computer and inclusive education. Lecture Notes in Computer Science, 2398:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980064.
pdf.

Watkins:2002:NIC

[946] Amanda Watkins and Harald Weber. National information and communication technology policies. Lecture Notes in Computer Science, 2398:68– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980068.
pdf.

Rainwald:2002:EVS

[947] Walter Rainwald. EUN virtual school special needs education. Lecture Notes in Computer Science, 2398:76– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980076.
                                        pdf.

Pieper:2002:TST

[948] M. Pieper. Tutorial systems to teach standard applications to the learning disabled. Lecture Notes in Computer Science, 2398:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980083.
                                        pdf.

Archambault:2002:VP

[949] Dominique Archambault and Dominique Burger. The Vickie Project. Lecture Notes in Computer Science, 2398:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980090.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980090.
                                        pdf.

Kurhila:2002:UAH

[950] Jaakko Kurhila and Hanna Varjola. Using adaptive hypermedia to evaluate basic arithmetic skills in special education. Lecture Notes in Computer Science, 2398:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980098.
pdf.

Feyerer:2002:IAT

[951] Ewald Feyerer, Klaus Miesenberger, and David Wohlhart. ICT and assistive technology in teachers education and training. Lecture Notes in Computer Science, 2398:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980107.
pdf.

Greve:2002:BIG

[952] Jørgen Greve and Harald Weber. Bridging the information gap between IST and SEN — A practical approach on European level. Lecture Notes in Computer Science, 2398:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980115.
pdf.

Mohamad:2002:EST

[953] Yehya Mohamad, S. Hammer, F. Haverkamp, M. Nöker, and H. Tebarth. Evaluational study: Training with animated pedagogical agents. Lecture Notes in Computer Science, 2398:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980117.
                                        pdf.

Prazak:2002:UMT

[954] Barbara Prazak and Mathilde Niehaus. Usage of modern technologies — A chance for people with special needs?!

— pedagogically relevant criteria for the review of educational software exemplified by “Step by Step 3”. Lecture Notes in Computer Science, 2398: 125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980125.
                                        pdf.

Edler:2002:SII

[955] Cordula Edler. Social innovation in the information society — take it easy. Lecture Notes in Computer Science, 2398: 128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980128.
                                        pdf.

Wohlhart:2002:OII

[956] David Wohlhart. ODL:Inclusive — improving inclusive teaching through online training. Lecture Notes in Computer Science, 2398:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980130.
pdf.

Coombs:2002:ALI

[957] Norman Coombs. Accessible Elearning: Infrastructure and content. Lecture Notes in Computer Science, 2398:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2398/23980133.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2398/23980133.
pdf.

Turrini:2002:AWY

[958] Giovanna Turrini, Laura Cignoni, and Alessandro Paccosi. Addizionario: Words in your pocket. Lecture Notes in Computer Science, 2398:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980136.
pdf.

Shopland:2002:VTT

[959] N. Shopland, J. Lewis, D. J. Brown, and H. M. Powell. Virtual travel training for people with learning disabilities accessing employment including the introduction to the special thematic session “Virtual Reality”. Lecture Notes in Computer Science, 2398: 140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980140.
pdf.

Alcantud:2002:AVR

[960] F. Alcantud, G. Herrera, G. Labajo, I. Dolz, C. Gayá, V. Avila, A. Blanquer, J. L. Cuesta, and J. Arnáiz. Assessing virtual reality as a tool for support imagination. Lecture Notes in Computer Science, 2398:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980143.
pdf.

Brown:2002:LNE

[961] D. J. Brown, M. Yazdanparast, J. Lewis, N. Shopland, and H. M. Powell. Literacy and numeracy edutainment packages for disaffected young learners. Lecture Notes in Computer Science, 2398:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980145.
pdf.

Standen:2002:CVE

[962] P. J. Standen, S. Battersby, and T. L. Lannen. Control of virtual environments for people with intellectual disabilities. Lecture Notes in Computer Science, 2398:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980147.
pdf.

Battersby:2002:OGP

[963] S. J. Battersby, N. Kelly, D. J. Brown, and H. M. Powell. Online gardening to promote social inclusion. Lecture Notes in Computer Science, 2398:150– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980150.
pdf.

Bauer:2002:SPS

[964] D. Bauer, A. Plinge, and M. Finke. Selective phoneme spotting for realization of an s, z, C, t transposer. Lecture Notes in Computer Science, 2398:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980153.
pdf.

Oliveira:2002:MBS

[965] Flavio Oliveira, Iwao Kobayashi, and Mitsumasa Sugawara. A multimedia based software for non-Japanese to learn Japanese sign language. Lecture Notes in Computer Science, 2398:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980162.
pdf.

Minagawa:2002:UID

[966] Hiroki Minagawa, Ichiro Naito, Nobuko Kato, Hiroshi Murakami, and Yasushi Ishihara. The user interface design for the sign language translator in a remote sign language interpretation system. Lecture Notes in Computer Science, 2398:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980164.
pdf.

Nishioka:2002:STH

[967] Tomoyuki Nishioka. The see-through head mount display as the information offering device for the hearing impaired students. Lecture Notes in Computer Science, 2398:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980166.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2398/23980166.
pdf.

Bauer:2002:DSS

[968] D. Bauer, A. Plinge, and M. Finke. Digital speech signal processing to compensate severe sensory hearing deficits: The /s, z, C, t/ transposer module in simulation — an overview and examples. Lecture Notes in Computer Science, 2398:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980168.
pdf.

Session:2002:CDY

[969] Introduction to the Special Thematic Session. Computers for the development of young disabled children. Lecture Notes in Computer Science, 2398: 170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980170.
pdf.

Buaud:2002:MGV

[970] Aurélie Buaud, Harry Svensson, Dominique Archambault, and Dominique Burger. Multimedia games for visually impaired children. Lecture Notes in Computer Science, 2398:173– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980173.
                                        pdf.

Panek:2002:TAM

[971] Paul Panek, Christian Beck, Stefan Mina, Gottfried Seisenbacher, and Wolfgang L. Zagler. Technical assistance for motor- and multiple disabled children — some long term experiences. Lecture Notes in Computer Science, 2398:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980181.
                                        pdf.

Chu:2002:DAW

[972] Chi Nung Chu, Tien Yu Li, and Ming Chung Chen. The design of an adaptive Web browser for young children with reading difficulties. Lecture Notes in Computer Science, 2398:189– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980189.
                                        pdf.

Hilden:2002:CAY

[973] Anita Hildén and Jenny Hammarlund. Can all young disabled children play at the computer? Lecture Notes in Computer Science, 2398:191– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980191.
pdf.

Dutot:2002:TLC

[974] Antoine Dutot, Damien Olivier, and Dominique Archambault. TL a language to create games for visually impaired children. Lecture Notes in Computer Science, 2398:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980193.
pdf.

Kawamoto:2002:PAS

[975] Hiroaki Kawamoto and Yoshiyuki Sankai. Power assist system HAL-3 for gait disorder person. Lecture Notes in Computer Science, 2398:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980196.
pdf.

Wu:2002:CAA

[976] Ting-Fang Wu, Ling-Fu Meng, Hwa-Pey Wang, Wu-Tien Wu, and Tien-Yu Li. Computer access assessment for persons with physical disabilities: A guide to assistive technology interventions. Lecture Notes in Computer Science, 2398:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980204.
pdf.

Ohyama:2002:PEM

[977] Junji Ohyama and Yoshiyuki Sankai. A proposal of effective motion learning method using sensory feedback for walk-disabled people. Lecture Notes in Computer Science, 2398:212– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980212.
pdf.

Nozawa:2002:CMW

[978] Masako Nozawa and Yoshiyuki Sankai. Control method of walking speed and step length for hybrid assistive leg. Lecture Notes in Computer Science, 2398: 220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980220.
pdf.

Ikeuchi:2002:DGT

[979] Hidetaka Ikeuchi, Satoshi Arakane, Kengo Ohnishi, Keiji Imado, Yukio Saito, and Hiroomi Miyagawa. The development of gait training system for computer-aided rehabilitation. Lecture Notes in Computer Science, 2398:228– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980228.
pdf.

Gips:2002:AIC

[980] James Gips, Philip A. DiMattia, Marialice Curran, Debra Lees, and Maureen Gates. Accessing Internet courses by eye movement and head movement. Lecture Notes in Computer Science, 2398:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980236.
pdf.

Li:2002:PIW

[981] Tien yu Li, Ling-Fu Meng, Chien-Huey Sophie Chang, Ming-Chung Chen, Chi-Nung Chu, Arr-Ming Chou, Tony Yang, Chin-Chen Hui, Ai-Chiao Ku, and Yun-Lung Lin. The program for improving the working interfaces and increasing the work competencies of people with severe physical disabilities: The evaluation, design, and training of the adaptive computer devices. Lecture Notes in Computer Science, 2398:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980238.
                                        pdf.

Session:2002:MTP

[982] Introduction to the Special Thematic Session. Matching typing persons and intelligent interfaces. Lecture Notes in Computer Science, 2398:241– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980241.
                                        pdf.

Matiasek:2002:FML

[983] Johannes Matiasek, Marco Baroni, and Harald Trost. FASTY — A multilingual approach to text prediction. Lecture Notes in Computer Science, 2398:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980243.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980243.
                                        pdf.

Willis:2002:IFA

[984] Tim Willis, Helen Pain, Shari Trewin, and Stephen Clark. Informing flexible abbreviation expansion for users with motor disabilities. Lecture Notes in Computer Science, 2398:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980251.
pdf.

Evreinov:2002:CIC

[985] Grigori Evreinov and Roope Raisamo. Cyclic input of characters through a single button manipulation. Lecture Notes in Computer Science, 2398:259– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980259.
pdf.

Lauruska:2002:DPC

[986] Vidas Lauruska and Tomas Musteikis. Developing of predictive communication system for the physically disabled people. Lecture Notes in Computer Science, 2398:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980267.
pdf.

Chen:2002:TAD

[987] Ming-Chung Chen, Hwa-Pey Wang, and Tien-Yu Li. Teaching adolescents with disabilities to learn Chinese keyboarding by using multimedia computer assisted input method learning system. Lecture Notes in Computer Science, 2398:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980271.
pdf.

Trewin:2002:AAS

[988] Shari Trewin, Patti Slobogin, and M. G. Power. Accelerating assessment with self-optimizing devices. Lecture Notes in Computer Science, 2398:273– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980273.
pdf.

Boissiere:2002:PEF

[989] Philippe Boissière and Daniel Dours. A proposal of an evaluation framework for writing assistance systems: Application to VITIPI. Lecture Notes in Computer Science, 2398:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980276.
pdf.

Clemens:2002:IAT

[990] Dirk Clemens, Helmut Heck, Michael Kühn, Olaf Perlick, and Frank Reins. Individually assisted text entry with situational and contextual prediction. Lecture Notes in Computer Science, 2398:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980279.
pdf.

Session:2002:ADM

[991] Introduction to the Special Thematic Session. Accessible digital media. Lecture Notes in Computer Science, 2398:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2398/23980282.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2398/23980282.
pdf.

Roberts:2002:MIE

[992] Vera Roberts and Deborah Fels. Methods for inclusion: Employing think aloud protocol with individuals who are deaf. Lecture Notes in Computer Science, 2398:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980284.
pdf.

Silverman:2002:ECD

[993] Charles Silverman and Deborah I. Fels. Emotive captioning in a digital world. Lecture Notes in Computer Science, 2398:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980292.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980292.
                                        pdf.

Verboom:2002:SSM

[994] Maarten Verboom, David Crombie, Evelien Dijk, and Mildred Theunisz. Spoken subtitles: Making subtitled TV programmes accessible. Lecture Notes in Computer Science, 2398:295– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980295.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980295.
                                        pdf.

Fels:2002:ADM

[995] Deborah I. Fels. Accessible digital media. Lecture Notes in Computer Science, 2398:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2398/23980303.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2398/23980303.pdf.

Session:2002:EVL

[996] Introduction to the Special Thematic Session. Electronic and virtual libraries: Access for print disabled people. Lecture Notes in Computer Science, 2398:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980308.
pdf.

Kerscher:2002:SRM

[997] George Kerscher. Semantically rich markup, should you be agnostic in your choice of a vocabulary. Lecture Notes in Computer Science, 2398:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980312.
pdf.

Kahlisch:2002:DGL

[998] Thomas Kahlisch. DZB — A German library for the blind goes digital. Lecture Notes in Computer Science, 2398: 316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980316.
pdf.

Hammerschmid:2002:HCL

[999] Ruth Hammerschmid, Klaus Miesenberger, and Bernhard Stöger. Harmonisation of the copyright law throughout the European Union — A challenge for all print disabled people. Lecture Notes in Computer Science, 2398:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980321.
pdf.

Moulin:2002:DDO

[1000] Claude Moulin, Sylvain Giroux, Dominique Archambault, Davide Carboni, and Dominique Burger. A distributed document oriented architecture for rendering services to visually impaired students. Lecture Notes in Computer Science, 2398:329– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980329.
pdf.

Taha:2002:AVL

[1001] Karim Taha and Lori Stefano Pertucci. ABA’s virtual library services for blind and visually impaired people. Lecture Notes in Computer Science, 2398:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980337.
pdf.

Sable:2002:LPL

[1002] Sébastien Sablé and Dominique Archambault. Libbraille: A portable library to easily access Braille displays. Lecture Notes in Computer Science, 2398:345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2398/23980345.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2398/23980345.
pdf.

Raynal:2002:CHB

[1003] Mathieu Raynal and Mathieu Serrurier. CYNTHIA: An HTML browser for visually handicapped people. Lecture Notes in Computer Science, 2398: 353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980353.
pdf.

Guillon:2002:SIS

[1004] Benoit Guillon, Dominique Burger, and Bruno Marmol. A secure Internet service for delivering documents for the blind. Lecture Notes in Computer Science, 2398:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980360.
pdf.

Kottstorfer:2002:VLI

[1005] Marco Köttstorfer and Klaus Miesenberger. Virtual libraries initiatives with usable results for print disabled people. Lecture Notes in Computer Science, 2398:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980366.
pdf.

Crombie:2002:UDB

[1006] David Crombie, Arne Leeman, Marian Oosting, and Maarten Verboom. Unlocking doors: Building an accessible online information node. Lecture Notes in Computer Science, 2398:374– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980374.
pdf.

Paepen:2002:UXR

[1007] Bert Paepen and Jan Engelen. Using XML as a reading enabler for visually impaired persons. Lecture Notes in Computer Science, 2398:382– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980382.
pdf.

Schmetzke:2002:AOL

[1008] Axel Schmetzke. The accessibility of online library resources for people with print disabilities: Research and strategies for change. Lecture Notes in Computer Science, 2398:390– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980390.
pdf.

Vincent:2002:QAD

[1009] Brasseur Vincent, Mohamed Slimane, Nicole Vincent, and Dominique Burger. Quick access to dictionaries for blind and visually impaired people. Lecture Notes in Computer Science, 2398:398– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980398.
pdf.

Jaworek:2002:PPA

[1010] Gerhard Jaworek. Perspectives and possibilities for the automated processing of scientific texts for blind and visually impaired people. Lecture Notes in Computer Science, 2398:401– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980401.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2398/23980401.
                                        pdf.

Patil:2002:UUI

[1011] Basawaraj Patil, Klaus Maetzel, and Erich J. Neuhold. Universal usability issues of textual information structures, commands, and languages of native visually challenged users: An inclusive design framework. Lecture Notes in Computer Science, 2398:403– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980403.
                                        pdf.

DeMeglio:2002:AID

[1012] Marisa DeMeglio, Markku T. Hakkinen, and Hiroshi Kawamura. Accessible interface design: Adaptive multimedia information system (AMIS). Lecture Notes in Computer Science, 2398:406–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980406.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980406.
                                        pdf.

Petrie:2002:RMD

[1013] Helen Petrie and Gerhard Weber. Reading multimedia documents. Lecture Notes in Computer Science, 2398: 413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2398/23980413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980413.
pdf.

Bellik:2002:CVI

[1014] Yacine Bellik and René Farcy. Comparison of various interface modalities for a locomotion assistance device. Lecture Notes in Computer Science, 2398: 421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980421.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980421.
pdf.

Guillet:2002:PHV

[1015] Vivien Guillet, Beatrice Rumpler, and Jean-Marie Pinon. Providing help for visually impaired people’s navigation in an urban environment using GPS. Lecture Notes in Computer Science, 2398: 429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980429.
pdf.

Al-Salman:2002:NTD

[1016] AbdulMalik S. Al-Salman. New trends and developments in computer and Internet technology for visually impaired people: A survey and usefulness in the Arab sub-region. Lecture Notes in Computer Science, 2398:437– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980437.
                                        pdf.

Schwende:2002:AEA

[1017] Harald Schwende. Auditory emotional access to visual information. Lecture Notes in Computer Science, 2398:445– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980445.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980445.
                                        pdf.

Nomura:2002:ASA

[1018] Yoshihiko Nomura, Sayuri Nishihama, Tokuhiro Sugiura, Hirokazu Matsui, and Norihiko Kato. Auditory sensation aided image perception: Integration of somesthetic sense of fingertip position. Lecture Notes in Computer Science, 2398:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980448.
                                        pdf.

Silva:2002:NAA

[1019] Luciano Silva and Olga Regina Pereira Bellon. A novel application to aid low vision computer users. Lecture Notes in Computer Science, 2398:455– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980455.
pdf.

Jung:2002:VIP

[1020] Martin Jung and Erdmuthe Meyer zu Bexten. Visually impaired persons using Microsoft PowerPoint presentations. Lecture Notes in Computer Science, 2398:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2398/23980463.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2398/23980463.pdf.

Gilden:2002:UMO

[1021] Deborah Gilden. Using MS Office features as low vision accessibility tools. Lecture Notes in Computer Science, 2398:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980469.
pdf.

Session:2002:AMB

[1022] Introduction to the Special Thematic Session. Access to mathematics by blind students. Lecture Notes in Computer Science, 2398:471– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980471.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980471.
                                        pdf.

Karshmer:2002:AAB

[1023] Arthur I. Karshmer, Gopal Gupta, and Doug Gillan. Architecting an auditory browser for navigating mathematical expressions. Lecture Notes in Computer Science, 2398:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980477.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980477.
                                        pdf.

Gaura:2002:RRE

[1024] Pavel Gaura. REMathEx — reader and editor of the mathematical expressions for blind students. Lecture Notes in Computer Science, 2398:486– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980486.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980486.
                                        pdf.

Fitzpatrick:2002:STD

[1025] D. Fitzpatrick. Speaking technical documents: Using prosody to convey textual and mathematical material. Lecture Notes in Computer Science, 2398: 494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980494.
pdf.

Gardner:2002:ABS

[1026] John A. Gardner. Access by blind students and professionals to mainstream math and science. Lecture Notes in Computer Science, 2398:502– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980502.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980502.
pdf.

MeyerzuBexten:2002:LUA

[1027] Erdmuthe Meyer zu Bexten and Martin Jung. L^{A}TEX at the University of Applied Sciences Giessen-Friedberg — experiences at the Institute for Visually Impaired Students. Lecture Notes in Computer Science, 2398:508– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980508.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980508.
pdf.

Kobolkova:2002:EAM

[1028] Martina Kobolkova and Peter Lecky.

Experience with access to mathematics for blind students in Slovakia. Lecture Notes in Computer Science, 2398: 510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980510.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980510.
                                        pdf.

Toshihiro:2002:IAT

[1029] Kanahori Toshihiro and Suzuki Masakazu. Infty alpha test site. Lecture Notes in Computer Science, 2398:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980512.
                                        pdf.

Chen:2002:DAP

[1030] Ming Chung Chen, Tien Yu Li, Chi Nung Chu, Ling Fu Meng, Chien-Huey Sophie Chang, Arr Mien Chou, Tony Yang, Chih Chen Hui, Ai Chiao Ku, and Yun Lung Lin. Designing an AutoMouse to promote computer accessibility for people with severe physical impairments. Lecture Notes in Computer Science, 2398:514– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980514.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980514.
                                        pdf.

Ault:2002:ELD

[1031] H. K. Ault, J. W. Deloge, R. W. Lapp, M. J. Morgan, and J. R. Barnett. Evaluation of long descriptions of statistical graphics for blind and low vision Web users. Lecture Notes in Computer Science, 2398:517– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980517.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980517.
pdf.

Kobayashi:2002:TDS

[1032] Makoto Kobayashi and Tetsuya Watanabe. A tactile display system equipped with a pointing device — MIMIZU. Lecture Notes in Computer Science, 2398:527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980527.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980527.
pdf.

Batusic:2002:PTG

[1033] Mario Batusic and Franz Urban.

Preparing tactile graphics for traditional Braille printers with Brl-GraphEditor. Lecture Notes in Computer Science, 2398:535–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2398/23980535.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2398/23980535.pdf.

Petrie:2002:TSP

[1034] Helen Petrie, Christoph Schlieder, Paul Blenkhorn, Gareth Evans, Alasdair King, Anne-Marie O’Neill, George T. Ioannidis, Blaithin Gallagher, David Crombie, Rolf Mager, and Maurizio Alafaci. TeDUB: A system for presenting and exploring technical drawings for blind people. Lecture Notes in Computer Science, 2398:537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980537.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980537.
                                        pdf.

Shimizu:2002:IUI

[1035] Yutaka Shimizu, Masami Shinohara, Hideji Nagaoka, and Yasushige Yonezawa. Improvement of user interface for blind PC users. Lecture Notes in Computer Science, 2398:540– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980540.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980540.
                                        pdf.

Graupp:2002:HPP

[1036] Helen Graupp, Keith Gladstone, and Leanne Thompson. Haptic perception of 2D pictures and 3D objects: Accurate mental representation as a function of visual status. Lecture Notes in Computer Science, 2398:543– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980543.
pdf.

Enikov:2002:MMS

[1037] Eniko T. Enikov, Kalin V. Lazarov, and Gilbert R. Gonzales. Microelectrical mechanical systems actuator array for tactile communication. Lecture Notes in Computer Science, 2398:551– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980551.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980551.
pdf.

Challis:2002:DIT

[1038] Ben Challis. Designing interactive tactile diagrams. Lecture Notes in Computer Science, 2398:559–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980559.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980559.
pdf.

Murphy:2002:AC

[1039] Damian T. Murphy, Michael C. Kelly, and Anthony I. Tew. 3D audio in the 21^{st} century. Lecture Notes in Computer Science, 2398:562–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980562.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980562.
                                        pdf.

Krueger:2002:GPK

[1040] Myron W. Krueger and Deborah Gilden. Going places with “KnowWare”: Virtual reality maps for blind people. Lecture Notes in Computer Science, 2398:565–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980565.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980565.
                                        pdf.

Kurniawan:2002:MMB

[1041] Sri Hastuti Kurniawan and Alistair Sutcliffe. Mental models of blind users in the Windows environment. Lecture Notes in Computer Science, 2398:568– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980568.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980568.
                                        pdf.

Evans:2002:APN

[1042] D. G. Evans, K. Polyzoaki, and P. Blenkhorn. An approach to producing new languages for talking applications for use by blind people. Lecture Notes in Computer Science, 2398: 575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980575.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980575.
pdf.

Fisher:2002:URT

[1043] Wendy Fisher and Helen Petrie. User requirements for technologies for personal communication and information use for deafblind people. Lecture Notes in Computer Science, 2398:583– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980583.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980583.
pdf.

Butorac:2002:PIW

[1044] Danko Butorac. Project IPSIS — Web portal and Linux for the blind. Lecture Notes in Computer Science, 2398: 585–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980585.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980585.
pdf.

Diggle:2002:ALE

[1045] T. Diggle, S. Kurniawan, D. G. Evans, and P. Blenkhorn. An analysis of layout errors in word processed documents produced by blind people. Lecture Notes in Computer Science, 2398:587– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980587.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980587.
                                        pdf.

Watanabe:2002:WSR

[1046] Tetsuya Watanabe, Chuji Sashida, and Shinichi Okada. Windows screen reader user survey in Japan. Lecture Notes in Computer Science, 2398:589– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980589.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980589.
                                        pdf.

Evans:2002:TCD

[1047] D. G. Evans and P. Blenkhorn. Tools for creating documents in ‘preferred format’ for visually impaired people. Lecture Notes in Computer Science, 2398:591–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980591.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980591.
                                        pdf.

Hara:2002:BPS

[1048] Shunsuke Hara, Ryoko Kawamitsu, Kanako Kusunoki, Hiroshi Sato, and Mamoru Fujiyoshi. Braille printing system. Lecture Notes in Computer Science, 2398:594–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980594.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980594.
pdf.

Blomquist:2002:BCW

[1049] Mats Blomquist. Braille contractions in WinBraille. Lecture Notes in Computer Science, 2398:602–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980602.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980602.
pdf.

Arato:2002:BTT

[1050] Andras Arato, Theresa Vaspori, Gareth Evans, and Paul Blenkhorn. Braille to text translation for Hungarian. Lecture Notes in Computer Science, 2398:610–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2398/23980610.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2398/23980610.pdf.

Blomquist:2002:WAP

[1051] Mats Blomquist and Per Burman.

The WinBraille approach to producing Braille quickly and effectively. Lecture Notes in Computer Science, 2398: 618–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980618.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980618.
                                        pdf.

Takahashi:2002:HDS

[1052] Yoshiyuki Takahashi, Takafumi Terada, Kaoru Inoue, Satoshi Sasada, and Takashi Komeda. Haptic device system for upper limb motor function and cognitive function rehabilitation. Lecture Notes in Computer Science, 2398: 620–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980620.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980620.
                                        pdf.

Mannikko-Barbutiu:2002:NAE

[1053] Sirkku Männikkö-Barbutiu. Networking for accessibility: Elderly people and new technology. Lecture Notes in Computer Science, 2398:628– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980628.
                                        pdf.

Panek:2002:THR

[1054] Paul Panek, Christian Beck, Andreas Hochgatterer, Stefan Mina, Barbara Prazak, Gottfried Seisenbacher, Mathijs Soede, and Wolfgang L. Zagler. Telehelp and remote service provision using RESORT prototype system. Lecture Notes in Computer Science, 2398: 635–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980635.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980635.
pdf.

Treviranus:2002:MYH

[1055] Jutta Treviranus. Making yourself at home — portable personal access preferences. Lecture Notes in Computer Science, 2398:643–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980643.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980643.
pdf.

Mayer:2002:EPS

[1056] Peter Mayer, Georg Edelmayer, and Wolfgang L. Zagler. EU project SILC — personal bio-sensor based alarm system. Lecture Notes in Computer Science, 2398:649–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980649.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980649.
pdf.

Tollefsen:2002:VUA

[1057] Morten Tollefsen and Magne Lunde.

Videophones: Useful assistive devices for visually impaired persons? Lecture Notes in Computer Science, 2398: 657–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980657.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980657.
                                        pdf.

Leloup:2002:HHM

[1058] Jérôme Leloup and Pierre Gaucher. HMPH — “Habitat Mobile pour Personnes Handicapées” — software for the design of an adapted living area. Lecture Notes in Computer Science, 2398:659–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980659.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980659.
                                        pdf.

James:2002:BIL

[1059] Diana P. James. “Bonkers”. increasing literacy skills: An eclectic electronic teaching approach gone “Bonkers”. Lecture Notes in Computer Science, 2398:662–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980662.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980662.
                                        pdf.

Session:2002:AM

[1060] Introduction to the Special Thematic Session. Access music. Lecture Notes in Computer Science, 2398:665– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980665.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980665.
pdf.

Crombie:2002:SME

[1061] David Crombie, Sijo Dijkstra, Edmar Schut, and Natasha Lindsay. Spoken music: Enhancing access to music for the print disabled. Lecture Notes in Computer Science, 2398:667– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980667.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980667.
pdf.

Ronch:2002:MSR

[1062] Angelo Da Ronch and Adriano Blason. A music score recogniser for the blind. Lecture Notes in Computer Science, 2398:675–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980675.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980675.
pdf.

Evreinova:2002:VMN

[1063] Tatiana Evreinova and Roope Raisamo.

Visualization of music notation for partially sighted persons. Lecture Notes in Computer Science, 2398:682– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980682.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980682.
                                        pdf.

Murphy:2002:ASS

[1064] Damian T. Murphy and Mark A. Hildred. Accessing the software studio. Lecture Notes in Computer Science, 2398:684–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980684.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980684.
                                        pdf.

Chamberlain:2002:ODI

[1065] Helen Chamberlain. Opening doors to information technology — section 508 of the US rehabilitation act of 1973. Lecture Notes in Computer Science, 2398:686–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980686.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980686.
                                        pdf.

Cameron:2002:ANA

[1066] Zachariah Cameron. Access: The North American experience of legislating attitudes. Lecture Notes in Computer Science, 2398:692–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980692.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980692.
pdf.

Viglas:2002:OMT

[1067] Constantinos Viglas and Georgios Kouroupetroglou. An open machine translation system for augmentative and alternative communication. Lecture Notes in Computer Science, 2398: 699–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980699.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980699.
pdf.

Kowalik:2002:CLD

[1068] Ryszard Kowalik. Capabilities and limitations of the disabled in remote access to information. Lecture Notes in Computer Science, 2398:707–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980707.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980707.
pdf.

Iversen:2002:IEI

[1069] Charmaine Mueller Iversen. IMPACT:

An example of information, awareness, and legislation, leading toward a more accessible workplace. Lecture Notes in Computer Science, 2398:709– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980709.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980709.
                                        pdf.

Session:2002:ICF

[1070] Introduction to the Special Thematic Session. IT-certificates to foster the vocational integration of people with disabilities: ECDL(R)PD. Lecture Notes in Computer Science, 2398: 712–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980712.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980712.
                                        pdf.

Tollefsen:2002:EVI

[1071] Morten Tollefsen and Magne Lunde. ECDL for visually impaired persons: Learning materials, individual training, and distance education. Lecture Notes in Computer Science, 2398:715– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980715.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980715.
                                        pdf.

Petz:2002:EPU

[1072] Andrea Petz and Klaus Miesenberger. ECDL(R) PD — using a well known standard to lift barriers on the labour market. Lecture Notes in Computer Science, 2398:723– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980723.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980723.
pdf.

Petz:2002:KKI

[1073] Andrea Petz. KISS — keep it short and simple? Lecture Notes in Computer Science, 2398:731–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980731.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980731.
pdf.

Lunde:2002:MWE

[1074] Magne Lunde and Morten Tollefsen. Many ways to ECDL. Lecture Notes in Computer Science, 2398:734– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980734.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980734.
pdf.

Session:2002:SAM

[1075] Introduction to the Special Thematic Session. Studying and academic mobility — IT support for students with disabilities. Lecture Notes in Computer Science, 2398:737– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980737.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980737.
                                        pdf.

Petz:2002:DAS

[1076] Andrea Petz, Klaus Miesenberger, and Bernhard Stöger. Developing academic skills among print disabled students: IT based Austrian-wide network for service provision. Lecture Notes in Computer Science, 2398:739– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980739.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980739.
                                        pdf.

Arion:2002:ALA

[1077] Mirela Arion, Radu Ciupa, and Marius Iulian Tutuianu. Access to learning for all. Lecture Notes in Computer Science, 2398:747–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980747.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980747.
                                        pdf.

Hecht:2002:EDP

[1078] Janice Hecht and Bogdan Zamfir. The evolving door: Providing technology access for post-secondary students in a constantly changing technology-based culture. Lecture Notes in Computer Science, 2398:749– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980749.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980749.
pdf.

Schmiedl:2002:TAT

[1079] Todd Schmiedl and S. L. Dyser.

Taking assistive technology to the highest degree. Lecture Notes in Computer Science, 2398:751–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980751.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980751.
pdf.

Katoh:2002:LEO

[1080] Hiroshi Katoh, Martin Pauly, Shunsuke Hara, Hideji Nagaoka, and Klaus Miesenberger. Learning from each other: Comparing the service provision for blind and visually-impaired students in post-secondary education in Japan and Europe. Lecture Notes in Computer Science, 2398:753–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980753.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2398/23980753.
                                        pdf.

Kurniawan:2002:HAW

[1081] Sri H. Kurniawan. How accessible are Web information resources for students with disabilities? Lecture Notes in Computer Science, 2398:756– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980756.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980756.
                                        pdf.

Roussel:2002:PAT

[1082] Calais Roussel. Providing assistive technology and alternate media in community colleges: A case study. Lecture Notes in Computer Science, 2398:758– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980758.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980758.
                                        pdf.

Scherer:2002:ICI

[1083] Marcia J. Scherer. International collaboration to improve assistive technology outcomes. Lecture Notes in Computer Science, 2398:761–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2398/23980761.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2398/23980761.
pdf.

Scherer:2002:SAT

[1084] Marcia J. Scherer. The study of assistive technology outcomes in the united states. Lecture Notes in Computer Science, 2398:764–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980764.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980764.
pdf.

deWitte:2002:FCO

[1085] Luc P. de Witte, G. J. Gelderblom, and R. D. Wessels. Future challenges in AT outcome measurement. Lecture Notes in Computer Science, 2398:772– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980772.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2398/23980772.
pdf.

Gelderblom:2002:EIA

[1086] G. J. Gelderblom, L. P. de Witte, and R. D. Wessels. European instruments for assistive technology outcome assessment. Lecture Notes in Computer Science, 2398:777–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2398/23980777.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2398/23980777.
                                        pdf.

Craddock:2002:IOM

[1087] Gerald Craddock. Implementing outcomes measurement in an assistive technology service delivery system. Lecture Notes in Computer Science, 2398:782–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2398/23980782.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2398/23980782.
                                        pdf.

Anonymous:2002:AIy

[1088] Anonymous. Author index. Lecture Notes in Computer Science, 2398:791– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2398/2398auth.pdf.

Schiper:2002:FDV

[1089] André Schiper. Failure detection vs group membership in faulttolerant distributed systems: Hidden trade-offs. Lecture Notes in Computer Science, 2399:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2399/23990001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2399/23990001.
                                        pdf.

Sands:2002:PTC

[1090] David Sands. Probability and timing: Challenges for secure programming (invited talk). Lecture Notes in Computer Science, 2399:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990016.
pdf.

Aldini:2002:SAP

[1091] Alessandro Aldini and Roberto Gorrieri. Security analysis of a probabilistic non-repudiation protocol. Lecture Notes in Computer Science, 2399:17– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990017.
pdf.

Bohnenkamp:2002:MVM

[1092] Henrik Bohnenkamp and Boudewijn Haverkort. The mean value of the maximum. Lecture Notes in Computer Science, 2399:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990037.
pdf.

DArgenio:2002:RRS

[1093] Pedro R. D’Argenio, Bertrand Jeannet, Henrik E. Jensen, and Kim G. Larsen. Reduction and refinement strategies for probabilistic analysis. Lecture Notes in Computer Science, 2399:57– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990057.
pdf.

Fecher:2002:ARP

[1094] Harald Fecher, Mila Majster-Cederbaum, and Jinzhao Wu. Action refinement for probabilistic processes with true concurrency models. Lecture Notes in Computer Science, 2399:77– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990077.
pdf.

Haar:2002:PUP

[1095] Stefan Haar. Probabilistic unfoldings and partial order fairness in petri nets. Lecture Notes in Computer Science, 2399:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990095.
pdf.

Huth:2002:PPA

[1096] Michael Huth. Possibilistic and probabilistic abstraction-based model checking. Lecture Notes in Computer Science, 2399:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2399/23990115.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2399/23990115.pdf.

Kwiatkowska:2002:CSL

[1097] Marta Kwiatkowska and Rashid Mehmood. Out-of-core solution of large linear systems of equations arising from stochastic modelling. Lecture Notes in Computer Science, 2399:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990135.
pdf.

Kwiatkowska:2002:MCC

[1098] Marta Kwiatkowska, Gethin Norman, and António Pacheco. Model checking CSL until formulae with random time bounds. Lecture Notes in Computer Science, 2399:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990152.
pdf.

Kwiatkowska:2002:PMC

[1099] Marta Kwiatkowska, Gethin Norman, and Jeremy Sproston. Probabilistic model checking of the IEEE 802.11 wireless local area network protocol. Lecture Notes in Computer Science, 2399:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2399/23990169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2399/23990169.
                                        pdf.

Kuntz:2002:DSR

[1100] Matthias Kuntz and Markus Siegle. Deriving symbolic representations from stochastic process algebras. Lecture Notes in Computer Science, 2399:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2399/23990188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2399/23990188.
                                        pdf.

Bournez:2002:GEP

[1101] Olivier Bournez. A generalization of equational proof theory? Lecture Notes in Computer Science, 2399:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2399/23990207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2399/23990207.
                                        pdf.

Bravetti:2002:IAS

[1102] Mario Bravetti. An integrated approach for the specification and analysis of stochastic real-time systems (short abstract). Lecture Notes in Computer Science, 2399:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990209.
pdf.

Pierro:2002:PAI

[1103] Alessandra Di Pierro and Herbert Wiklicky. Probabilistic abstract interpretation and statistical testing (extended abstract). Lecture Notes in Computer Science, 2399:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990211.
pdf.

Lassaigne:2002:AVP

[1104] Richard Lassaigne and Sylvain Peyronnet. Approximate verification of probabilistic systems. Lecture Notes in Computer Science, 2399:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2399/23990213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2399/23990213.
pdf.

Anonymous:2002:AIz

[1105] Anonymous. Author index. Lecture Notes in Computer Science, 2399:215– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2399/23990215.pdf.

Choi:2002:OCW

[1106] Y. r. Choi, A. Garg, S. Rai, J. Misra, and H. Vin. Orchestrating computations on the World-Wide Web. Lecture Notes in Computer Science, 2400: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000001.
                                        pdf.

Chalmers:2002:RRR

[1107] A. Chalmers and K. Cater. Realistic rendering in real-time. Lecture Notes in Computer Science, 2400:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000021.
                                        pdf.

Benveniste:2002:NMN

[1108] A. Benveniste. Non-massive, nonhigh performance, distributed computing: Selected issues. Lecture Notes in Computer Science, 2400:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000029.
pdf.

Feitelson:2002:FFF

[1109] D. G. Feitelson. The forgotten factor:

Facts on performance evaluation and its dependence on workloads. Lecture Notes in Computer Science, 2400:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000049.
pdf.

Khosla:2002:SNP

[1110] P. K. Khosla. Sensor networks — promise and challenges. Lecture Notes in Computer Science, 2400:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000061.
pdf.

Reinefeld:2002:CTW

[1111] A. Reinefeld and F. Schintke. Concepts and technologies for a worldwide Grid infrastructure. Lecture Notes in Computer Science, 2400:62– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000062.
                                        pdf.

Bubak:2002:STE

[1112] M. Bubak and T. Ludwig. Support tools and environments. Lecture Notes in Computer Science, 2400:73– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000073.
                                        pdf.

Truong:2002:SPA

[1113] H.-L. Truong and T. Fahringer. SCALEA: A performance analysis tool for distributed and parallel programs. Lecture Notes in Computer Science, 2400:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000075.
                                        pdf.

Roth:2002:DSH

[1114] P. C. Roth and B. P. Miller. Deep start: A hybrid strategy for automated performance problem searches. Lecture Notes in Computer Science, 2400:86– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2400/24000086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000086.
pdf.

Freitag:2002:STM

[1115] F. Freitag, J. Caubet, and J. Labarta.

On the scalability of tracing mechanisms. Lecture Notes in Computer Science, 2400:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2400/24000097.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2400/24000097.pdf.

Hey:2002:CBP

[1116] A. J. G. Hey, J. Papay, A. J. Keane, and S. J. Cox. Component based problem solving environment. Lecture Notes in Computer Science, 2400:105– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000105.
pdf.

Kovacs:2002:ITA

[1117] J. Kovacs, G. Kusper, R. Lovas, and W. Schreiner. Integrating temporal assertions into a parallel debugger. Lecture Notes in Computer Science, 2400: 113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000113.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2400/24000113.
                                        pdf.

Nonaka:2002:LCH

[1118] J. Nonaka, G. H. Pfitscher, K. Onisi, and H. Nakano. Low-cost hybrid internal clock synchronization mechanism for COTS PC cluster (research note). Lecture Notes in Computer Science, 2400:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000121.
                                        pdf.

Nebro:2002:NPI

[1119] A. J. Nebro, E. Alba, F. Luna, and J. M. Troya. .NET as a platform for implementing concurrent objects (research note). Lecture Notes in Computer Science, 2400:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000125.
                                        pdf.

Miller:2002:PEA

[1120] B. P. Miller, J. Labarta, F. Schintke, and J. Simon. Performance evaluation, analysis and optimization. Lecture Notes in Computer Science, 2400: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2400/24000131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000131.
pdf.

Dementiev:2002:PMS

[1121] R. Dementiev, M. Klein, and W. J.

Paul. Performance of MP3D on the SB-PRAM prototype (research note). Lecture Notes in Computer Science, 2400: 132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000132.
pdf.

Cohen:2002:MPP

[1122] A. Cohen, D. Genius, A. Kortebi, Z. Chamski, M. Duranton, and P. Feautrier. Multi-periodic process networks: Prototyping and verifying stream-processing systems. Lecture Notes in Computer Science, 2400:137– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000137.
pdf.

vanGemund:2002:SCE

[1123] A. J. C. van Gemund. Symbolic cost estimation of parallel applications. Lecture Notes in Computer Science, 2400: 147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2400/24000147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000147.
                                        pdf.

Baker:2002:PMI

[1124] Z. K. Baker and V. K. Prasanna. Performance modeling and interpretive simulation of PIM architectures and applications (research note). Lecture Notes in Computer Science, 2400:157– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000157.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000157.
                                        pdf.

Bane:2002:EOA

[1125] M. K. Bane and G. D. Riley. Extended overhead analysis for OpenMP (research note). Lecture Notes in Computer Science, 2400:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000162.
                                        pdf.

DeRose:2002:CCG

[1126] L. DeRose and F. Wolf. CATCH — A call-graph based automatic tool for capture of hardware performance metrics for MPI and OpenMP applications. Lecture Notes in Computer Science, 2400:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000167.
pdf.

Berg:2002:SPT

[1127] E. Berg and E. Hagersten. SIP:

Performance tuning through source code interdependence. Lecture Notes in Computer Science, 2400:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000177.
pdf.

Drozdowski:2002:SLB

[1128] M. Drozdowski, I. Milis, L. Rudolph, and D. Trystram. Scheduling and load balancing. Lecture Notes in Computer Science, 2400:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000187.
pdf.

Lowe:2002:STG

[1129] W. Löwe and W. Zimmermann.

On scheduling task-graphs to LogPmachines with disturbances. Lecture Notes in Computer Science, 2400:189– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000189.
                                        pdf.

Altilar:2002:OSA

[1130] D. T. Altilar and Y. Paker. Optimal scheduling algorithms for communication constrained parallel processing. Lecture Notes in Computer Science, 2400:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2400/24000197.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000197.
                                        pdf.

Krevat:2002:JSB

[1131] E. Krevat, J. G. Castaños, and J. E. Moreira. Job scheduling for the Blue-Gene/L system (research note). Lecture Notes in Computer Science, 2400: 207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000207.
                                        pdf.

Solar:2002:ASP

[1132] M. Solar and M. Inostroza. An automatic scheduler for parallel machines (research note). Lecture Notes in Computer Science, 2400:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2400/24000212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000212.
pdf.

Angel:2002:NAR

[1133] E. Angel, E. Bampis, and R. Giroudeau.

Non-approximability results for the hierarchical communication problem with a bounded number of clusters. Lecture Notes in Computer Science, 2400:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2400/24000217.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2400/24000217.
pdf.

Fujimoto:2002:NAB

[1134] N. Fujimoto and K. Hagihara. Nonapproximability of the bulk synchronous task scheduling problem. Lecture Notes in Computer Science, 2400: 225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000225.
pdf.

Gine:2002:ATS

[1135] F. Giné, F. Solsona, P. Hernández, and E. Luque. Adjusting time slices to apply coscheduling techniques in a nondedicated NOW (research note). Lecture Notes in Computer Science, 2400: 234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000234.
                                        pdf.

Fujita:2002:SDM

[1136] S. Fujita. A semi-dynamic multiprocessor scheduling algorithm with an asymptotically optimal competitive ratio. Lecture Notes in Computer Science, 2400:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2400/24000240.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2400/24000240.pdf.

Yuan:2002:AGP

[1137] X. Yuan, C. Roig, A. Ripoll, M. A. Senar, F. Guirado, and E. Luque. AMEEDA: A general-purpose mapping tool for parallel applications on dedicated clusters (research note). Lecture Notes in Computer Science, 2400:248– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000248.
                                        pdf.

Griebl:2002:CHP

[1138] M. Griebl. Compilers for high performance (compilation and parallelization techniques). Lecture Notes in Computer Science, 2400:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000253.
pdf.

Bouchebaba:2002:TMR

[1139] Y. Bouchebaba and F. Coelho. Tiling and memory reuse for sequences of nested loops. Lecture Notes in Computer Science, 2400:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000255.
pdf.

Beyls:2002:RDB

[1140] K. Beyls and E. H. D’Hollander. Reuse distance-based cache hint selection. Lecture Notes in Computer Science, 2400:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2400/24000265.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2400/24000265.
pdf.

Martin:2002:ILP

[1141] M. J. Martín, D. E. Singh, J. Touriño, and F. F. Rivera. Improving locality in the parallelization of doacross loops (research note). Lecture Notes in Computer Science, 2400:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000275.
                                        pdf.

Thiyagalingam:2002:MLC

[1142] J. Thiyagalingam and P. H. J. Kelly. Is morton layout competitive for large two-dimensional arrays? Lecture Notes in Computer Science, 2400:280– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000280.
                                        pdf.

Arenaz:2002:TDC

[1143] M. Arenaz, J. Touriño, and R. Doallo. Towards detection of coarse-grain looplevel parallelism in irregular computations. Lecture Notes in Computer Science, 2400:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2400/24000289.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2400/24000289.pdf.

Vivien:2002:OFS

[1144] F. Vivien. On the optimality of Feautrier’s scheduling algorithm. Lecture Notes in Computer Science, 2400: 299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000299.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2400/24000299.
pdf.

Barthou:2002:ETS

[1145] D. Barthou, P. Feautrier, and X. Redon. On the equivalence of two systems of affine recurrence equations (research note). Lecture Notes in Computer Science, 2400:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000309.
pdf.

Diessel:2002:THL

[1146] O. Diessel, U. Malik, and K. So. Towards high-level specification, synthesis, and virtualization of programmable logic designs (research note). Lecture Notes in Computer Science, 2400:314– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000314.
pdf.

Kosch:2002:PDD

[1147] H. Kosch, D. Skilicorn, and D. Talia.

Parallel and distributed databases, data mining and knowledge discovery. Lecture Notes in Computer Science, 2400:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2400/24000319.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000319.
                                        pdf.

Martens:2002:DQS

[1148] H. Märtens, E. Rahm, and T. Stöhr. Dynamic query scheduling in parallel data warehouses. Lecture Notes in Computer Science, 2400:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000321.
                                        pdf.

Smith:2002:SNR

[1149] J. Smith, P. Watson, S. de F. Mendes Sampaio, and N. W. Paton. Speeding up navigational requests in a parallel object database system. Lecture Notes in Computer Science, 2400:332– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000332.
                                        pdf.

Bretschneider:2002:RMS

[1150] T. Bretschneider and O. Kao. Retrieval of multispectral satellite imagery on cluster architectures (research note). Lecture Notes in Computer Science, 2400:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2400/24000342.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2400/24000342.
pdf.

Jin:2002:SMP

[1151] R. Jin and G. Agrawal. Shared memory parallelization of decision tree construction using a general data mining middleware. Lecture Notes in Computer Science, 2400:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000346.
pdf.

Zhang:2002:CSD

[1152] Y. Zhang, A. Sivasubramaniam, J. Zhang, S. Nagar, and H. Franke. Characterizing the scalability of decision-support workloads on clusters and SMP systems. Lecture Notes in Computer Science, 2400:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000355.
pdf.

Kwok:2002:PFC

[1153] T. Kwok, K. Smith, S. Lozano, and D. Taniar. Parallel fuzzy c-means clustering for large data sets. Lecture Notes in Computer Science, 2400:365– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000365.
                                        pdf.

Orlando:2002:SHP

[1154] S. Orlando, P. Palmerini, R. Perego, and F. Silvestri. Scheduling high performance data mining tasks on a data grid environment. Lecture Notes in Computer Science, 2400:375– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000375.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000375.
                                        pdf.

Boukerche:2002:DIR

[1155] A. Boukerche and T. Tuck. A delayed-initiation risk-free multiversion temporally correct algorithm (research note). Lecture Notes in Computer Science, 2400:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000385.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000385.
                                        pdf.

Mayr:2002:CTA

[1156] E. W. Mayr. Complexity theory and algorithms. Lecture Notes in Computer Science, 2400:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000391.
pdf.

Tiskin:2002:PCH

[1157] A. Tiskin. Parallel convex hull computation by generalised regular sampling. Lecture Notes in Computer Science, 2400:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000392.
pdf.

Niculescu:2002:PAF

[1158] V. Niculescu. Parallel algorithms for fast Fourier transformation using PowerList, ParList and PList theories (research note). Lecture Notes in Computer Science, 2400:400–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000400.
pdf.

Han:2002:BBAa

[1159] J. Han, G. McMahon, and S. Sugden. A branch and bound algorithm for capacitated minimum spanning tree problem (research note). Lecture Notes in Computer Science, 2400: 404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000404.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000404.
                                        pdf.

Kumar:2002:AHP

[1160] V. Kumar, F.-J. Pfreundt, H. Burkhard, and J. Laghina Palma. Applications on high performance computers. Lecture Notes in Computer Science, 2400: 409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000409.
                                        pdf.

Plachetka:2002:PLB

[1161] T. Plachetka. Perfect load balancing for demand-driven parallel ray tracing. Lecture Notes in Computer Science, 2400:410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000410.
                                        pdf.

Lorenz:2002:PCC

[1162] U. Lorenz. Parallel controlled conspiracy number search. Lecture Notes in Computer Science, 2400:420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2400/24000420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000420.
pdf.

Svolos:2002:PST

[1163] A. I. Svolos, C. Konstantopoulos, and C. Kaklamanis. A parallel solution in texture analysis employing a massively parallel processor (research note). Lecture Notes in Computer Science, 2400: 431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000431.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000431.
pdf.

Langlais:2002:SSM

[1164] M. Langlais, G. Latu, J. Roman, and P. Silan. Stochastic simulation of a marine host-parasite system using a hybrid MPI/OpenMP programming. Lecture Notes in Computer Science, 2400: 436–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000436.
pdf.

Abdalhaq:2002:OFP

[1165] B. Abdalhaq, A. Cortés, T. Margalef, and E. Luque. Optimization of fire propagation model inputs: A grand challenge application on metacomputers (research note). Lecture Notes in Computer Science, 2400:447– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000447.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000447.
                                        pdf.

Webster:2002:PNS

[1166] S. G. Webster, M. K. Gobbert, J.- F. Remacle, and T. S. Cale. Parallel numerical solution of the Boltzmann equation for atomic layer deposition (research note). Lecture Notes in Computer Science, 2400:452– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000452.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000452.
                                        pdf.

Gaudiot:2002:PCA

[1167] J.-L. Gaudiot. Parallel computer architecture and instruction-level parallelism. Lecture Notes in Computer Science, 2400:457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2400/24000457.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2400/24000457.pdf.

Desmet:2002:IHC

[1168] V. Desmet, B. Goeman, and K. De Bosschere. Independent hashing as confidence mechanism for value predictors in microprocessors. Lecture Notes in Computer Science, 2400:458– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000458.
pdf.

Sendag:2002:EPE

[1169] R. Sendag, D. J. Lilja, and S. R.

Kunkel. Exploiting the prefetching effect provided by executing mispredicted load instructions. Lecture Notes in Computer Science, 2400:468– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000468.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000468.
pdf.

Yi:2002:IIL

[1170] J. J. Yi, R. Sendag, and D. J.

Lilja. Increasing instruction-level parallelism with instruction precomputation (research note). Lecture Notes in Computer Science, 2400:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000481.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000481.
pdf.

Chi:2002:RAS

[1171] C.-H. Chi and J. Yuan. Runtime association of software prefetch control to memory access instructions (research note). Lecture Notes in Computer Science, 2400:486–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000486.
pdf.

Uht:2002:RHI

[1172] A. Uht, A. Khalafi, D. Morano, M. de Alba, and D. Kaeli. Realizing high IPC using time-tagged resource-flow computing. Lecture Notes in Computer Science, 2400:490–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000490.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000490.
pdf.

Kailas:2002:RFA

[1173] K. Kailas, M. Franklin, and K. Ebcioğlu. A register file architecture and compilation scheme for clustered ILP processors. Lecture Notes in Computer Science, 2400:500–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000500.
pdf.

Vandierendonck:2002:CSR

[1174] H. Vandierendonck, A. Ramírez, K. De Bosschere, and M. Valero. A comparative study of redundancy in trace caches (research note). Lecture Notes in Computer Science, 2400:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000512.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000512.
pdf.

Moure:2002:STA

[1175] J. C. Moure, D. I. Rexachs, and E. Luque. Speeding up target address generation using a self-indexed FTB (research note). Lecture Notes in Computer Science, 2400:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000517.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000517.
pdf.

Paul:2002:RPP

[1176] W. J. Paul, P. Bach, M. Bosch, J. Fischer, C. Lichtenau, and J. Röhrig. Real PRAM programming. Lecture Notes in Computer Science, 2400:522– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000522.
pdf.

Trancoso:2002:MPD

[1177] P. Trancoso. In-memory parallelism for database workloads. Lecture Notes in Computer Science, 2400:532–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000532.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000532.
                                        pdf.

Sarojadevi:2002:ECC

[1178] H. Sarojadevi, S. K. Nandy, and S. Balakrishnan. Enforcing cache coherence at data sharing boundaries without global control: A hardware-software approach (research note). Lecture Notes in Computer Science, 2400:543– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000543.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000543.
                                        pdf.

Verdoscia:2002:CPD

[1179] L. Verdoscia. CODACS project: A demand-data driven reconfigurable architecture (research note). Lecture Notes in Computer Science, 2400:547– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000547.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000547.
                                        pdf.

Mavronicolas:2002:DSA

[1180] M. Mavronicolas and A. Schiper. Distributed systems and algorithms. Lecture Notes in Computer Science, 2400: 551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000551.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000551.
pdf.

Datta:2002:SST

[1181] A. K. Datta, R. Hadid, and V. Villain. A self-stabilizing token-based kout-of-ℓ exclusion algorithm. Lecture Notes in Computer Science, 2400:553– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000553.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000553.
pdf.

Ruiz:2002:AEF

[1182] D. Ruiz, R. Corchuelo, J. A. Pérez, and M. Toro. An algorithm for ensuring fairness and liveness in nondeterministic systems based on multiparty interactions. Lecture Notes in Computer Science, 2400:563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000563.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000563.
pdf.

Jelasity:2002:OGI

[1183] M. Jelasity and M. Preuß. On obtaining global information in a peer-to-peer fully distributed environment (research note). Lecture Notes in Computer Science, 2400:573–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000573.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000573.
                                        pdf.

Baldoni:2002:FTS

[1184] R. Baldoni, C. Marchetti, and S. Tucci Piergiovanni. A faulttolerant sequencer for timed asynchronous systems. Lecture Notes in Computer Science, 2400:578–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000578.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000578.
                                        pdf.

Gallard:2002:DRM

[1185] P. Gallard, C. Morin, and R. Lottiaux. Dynamic resource management in a cluster for high-availability (research note). Lecture Notes in Computer Science, 2400:589–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000589.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2400/24000589.
pdf.

Renault:2002:PIS

[1186] É. Renault and D. Millot. Progressive introduction of security in remotewrite communications with no performance sacrifice (research note). Lecture Notes in Computer Science, 2400: 593–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000593.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000593.
pdf.

Suppi:2002:PDP

[1187] R. Suppi, M. Solsona, and E. Luque.

Parasite: Distributing processing using Java applets (research note). Lecture Notes in Computer Science, 2400: 598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000598.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000598.
pdf.

Hammond:2002:PPM

[1188] K. Hammond. Parallel programming: Models, methods and programming languages. Lecture Notes in Computer Science, 2400:603–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000603.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2400/24000603.
                                        pdf.

Bouge:2002:IRE

[1189] L. Bougé, V. Danjean, and R. Namyst. Improving reactivity to I/O events in multithreaded environments using a uniform, scheduler-centric API. Lecture Notes in Computer Science, 2400: 605–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000605.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000605.
                                        pdf.

Hawkins:2002:OSD

[1190] J. Hawkins and A. E. Abdallah. An overview of systematic development of parallel systems for reconfigurable hardware (research note). Lecture Notes in Computer Science, 2400:615– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000615.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000615.
                                        pdf.

Kuchen:2002:SL

[1191] H. Kuchen. A skeleton library. Lecture Notes in Computer Science, 2400: 620–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000620.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2400/24000620.
pdf.

Field:2002:OSR

[1192] A. J. Field, P. H. J. Kelly, and T. L.

Hansen. Optimising shared reduction variables in MPI programs. Lecture Notes in Computer Science, 2400:630– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000630.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000630.
pdf.

Bischof:2002:DSI

[1193] H. Bischof and S. Gorlatch. Doublescan: Introducing and implementing a new data-parallel skeleton. Lecture Notes in Computer Science, 2400:640– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000640.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000640.
pdf.

Pedicini:2002:SVC

[1194] M. Pedicini and F. Quaglia. Scheduling vs communication in PELCR. Lecture Notes in Computer Science, 2400: 648–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000648.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000648.
pdf.

Keen:2002:EHD

[1195] A. W. Keen and R. A. Olsson. Exception handling during asynchronous method invocation (research note). Lecture Notes in Computer Science, 2400:656–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000656.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000656.
pdf.

Sobral:2002:DSO

[1196] J. L. Sobral and A. J. Proença. Designing scalable object oriented parallel applications (research note). Lecture Notes in Computer Science, 2400: 661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000661.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000661.
pdf.

Liniker:2002:DES

[1197] P. Liniker, O. Beckmann, and P. H. J. Kelly. Delayed evaluation, selfoptimising software components as a programming model. Lecture Notes in Computer Science, 2400:666–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000666.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000666.
pdf.

Duff:2002:NA

[1198] I. S. Duff, W. Borchers, L. Giraud, and H. A. van der Vorst. Numerical algorithms. Lecture Notes in Computer Science, 2400:675–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000675.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000675.
pdf.

daCunha:2002:NPR

[1199] R. Dias da Cunha, D. Becker, and J. Carlton Patterson. New parallel (rank-revealing) QR factorization algorithms. Lecture Notes in Computer Science, 2400:677–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000677.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000677.
pdf.

Badia:2002:SLS

[1200] J. M. Badía, P. Benner, R. Mayo, and E. S. Quintana-Ortí. Solving large sparse Lyapunov equations on parallel computers (research note). Lecture Notes in Computer Science, 2400:687– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000687.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000687.
pdf.

Takahashi:2002:BAP

[1201] D. Takahashi, T. Boku, and M. Sato. A blocking algorithm for parallel 1-D FFT on clusters of PCs. Lecture Notes in Computer Science, 2400:691– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000691.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000691.
                                        pdf.

Buijssen:2002:SPI

[1202] S. H. M. Buijssen and S. Turek. Sources of parallel inefficiency for incompressible CFD simulations (research note). Lecture Notes in Computer Science, 2400:701–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2400/24000701.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000701.
                                        pdf.

Graham:2002:PIM

[1203] I. G. Graham, A. Spence, and E. Vainikko. Parallel iterative methods for Navier–Stokes equations and application to stability assessment (distinguished paper). Lecture Notes in Computer Science, 2400:705–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000705.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000705.
                                        pdf.

Boufflet:2002:MDP

[1204] J.-P. Boufflet, P. Breitkopf, A. Rassineux, and P. Villon. A modular design for a parallel multifrontal mesh generator. Lecture Notes in Computer Science, 2400:715–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000715.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000715.
pdf.

Korch:2002:PLI

[1205] M. Korch, T. Rauber, and G. Rünger.

Pipelining for locality improvement in RK methods. Lecture Notes in Computer Science, 2400:724–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000724.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000724.
pdf.

Flammini:2002:RCI

[1206] M. Flammini, B. Maggs, J. Sibeyn, and B. Vöcking. Routing and communication in interconnection networks. Lecture Notes in Computer Science, 2400: 735–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000735.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000735.
pdf.

Bang:2002:MMC

[1207] Y.-C. Bang and H. Choo. On multicasting with minimum costs for the Internet topology. Lecture Notes in Computer Science, 2400:736–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000736.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000736.
                                        pdf.

Jin:2002:SOU

[1208] H.-W. Jin, C. Yoo, and S.-K. Park. Stepwise optimizations of UDP/IP on a gigabit network (research note). Lecture Notes in Computer Science, 2400: 745–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000745.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000745.
                                        pdf.

Chen:2002:SID

[1209] Y. Chen, A. K. Datta, and S. Tixeuil. Stabilizing inter-domain routing in the Internet (research note). Lecture Notes in Computer Science, 2400:749– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000749.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000749.
                                        pdf.

Jegou:2002:PAC

[1210] Y. Jégou. Performance analysis of code coupling on long distance high bandwidth network (research note). Lecture Notes in Computer Science, 2400:753– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000753.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000753.
pdf.

Wang:2002:APB

[1211] C.-M. Wang, Y. Hou, and L.-H.

Hsu. Adaptive path-based multicast on wormhole-routed hypercubes. Lecture Notes in Computer Science, 2400: 757–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000757.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000757.
pdf.

Barth:2002:MDC

[1212] D. Barth, P. Berthomé, T. Czarchoski, J. M. Fourneau, C. Laforest, and S. Vial. A mixed deflection and convergence routing algorithm: Design and performance. Lecture Notes in Computer Science, 2400: 767–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000767.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000767.
pdf.

Gomez:2002:ERA

[1213] M. E. Gómez, J. Flich, A. Robles, P. López, and J. Duato. Evaluation of routing algorithms for Infini-Band networks (research note). Lecture Notes in Computer Science, 2400: 775–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000775.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000775.
                                        pdf.

Baydal:2002:CCB

[1214] E. Baydal, P. López, and J. Duato. Congestion control based on transmission times. Lecture Notes in Computer Science, 2400:781–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000781.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000781.
                                        pdf.

Park:2002:DLT

[1215] Jihoon Park, Jonggyu Park, Ilsuk Han, and Hagbae Kim. A dual-LAN topology with the dual-path Ethernet module (research note). Lecture Notes in Computer Science, 2400:791– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000791.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000791.
                                        pdf.

Fujita:2002:FBS

[1216] S. Fujita and S. Tagashira. A fast barrier synchronization protocol for broadcast networks based on a dynamic access control (research note). Lecture Notes in Computer Science, 2400:795– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000795.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000795.
pdf.

Sanders:2002:HFA

[1217] P. Sanders and J. Larsson Träff. The hierarchical factor algorithm for all-to-all communication (research note). Lecture Notes in Computer Science, 2400: 799–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000799.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000799.
pdf.

Uhl:2002:AAM

[1218] A. Uhl. Architectures and algorithms for multimedia applications. Lecture Notes in Computer Science, 2400:805– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000805.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000805.
pdf.

Mourlas:2002:DSC

[1219] C. Mourlas. Deterministic scheduling of CBR and VBR media flows on parallel media servers. Lecture Notes in Computer Science, 2400:807– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000807.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000807.
pdf.

Cores:2002:DPT

[1220] F. Cores, A. Ripoll, and E. Luque. Double P-tree: A distributed architecture for large-scale video-on-demand. Lecture Notes in Computer Science, 2400:816–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2400/24000816.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2400/24000816.
pdf.

Polak:2002:MPX

[1221] S. Polak, R. Slota, and J. Kitowski. Message passing in XML-based language for creating multimedia presentations (research note). Lecture Notes in Computer Science, 2400:826– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000826.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000826.
pdf.

Fernandez:2002:PIH

[1222] J. C. Fernández and M. P. Malumbres.

A parallel implementation of H.26L video encoder (research note). Lecture Notes in Computer Science, 2400:830– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000830.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000830.
pdf.

Paar:2002:NPS

[1223] A. Paar, M. L. Anido, and N. Bagherzadeh.

A novel predication scheme for a SIMD system-on-chip. Lecture Notes in Computer Science, 2400:834–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000834.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000834.
pdf.

Parizi:2002:MCG

[1224] H. Parizi, A. Niktash, N. Bagherzadeh, and F. Kurdahi. MorphoSys: A coarse grain reconfigurable architecture for multimedia applications (research note). Lecture Notes in Computer Science, 2400:844–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000844.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000844.
pdf.

Cheresiz:2002:PSM

[1225] D. Cheresiz, B. Juurlink, S. Vassiliadis, and H. Wijshoff. Performance scalability of multimedia instruction set extensions. Lecture Notes in Computer Science, 2400:849–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000849.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000849.
                                        pdf.

Cosnard:2002:MGC

[1226] M. Cosnard and A. Merzky. Metaand grid-computing. Lecture Notes in Computer Science, 2400:861–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000861.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000861.
                                        pdf.

Kelly:2002:IAC

[1227] P. H. J. Kelly, S. Pelagatti, and M. Rossiter. Instant-access cyclestealing for parallel applications requiring interactive response. Lecture Notes in Computer Science, 2400:863– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000863.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000863.
                                        pdf.

Nikolow:2002:ATE

[1228] D. Nikolow, R. Słota, M. Dziewierz, and J. Kitowski. Access time estimation for tertiary storage systems. Lecture Notes in Computer Science, 2400: 873–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000873.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000873.
pdf.

Pytlinski:2002:BUP

[1229] J. Pytliński, Ł. Skorwider, P. Bała, M. Nazaruk, and K. Wawruch. BioGRID — uniform platform for biomoleculago nazar applications (research note). Lecture Notes in Computer Science, 2400:881–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2400/24000881.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2400/24000881.pdf.

Stanton:2002:ISV

[1230] J. Stanton, S. Newhouse, and J. Darlington. Implementing a scientific visualisation capability within a grid enabled component framework (research note). Lecture Notes in Computer Science, 2400:885–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000885.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2400/24000885.
                                        pdf.

Dialani:2002:TFT

[1231] V. Dialani, S. Miles, L. Moreau, D. De Roure, and M. Luck. Transparent fault tolerance for Web services based architectures. Lecture Notes in Computer Science, 2400:889– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000889.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000889.
                                        pdf.

Alt:2002:ADP

[1232] M. Alt, H. Bischof, and S. Gorlatch. Algorithm design and performance prediction in a Java-based grid system with skeletons. Lecture Notes in Computer Science, 2400:899– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000899.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000899.
                                        pdf.

Caron:2002:SAN

[1233] E. Caron, F. Desprez, F. Lombard, J.-M. Nicod, L. Philippe, M. Quinson, and F. Suter. A scalable approach to network enabled servers (research note). Lecture Notes in Computer Science, 2400:907–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2400/24000907.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000907.
pdf.

Feldmann:2002:DO

[1234] R. Feldmann and C. Roucairol. Discrete optimization. Lecture Notes in Computer Science, 2400:911–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000911.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000911.
pdf.

Gebremedhin:2002:PDC

[1235] A. H. Gebremedhin, F. Manne, and A. Pothen. Parallel distancek coloring algorithms for numerical optimization. Lecture Notes in Computer Science, 2400:912–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000912.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000912.
pdf.

Ribeiro:2002:PGH

[1236] C. C. Ribeiro and I. Rosseti.

A parallel GRASP heuristic for the 2-path network design problem (research note). Lecture Notes in Computer Science, 2400:922–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000922.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000922.
                                        pdf.

Alba:2002:MLS

[1237] E. Alba, F. Almeida, M. Blesa, J. Cabeza, C. Cotta, M. Díaz, I. Dorta, J. Gabarró, C. León, J. Luna, L. Moreno, C. Pablos, J. Petit, A. Rojas, and F. Xhafa. MALLBA: A library of skeletons for combinatorial optimisation (research note). Lecture Notes in Computer Science, 2400:927– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000927.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000927.
                                        pdf.

aufderHeide:2002:MCM

[1238] F. Meyer auf der Heide, M. Kumar, S. Nikoletseas, and P. Spirakis. Mobile computing, mobile networks. Lecture Notes in Computer Science, 2400: 933–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2400/24000933.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2400/24000933.
                                        pdf.

Grunewald:2002:DMR

[1239] M. Grünewald, T. Lukovszki, C. Schindelhauer, and K. Volbert. Distributed maintenance of resource efficient wireless network topologies (distinguished paper). Lecture Notes in Computer Science, 2400:935–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000935.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000935.
pdf.

Clematis:2002:LDA

[1240] A. Clematis, D. D’Agostino, and V. Gianuzzi. A local decision algorithm for maximum lifetime in ad hoc networks. Lecture Notes in Computer Science, 2400:947–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000947.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000947.
pdf.

Boukerche:2002:PSD

[1241] A. Boukerche, J. Linus, and A. Saurabha.

A performance study of distance source routing based protocols for mobile and wireless ad hoc networks. Lecture Notes in Computer Science, 2400:957– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000957.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000957.
pdf.

Jurdzinski:2002:WCR

[1242] T. Jurdziński, M. Kutylowski, and J. Zatopiański. Weak communication in radio networks. Lecture Notes in Computer Science, 2400:965–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000965.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000965.
pdf.

Zaini:2002:CMI

[1243] N. Zaini and L. Moreau. Coordination of mobile intermediaries acting on behalf of mobile users (research note). Lecture Notes in Computer Science, 2400:973–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000973.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000973.
pdf.

Lin:2002:ETB

[1244] C.-Y. Lin, S.-C. Wang, and S.-Y. Kuo. An efficient time-based checkpointing protocol for mobile computing systems over wide area networks (research note). Lecture Notes in Computer Science, 2400:978–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000978.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000978.
pdf.

Hwang:2002:DCR

[1245] S.-H. Hwang and K.-J. Han. Discriminative collision resolution algorithm for wireless MAC protocol (research note). Lecture Notes in Computer Science, 2400:983–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2400/24000983.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2400/24000983.
pdf.

Anonymous:2002:AIba

[1246] Anonymous. Author index. Lecture Notes in Computer Science, 2400:989– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2400/2400auth.pdf.

Bae:2002:NSA

[1247] Sungil Bae, Daesik Seo, Gilyoung Kang, and Sungho Kang. A new survival architecture for network processors. Lecture Notes in Computer Science, 2402:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020001.
pdf.

White:2002:MTU

[1248] Allan L. White. Mean time to unsafe failure. Lecture Notes in Computer Science, 2402:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020011.
pdf.

Kim:2002:PAA

[1249] Youngseok Kim, Seunghoon Hwang, Dongkyoon Cho, and Keumchan Whang. Performance of antenna arrays with reverse-link synchronous transmission technique for DS-CDMA system in multipath fading channels. Lecture Notes in Computer Science, 2402: 21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020021.
pdf.

Kim:2002:NRA

[1250] Youngyong Kim. A noble routing algorithm for the Internet computational GRID. Lecture Notes in Computer Science, 2402:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020033.
pdf.

Lee:2002:GRA

[1251] Sangmoon Lee, Daekeun Moon, Hagbae Kim, and Whie Chang. A genetic routing algorithm for a 2D-meshed fault-tolerant network system. Lecture Notes in Computer Science, 2402:39– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020039.
pdf.

Na:2002:APD

[1252] Songun Na and Seungwha Yoo. Allowable propagation delay for VoIP calls of acceptable quality. Lecture Notes in Computer Science, 2402:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020047.
pdf.

Lee:2002:RNS

[1253] Dongho Lee, Jungsup Kim, Sooduk Kim, Kichang Kim, and Jaehyun Park. Robustness of a neighbor selection Markov chain in prefetching tiled Web data. Lecture Notes in Computer Science, 2402:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020057.
pdf.

Shankar:2002:QSP

[1254] Sai Shankar and Sunghyun Choi. QoS signaling for parameterized traffic in IEEE 802.11e wireless LANs. Lecture Notes in Computer Science, 2402:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020067.
                                        pdf.

Kim:2002:CAR

[1255] Hagyoung Kim, Sungin Jung, and Cheolhoon Lee. Content-adaptive request distribution strategy for Internet server cluster. Lecture Notes in Computer Science, 2402:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020085.
                                        pdf.

Yoon:2002:NDA

[1256] Sungchul Yoon, Sangwook Kim, Jaeseuk Oh, and Sungho Kang. A new DSP architecture for correcting errors using Viterbi algorithm. Lecture Notes in Computer Science, 2402:95– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020095.
                                        pdf.

Kang:2002:AWR

[1257] Moonsik Kang, Junho Lee, and Sangmin Lee. An adapting weight rerouting algorithm for handoff control in wireless ATM networks. Lecture Notes in Computer Science, 2402:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020103.
pdf.

Hong:2002:OBS

[1258] Taechul Hong and Youngyong Kim.

Optimal buffering strategy for streaming service in time varying wireless environment. Lecture Notes in Computer Science, 2402:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020115.
pdf.

Lee:2002:OSR

[1259] Heeyong Lee, Gilyoung Kong, Sihwa Kim, Cholseong Kim, and Jaechul Lee. Optimum ship routing and it’s implementation on the Web. Lecture Notes in Computer Science, 2402:125– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020125.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2402/24020125.
                                        pdf.

Park:2002:DRG

[1260] Jongkyu Park, Ilseok Han, Jinhyuck Kwon, Jun Hwang, Hagbae Kim, Sangtae Ahn, and Whie Chang. Development of a residential gateway and a service server for home automation. Lecture Notes in Computer Science, 2402: 137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020137.
                                        pdf.

Pak:2002:PMT

[1261] Uiyoung Pak, Joonseok Maeng, Iksoo Jin, and Keumchan Whang. Performance of multiple TCM codes with transmit antenna diversity. Lecture Notes in Computer Science, 2402:151– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020151.
                                        pdf.

Kim:2002:SML

[1262] Sangyoub Kim. On the study of MAC layer for cdma2000 wireless packet data service. Lecture Notes in Computer Science, 2402:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2402/24020161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020161.
pdf.

Jeon:2002:ITO

[1263] Taewoong Jeon, Sungyoung Lee, and Hyonwoo Seung. Increasing the testability of object-oriented frameworks with built-in tests. Lecture Notes in Computer Science, 2402:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020169.
pdf.

Jung:2002:GCS

[1264] Sungin Jung, Hagyoung Kim, and Cheolhoon Lee. Guaranteeing the continuous stream service in cluster media servers. Lecture Notes in Computer Science, 2402:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020183.
pdf.

Lee:2002:SBF

[1265] Changho Lee, Jaehyun Park, and Yoosung Kim. A SOAP-based framework for the Internetworked distributed control systems. Lecture Notes in Computer Science, 2402:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020195.
                                        pdf.

Bae:2002:ELM

[1266] Sangmin Bae, Dongsup Song, Jihye Kim, and Sungho Kang. An efficient on-line monitoring BIST for remote service system. Lecture Notes in Computer Science, 2402:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020205.
                                        pdf.

Roh:2002:TPE

[1267] Inho Roh and Youngyong Kim. TCP performance enhancement using FG-LIW (fairness guaranteed — larger initial window) scheme in the wired-cum-wireless network environment. Lecture Notes in Computer Science, 2402:215– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020215.
                                        pdf.

Yun:2002:CSP

[1268] Yongchul Yun, Jinwoo Choe, and Wonjin Sung. A comparative study on the performance of detectors and interference models for OCDMA LANs. Lecture Notes in Computer Science, 2402: 223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020223.
pdf.

Baek:2002:TMO

[1269] Yuncheol Baek. A tool measuring operating system supports for squid proxy cache. Lecture Notes in Computer Science, 2402:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020239.
pdf.

Jeon:2002:DIO

[1270] Taewoong Jeon, Sunghwan Roh, Hyonwoo Seung, and Sungyoung Lee. The design and implementation of an object-oriented process control loop framework. Lecture Notes in Computer Science, 2402:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020249.
pdf.

Lee:2002:PIL

[1271] Sangmoon Lee, Sinjun Kang, Byungseok Min, and Hagbae Kim. A personal Internet live-broadcasting system. Lecture Notes in Computer Science, 2402: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020263.
                                        pdf.

Kong:2002:DIS

[1272] Gilyoung Kong, Chungro Lee, Soonkap Kim, Gilsoo Kim, and Cholseong Kim. On the development of an information service system for navigational safety of seagoing vessels. Lecture Notes in Computer Science, 2402:273– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020273.
                                        pdf.

Kim:2002:OCW

[1273] Taewon Kim, Wonjin Sung, and Jinwoo Choe. Optimal code weight selection for 2-D optical CDMA systems. Lecture Notes in Computer Science, 2402: 285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2402/24020285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2402/24020285.
                                        pdf.

Lee:2002:RCI

[1274] Sangyeob Lee and Whoiyul Kim. Robust character image retrieval method using bipartite matching and pseudobipartite matching. Lecture Notes in Computer Science, 2402:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2402/24020295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2402/24020295.
pdf.

Anonymous:2002:AIbb

[1275] Anonymous. Author index. Lecture Notes in Computer Science, 2402:307– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2402/2402auth.pdf.

Decker:2002:CIA

[1276] Keith Decker. Coordinating intelligent agents. Lecture Notes in Computer Science, 2403:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2403/24030001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2403/24030001.
pdf.

Durfee:2002:SDC

[1277] Edmund H. Durfee. Strategies for discovering coordination needs in MultiAgent systems. Lecture Notes in Computer Science, 2403:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2403/24030019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030019.
                                        pdf.

Sierra:2002:AMI

[1278] Carles Sierra and Pablo Noriega. Agent-mediated interaction. from auctions to negotiation and argumentation. Lecture Notes in Computer Science, 2403:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2403/24030027.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2403/24030027.pdf.

Tennenholtz:2002:GTA

[1279] Moshe Tennenholtz. Game theory and artificial intelligence. Lecture Notes in Computer Science, 2403:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2403/24030049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030049.
                                        pdf.

Alonso:2002:RMA

[1280] Eduardo Alonso. Rights for multiagent systems. Lecture Notes in Computer Science, 2403:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2403/24030059.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2403/24030059.
pdf.

Ashri:2002:ISA

[1281] Ronald Ashri, Michael Luck, and Mark d’Inverno. Infrastructure support for agent-based development. Lecture Notes in Computer Science, 2403:73– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2403/24030073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2403/24030073.
pdf.

Bordini:2002:AAD

[1282] Rafael H. Bordini, Renata Vieira, and John A. Campbell. An anthropological approach to the discovery of ontologies in multi-agent societies. Lecture Notes in Computer Science, 2403: 89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2403/24030089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2403/24030089.
pdf.

Buckle:2002:SMA

[1283] Phil Buckle, Tom Moore, Steve Robertshaw, Alan Treadway, Sasu Tarkoma, and Stefan Poslad. Scalability in multi-agent systems: The FIPA-OS perspective. Lecture Notes in Computer Science, 2403:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2403/24030110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030110.
                                        pdf.

Doran:2002:AS

[1284] Jim Doran. Agents and MAS in STaMs. Lecture Notes in Computer Science, 2403:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2403/24030131.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030131.
                                        pdf.

vanEijk:2002:SAC

[1285] Rogier M. van Eijk. Semantics of agent communication: An introduction. Lecture Notes in Computer Science, 2403: 152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2403/24030152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030152.
                                        pdf.

Fisher:2002:ABT

[1286] Michael Fisher and Chiara Ghidini. Agents with bounded temporal resources. Lecture Notes in Computer Science, 2403:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2403/24030169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030169.
                                        pdf.

Norman:2002:MDM

[1287] Timothy J. Norman and Chris Reed. A model of delegation for multi-agent systems. Lecture Notes in Computer Science, 2403:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2403/24030185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2403/24030185.
pdf.

Parsons:2002:ASU

[1288] Simon Parsons, Nicholas R. Jennings, Jordi Sabater, and Carles Sierra. Agent specification using multi-context systems. Lecture Notes in Computer Science, 2403:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2403/24030205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2403/24030205.
pdf.

Preist:2002:ACM

[1289] Chris Preist and Siani Pearson. An adaptive choice of messaging protocol in multi agent systems. Lecture Notes in Computer Science, 2403:227– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2403/24030227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2403/24030227.
pdf.

Schut:2002:POM

[1290] Martijn Schut, Michael Wooldridge, and Simon Parsons. On partially observable MDPs and BDI models. Lecture Notes in Computer Science, 2403: 243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2403/24030243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2403/24030243.
                                        pdf.

Anonymous:2002:AIbc

[1291] Anonymous. Author index. Lecture Notes in Computer Science, 2403:261– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2403/2403auth.pdf.

Hull:2002:IYW

[1292] Richard Hull, Bharat Kumar, Arnaud Sahuguet, and Ming Xiong. Have it your way: Personalization of network-hosted services. Lecture Notes in Computer Science, 2405:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2405/24050001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2405/24050001.
                                        pdf.

Gounaris:2002:AQP

[1293] Anastasios Gounaris, Norman W. Paton, Alvaro A. A. Fernandes, and Rizos Sakellariou. Adaptive query processing: A survey. Lecture Notes in Computer Science, 2405:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050011.
pdf.

Giannella:2002:IQE

[1294] Chris M. Giannella, Mehmet M. Dalkilic, Dennis P. Groth, and Edward L. Robertson. Improving query evaluation with approximate functional dependency based decompositions. Lecture Notes in Computer Science, 2405: 26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050026.
pdf.

Boyd:2002:ASI

[1295] Michael Boyd, Peter McBrien, and Nerissa Tong. The AutoMed schema integration repository. Lecture Notes in Computer Science, 2405:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050042.
pdf.

Jasper:2002:GQP

[1296] Edgar Jasper. Global query processing in the AutoMed heterogeneous database environment. Lecture Notes in Computer Science, 2405:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050046.
pdf.

Fan:2002:TDL

[1297] Hao Fan. Tracing data lineage using automed schema transformation pathways. Lecture Notes in Computer Science, 2405:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050050.
pdf.

Gupta:2002:SMA

[1298] Amarnath Gupta, Bertram Ludäscher, Maryann E. Martone, Xufei Qian, Edward Ross, Joshua Tran, and Ilya Zaslavsky. A system for managing alternate models in modelbased mediation. Lecture Notes in Computer Science, 2405:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050054.
pdf.

Ram:2002:CMB

[1299] Sudha Ram, Jinsoo Park, and Yousub Hwang. CREAM: A mediator based environment for modeling and accessing distributed information on the Web. Lecture Notes in Computer Science, 2405:58– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050058.
pdf.

Shou:2002:IAH

[1300] Xiao Mang Shou and Siobhán North.

An integrated approach to handling collaborative diagram databases on the WWW. Lecture Notes in Computer Science, 2405:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050062.
pdf.

Medina:2002:HCM

[1301] Enrique Medina, Sergio Luján-Mora, and Juan Trujillo. Handling conceptual multidimensional models using XML through DTDs. Lecture Notes in Computer Science, 2405:66– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050066.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2405/24050066.
                                        pdf.

Lepinioti:2002:IDM

[1302] Konstantina Lepinioti and Stephen McKearney. Implementing data mining in a DBMS. Lecture Notes in Computer Science, 2405:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2405/24050070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2405/24050070.
                                        pdf.

Lodi:2002:FDC

[1303] Stefano Lodi. Fully dynamic clustering of metric data sets. Lecture Notes in Computer Science, 2405:73– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2405/24050073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2405/24050073.
                                        pdf.

Veloso:2002:RWA

[1304] Adriano Veloso, Bruno Rocha, Márcio de Carvalho, and Wagner Meira Jr. Real world association rule mining. Lecture Notes in Computer Science, 2405:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2405/24050077.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2405/24050077.
                                        pdf.

Engstrom:2002:ICE

[1305] Henrik Engström, Sharma Chakravarthy, and Brian Lings. Implementation and comparative evaluation of maintenance policies in a data warehouse environment. Lecture Notes in Computer Science, 2405:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050090.
pdf.

Burgess:2002:ETQ

[1306] Mikhaila Burgess, W. Alex Gray, and Nick Fiddian. Establishing a taxonomy of quality for use in information filtering. Lecture Notes in Computer Science, 2405:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050103.
pdf.

Couchot:2002:IRT

[1307] Alain Couchot. Improving the refined triggering graph method for active rules termination analysis. Lecture Notes in Computer Science, 2405: 114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050114.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2405/24050114.
                                        pdf.

Fu:2002:ADR

[1308] Gaihua Fu, Jianhua Shao, Suzanne M. Embury, and W. Alex Gray. An algorithm for determining related constraints. Lecture Notes in Computer Science, 2405:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2405/24050134.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2405/24050134.pdf.

vanZwol:2002:RPE

[1309] Roelof van Zwol and Peter M. G. Apers. Retrieval performance experiment with the Webspace method. Lecture Notes in Computer Science, 2405: 150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2405/24050150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2405/24050150.
                                        pdf.

Theodoratos:2002:SIQ

[1310] Dimitri Theodoratos. Semantic integration and querying of heterogeneous data sources using a hypergraph data model. Lecture Notes in Computer Science, 2405:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2405/24050166.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2405/24050166.
pdf.

Brisaboa:2002:DDQ

[1311] Nieves R. Brisaboa, Miguel R. Penabad, Ángeles S. Places, and Francisco J. Rodríguez. A document database query language. Lecture Notes in Computer Science, 2405:183– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2405/24050183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2405/24050183.
pdf.

Anonymous:2002:AIbd

[1312] Anonymous. Author index. Lecture Notes in Computer Science, 2405:199– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2405/24050199.pdf.

Peters:2002:I

[1313] Carol Peters. Introduction. Lecture Notes in Computer Science, 2406:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060001.
pdf.

Braschler:2002:COR

[1314] Martin Braschler. CLEF 2001 — overview of results. Lecture Notes in Computer Science, 2406:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060009.
pdf.

Savoy:2002:RCE

[1315] Jacques Savoy. Report on CLEF-2001 experiments: Effective combined query-translation approach. Lecture Notes in Computer Science, 2406:27– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060027.
pdf.

Chen:2002:MIR

[1316] Aitao Chen. Multilingual information retrieval using English and Chinese queries. Lecture Notes in Computer Science, 2406:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060044.
pdf.

Jones:2002:ECE

[1317] Gareth J. F. Jones and Adenike M. Lam-Adesina. Exeter at CLEF 2001: Experiments with machine translation for bilingual retrieval. Lecture Notes in Computer Science, 2406:59– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060059.
pdf.

Kraaij:2002:TCC

[1318] Wessel Kraaij. TNO at CLEF-2001:

Comparing translation resources. Lecture Notes in Computer Science, 2406: 78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060078.
pdf.

Bertoldi:2002:IIC

[1319] Nicola Bertoldi and Marcello Federico.

ITC-irst at CLEF 2001: Monolingual and bilingual tracks. Lecture Notes in Computer Science, 2406:94– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060094.
pdf.

Braschler:2002:EER

[1320] Martin Braschler, Bärbel Ripplinger, and Peter Schäuble. Experiments with the eurospider retrieval system for CLEF 2001. Lecture Notes in Computer Science, 2406:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060102.
                                        pdf.

Lin:2002:UCO

[1321] Wen-Cheng Lin and Hsin-Hsi Chen. Using co-occurrence, augmented restrictions, and C-E WordNet for Chinese–English cross-language information retrieval at CLEF 2001. Lecture Notes in Computer Science, 2406: 111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060111.
                                        pdf.

Hedlund:2002:UCE

[1322] Turid Hedlund, Heikki Keskustalo, Ari Pirkola, Eija Airio, and Kalervo Järvelin. Utaclir @ CLEF 2001 — effects of compound splitting and N-gram techniques. Lecture Notes in Computer Science, 2406:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060118.
                                        pdf.

Nie:2002:UST

[1323] Jian-Yun Nie and Michel Simard. Using statistical translation models for bilingual IR. Lecture Notes in Computer Science, 2406:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060137.
pdf.

Rogati:2002:CLP

[1324] Monica Rogati and Yiming Yang.

Cross-lingual pseudo-relevance feedback using a comparable corpus. Lecture Notes in Computer Science, 2406: 151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060151.
pdf.

Boughanem:2002:IDC

[1325] M. Boughanem, C. Chrisment, and N. Nassr. Investigation on disambiguation in CLIR: Aligned corpus and bi-directional translationbased strategies. Lecture Notes in Computer Science, 2406:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060158.
pdf.

Sahlgren:2002:VBS

[1326] Magnus Sahlgren and Jussi Karlgren. Vector-based semantic analysis using random indexing for crosslingual query expansion. Lecture Notes in Computer Science, 2406:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060169.
pdf.

Sadat:2002:QET

[1327] Fatiha Sadat, Akira Maeda, Masatoshi Yoshikawa, and Shunsuke Uemura. Query expansion techniques for the CLEF bilingual track. Lecture Notes in Computer Science, 2406:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060177.
pdf.

Martinez-Santiago:2002:IIA

[1328] Fernando Martínez-Santiago, L. Alfonso Ureña-López, M. Carlos Díaz-Galiano, Maite Martín-Valdivia, and Manuel García-Vega. Intelligent information access systems (SINAI) at CLEF 2001: Calculating translation probabilities with SemCor. Lecture Notes in Computer Science, 2406:185– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060185.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2406/24060185.
pdf.

McNamee:2002:JAE

[1329] Paul McNamee and James Mayfield. JHU/APL experiments at CLEF: Translation resources and score normalization. Lecture Notes in Computer Science, 2406:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060193.
pdf.

Chuleerat:2002:DBT

[1330] Jaruskulchai Chuleerat. Dictionarybased Thai CLIR: An experimental survey of Thai CLIR. Lecture Notes in Computer Science, 2406:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060209.
pdf.

Adriani:2002:EDC

[1331] Mirna Adriani. English-dutch CLIR using query translation techniques. Lecture Notes in Computer Science, 2406: 219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060219.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2406/24060219.
pdf.

Molina-Salgado:2002:TLR

[1332] Hugo Molina-Salgado, Isabelle Moulinier, Mark Knudson, Elizabeth Lund, and Kirat Sekhon. Thomson legal and regulatory at CLEF 2001: Monolingual and bilingual experiments. Lecture Notes in Computer Science, 2406: 226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060226.
pdf.

Gey:2002:WRQ

[1333] Fredric C. Gey, Hailing Jiang, and Natalia Perelman. Working with Russian queries for the GIRT, bilingual, and multilingual CLEF tasks. Lecture Notes in Computer Science, 2406:235– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060235.
pdf.

Llopis:2002:IPR

[1334] Fernando Llopis and José L. Vicedo. IR-n: A passage retrieval system at CLEF-2001. Lecture Notes in Computer Science, 2406:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2406/24060244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060244.
pdf.

Figuerola:2002:SMT

[1335] Carlos G. Figuerola, Raquel Gómez, Angel F. Zazo Rodríguez, and José Luis Alonso Berrocal. Spanish monolingual track: The impact of stemming on retrieval. Lecture Notes in Computer Science, 2406:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060253.
pdf.

Monz:2002:SMA

[1336] Christof Monz and Maarten de Rijke. Shallow morphological analysis in monolingual information retrieval for dutch, german, and italian. Lecture Notes in Computer Science, 2406:262– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060262.
pdf.

Tomlinson:2002:SEL

[1337] Stephen Tomlinson. Stemming evaluated in 6 languages by hummingbird SearchServer^{T M} at CLEF 2001. Lecture Notes in Computer Science, 2406: 278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060278.
                                        pdf.

Frizzarin:2002:MTR

[1338] Teresita Frizzarin and Annius Groenink. Minimalistic test runs of the eidetica indexer. Lecture Notes in Computer Science, 2406:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060288.
                                        pdf.

Matoyo:2002:ABC

[1339] Eugenia Matoyo and Tony Valsamidis. Across the bridge: CLEF 2001 — non-English monolingual retrieval. the French task. Lecture Notes in Computer Science, 2406:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060291.
                                        pdf.

Ripplinger:2002:MIC

[1340] Bärbel Ripplinger. Mpro-IR in CLEF 2001. Lecture Notes in Computer Science, 2406:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/2406/24060300.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2406/24060300.pdf.

Klok:2002:STM

[1341] Jakob Klok, Samuel Driessen, and Marvin Brunner. Some terms are more interchangeable than others. Lecture Notes in Computer Science, 2406:303– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060303.
pdf.

Oard:2002:CIT

[1342] Douglas W. Oard and Julio Gonzalo.

The CLEF 2001 interactive track. Lecture Notes in Computer Science, 2406: 308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060308.
pdf.

Lopez-Ostenero:2002:NPT

[1343] Fernando López-Ostenero, Julio Gonzalo, Anselmo Peñas, and Felisa Verdejo. Noun phrase translations for cross-language document selection. Lecture Notes in Computer Science, 2406:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2406/24060320.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060320.
                                        pdf.

Sanderson:2002:IS

[1344] Mark Sanderson and Zoë Bathie. iCLEF at sheffield. Lecture Notes in Computer Science, 2406:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060332.
                                        pdf.

Wang:2002:IMC

[1345] Jianqiang Wang and Douglas W. Oard. iCLEF 2001 at maryland: Comparing term-for-term gloss and MT. Lecture Notes in Computer Science, 2406:336– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060336.
                                        pdf.

Voorhees:2002:PIR

[1346] Ellen M. Voorhees. The philosophy of information retrieval evaluation. Lecture Notes in Computer Science, 2406: 355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2406/24060355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2406/24060355.
                                        pdf.

Kando:2002:CSE

[1347] Noriko Kando. CLIR system evaluation at the second NTCIR workshop. Lecture Notes in Computer Science, 2406: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060371.
pdf.

Womser-Hacker:2002:MTG

[1348] Christa Womser-Hacker. Multilingual topic generation within the CLEF 2001 experiments. Lecture Notes in Computer Science, 2406:389–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060389.
pdf.

Braschler:2002:CMM

[1349] Martin Braschler and Carol Peters.

CLEF methodology and metrics. Lecture Notes in Computer Science, 2406: 394–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2406/24060394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2406/24060394.
pdf.

Anonymous:2002:AIbe

[1350] Anonymous. Author index. Lecture Notes in Computer Science, 2406:601– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2406/2406auth.pdf.

Bruynooghe:2002:PSC

[1351] Maurice Bruynooghe, Luís Moniz Pereira, Jörg H. Siekmann, and Maarten van Emden. A portrait of a scientist as a computational logician. Lecture Notes in Computer Science, 2407:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070001.
pdf.

Sergot:2002:BKP

[1352] Marek Sergot. Bob Kowalski: A portrait. Lecture Notes in Computer Science, 2407:5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2407/24070005.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2407/24070005.pdf.

Kowalski:2002:DLP

[1353] Robert A. Kowalski. Directions for logic programming. Lecture Notes in Computer Science, 2407:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070026.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2407/24070026.
pdf.

Clark:2002:AMT

[1354] Keith Clark and Peter J. Robinson. Agents as multi-threaded logical objects. Lecture Notes in Computer Science, 2407:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070033.
pdf.

Davison:2002:LPL

[1355] Andrew Davison. Logic programming languages for the Internet. Lecture Notes in Computer Science, 2407:66– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070066.
pdf.

Lloyd:2002:HOC

[1356] John W. Lloyd. Higher-order computational logic. Lecture Notes in Computer Science, 2407:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070105.
pdf.

Ueda:2002:PMI

[1357] Kazunori Ueda. A pure metainterpreter for flat GHC, a concurrent constraint language. Lecture Notes in Computer Science, 2407:138– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070138.
                                        pdf.

Bossi:2002:TSN

[1358] Annalisa Bossi, Nicoletta Cocco, and Sandro Etalle. Transformation systems and nondeclarative properties. Lecture Notes in Computer Science, 2407:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070162.
                                        pdf.

Schreye:2002:AGO

[1359] Danny De Schreye and Alexander Serebrenik. Acceptability with general orderings. Lecture Notes in Computer Science, 2407:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070187.
                                        pdf.

Gupta:2002:SIV

[1360] Gopal Gupta and Enrico Pontelli. Specification, implementation, and verification of domain specific languages: A logic programmingbased approach. Lecture Notes in Computer Science, 2407:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070211.
pdf.

Mancarella:2002:NFT

[1361] Paolo Mancarella, Dino Pedreschi, and Salvatore Ruggieri. Negation as failure through abduction: Reasoning about termination. Lecture Notes in Computer Science, 2407:240– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070240.
pdf.

Pettorossi:2002:PDR

[1362] Alberto Pettorossi and Maurizio Proietti. Program derivation = rules + strategies. Lecture Notes in Computer Science, 2407:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070273.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2407/24070273.
                                        pdf.

Flener:2002:APP

[1363] Pierre Flener. Achievements and prospects of program synthesis. Lecture Notes in Computer Science, 2407: 310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070310.
                                        pdf.

Lau:2002:LCB

[1364] Kung-Kiu Lau and Mario Ornaghi. Logic for component-based software development. Lecture Notes in Computer Science, 2407:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070347.
                                        pdf.

Sterling:2002:PPP

[1365] Leon Sterling. Patterns for Prolog programming. Lecture Notes in Computer Science, 2407:374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070374.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070374.
                                        pdf.

Denecker:2002:ALP

[1366] Marc Denecker and Antonis Kakas. Abduction in logic programming. Lecture Notes in Computer Science, 2407:402– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070402.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070402.
pdf.

Flach:2002:LCL

[1367] Peter Flach and Nada Lavrac. Learning in clausal logic: A perspective on inductive logic programming. Lecture Notes in Computer Science, 2407:437– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070437.
pdf.

Minker:2002:DLP

[1368] Jack Minker and Dietmar Seipel. Disjunctive logic programming: A survey and assessment. Lecture Notes in Computer Science, 2407:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070472.
pdf.

Wallace:2002:CLP

[1369] Mark Wallace. Constraint logic programming. Lecture Notes in Computer Science, 2407:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2407/24070512.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2407/24070512.pdf.

Aiello:2002:PAS

[1370] Luigia Carlucci Aiello and Fabio Massacci. Planning attacks to security protocols: Case studies in logic programming. Lecture Notes in Computer Science, 2407:533–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070533.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070533.
                                        pdf.

Buccafurri:2002:MCJ

[1371] Francesco Buccafurri and Georg Gottlob. Multiagent compromises, joint fixpoints, and stable models. Lecture Notes in Computer Science, 2407:561– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2407/24070561.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2407/24070561.
                                        pdf.

Eiter:2002:ETA

[1372] Thomas Eiter, Viviana Mascardi, and V. S. Subrahmanian. Errortolerant agents. Lecture Notes in Computer Science, 2407:586–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070586.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070586.
pdf.

Jung:2002:LBH

[1373] Christoph G. Jung and Klaus Fischer.

Logic-based hybrid agents. Lecture Notes in Computer Science, 2407:626– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070626.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070626.
pdf.

Sjoland:2002:HSR

[1374] Thomas Sjöland, Per Kreuger, and Martin Aronsson. Heterogeneous scheduling and rotation. Lecture Notes in Computer Science, 2407:655– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2407/24070655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2407/24070655.
pdf.

Anonymous:2002:AIbf

[1375] Anonymous. Author index. Lecture Notes in Computer Science, 2407:677– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2407/2407indx.pdf.

Baldan:2002:MLT

[1376] Paolo Baldan, Paolo Mancarella, Alessandra Raffaetà, and Franco Turini. MuTACLP: A language for temporal reasoning with multiple theories. Lecture Notes in Computer Science, 2408:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2408/24080001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2408/24080001.pdf.

Calvanese:2002:DLI

[1377] Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. Description logics for information integration. Lecture Notes in Computer Science, 2408:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2408/24080041.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080041.
                                        pdf.

Greco:2002:SOP

[1378] Sergio Greco and Domenico Saccà. Search and optimization problems in datalog. Lecture Notes in Computer Science, 2408:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2408/24080061.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2408/24080061.pdf.

Mascellani:2002:DSM

[1379] Paolo Mascellani and Dino Pedreschi.

The declarative side of magic. Lecture Notes in Computer Science, 2408:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080083.
pdf.

Zaniolo:2002:KCM

[1380] Carlo Zaniolo. Key constraints and monotonic aggregates in deductive databases. Lecture Notes in Computer Science, 2408:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080109.
pdf.

Broda:2002:DCS

[1381] Krysia Broda. A decidable CLDS for some propositional resource logics. Lecture Notes in Computer Science, 2408: 135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2408/24080135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080135.
                                        pdf.

Bundy:2002:CPP

[1382] Alan Bundy. A critique of proof planning. Lecture Notes in Computer Science, 2408:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2408/24080160.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2408/24080160.pdf.

Hasegawa:2002:MGB

[1383] Ryuzo Hasegawa, Hiroshi Fujita, Miyuki Koshimura, and Yasuyuki Shirai. A model generation based theorem prover MGTP for first-order logic. Lecture Notes in Computer Science, 2408: 178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2408/24080178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080178.
                                        pdf.

Omodeo:2002:TMP

[1384] Eugenio G. Omodeo and Jacob T. Schwartz. A ‘theory’ mechanism for a proof-verifier based on firstorder set theory. Lecture Notes in Computer Science, 2408:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2408/24080214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080214.
pdf.

Siekmann:2002:ORP

[1385] Jörg Siekmann and Graham Wrightson. An open research problem: Strong completeness of R. kowalski’s connection graph proof procedure. Lecture Notes in Computer Science, 2408:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080231.
pdf.

Costantini:2002:MRS

[1386] Stefania Costantini. Meta-reasoning:

A survey. Lecture Notes in Computer Science, 2408:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2408/24080253.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2408/24080253.pdf.

Dung:2002:ABP

[1387] Phan Minh Dung, Paolo Mancarella, and Francesca Toni. Argumentationbased proof procedures for credulous and sceptical non-monotonic reasoning. Lecture Notes in Computer Science, 2408:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2408/24080289.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080289.
                                        pdf.

Inoue:2002:AA

[1388] Katsumi Inoue. Automated abduction. Lecture Notes in Computer Science, 2408:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2408/24080311.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080311.
                                        pdf.

Prakken:2002:RLC

[1389] Henry Prakken and Giovanni Sartor. The role of logic in computational models of legal argument: A critical survey. Lecture Notes in Computer Science, 2408:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2408/24080342.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080342.
                                        pdf.

Alferes:2002:LPU

[1390] José Júlio Alferes and Luís Moniz Pereira. Logic programming updating — A guided approach. Lecture Notes in Computer Science, 2408:382– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2408/24080382.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2408/24080382.
pdf.

Gelfond:2002:RKP

[1391] Michael Gelfond. Representing knowledge in A-prolog. Lecture Notes in Computer Science, 2408:413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080413.
pdf.

Miller:2002:SAF

[1392] Rob Miller and Murray Shanahan.

Some alternative formulations of the event calculus. Lecture Notes in Computer Science, 2408:452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080452.
pdf.

Cussens:2002:ILL

[1393] James Cussens. Issues in learning language in logic. Lecture Notes in Computer Science, 2408:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080491.
pdf.

Dahl:2002:IM

[1394] Veronica Dahl. On implicit meanings. Lecture Notes in Computer Science, 2408:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2408/24080506.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080506.
                                        pdf.

Raedt:2002:DMC

[1395] Luc De Raedt. Data mining as constraint logic programming. Lecture Notes in Computer Science, 2408:526– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2408/24080526.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080526.
                                        pdf.

Mellish:2002:DPD

[1396] Chris Mellish. DCGs: Parsing as deduction? Lecture Notes in Computer Science, 2408:548–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2408/24080548.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2408/24080548.
                                        pdf.

Sato:2002:SAT

[1397] Taisuke Sato and Yoshitaka Kameya. Statistical abduction with tabulation. Lecture Notes in Computer Science, 2408:567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2408/24080567.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2408/24080567.
pdf.

Gillies:2002:LDC

[1398] Donald Gillies. Logicism and the development of computer science. Lecture Notes in Computer Science, 2408: 588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080588.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080588.
pdf.

Psillos:2002:SBC

[1399] Stathis Psillos. Simply the best: A case for abduction. Lecture Notes in Computer Science, 2408:605–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2408/24080605.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2408/24080605.
pdf.

Anonymous:2002:AIbg

[1400] Anonymous. Author index. Lecture Notes in Computer Science, 2408:627– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2408/2408indx.pdf.

deAragao:2002:IMB

[1401] Marcus Poggi de Aragão and Renato F. Werneck. On the implementation of MST-based heuristics for the Steiner problem in graphs. Lecture Notes in Computer Science, 2409:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090001.
pdf.

Phan:2002:TSS

[1402] Vinhthuy Phan, Pavel Sumazin, and Steven Skiena. A time-sensitive system for black-box combinatorial optimization. Lecture Notes in Computer Science, 2409:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090016.
pdf.

Motter:2002:CBF

[1403] DoRon B. Motter and Igor L. Markov. A compressed breadth-first search for satisfiability. Lecture Notes in Computer Science, 2409:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090029.
pdf.

Schulz:2002:UML

[1404] Frank Schulz, Dorothea Wagner, and Christos Zaroliagis. Using multilevel graphs for timetable information in railway systems. Lecture Notes in Computer Science, 2409:43– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090043.
pdf.

Pruhs:2002:ELR

[1405] Kirk Pruhs and Eric Wiewiora. Evaluating the local ratio algorithm for dynamic storage allocation. Lecture Notes in Computer Science, 2409:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090060.
pdf.

Curcio:2002:ESP

[1406] Massimiliano Curcio, Stefano Leonardi, and Andrea Vitaletti. An experimental study of prefetching and caching algorithms for the World Wide Web. Lecture Notes in Computer Science, 2409: 71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090071.
pdf.

Gustedt:2002:TJP

[1407] Jens Gustedt, Ole A. Mæhle, and Jan Arne Telle. The treewidth of Java programs. Lecture Notes in Computer Science, 2409:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2409/24090086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2409/24090086.
                                        pdf.

Aleksandrov:2002:PPG

[1408] Lyudmil Aleksandrov, Hristo Djidjev, Hua Guo, and Anil Maheshwari. Partitioning planar graphs with costs and weights. Lecture Notes in Computer Science, 2409:98– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2409/24090098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2409/24090098.
                                        pdf.

Cattaneo:2002:MDM

[1409] Giuseppe Cattaneo, Pompeo Faruolo, Umberto Ferraro Petrillo, and Giuseppe F. Italiano. Maintaining dynamic minimum spanning trees: An experimental study. Lecture Notes in Computer Science, 2409:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2409/24090111.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2409/24090111.
pdf.

Pettie:2002:EEN

[1410] Seth Pettie, Vijaya Ramachandran, and Srinath Sridhar. Experimental evaluation of a new shortest path algorithm. Lecture Notes in Computer Science, 2409:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090126.
pdf.

Chaudhry:2002:GMC

[1411] Geeta Chaudhry and Thomas H.

Cormen. Getting more from outof-core columnsort. Lecture Notes in Computer Science, 2409:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2409/24090143.
pdf.

Rafalin:2002:TSD

[1412] Eynat Rafalin, Diane Souvaine, and Ileana Streinu. Topological sweep in degenerate cases. Lecture Notes in Computer Science, 2409:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2409/24090155.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2409/24090155.
                                        pdf.

Phillips:2002:AKM

[1413] Steven J. Phillips. Acceleration of K-means and related clustering algorithms. Lecture Notes in Computer Science, 2409:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2409/24090166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2409/24090166.
                                        pdf.

Procopiuc:2002:STE

[1414] Cecilia M. Procopiuc, Pankaj K. Agarwal, and Sariel Har-Peled. STAR-Tree: An efficient self-adjusting index for moving objects. Lecture Notes in Computer Science, 2409:178– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2409/24090178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2409/24090178.
                                        pdf.

Chen:2002:ITS

[1415] Jingchao Chen. An improvement on tree selection sort. Lecture Notes in Computer Science, 2409:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2409/24090194.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2409/24090194.
pdf.

Anonymous:2002:AIbh

[1416] Anonymous. Author index. Lecture Notes in Computer Science, 2409:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2409/2409auth.pdf.

Butler:2002:FMN

[1417] Ricky Butler. Formal methods at NASA langley. Lecture Notes in Computer Science, 2410:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100001.
pdf.

Huet:2002:HOU

[1418] Gérard Huet. Higher order unification 30 years later (extended abstract). Lecture Notes in Computer Science, 2410: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100003.
pdf.

Ambler:2002:CHO

[1419] Simon J. Ambler, Roy L. Crole, and Alberto Momigliano. Combining higher order abstract syntax with tactical theorem proving and (co)induction. Lecture Notes in Computer Science, 2410: 13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100013.
pdf.

Barthe:2002:ERA

[1420] Gilles Barthe and Pierre Courtieu. Efficient reasoning about executable specifications in Coq. Lecture Notes in Computer Science, 2410:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100031.
pdf.

Basin:2002:VBM

[1421] David Basin, Stefan Friedrich, and Marek Gawkowski. Verified bytecode model checkers. Lecture Notes in Computer Science, 2410:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100047.
pdf.

Bauer:2002:CTI

[1422] Gertrud Bauer and Tobias Nipkow. The 5 colour theorem in Isabelle/Isar. Lecture Notes in Computer Science, 2410:67–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2410/24100067.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2410/24100067.
pdf.

Bertot:2002:TTF

[1423] Yves Bertot, Venanzio Capretta, and Kuntal Das Barman. Type-theoretic functional semantics. Lecture Notes in Computer Science, 2410:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100083.
pdf.

Brucker:2002:PFO

[1424] Achim D. Brucker and Burkhart Wolff.

A proposal for a formal OCL semantics in Isabelle/HOL. Lecture Notes in Computer Science, 2410:99– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100099.
pdf.

Courant:2002:EUC

[1425] Judicaël Courant. Explicit universes for the calculus of constructions. Lecture Notes in Computer Science, 2410: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2410/24100115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2410/24100115.
                                        pdf.

Dawson:2002:FCA

[1426] Jeremy E. Dawson and Rajeev Goré. Formalised cut admissibility for display logic. Lecture Notes in Computer Science, 2410:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2410/24100131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2410/24100131.
                                        pdf.

Dehlinger:2002:FTT

[1427] Christophe Dehlinger and Jean-François Dufourd. Formalizing the trading theorem for the classification of surfaces. Lecture Notes in Computer Science, 2410:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2410/24100148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2410/24100148.
                                        pdf.

Delahaye:2002:FST

[1428] David Delahaye. Free-style theorem proving. Lecture Notes in Computer Science, 2410:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2410/24100164.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2410/24100164.pdf.

Dennis:2002:CTP

[1429] Louise A. Dennis and Alan Bundy.

A comparison of two proof critics: Power vs. robustness. Lecture Notes in Computer Science, 2410:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100182.
pdf.

Felty:2002:TLM

[1430] Amy P. Felty. Two-level metareasoning in Coq. Lecture Notes in Computer Science, 2410:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100198.
pdf.

Gordon:2002:PEP

[1431] Michael J. C. Gordon. PuzzleTool:

An example of programming computation and deduction. Lecture Notes in Computer Science, 2410:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100214.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2410/24100214.
pdf.

Hurd:2002:FAP

[1432] Joe Hurd. A formal approach to probabilistic termination. Lecture Notes in Computer Science, 2410:230– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100230.
pdf.

Mayero:2002:UTP

[1433] Micaela Mayero. Using theorem proving for numerical analysis (correctness proof of an automatic differentiation algorithm). Lecture Notes in Computer Science, 2410:246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100246.
pdf.

Nogin:2002:QTM

[1434] Aleksey Nogin. Quotient types: A modular approach. Lecture Notes in Computer Science, 2410:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100263.
pdf.

Nogin:2002:SSD

[1435] Aleksey Nogin and Jason Hickey. Sequent schema for derived rules. Lecture Notes in Computer Science, 2410: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100281.
pdf.

Prevosto:2002:ASD

[1436] Virgile Prevosto, Damien Doligez, and Thérèse Hardin. Algebraic structures and dependent records. Lecture Notes in Computer Science, 2410:298– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100298.
pdf.

Schneider:2002:PEM

[1437] Klaus Schneider. Proving the equivalence of microstep and macrostep semantics. Lecture Notes in Computer Science, 2410:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2410/24100314.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2410/24100314.pdf.

Zhang:2002:WPG

[1438] Xingyuan Zhang, Malcolm Munro, Mark Harman, and Lin Hu. Weakest precondition for general recursive programs formalized in Coq. Lecture Notes in Computer Science, 2410:332– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2410/24100332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2410/24100332.
pdf.

Anonymous:2002:AIbi

[1439] Anonymous. Author index. Lecture Notes in Computer Science, 2410:349– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2410/24100349.pdf.

Capelle:2002:MFS

[1440] Matthieu Capelle, Cyrille Masson, and Jean-François Boulicaut. Mining frequent sequential patterns under a similarity constraint. Lecture Notes in Computer Science, 2412:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120001.
pdf.

Bramer:2002:PPC

[1441] Max Bramer. Pre-pruning classification trees to reduce overfitting in noisy domains. Lecture Notes in Computer Science, 2412:7– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120007.
pdf.

Smith:2002:DMF

[1442] James F. Smith III. Data mining for fuzzy decision tree structure with a genetic program. Lecture Notes in Computer Science, 2412:13– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120013.
pdf.

Smith:2002:CED

[1443] James F. Smith III. Co-evolutionary data mining to discover rules for fuzzy resource management. Lecture Notes in Computer Science, 2412:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120019.
pdf.

Karimi:2002:DTR

[1444] Kamran Karimi and Howard J. Hamilton. Discovering temporal rules from temporally ordered data. Lecture Notes in Computer Science, 2412:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120025.
                                        pdf.

Hiltunen:2002:API

[1445] Yrjö Hiltunen and Mika Lappalainen. Automated personalisation of Internet users using self-organising maps. Lecture Notes in Computer Science, 2412: 31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120031.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120031.
                                        pdf.

Narita:2002:DAN

[1446] Masaaki Narita, Makoto Haraguchi, and Yoshiaki Okubo. Data abstractions for numerical attributes in data mining. Lecture Notes in Computer Science, 2412:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120035.
                                        pdf.

Bhutta:2002:CAR

[1447] Kashif Bhutta. Calculating aggregates with range-encoded bit-sliced index. Lecture Notes in Computer Science, 2412:43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2412/24120043.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2412/24120043.
pdf.

Tjortjis:2002:TCA

[1448] Christos Tjortjis and John Keane.

T3: A classification algorithm for data mining. Lecture Notes in Computer Science, 2412:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2412/24120050.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2412/24120050.pdf.

Hayashi:2002:HMS

[1449] Tomofumi Hayashi, Akio Sato, and Nadia Berthouze. A hierarchical model to support Kansei mining process. Lecture Notes in Computer Science, 2412: 56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120056.
pdf.

Salim:2002:ESQ

[1450] Majid Salim and Xin Yao. Evolving SQL queries for data mining. Lecture Notes in Computer Science, 2412:62– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120062.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2412/24120062.
pdf.

Jin:2002:IML

[1451] Xiaoming Jin, Likun Wang, Yuchang Lu, and Chunyi Shi. Indexing and mining of the local patterns in sequence database. Lecture Notes in Computer Science, 2412:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120068.
pdf.

Sornkaew:2002:KDF

[1452] Thanakorn Sornkaew and Yasuo Yamashita. A knowledge discovery by fuzzy rule based Hopfield network. Lecture Notes in Computer Science, 2412: 74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120074.
pdf.

Rossiter:2002:FPI

[1453] Jonathan Rossiter. Fusing partially inconsistent expert and learnt knowledge in uncertain hierarchies. Lecture Notes in Computer Science, 2412:80– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120080.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2412/24120080.
pdf.

Moreale:2002:OIM

[1454] Emanuela Moreale and Stuart Watt.

Organisational information management and knowledge discovery in email within mailing lists. Lecture Notes in Computer Science, 2412:87– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120087.
pdf.

Klene:2002:DMD

[1455] G. Klene, A. Grauel, H. J. Convey, and A. J. Hartley. Design of multi-drilling gear machines by knowledge processing and machine simulation. Lecture Notes in Computer Science, 2412:93– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120093.
pdf.

Zhdanova:2002:CEQ

[1456] Anna V. Zhdanova and Denis V.

Shishkin. Classification of email queries by topic: Approach based on hierarchically structured subject domain. Lecture Notes in Computer Science, 2412: 99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2412/24120099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120099.
                                        pdf.

Yang:2002:KBI

[1457] Jaeyoung Yang, Heekuck Oh, Kyung-Goo Doh, and Joongmin Choi. A knowledge-based information extraction system for semi-structured labeled documents. Lecture Notes in Computer Science, 2412:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120105.
                                        pdf.

Li:2002:MSS

[1458] Yuhua Li, Zuhair Bandar, and David Mclean. Measuring semantic similarity between words using lexical knowledge and neural networks. Lecture Notes in Computer Science, 2412:111– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120111.
                                        pdf.

Carchiolo:2002:EHS

[1459] Vincenza Carchiolo, Alessandro Longheu, and Michele Malgeri. Extraction of hidden semantics from Web pages. Lecture Notes in Computer Science, 2412:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120117.
pdf.

Freeman:2002:SOM

[1460] Richard Freeman and Hujun Yin. Selforganising maps for hierarchical tree view document clustering using contextual information. Lecture Notes in Computer Science, 2412:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120123.
pdf.

He:2002:SDS

[1461] Jianwen He. Schema discovery of the semi-structured and hierarchical data. Lecture Notes in Computer Science, 2412:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2412/24120129.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2412/24120129.
pdf.

Marir:2002:RIR

[1462] Farhi Marir and Kamel Houam.

RSTIndex: Indexing and retrieving Web document using computational and linguistic techniques. Lecture Notes in Computer Science, 2412:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120135.
                                        pdf.

Umehara:2002:CBR

[1463] Masayuki Umehara, Koji Iwanuma, and Hidetomo Nabeshima. A casebased recognition of semantic structures in HTML documents. Lecture Notes in Computer Science, 2412:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120141.
                                        pdf.

Yeow:2002:EXP

[1464] Kelvin Yeow, R. Nigel Horspool, and Michael R. Levy. Expeditious XML processing. Lecture Notes in Computer Science, 2412:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120148.
                                        pdf.

Russell:2002:DCU

[1465] Ben Russell, Hujun Yin, and Nigel M. Allinson. Document clustering using the 1 + 1 dimensional selforganising map. Lecture Notes in Computer Science, 2412:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120154.
pdf.

Kim:2002:NLP

[1466] Sanghee Kim, Wendy Hall, and Andy Keane. Natural language processing for expertise modelling in Email communication. Lecture Notes in Computer Science, 2412:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120161.
pdf.

Han:2002:BBAb

[1467] Jun Han, Graham McMahon, and Stephen Sugden. A branch and bound algorithm for minimum cost network flow problem. Lecture Notes in Computer Science, 2412:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120167.
pdf.

Durand:2002:SRU

[1468] Nicolas Durand and Luigi Lancieri.

Study of the regularity of the users’ Internet accesses. Lecture Notes in Computer Science, 2412:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120173.
                                        pdf.

Jang:2002:IMC

[1469] Sera Jang and Eunseok Lee. An intelligent mobile commerce system with dynamic contents builder and mobile products browser. Lecture Notes in Computer Science, 2412:179– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120179.
                                        pdf.

Kononen:2002:FCU

[1470] Ville Könönen. Focused crawling using fictitious play. Lecture Notes in Computer Science, 2412:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120186.
                                        pdf.

Lee:2002:UAM

[1471] Eunseok Lee and Sera Jang. A user adaptive mobile commerce system with a middlet application. Lecture Notes in Computer Science, 2412: 193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120193.
pdf.

Prasad:2002:WVB

[1472] Bhanu Prasad. Weight-vector based approach for product recommendation in E-commerce. Lecture Notes in Computer Science, 2412:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120200.
pdf.

Huang:2002:DXB

[1473] Shi-Ming Huang and Chun-Hao Su.

The development of an XML-based data warehouse system. Lecture Notes in Computer Science, 2412:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120206.
pdf.

Koncilia:2002:IDS

[1474] Christian Koncilia and Heinz Pozewaunig. Identifying data sources for data warehouses. Lecture Notes in Computer Science, 2412:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120213.
                                        pdf.

Lynden:2002:CLA

[1475] Steven Lynden and Omer F. Rana. Coordinating learning agents via utility assignment. Lecture Notes in Computer Science, 2412:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120219.
                                        pdf.

Dessi:2002:AAA

[1476] Nicoletta Dessì. AGILE: An agentassisted infrastructure to support learning environments. Lecture Notes in Computer Science, 2412:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120225.
                                        pdf.

Smith:2002:MAF

[1477] James F. Smith III. Multi-agent fuzzy logic resource manager. Lecture Notes in Computer Science, 2412:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2412/24120231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120231.
pdf.

Ye:2002:TMA

[1478] Xinfeng Ye, John Keane, and Guoqing Zhang. Transactional multiple agents. Lecture Notes in Computer Science, 2412:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120237.
pdf.

Kasiran:2002:IMM

[1479] Mohd Khairudin Kasiran and Farid Meziane. An information model for a merchant trust agent in electronic commerce. Lecture Notes in Computer Science, 2412:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120243.
pdf.

Trajkovski:2002:MCS

[1480] Goran Trajkovski. MASIVE: A case study in multiagent systems. Lecture Notes in Computer Science, 2412:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120249.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120249.
                                        pdf.

Wright:2002:LMA

[1481] W. Andy Wright. Learning multiagent strategies in multi-stage collaborative games. Lecture Notes in Computer Science, 2412:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120255.
                                        pdf.

Li:2002:ESS

[1482] Ling Li, Alcherio Martinoli, and Yaser S. Abu-Mostafa. Emergent specialization in swarm systems. Lecture Notes in Computer Science, 2412:261– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120261.
                                        pdf.

Feng:2002:DMC

[1483] J. Q. Feng, D. P. Buse, Q. H. Wu, and J. Fitch. Distributed mobile communication base station diagnosis and monitoring using multi-agents. Lecture Notes in Computer Science, 2412:267– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120267.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2412/24120267.
pdf.

Perrone:2002:AAB

[1484] Alessandro Perrone and Gianluigi Ferraris. ABBA — agent based beaver application — busy beaver in swarm. Lecture Notes in Computer Science, 2412: 273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120273.
pdf.

Gabbai:2002:CDO

[1485] J. M. E. Gabbai, W. A. Wright, and N. M. Allinson. Centralised and distributed organisational control. Lecture Notes in Computer Science, 2412: 279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120279.
pdf.

Xu:2002:MDS

[1486] Lei Xu. Mining dependence structures from statistical learning perspective. Lecture Notes in Computer Science, 2412:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2412/24120285.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2412/24120285.
pdf.

Cheung:2002:MGM

[1487] Yiu ming Cheung. k-means — A generalized k-means clustering algorithm with unknown cluster number. Lecture Notes in Computer Science, 2412:307– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120307.
pdf.

Jin:2002:MSM

[1488] Xiaolong Jin and Jiming Liu. Multiagent SAT (MASSAT): Autonomous pattern search in constrained domains. Lecture Notes in Computer Science, 2412:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2412/24120318.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2412/24120318.
pdf.

Zhong:2002:TMA

[1489] Ning Zhong, Takahisa Matsunaga, and Chunnian Liu. A text mining agents based architecture for personal E-mail filtering and management. Lecture Notes in Computer Science, 2412:329– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120329.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2412/24120329.
pdf.

Zhong:2002:FMA

[1490] Ning Zhong, Yasuaki Matsui, Tomohiro Okuno, and Chunnian Liu. Framework of a multi-agent KDD system. Lecture Notes in Computer Science, 2412:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2412/24120337.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2412/24120337.
pdf.

Dempster:2002:IFT

[1491] M. A. H. Dempster and Y. S. Romahi.

Intraday FX trading: An evolutionary reinforcement learning approach. Lecture Notes in Computer Science, 2412: 347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120347.
pdf.

Baek:2002:TDU

[1492] Jinwoo Baek and Sungzoon Cho. An up-trend detection using an autoassociative neural network: KOSPI 200 futures. Lecture Notes in Computer Science, 2412:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2412/24120359.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2412/24120359.pdf.

Chiu:2002:SPI

[1493] Kai Chun Chiu and Lei Xu. Stock price and index forecasting by arbitrage pricing theory-based Gaussian TFA learning. Lecture Notes in Computer Science, 2412:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120366.
pdf.

Liu:2002:CST

[1494] Zhiyong Liu and Lei Xu. A comparative study on three MAP factor estimate approaches for NFA. Lecture Notes in Computer Science, 2412:372– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120372.
pdf.

Kim:2002:NCF

[1495] Min-Jung Kim and Taek-Soo Kim. A neural classifier with fraud density map for effective credit card fraud detection. Lecture Notes in Computer Science, 2412:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120378.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2412/24120378.
pdf.

Lora:2002:CTT

[1496] Alicia Troncoso Lora, Jesús Riquelme Santos, José RiquelmeSantos, Antonio Gómez Expósito, and José Luís Martínez Ramos. A comparison of two techniques for next-day electricity price forecasting. Lecture Notes in Computer Science, 2412:384– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120384.
pdf.

Yang:2002:SVM

[1497] Haiqin Yang, Laiwan Chan, and Irwin King. Support vector machine regression for volatile stock market prediction. Lecture Notes in Computer Science, 2412:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120391.
pdf.

Han:2002:CPF

[1498] Ying Han and Colin Fyfe. Complexity pursuit for financial prediction. Lecture Notes in Computer Science, 2412: 397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120397.
                                        pdf.

Chan:2002:AIP

[1499] Man-Chung Chan, Chi-Cheong Wong, W. F. Tse, Bernard K.-S. Cheung, and Gordon Y.-N. Tang. Artificial intelligence in portfolio management. Lecture Notes in Computer Science, 2412: 403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120403.
                                        pdf.

Magdon-Ismail:2002:MCP

[1500] Malik Magdon-Ismail, Hung-Ching (Justin) Chen, and Yaser S. Abu-Mostafa. The multilevel classification problem and a monotonicity hint. Lecture Notes in Computer Science, 2412:410– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120410.
                                        pdf.

Lynch:2002:AFG

[1501] Paul E. Lynch and Nigel M. Allinson. Adaptive filtering for GARCH models. Lecture Notes in Computer Science, 2412:416–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2412/24120416.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2412/24120416.
pdf.

Pulkkinen:2002:ASO

[1502] Juhani Pulkkinen, Mika Lappalainen, Anna-Maija Häkkinen, Nina Lundbom, Risto A. Kauppinen, and Yrjö Hiltunen. Application of self-organising maps in automated chemical shift correction of in vivo H MR spectra. Lecture Notes in Computer Science, 2412: 423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120423.
pdf.

Spasic:2002:SLT

[1503] Irena Spasić, Goran Nenadić, Kostas Manios, and Sophia Ananiadou. Supervised learning of term similarities. Lecture Notes in Computer Science, 2412: 429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120429.
pdf.

Lopez-Alonso:2002:BKE

[1504] Victoria López-Alonso, Lucia Moreno, Guillermo López-Campos, Victor Maojo, and Fernando Martín-Sanchez. BIKMAS: A knowledge engineering system for bioinformatics. Lecture Notes in Computer Science, 2412:435– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120435.
                                        pdf.

Ladroue:2002:UFE

[1505] C. Ladroue, A. R. Tate, F. A. Howe, and J. R. Griffiths. Unsupervised feature extraction of in vivo magnetic resonance spectra of brain tumours using independent component analysis. Lecture Notes in Computer Science, 2412: 441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120441.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120441.
                                        pdf.

Supekar:2002:FRB

[1506] K. Supekar, A. Marwadi, Y. Lee, and D. Medhi. Fuzzy rule-based framework for medical record validation. Lecture Notes in Computer Science, 2412:447– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120447.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120447.
                                        pdf.

Hwang:2002:CLD

[1507] Grace J. Hwang and Chun-Chan Tung. Classification learning by decomposition of numerical datasets. Lecture Notes in Computer Science, 2412:454– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120454.
pdf.

Bao:2002:CFS

[1508] Yongguang Bao, Xiaoyong Du, and Naohiro Ishii. Combining feature selection with feature weighting for k-NN classifier. Lecture Notes in Computer Science, 2412:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120461.
pdf.

Shin:2002:PSS

[1509] Hyunjung Shin and Sungzoon Cho.

Pattern selection for support vector classifiers. Lecture Notes in Computer Science, 2412:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2412/24120469.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2412/24120469.pdf.

Chang:2002:GFS

[1510] Yuan chin Ivan Chang, Haoran Hsu, and Lin-Yi Chou. Graphical features selection method. Lecture Notes in Computer Science, 2412:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120475.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120475.
                                        pdf.

Crockett:2002:FNI

[1511] Keeley Crockett, Zuhair Bandar, and James O’Shea. Fuzzy-neural inference in decision trees. Lecture Notes in Computer Science, 2412:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120481.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120481.
                                        pdf.

Yook:2002:DTB

[1512] Dongsuk Yook. Decision tree based clustering. Lecture Notes in Computer Science, 2412:487–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120487.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120487.
                                        pdf.

Levashenko:2002:UNI

[1513] Vitaly G. Levashenko and Elena N. Zaitseva. Usage of new information estimations for induction of fuzzy decision trees. Lecture Notes in Computer Science, 2412:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120493.
pdf.

Bin:2002:GAB

[1514] Bin LI and Zhen quan Zhuang. Genetic algorithm based-on the quantum probability representation. Lecture Notes in Computer Science, 2412: 500–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120500.
pdf.

Hwang:2002:DMD

[1515] Grace J. Hwang and Fumin Li. A dynamic method for discretization of continuous attributes. Lecture Notes in Computer Science, 2412:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120506.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120506.
pdf.

Fyfe:2002:NNI

[1516] Colin Fyfe and Emilio Corchado. A new neural implementation of exploratory projection pursuit. Lecture Notes in Computer Science, 2412:512– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120512.
                                        pdf.

Kaban:2002:GFP

[1517] Ata Kabán, Peter Tiňo, and Mark Girolami. A general framework for a principled hierarchical visualization of multivariate data. Lecture Notes in Computer Science, 2412:518–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120518.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120518.
                                        pdf.

Singh:2002:CCR

[1518] Sameer Singh, Adnan Amin, and K. C. Sum. Chinese character recognition — comparison of classification methodologies. Lecture Notes in Computer Science, 2412:524–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120524.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120524.
                                        pdf.

Iwata:2002:LZC

[1519] Kazunori Iwata and Naohiro Ishii. Lempel–Ziv coding in reinforcement learning. Lecture Notes in Computer Science, 2412:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120531.
pdf.

Kim:2002:EFE

[1520] Jong-Bae Kim, Chae-Hyun Moon, and Hang-Joon Kim. Efficient face extraction using skin-color model and a neural network. Lecture Notes in Computer Science, 2412:538–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120538.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120538.
pdf.

Ding:2002:FWD

[1521] Shan Ding and Naohiro Ishii. Feature weights determining of pattern classification by using a rough genetic algorithm with fuzzy similarity measure. Lecture Notes in Computer Science, 2412:544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120544.
pdf.

Dokur:2002:RFD

[1522] Zümray Dokur and Tamer Ölmez. Recursive form of the discrete Fourier transform for two-dimensional signals. Lecture Notes in Computer Science, 2412:551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2412/24120551.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120551.
                                        pdf.

Lee:2002:VRE

[1523] Soonkyu Lee and Dongsuk Yook. Viseme recognition experiment using context dependent hidden Markov models. Lecture Notes in Computer Science, 2412:557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2412/24120557.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2412/24120557.pdf.

Miyao:2002:SEP

[1524] Hidetoshi Miyao. Stave extraction for printed music scores. Lecture Notes in Computer Science, 2412:562– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120562.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120562.
                                        pdf.

Jin:2002:SMB

[1525] Huidong Jin, Kwong-Sak Leung, and Man-Leung Wong. Scaling-up modelbased clustering algorithm by working on clustering features. Lecture Notes in Computer Science, 2412:569– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120569.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120569.
pdf.

Liu:2002:NAH

[1526] Yang Liu, Huanqiang Zhang, and Zhimei Wu. A new approach to hierarchically retrieve MPEG video. Lecture Notes in Computer Science, 2412: 576–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120576.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120576.
pdf.

Parkins:2002:ABS

[1527] Paul Parkins and John A. Keane.

Alpha-beta search revisited. Lecture Notes in Computer Science, 2412:582– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2412/24120582.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2412/24120582.
pdf.

Wang:2002:QRI

[1528] Wenjia Wang. Quantifying relevance of input features. Lecture Notes in Computer Science, 2412:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2412/24120588.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2412/24120588.
                                        pdf.

Anonymous:2002:AIbj

[1529] Anonymous. Author index. Lecture Notes in Computer Science, 2412:595– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2412/2412auth.pdf.

Boella:2002:ANR

[1530] Guido Boella and Rossana Damiano. An architecture for normative reactive agents. Lecture Notes in Computer Science, 2413:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2413/24130001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130001.
                                        pdf.

Lee:2002:RTA

[1531] Jimmy H. M. Lee and Lei Zhao. A real-time agent architecture: Design, implementation and evaluation. Lecture Notes in Computer Science, 2413: 18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2413/24130018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130018.
                                        pdf.

Takashina:2002:ILA

[1532] Tomomi Takashina, Kazuhide Tanaka, and Shigeyoshi Watanabe. Individual level analysis using decision making features in multiagent based simulation. Lecture Notes in Computer Science, 2413:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2413/24130033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2413/24130033.
pdf.

Terada:2002:FNP

[1533] Kenji Terada and Makoto Yokoo.

False-name-proof multi-unit auction protocol utilizing greedy allocation based on approximate evaluation values. Lecture Notes in Computer Science, 2413:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2413/24130048.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2413/24130048.pdf.

Chen:2002:DAM

[1534] Wei Chen and Keith Decker. Developing alternative mechanisms for multiagent coordination. Lecture Notes in Computer Science, 2413:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2413/24130063.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2413/24130063.
                                        pdf.

Jmaiel:2002:OSN

[1535] Mohamed Jmaiel and Ahmed Hadj Kacem. An operational semantics for negotiating agents. Lecture Notes in Computer Science, 2413:77– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2413/24130077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130077.
                                        pdf.

Serrano:2002:OMD

[1536] Juan Manuel Serrano and Sascha Ossowski. An organizational metamodel for the design of catalogues of communicative actions. Lecture Notes in Computer Science, 2413:92– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2413/24130092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130092.
                                        pdf.

Mathieu:2002:PDM

[1537] Philippe Mathieu, Jean-Christophe Routier, and Yann Secq. Principles for dynamic multi-agent organizations. Lecture Notes in Computer Science, 2413:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2413/24130109.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2413/24130109.
pdf.

Soo:2002:CDD

[1538] Von-Wun Soo and Hai-Long Cheng.

Conducting the disambiguation dialogues between software agent sellers and human buyers. Lecture Notes in Computer Science, 2413:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2413/24130123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2413/24130123.
pdf.

Yamada:2002:MLM

[1539] Seiji Yamada and Tomohiro Yamaguchi. Mutual learning of mind reading between a human and a life-like agent. Lecture Notes in Computer Science, 2413:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2413/24130138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2413/24130138.
pdf.

Shim:2002:ASS

[1540] Yunju Shim and Minkoo Kim. Automatic short story generator based on autonomous agents. Lecture Notes in Computer Science, 2413:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2413/24130151.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130151.
                                        pdf.

Gao:2002:MLA

[1541] Zhiqiang Gao, Tomoyuki Kawasoe, Akishige Yamamoto, and Toru Ishida. Meta-level architecture for executing multi-agent scenarios. Lecture Notes in Computer Science, 2413:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2413/24130163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130163.
                                        pdf.

Kitagata:2002:AOF

[1542] Gen Kitagata, Takuo Suganuma, and Tetsuo Kinoshita. Applicationoriented flow control in agent-based network middleware. Lecture Notes in Computer Science, 2413:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2413/24130178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2413/24130178.
                                        pdf.

Sawamoto:2002:CTD

[1543] Jun Sawamoto, Hidekazu Tsuji, and Hisao Koizumi. Continuous truck delivery scheduling and execution system with multiple agents. Lecture Notes in Computer Science, 2413:190– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2413/24130190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2413/24130190.
pdf.

Liu:2002:DAC

[1544] Jyi-Shane Liu. Designing agents for context-rich textual information tasks. Lecture Notes in Computer Science, 2413:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2413/24130205.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2413/24130205.
pdf.

Anonymous:2002:AIbk

[1545] Anonymous. Author index. Lecture Notes in Computer Science, 2413:221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2413/2413auth.pdf.

Raman:2002:SMS

[1546] Bhaskaran Raman, Sharad Agarwal, Yan Chen, Matthew Caesar, Weidong Cui, Per Johansson, Kevin Lai, Tal Lavian, Sridhar Machiraju, Z. Morley Mao, George Porter, Timothy Roscoe, Mukund Seshadri, Jimmy Shih, Keith Sklower, Lakshminarayanan Subramanian, Takashi Suzuki, Shelley Zhuang, Anthony D. Joseph, Randy H. Katz, and Ion Stoica. The SAHARA model for service composition across multiple providers. Lecture Notes in Computer Science, 2414: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140001.
                                        pdf.

Herrtwich:2002:UCA

[1547] Ralf G. Herrtwich. Ubiquitous computing in the automotive domain (abstract). Lecture Notes in Computer Science, 2414:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140015.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140015.
                                        pdf.

Hess:2002:BAU

[1548] Christopher K. Hess, Manuel Román, and Roy H. Campbell. Building applications for ubiquitous computing environments. Lecture Notes in Computer Science, 2414:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140016.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140016.
                                        pdf.

Arnstein:2002:SSU

[1549] Larry Arnstein, Robert Grimm, Chia-Yang Hung, Jong Hee Kang, Anthony LaMarca, Gary Look, Stefan B. Sigurdsson, Jing Su, and Gaetano Borriello. Systems support for ubiquitous computing: A case study of two implementations of labscape. Lecture Notes in Computer Science, 2414:30– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140030.
pdf.

Lueg:2002:GBV

[1550] Christopher Lueg. On the gap between vision and feasibility. Lecture Notes in Computer Science, 2414:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140045.
pdf.

Levy:2002:FKP

[1551] David Levy. The fastap keypad and pervasive computing. Lecture Notes in Computer Science, 2414:58– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140058.
pdf.

Bayon:2002:GBS

[1552] Victor Bayon, Tom Rodden, Chris Greenhalgh, and Steve Benford. Going back to school: Putting a pervasive environment into the real world. Lecture Notes in Computer Science, 2414: 69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140069.
                                        pdf.

Ferscha:2002:PWA

[1553] Alois Ferscha and Simon Vogl. Pervasive Web access via public communication walls. Lecture Notes in Computer Science, 2414:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140084.
                                        pdf.

Vogt:2002:EOI

[1554] Harald Vogt. Efficient object identification with passive RFID tags. Lecture Notes in Computer Science, 2414:98– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140098.
                                        pdf.

Clarke:2002:UCP

[1555] Dwaine Clarke, Blaise Gassend, Thomas Kotwal, Matt Burnside, Marten van Dijk, Srinivas Devadas, and Ronald Rivest. The untrusted computer problem and camerabased authentication. Lecture Notes in Computer Science, 2414:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140114.
pdf.

Tuulari:2002:SPU

[1556] Esa Tuulari and Arto Ylisaukko-oja.

SoapBox: A platform for ubiquitous computing research and applications. Lecture Notes in Computer Science, 2414:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2414/24140125.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2414/24140125.
pdf.

Lifton:2002:PCS

[1557] Joshua Lifton, Deva Seetharam, Michael Broxton, and Joseph Paradiso. Pushpin computing system overview: A platform for distributed, embedded, ubiquitous sensor networks. Lecture Notes in Computer Science, 2414:139– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140139.
pdf.

LaMarca:2002:MSN

[1558] Anthony LaMarca, Waylon Brunette, David Koizumi, Matthew Lease, Stefan B. Sigurdsson, Kevin Sikorski, Dieter Fox, and Gaetano Borriello. Making sensor networks practical with robots. Lecture Notes in Computer Science, 2414:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140152.
pdf.

Henricksen:2002:MCI

[1559] Karen Henricksen, Jadwiga Indulska, and Andry Rakotonirainy. Modeling context information in pervasive computing systems. Lecture Notes in Computer Science, 2414:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140167.
pdf.

Schubiger-Banz:2002:MSC

[1560] Simon Schubiger-Banz and Béat Hirsbrunner. A model for software configuration in ubiquitous computing environments. Lecture Notes in Computer Science, 2414:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140181.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2414/24140181.
pdf.

Balazinska:2002:ITS

[1561] Magdalena Balazinska, Hari Balakrishnan, and David Karger. INS/Twine: A scalable peer-to-peer architecture for intentional resource discovery. Lecture Notes in Computer Science, 2414:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140195.
pdf.

Vildjiounaite:2002:LEI

[1562] Elena Vildjiounaite, Esko-Juhani Malm, Jouni Kaartinen, and Petteri Alahuhta. Location estimation indoors by means of small computing power devices, accelerometers, magnetic sensors, and map knowledge. Lecture Notes in Computer Science, 2414:211– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140211.
pdf.

Kubach:2002:EBL

[1563] Uwe Kubach and Kurt Rothermel.

Estimating the benefit of locationawareness for mobile data management mechanisms. Lecture Notes in Computer Science, 2414:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140225.
                                        pdf.

Nakanishi:2002:IMC

[1564] Yasuto Nakanishi, Kazunari Takahashi, Takayuki Tsuji, and Katsuya Hakozaki. iCAMS: A mobile communication tool using location and schedule information. Lecture Notes in Computer Science, 2414:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140239.
                                        pdf.

Song:2002:BSR

[1565] Henry Song, Hao hua Chu, Nayeem Islam, Shoji Kurakake, and Masaji Katagiri. Browser state repository service. Lecture Notes in Computer Science, 2414:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2414/24140253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2414/24140253.
                                        pdf.

Hori:2002:ATA

[1566] Masahiro Hori, Kouichi Ono, Teruo Koyanagi, and Mari Abe. Annotation by transformation for the automatic generation of content customization metadata. Lecture Notes in Computer Science, 2414:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140267.
pdf.

Chen:2002:SDS

[1567] Yan Chen, Randy H. Katz, and John D. Kubiatowicz. SCAN: A dynamic, scalable, and efficient content distribution network. Lecture Notes in Computer Science, 2414:282– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2414/24140282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2414/24140282.
pdf.

Anonymous:2002:AIbl

[1568] Anonymous. Author index. Lecture Notes in Computer Science, 2414:297– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2414/2414auth.pdf.

Deco:2002:NTV

[1569] Gustavo Deco and Edmund Rolls.

A neurodynamical theory of visual attention: Comparisons with fMRIand single-neuron data. Lecture Notes in Computer Science, 2415:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150003.
                                        pdf.

Molina-Vilaplana:2002:NMS

[1570] Javier Molina-Vilaplana, Jorge Feliu Batlle, and Juan López Coronado. A neural model of spatio temporal coordination in prehension. Lecture Notes in Computer Science, 2415:9– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150009.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150009.
                                        pdf.

Thiel:2002:SDP

[1571] Andreas Thiel, Christian W. Eurich, and Helmut Schwegler. Stabilized dynamics in physiological and neural systems despite strongly delayed feedback. Lecture Notes in Computer Science, 2415:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150015.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150015.
                                        pdf.

Einhauser:2002:LMF

[1572] Wolfgang Einhäuser, Christoph Kayser, Konrad P. Körding, and Peter König. Learning multiple feature representations from natural image sequences. Lecture Notes in Computer Science, 2415:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150021.
pdf.

Aguirre:2002:ABI

[1573] Carlos Aguirre, Ramón Huerta, Fernando Corbacho, and Pedro Pascual. Analysis of biologically inspired small-world networks. Lecture Notes in Computer Science, 2415:27– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150027.
pdf.

Hurri:2002:RFS

[1574] Jarmo Hurri and Aapo Hyvärinen.

Receptive fields similar to simple cells maximize temporal coherence in natural video. Lecture Notes in Computer Science, 2415:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150033.
pdf.

Marinaro:2002:NIS

[1575] Maria Marinaro and Silvia Scarpetta. Noise induces spontaneous synchronous aperiodic activity in EI neural networks. Lecture Notes in Computer Science, 2415:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150039.
pdf.

Graham:2002:MFA

[1576] Bruce Graham. Multiple forms of activity-dependent plasticity enhance information transfer at a dynamic synapse. Lecture Notes in Computer Science, 2415:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150045.
pdf.

Caputo:2002:SCK

[1577] Barbara Caputo and Heinrich Niemann. Storage capacity of kernel associative memories. Lecture Notes in Computer Science, 2415:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150051.
pdf.

Lucke:2002:MDU

[1578] Jörg Lücke, Christoph von der Malsburg, and Rolf P. Würtz. Macrocolumns as decision units. Lecture Notes in Computer Science, 2415:57– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150057.
pdf.

Wennekers:2002:NAS

[1579] Thomas Wennekers. Nonlinear analysis of simple cell tuning in visual cortex. Lecture Notes in Computer Science, 2415:63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24150063.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24150063.
pdf.

Rowcliffe:2002:CWI

[1580] Phill Rowcliffe, Jianfeng Feng, and Hilary Buxton. Clustering within integrate-and-fire neurons for image segmentation. Lecture Notes in Computer Science, 2415:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150069.
pdf.

deOliveira:2002:SDU

[1581] Rogério de Oliveira and Luiz Henrique Alves Monteiro. Symmetry detection using global-locally coupled maps. Lecture Notes in Computer Science, 2415:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150075.
pdf.

Berkes:2002:ASF

[1582] Pietro Berkes and Laurenz Wiskott. Applying slow feature analysis to image sequences yields a rich repertoire of complex cell properties. Lecture Notes in Computer Science, 2415:81– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150081.
pdf.

Strosslin:2002:CMS

[1583] Thomas Strösslin, Christophe Krebser, Angelo Arleo, and Wulfram Gerstner. Combining multimodal sensory input for spatial learning. Lecture Notes in Computer Science, 2415:87– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150087.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2415/24150087.
pdf.

Kupper:2002:NNM

[1584] Rüdiger Kupper and Reinhard Eckhorn. A neural network model generating invariance for visual distance. Lecture Notes in Computer Science, 2415: 93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150093.
pdf.

Rybak:2002:MNC

[1585] Ilya A. Rybak, Dmitry G. Ivashko, Boris I. Prilutsky, M. Anthony Lewis, and John K. Chapin. Modeling neural control of locomotion: Integration of reflex circuits with CPG. Lecture Notes in Computer Science, 2415:99– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150099.
pdf.

Nevado:2002:CIE

[1586] Angel Nevado, Malcolm P. Young, and Stefano Panzeri. Comparing the information encoded by different brain areas with functional imaging techniques. Lecture Notes in Computer Science, 2415:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150105.
                                        pdf.

Mattia:2002:MFP

[1587] Maurizio Mattia and Paolo Del Giudice. Mean-field population dynamics of spiking neurons with random synaptic delays. Lecture Notes in Computer Science, 2415:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150111.
                                        pdf.

Mtetwa:2002:SRF

[1588] Nhamo Mtetwa, Leslie S. Smith, and Amir Hussain. Stochastic resonance and finite resolution in a network of leaky integrate-and-fire neurons. Lecture Notes in Computer Science, 2415: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150117.
                                        pdf.

Auer:2002:RCD

[1589] Peter Auer, Harald Burgsteiner, and Wolfgang Maass. Reducing communication for distributed learning in neural networks. Lecture Notes in Computer Science, 2415:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150123.
pdf.

Dominguez:2002:FDQ

[1590] David R. C. Dominguez, E. Korutcheva, W. K. Theumann, and R. Erichsen Jr. Flow diagrams of the quadratic neural network. Lecture Notes in Computer Science, 2415:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150129.
pdf.

Mongillo:2002:DPC

[1591] Gianluigi Mongillo and Daniel J.

Amit. Dynamics of a plastic cortical network. Lecture Notes in Computer Science, 2415:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150135.
pdf.

Giugliano:2002:NMC

[1592] Michele Giugliano, Giancarlo La Camera, Alexander Rauch, Hans-Rudolf Lüscher, and Stefano Fusi. Nonmonotonic current-to-rate response function in a novel integrate-and-fire model neuron. Lecture Notes in Computer Science, 2415:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150141.
                                        pdf.

Meyer:2002:SWE

[1593] Kris De Meyer, J. Mark Bishop, and Slawomir J. Nasuto. Smallworld effects in lattice stochastic diffusion search. Lecture Notes in Computer Science, 2415:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150147.
                                        pdf.

Iske:2002:DSN

[1594] Burkhard Iske, Axel Löffler, and Ulrich Rückert. A direction sensitive network based on a biophysical neurone model. Lecture Notes in Computer Science, 2415:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150153.
                                        pdf.

Latorre:2002:CTR

[1595] Roberto Latorre, Francisco B. Rodríguez, and Pablo Varona. Characterization of triphasic rhythms in central pattern generators (I): Interspike interval analysis. Lecture Notes in Computer Science, 2415:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150160.
pdf.

Rodriguez:2002:CTR

[1596] Francisco B. Rodríguez, Roberto Latorre, and Pablo Varona. Characterization of triphasic rhythms in central pattern generators (II): Burst information analysis. Lecture Notes in Computer Science, 2415:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150167.
pdf.

Ferrandez:2002:NCA

[1597] J. M. Ferrández, M. Bongard, F. García de Quirós, J. A. Bolea, and E. Fernández. Neural coding analysis in retinal ganglion cells using information theory. Lecture Notes in Computer Science, 2415:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150174.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2415/24150174.
                                        pdf.

Camera:2002:FRA

[1598] Giancarlo La Camera, Alexander Rauch, Walter Senn, Hans-R. Lüscher, and Stefano Fusi. Firing rate adaptation without losing sensitivity to input fluctuations. Lecture Notes in Computer Science, 2415:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150180.
                                        pdf.

Sterratt:2002:DMI

[1599] David C. Sterratt and Arjen van Ooyen. Does morphology influence temporal plasticity? Lecture Notes in Computer Science, 2415:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150186.
                                        pdf.

Johansson:2002:ANN

[1600] Christopher Johansson, Anders Sandberg, and Anders Lansner. Attractor neural networks with hypercolumns. Lecture Notes in Computer Science, 2415:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2415/24150192.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2415/24150192.
pdf.

Sanchez:2002:EDM

[1601] Eduardo Sánchez, S. Barro, and A. Canedo. Edge detection and motion discrimination in the cuneate nucleus. Lecture Notes in Computer Science, 2415:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24150198.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24150198.
pdf.

Viretta:2002:ETS

[1602] Alessandro Usseglio Viretta, Stefano Fusi, and Shih-Chii Liu. Encoding the temporal statistics of Markovian sequences of stimuli in recurrent neuronal networks. Lecture Notes in Computer Science, 2415:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150204.
pdf.

Charles:2002:MSE

[1603] Darryl Charles, Jos Koetsier, Donald MacDonald, and Colin Fyfe. Multistream exploratory projection pursuit for the formation of complex cells similar to visual cortical neurons. Lecture Notes in Computer Science, 2415:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150210.
                                        pdf.

Garcia-Cordova:2002:CNC

[1604] Francisco García-Córdova, Javier Molina-Vilaplana, and Juan López-Coronado. A corticospinal network for control of voluntary movements of a physiologically based experimental platform. Lecture Notes in Computer Science, 2415:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150216.
                                        pdf.

Moreno:2002:FRG

[1605] Rubén Moreno and Néstor Parga. Firing rate for a generic integrateand-fire neuron with exponentially correlated input. Lecture Notes in Computer Science, 2415:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150223.
                                        pdf.

Otterpohl:2002:IPD

[1606] Jens R. Otterpohl and K. Pawelzik. Iterative population decoding based on prior beliefs. Lecture Notes in Computer Science, 2415:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150229.
pdf.

Camera:2002:WNR

[1607] Giancarlo La Camera, Stefano Fusi, Walter Senn, Alexander Rauch, and Hans-R. Lüscher. When NMDA receptor conductances increase interspike interval variability. Lecture Notes in Computer Science, 2415:235– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150235.
pdf.

Fusi:2002:SDS

[1608] Stefano Fusi. Spike-driven synaptic plasticity for learning correlated patterns of asynchronous activity. Lecture Notes in Computer Science, 2415:241– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150241.
pdf.

Sanchez-Montanes:2002:MHC

[1609] Manuel Sánchez-Montañés, Luis F.

Lago-Fernández, Nazareth P. Castellanos, Ángel Merchán-Pérez, Jon I. Arellano, and Javier DeFelipe. A model of human cortical microcircuits for the study of the development of epilepsy. Lecture Notes in Computer Science, 2415:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150248.
pdf.

Maass:2002:CPN

[1610] Wolfgang Maass. On the computational power of neural microcircuit models: Pointers to the literature. Lecture Notes in Computer Science, 2415: 254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150254.
pdf.

Gelenbe:2002:NCP

[1611] Erol Gelenbe, Ricardo Lent, and Zhiguang Xu. Networking with cognitive packets. Lecture Notes in Computer Science, 2415:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150259.
pdf.

Wallace:2002:EMC

[1612] J. G. Wallace and K. Bluff. Episodic memory: A connectionist interpretation. Lecture Notes in Computer Science, 2415:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2415/24150265.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2415/24150265.pdf.

Frezza-Buet:2002:ASS

[1613] Hervé Frezza-Buet. Action scheme scheduling with a neural architecture: A prefrontal cortex approach. Lecture Notes in Computer Science, 2415:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150271.
pdf.

Stoianov:2002:AAB

[1614] Ivilin Stoianov, Marco Zorzi, Suzanna Becker, and Carlo Umilta. Associative arithmetic with Boltzmann machines: The role of number representations. Lecture Notes in Computer Science, 2415:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150277.
pdf.

Eck:2002:LLT

[1615] Douglas Eck and Jürgen Schmidhuber. Learning the long-term structure of the blues. Lecture Notes in Computer Science, 2415:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150284.
                                        pdf.

Bua:2002:RNN

[1616] Antonella Bua, Marco Gori, and Fabrizio Santini. Recursive neural networks applied to discourse representation theory. Lecture Notes in Computer Science, 2415:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150290.
                                        pdf.

Garfield:2002:RNL

[1617] Sheila Garfield and Stefan Wermter. Recurrent neural learning for helpdesk call routing. Lecture Notes in Computer Science, 2415:296–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/2415/24150296.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2415/24150296.pdf.

Korczak:2002:AEM

[1618] Jerzy Korczak and Emmanuel Blindauer. An approach to encode multilayer perceptrons. Lecture Notes in Computer Science, 2415:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150302.
pdf.

Bishop:2002:DKR

[1619] J. Mark Bishop, Slawomir J. Nasuto, and Kris De Meyer. Dynamic knowledge representation in connectionist systems. Lecture Notes in Computer Science, 2415:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150308.
pdf.

Gutierrez:2002:GCC

[1620] Germán Gutiérrez, Inés M. Galván, José M. Molina, and Araceli Sanchis. Generative capacities of cellular automata codification for evolution of NN codification. Lecture Notes in Computer Science, 2415:314– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150314.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2415/24150314.
                                        pdf.

Buldain:2002:EMR

[1621] J. David Buldain. Entropic measures with radial basis units. Lecture Notes in Computer Science, 2415:323– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150323.
                                        pdf.

Sonnenburg:2002:NMS

[1622] Sören Sonnenburg, Gunnar Rätsch, Arun Jagota, and Klaus-Robert Müller. New methods for splice site recognition. Lecture Notes in Computer Science, 2415:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150329.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150329.
                                        pdf.

Ito:2002:WCL

[1623] Yoshifusa Ito. A weak condition on linear independence of unscaled shifts of a function and finite mappings by neural networks. Lecture Notes in Computer Science, 2415:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150337.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2415/24150337.
pdf.

Azhar:2002:IWM

[1624] Ali Syed Saad Azhar and Hussain N. Al-Duwaish. Identification of Wiener model using radial basis functions neural networks. Lecture Notes in Computer Science, 2415:344– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150344.
pdf.

Welling:2002:NLA

[1625] Max Welling and Geoffrey E. Hinton.

A new learning algorithm for mean field Boltzmann machines. Lecture Notes in Computer Science, 2415:351– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150351.
pdf.

Chen:2002:CRB

[1626] Hsin Chen and Alan Murray. A continuous restricted Boltzmann machine with a hardware-amenable learning algorithm. Lecture Notes in Computer Science, 2415:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150358.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150358.
                                        pdf.

Herrero-Jaraba:2002:HRG

[1627] J. Elías Herrero-Jaraba, Carlos Orrite-Uruñuela, David Buldain, and Armando Roy-Yarza. Human recognition by gait analysis using neural networks. Lecture Notes in Computer Science, 2415:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2415/24150364.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150364.
                                        pdf.

Hammer:2002:LVQ

[1628] Barbara Hammer, Marc Strickert, and Thomas Villmann. Learning vector quantization for multimodal data. Lecture Notes in Computer Science, 2415: 370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150370.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150370.
                                        pdf.

Galicki:2002:LDN

[1629] Mirosław Galicki, Lutz Leistritz, and Herbert Witte. Learning the dynamic neural networks with the improvement of generalization capabilities. Lecture Notes in Computer Science, 2415:377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2415/24150377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150377.
pdf.

Bakker:2002:MCN

[1630] Bart Bakker and Tom Heskes.

Model clustering for neural network ensembles. Lecture Notes in Computer Science, 2415:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150383.
pdf.

Martin:2002:DCP

[1631] José Luis Fernández-Villacañas Martín and Mónica Sierra Sánchez. Does crossover probability depend on fitness and Hamming differences in genetic algorithms? Lecture Notes in Computer Science, 2415:389–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150389.
pdf.

Besada-Juez:2002:EFR

[1632] Jesús Manuel Besada-Juez and Miguel A.

Sanz-Bobi. Extraction of fuzzy rules using sensibility analysis in a neural network. Lecture Notes in Computer Science, 2415:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150395.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150395.
                                        pdf.

Albrecht:2002:SAR

[1633] Andreas A. Albrecht, Staal A. Vinterbo, C. K. Wong, and Lucila Ohno-Machado. A simulated annealing and resampling method for training perceptrons to classify geneexpression data. Lecture Notes in Computer Science, 2415:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150401.
                                        pdf.

Alaiz-Rodriguez:2002:NMC

[1634] Rocío Alaiz-Rodríguez and Jesús Cid-Sueiro. Neural minimax classifiers. Lecture Notes in Computer Science, 2415: 408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150408.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150408.
                                        pdf.

Lauer:2002:SPE

[1635] Martin Lauer. Sampling parameters to estimate a mixture distribution with unknown size. Lecture Notes in Computer Science, 2415:414– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150414.
pdf.

Verikas:2002:SNN

[1636] Antanas Verikas, Arunas Lipnickas, and Kerstin Malmqvist. Selecting neural networks for making a committee decision. Lecture Notes in Computer Science, 2415:420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150420.
pdf.

Fleury:2002:HAM

[1637] Patrice Fleury, Alan F. Murray, and Martin Reekie. High-accuracy mixedsignal VLSI for weight modification in contrastive divergence learning. Lecture Notes in Computer Science, 2415: 426–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150426.
pdf.

Weng:2002:DDG

[1638] Sebastian Weng and Jochen J. Steil.

Data driven generation of interactions for feature binding and relaxation labeling. Lecture Notes in Computer Science, 2415:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150432.
                                        pdf.

Bote-Lorenzo:2002:HTS

[1639] Miguel L. Bote-Lorenzo, Yannis A. Dimitriadis, and Eduardo Gómez-Sánchez. A hybrid two-stage fuzzy ARTMAP and LVQ neurofuzzy system for on-line handwriting recognition. Lecture Notes in Computer Science, 2415:438–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150438.
                                        pdf.

Ferrari-Trecate:2002:NLM

[1640] Giancarlo Ferrari-Trecate and Marco Muselli. A new learning method for piecewise linear regression. Lecture Notes in Computer Science, 2415:444– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150444.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150444.
                                        pdf.

Graepel:2002:SAM

[1641] Thore Graepel and Nicol N. Schraudolph. Stable adaptive momentum for rapid online learning in nonlinear systems. Lecture Notes in Computer Science, 2415:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150450.
pdf.

Erdogmus:2002:PEP

[1642] Deniz Erdogmus, Jose C. Principe, Luis Vielva, and David Luengo. Potential energy and particle interaction approach for learning in adaptive systems. Lecture Notes in Computer Science, 2415:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150456.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150456.
pdf.

Preciado:2002:PLA

[1643] Víctor M. Preciado. Piecewiselinear approximation of any smooth output function on the cellular neural network. Lecture Notes in Computer Science, 2415:462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150462.
pdf.

Anguita:2002:MBM

[1644] Davide Anguita and Matteo Gagliolo. MDL based model selection for relevance vector regression. Lecture Notes in Computer Science, 2415:468– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150468.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150468.
                                        pdf.

Koppen:2002:TKN

[1645] Mario Köppen. On the training of a Kolmogorov network. Lecture Notes in Computer Science, 2415:474– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150474.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150474.
                                        pdf.

Kwak:2002:NMF

[1646] Nojun Kwak and Chong-Ho Choi. A new method of feature extraction and its stability. Lecture Notes in Computer Science, 2415:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150480.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150480.
                                        pdf.

Benabdeslem:2002:VAW

[1647] Khalid Benabdeslem, Younes Bennani, and Eric Janvier. Visualization and analysis of Web navigation data. Lecture Notes in Computer Science, 2415: 486–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150486.
pdf.

Oba:2002:MVE

[1648] Shigeyuki Oba, Masa aki Sato, Ichiro Takemasa, Morito Monden, Ken ichi Matsubara, and Shin Ishii. Missing value estimation using mixture of PCAs. Lecture Notes in Computer Science, 2415:492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150492.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150492.
pdf.

Svensson:2002:HPM

[1649] Jakob Svensson, Manfred von Hellermann, and Ralf König. High precision measurement of fuel density profiles in nuclear fusion plasmas. Lecture Notes in Computer Science, 2415: 498–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150498.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2415/24150498.
                                        pdf.

Grabczewski:2002:HFD

[1650] Krzysztof Grabczewski and Wlodzislaw Duch. Heterogeneous forests of decision trees. Lecture Notes in Computer Science, 2415:504–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150504.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150504.
                                        pdf.

Bounkong:2002:ICA

[1651] Stéphane Bounkong, David Saad, and David Lowe. Independent component analysis for domain independent watermarking. Lecture Notes in Computer Science, 2415:510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150510.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150510.
                                        pdf.

Ragg:2002:AML

[1652] Thomas Ragg, Frank Padberg, and Ralf Schoknecht. Applying machine learning to solve an estimation problem in software inspections. Lecture Notes in Computer Science, 2415:516– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150516.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2415/24150516.
pdf.

Yoshioka:2002:CGE

[1653] Taku Yoshioka, Ryouko Morioka, Kazuo Kobayashi, Shigeyuki Oba, Naotake Ogawsawara, and Shin Ishii. Clustering of gene expression data by mixture of PCA models. Lecture Notes in Computer Science, 2415:522– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150522.
pdf.

Sugiyama:2002:SRP

[1654] Masashi Sugiyama and Klaus-Robert Müller. Selecting ridge parameters in infinite dimensional hypothesis spaces. Lecture Notes in Computer Science, 2415:528–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150528.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150528.
pdf.

Kanamori:2002:NSA

[1655] Takafumi Kanamori. A new sequential algorithm for regression problems by using mixture distribution. Lecture Notes in Computer Science, 2415:535– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150535.
                                        pdf.

Lopez:2002:NBC

[1656] Damián López and María José Castro. Neural-based classification of blocks from documents. Lecture Notes in Computer Science, 2415:541–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150541.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150541.
                                        pdf.

Salcedo-Sanz:2002:FSG

[1657] Sancho Salcedo-Sanz, Mario Prado-Cumplido, Fernando Pérez-Cruz, and Carlos Bousoño-Calzón. Feature selection via genetic optimization. Lecture Notes in Computer Science, 2415:547– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150547.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150547.
                                        pdf.

Gonzalez:2002:NNC

[1658] Jesús González, Ignacio Rojas, and Héctor Pomares. Neural networks, clustering techniques, and function approximation problems. Lecture Notes in Computer Science, 2415:553–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150553.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150553.
pdf.

Gonzalez:2002:ETN

[1659] Jésus González, Ignacio Rojas, Hector Pomares, Alberto Prieto, and K. Goser. Evolutionary training of neuro-fuzzy patches for function approximation. Lecture Notes in Computer Science, 2415:559–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150559.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150559.
pdf.

Martinez:2002:URN

[1660] César Martínez, Alfons Juan, and Francisco Casacuberta. Using recurrent neural networks for automatic chromosome classification. Lecture Notes in Computer Science, 2415:565– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150565.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150565.
pdf.

Dimitriadou:2002:MEA

[1661] Evgenia Dimitriadou, Andreas Weingessel, and Kurt Hornik. A mixed ensemble approach for the semisupervised problem. Lecture Notes in Computer Science, 2415:571–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150571.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150571.
pdf.

Mateos:2002:UPS

[1662] Alvaro Mateos, Javier Herrero, and Joaquín Dopazo. Using perceptrons for supervised classification of DNA microarray samples: Obtaining the optimal level of information and finding differentially expressed genes. Lecture Notes in Computer Science, 2415:577– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150577.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150577.
pdf.

Gavin:2002:LBT

[1663] Gerald Gavin and Olivier Teytaud. Lower bounds for training and leave-one-out estimates of the generalization error. Lecture Notes in Computer Science, 2415:583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150583.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150583.
pdf.

Salmeron:2002:SSQ

[1664] Moisés Salmerón, Julio Ortega, Carlos García Puntonet, Alberto Prieto, and Ignacio Rojas. SSA, SVD, QRcp, and RBF model reduction. Lecture Notes in Computer Science, 2415: 589–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150589.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150589.
pdf.

Leisink:2002:LAB

[1665] Martijn A. R. Leisink, Hilbert J. Kappen, and Han G. Brunner. Linkage analysis: A Bayesian approach. Lecture Notes in Computer Science, 2415: 595–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150595.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150595.
pdf.

Sperduti:2002:LSS

[1666] Alessandro Sperduti. On linear separability of sequences and structures. Lecture Notes in Computer Science, 2415: 601–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150601.
pdf.

Roth:2002:SBM

[1667] Volker Roth, Mikio L. Braun, Tilman Lange, and Joachim M. Buhmann. Stability-based model order selection in clustering with applications to gene expression data. Lecture Notes in Computer Science, 2415:607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150607.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150607.
pdf.

Zufiria:2002:EBR

[1668] Pedro J. Zufiria and Carlos Rivero. EM-based radial basis function training with partial information. Lecture Notes in Computer Science, 2415:613– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150613.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150613.
pdf.

Giudici:2002:SSL

[1669] Matteo Giudici, Filippo Queirolo, and Maurizio Valle. Stochastic supervised learning algorithms with local and adaptive learning rate for recognising hand-written characters. Lecture Notes in Computer Science, 2415:619– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150619.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2415/24150619.
pdf.

Sierra:2002:IOF

[1670] Alejandro Sierra and Fernando Corbacho. Input and output feature selection. Lecture Notes in Computer Science, 2415:625–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24150625.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24150625.
pdf.

Lago-Fernandez:2002:OEH

[1671] Luis F. Lago-Fernández and Fernando Corbacho. Optimal extraction of hidden causes. Lecture Notes in Computer Science, 2415:631–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150631.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150631.
pdf.

Sanchez-Montanes:2002:TNI

[1672] Manuel A. Sánchez-Montañés and Fernando J. Corbacho. Towards a new information processing measure for neural computation. Lecture Notes in Computer Science, 2415:637–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150637.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2415/24150637.
pdf.

Stensmo:2002:SEP

[1673] Magnus Stensmo. A scalable and efficient probabilistic information retrieval and text mining system. Lecture Notes in Computer Science, 2415: 643–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150643.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150643.
pdf.

MacDonald:2002:MML

[1674] Donald MacDonald, Emilio Corchado, Colin Fyfe, and Erzsebet Merenyi. Maximum and minimum likelihood Hebbian learning for exploratory projection pursuit. Lecture Notes in Computer Science, 2415:649– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150649.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150649.
pdf.

Gers:2002:LCS

[1675] Felix A. Gers, Juan Antonio Pérez-Ortiz, Douglas Eck, and Jürgen Schmidhuber. Learning context sensitive languages with LSTM trained with Kalman filters. Lecture Notes in Computer Science, 2415:655–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2415/24150655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150655.
pdf.

Yoshimoto:2002:HMS

[1676] Junichiro Yoshimoto, Shin Ishii, and Masa aki Sato. Hierarchical model selection for NGnet based on variational Bayes inference. Lecture Notes in Computer Science, 2415:661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150661.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150661.
pdf.

Conan-Guez:2002:MLP

[1677] Brieuc Conan-Guez and Fabrice Rossi.

Multi-layer perceptrons for functional data analysis: A projection based approach. Lecture Notes in Computer Science, 2415:667–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150667.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150667.
pdf.

Dorronsoro:2002:NGM

[1678] José R. Dorronsoro and Ana González.

Natural gradient and multiclass NLDA networks. Lecture Notes in Computer Science, 2415:673–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150673.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150673.
                                        pdf.

Cawley:2002:GTA

[1679] Gavin C. Cawley and Nicola L. C. Talbot. A greedy training algorithm for sparse least-squares support vector machines. Lecture Notes in Computer Science, 2415:681–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150681.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150681.
                                        pdf.

Cherkassky:2002:SMP

[1680] Vladimir Cherkassky and Yunqian Ma. Selection of meta-parameters for support vector regression. Lecture Notes in Computer Science, 2415:687– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150687.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150687.
                                        pdf.

Graepel:2002:KMC

[1681] Thore Graepel. Kernel matrix completion by semidefinite programming. Lecture Notes in Computer Science, 2415: 694–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2415/24150694.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150694.
pdf.

Sato:2002:ISK

[1682] Masa aki Sato and Shigeyuki Oba. Incremental sparse kernel machine. Lecture Notes in Computer Science, 2415: 700–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150700.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150700.
pdf.

Rakotomamonjy:2002:FKL

[1683] Alain Rakotomamonjy and Stéphane Canu. Frame kernels for learning. Lecture Notes in Computer Science, 2415: 707–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150707.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150707.
pdf.

Brabanter:2002:RCV

[1684] Jos De Brabanter, Kristiaan Pelckmans, Johan A. K. Suykens, and Joos Vandewalle. Robust crossvalidation score function for non-linear function estimation. Lecture Notes in Computer Science, 2415:713–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2415/24150713.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150713.
                                        pdf.

Hamers:2002:CSR

[1685] Bart Hamers, Johan A. K. Suykens, and Bart De Moor. Compactly supported RBF kernels for sparsifying the gram matrix in LS-SVM regression models. Lecture Notes in Computer Science, 2415:720–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150720.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150720.
                                        pdf.

Tsuda:2002:LOK

[1686] Koji Tsuda and Motoaki Kawanabe. The leave-one-out kernel. Lecture Notes in Computer Science, 2415:727– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150727.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150727.
                                        pdf.

Borer:2002:SVR

[1687] Silvio Borer and Wulfram Gerstner. Support vector representation of multi-categorical data. Lecture Notes in Computer Science, 2415:733– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2415/24150733.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150733.
pdf.

Takahashi:2002:RNK

[1688] Takashi Takahashi and Takio Kurita.

Robust de-noising by kernel PCA. Lecture Notes in Computer Science, 2415: 739–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150739.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150739.
pdf.

Meinicke:2002:MCC

[1689] Peter Meinicke, Thorsten Twellmann, and Helge Ritter. Maximum contrast classifiers. Lecture Notes in Computer Science, 2415:745–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150745.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150745.
pdf.

Perez-Cruz:2002:PMC

[1690] Fernando Pérez-Cruz and Antonio Artés-Rodríguez. Puncturing multiclass support vector machines. Lecture Notes in Computer Science, 2415:751– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150751.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2415/24150751.
                                        pdf.

Perez-Cruz:2002:MDF

[1691] Fernando Pérez-Cruz, Gustavo Camps-Valls, Emilio Soria-Olivas, Juan José Pérez-Ruixo, Aníbal R. Figueiras-Vidal, and Antonio Artés-Rodríguez. Multi-dimensional function approximation and regression estimation. Lecture Notes in Computer Science, 2415:757– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150757.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150757.
                                        pdf.

Moguerza:2002:DNC

[1692] Javier M. Moguerza, Alberto Muñoz, and Manuel Martín-Merino. Detecting the number of clusters using a support vector machine approach. Lecture Notes in Computer Science, 2415:763– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150763.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150763.
                                        pdf.

Siolas:2002:MPP

[1693] George Siolas and Florence d’Alché Buc. Mixtures of probabilistic PCAs and Fisher kernels for word and document modeling. Lecture Notes in Computer Science, 2415:769–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150769.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150769.
pdf.

Sato:2002:RLB

[1694] Masa aki Sato, Yutaka Nakamura, and Shin Ishii. Reinforcement learning for biped locomotion. Lecture Notes in Computer Science, 2415:777– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150777.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150777.
pdf.

Hulse:2002:DNS

[1695] Martin Hülse and Frank Pasemann.

Dynamical neural schmitt trigger for robot control. Lecture Notes in Computer Science, 2415:783–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150783.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150783.
pdf.

McMinn:2002:EAN

[1696] David McMinn, Grant Maxwell, and Christopher MacLeod. Evolutionary artificial neural networks for quadruped locomotion. Lecture Notes in Computer Science, 2415:789–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150789.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150789.
                                        pdf.

Conradt:2002:SMO

[1697] Jörg Conradt, Pascal Simon, Michel Pescatore, and Paul F. M. J. Verschure. Saliency maps operating on stereo images detect landmarks and their distance. Lecture Notes in Computer Science, 2415:795–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150795.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150795.
                                        pdf.

Herzallah:2002:NAM

[1698] Randa Herzallah and David Lowe. A novel approach to modelling and exploiting uncertainty in stochastic control systems. Lecture Notes in Computer Science, 2415:801–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150801.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150801.
                                        pdf.

Haber:2002:TWP

[1699] Rodolfo E. Haber, A. Alique, and J. R. Alique. Tool wear prediction in milling using neural networks. Lecture Notes in Computer Science, 2415:807– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150807.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150807.
pdf.

Schoknecht:2002:SRL

[1700] Ralf Schoknecht and Martin Riedmiller. Speeding-up reinforcement learning with multi-step actions. Lecture Notes in Computer Science, 2415: 813–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150813.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150813.
pdf.

Todorovic:2002:EKF

[1701] Branimir Todorovic, Miomir Stankovic, and Claudio Moraga. Extended Kalman filter trained recurrent radial basis function network in nonlinear system identification. Lecture Notes in Computer Science, 2415:819– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150819.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150819.
pdf.

Hubner:2002:IMP

[1702] Wolfgang Hübner and Hanspeter A.

Mallot. Integration of metric place relations in a landmark graph. Lecture Notes in Computer Science, 2415:825– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150825.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150825.
                                        pdf.

Simon:2002:HOC

[1703] Steffen Simon, Friedhelm Schwenker, Hans A. Kestler, Gerhard Kraetzschmar, and Günther Palm. Hierarchical object classification for autonomous mobile robots. Lecture Notes in Computer Science, 2415:831– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150831.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150831.
                                        pdf.

Becerra:2002:SPG

[1704] J. A. Becerra, R. J. Duro, and J. Santos. Self pruning Gaussian synapse networks for behavior based robots. Lecture Notes in Computer Science, 2415: 837–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150837.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150837.
                                        pdf.

Benzaquen:2002:SOC

[1705] Samuel Benzaquen and Carolina Chang. Second-order conditioning in mobile robots. Lecture Notes in Computer Science, 2415:844–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150844.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150844.
pdf.

Lichtensteiger:2002:OSM

[1706] Lukas Lichtensteiger and Rolf Pfeifer.

An optimal sensor morphology improves adaptability of neural network controllers. Lecture Notes in Computer Science, 2415:850–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150850.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150850.
pdf.

deAngulo:2002:LIK

[1707] Vicente Ruiz de Angulo and Carme Torras. Learning inverse kinematics via cross-point function decomposition. Lecture Notes in Computer Science, 2415:856–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24150856.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24150856.
pdf.

Lopez-Rubio:2002:PCA

[1708] Ezequiel López-Rubio, José Muñoz-Pérez, and José Antonio Gómez-Ruiz.

The principal components analysis self-organizing map. Lecture Notes in Computer Science, 2415:865–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150865.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150865.
pdf.

Pampalk:2002:USD

[1709] Elias Pampalk, Andreas Rauber, and Dieter Merkl. Using smoothed data histograms for cluster visualization in self-organizing maps. Lecture Notes in Computer Science, 2415:871–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150871.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150871.
pdf.

Hammer:2002:RES

[1710] Barbara Hammer, Andreas Rechtien, Marc Strickert, and Thomas Villmann. Rule extraction from selforganizing networks. Lecture Notes in Computer Science, 2415:877–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150877.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150877.
pdf.

Hirose:2002:PSO

[1711] Akira Hirose and Tomoyuki Nagashima. Predictive self-organizing map for vector quantization of migratory signals. Lecture Notes in Computer Science, 2415:884–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150884.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150884.
pdf.

Lebbah:2002:CTM

[1712] Mustapha Lebbah, Christian Chabanon, Fouad Badran, and Sylvie Thiria. Categorical topological map. Lecture Notes in Computer Science, 2415:890–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24150890.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24150890.
pdf.

Panchev:2002:STD

[1713] Christo Panchev, Stefan Wermter, and Huixin Chen. Spike-timing dependent competitive learning of integrate-and-fire neurons with active dendrites. Lecture Notes in Computer Science, 2415: 896–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150896.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150896.
pdf.

Saalbach:2002:PSO

[1714] Axel Saalbach, Gunther Heidemann, and Helge Ritter. Parametrized SOMs for object recognition and pose estimation. Lecture Notes in Computer Science, 2415:902–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150902.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150902.
                                        pdf.

Plebe:2002:ETS

[1715] Alessio Plebe. An effective traveling salesman problem solver based on self-organizing map. Lecture Notes in Computer Science, 2415:908–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150908.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150908.
                                        pdf.

Verbeek:2002:CPC

[1716] Jakob J. Verbeek, Nikos Vlassis, and Ben Kröse. Coordinating principal component analyzers. Lecture Notes in Computer Science, 2415:914–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150914.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150914.
                                        pdf.

Viviani:2002:LIS

[1717] Roberto Viviani. Lateral interactions in self-organizing maps. Lecture Notes in Computer Science, 2415:920– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150920.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150920.
pdf.

Lensu:2002:CSS

[1718] Anssi Lensu and Pasi Koikkalainen.

Complexity selection of the selforganizing map. Lecture Notes in Computer Science, 2415:927–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150927.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150927.
pdf.

Lee:2002:NPI

[1719] John A. Lee and Michel Verleysen. Nonlinear projection with the isotop method. Lecture Notes in Computer Science, 2415:933–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150933.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150933.
pdf.

Claussen:2002:ALD

[1720] Jens Christian Claussen and Heinz Georg Schuster. Asymptotic level density of the elastic net self-organizing feature map. Lecture Notes in Computer Science, 2415:939–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2415/24150939.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2415/24150939.pdf.

Fontenla-Romero:2002:LMU

[1721] Oscar Fontenla-Romero, Amparo Alonso-Betanzos, Enrique Castillo, Jose C. Principe, and Bertha Guijarro-Berdiñas. Local modeling using self-organizing maps and single layer neural networks. Lecture Notes in Computer Science, 2415:945–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150945.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150945.
pdf.

Vesanto:2002:DMB

[1722] Juha Vesanto and Mika Sulkava. Distance matrix based clustering of the self-organizing map. Lecture Notes in Computer Science, 2415:951–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150951.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2415/24150951.
pdf.

Vogiatzis:2002:MGN

[1723] Dimitrios Vogiatzis and Andreas Stafylopatis. Mapping the growing neural gas to situation calculus. Lecture Notes in Computer Science, 2415:957– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150957.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150957.
pdf.

Chiarantoni:2002:RUC

[1724] Ernesto Chiarantoni, Giuseppe Acciani, Girolamo Fornarelli, and Silvano Vergura. Robust unsupervised competitive neural network by local competitive signals. Lecture Notes in Computer Science, 2415:963–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150963.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150963.
pdf.

Panangadan:2002:GSC

[1725] Anand Panangadan and Michael G.

Dyer. Goal sequencing for construction agents in a simulated environment. Lecture Notes in Computer Science, 2415: 969–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150969.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150969.
                                        pdf.

Barreto:2002:NMD

[1726] Guilherme de A. Barreto and Aluizio F. R. Araújo. Nonlinear modeling of dynamic systems with the self-organizing map. Lecture Notes in Computer Science, 2415:975–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150975.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150975.
                                        pdf.

Koskela:2002:IRF

[1727] Markus Koskela, Jorma Laaksonen, and Erkki Oja. Implementing relevance feedback as convolutions of local neighborhoods on self-organizing maps. Lecture Notes in Computer Science, 2415: 981–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24150981.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24150981.
                                        pdf.

Hunter:2002:PSO

[1728] Andrew Hunter and Richard Lee Kennedy. A Pareto self-organizing map. Lecture Notes in Computer Science, 2415:987–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2415/24150987.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2415/24150987.pdf.

Georgakis:2002:SVB

[1729] Apostolos Georgakis, Costas Kotropoulos, and Ioannis Pitas. A SOM variant based on the Wilcoxon test for document organization and retrieval. Lecture Notes in Computer Science, 2415: 993–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150993.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150993.
pdf.

Peltonen:2002:LMA

[1730] Jaakko Peltonen, Arto Klami, and Samuel Kaski. Learning more accurate metrics for self-organizing maps. Lecture Notes in Computer Science, 2415: 999–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24150999.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24150999.
pdf.

Blanco:2002:CVH

[1731] Ignacio Díaz Blanco, Abel A. Cuadrado Vega, and Alberto B. Diez González. Correlation visualization of high dimensional data using topographic maps. Lecture Notes in Computer Science, 2415:1005–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/2415/24151005.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2415/24151005.pdf.

Flexer:2002:CUS

[1732] Arthur Flexer, Georg Gruber, and Georg Dorffner. Continuous unsupervised sleep staging based on a single EEG signal. Lecture Notes in Computer Science, 2415:1013–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151013.
                                        pdf.

Chiu:2002:FAB

[1733] Kai Chun Chiu and Lei Xu. Financial APT-based Gaussian TFA learning for adaptive portfolio management. Lecture Notes in Computer Science, 2415:1019–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2415/24151019.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151019.
                                        pdf.

Liu:2002:CIF

[1734] Zhiyong Liu and Lei Xu. On convergence of an iterative factor estimate algorithm for the NFA model. Lecture Notes in Computer Science, 2415:1025– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2415/24151025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151025.
pdf.

Foxall:2002:EFP

[1735] Robert J. Foxall, Gavin C. Cawley, Stephen R. Dorling, and Danilo P. Mandic. Error functions for prediction of episodes of poor air quality. Lecture Notes in Computer Science, 2415:1031– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151031.
pdf.

Sanz-Gonzalez:2002:AIS

[1736] José L. Sanz-González and Francisco Álvarez-Vaquero. Adaptive importance sampling technique for neural detector training. Lecture Notes in Computer Science, 2415:1037–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151037.
pdf.

vanLint:2002:SSN

[1737] Hans van Lint, Serge P. Hoogendoorn, and Henk J. van Zuylen. State space neural networks for freeway travel time prediction. Lecture Notes in Computer Science, 2415:1043– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151043.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151043.
                                        pdf.

Theis:2002:OIG

[1738] Fabian J. Theis, Elmar W. Lang, Tobias Westenhuber, and Carlos G. Puntonet. Overcomplete ICA with a geometric algorithm. Lecture Notes in Computer Science, 2415:1049–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151049.
                                        pdf.

Perez-Ortiz:2002:ILT

[1739] Juan A. Pérez-Ortiz, Jürgen Schmidhuber, Felix A. Gers, and Douglas Eck. Improving long-term online prediction with decoupled extended Kalman filters. Lecture Notes in Computer Science, 2415:1055–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151055.
                                        pdf.

Zimmermann:2002:MMB

[1740] Georg Zimmermann, Ralph Grothmann, Christoph Tietz, and Ralph Neuneier. Market modeling based on cognitive agents. Lecture Notes in Computer Science, 2415:1061–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151061.
pdf.

Kurimo:2002:EFL

[1741] Mikko Kurimo and Krista Lagus.

An efficiently focusing large vocabulary language model. Lecture Notes in Computer Science, 2415:1068–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151068.
pdf.

Teranishi:2002:NCB

[1742] Masaru Teranishi, Sigeru Omatu, and Toshihisa Kosaka. Neuro-classification of bill fatigue levels based on acoustic wavelet components. Lecture Notes in Computer Science, 2415:1074–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151074.
pdf.

Allende:2002:REL

[1743] Héctor Allende, Claudio Moraga, and Rodrigo Salas. Robust estimator for the learning process in neural networks applied in time series. Lecture Notes in Computer Science, 2415:1080– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151080.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151080.
                                        pdf.

Blaschke:2002:ICB

[1744] Tobias Blaschke and Laurenz Wiskott. An improved cumulant based method for independent component analysis. Lecture Notes in Computer Science, 2415:1087–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2415/24151087.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151087.
                                        pdf.

Kindermann:2002:FOC

[1745] Lars Kindermann, Achim Lewandowski, and Peter Protzel. Finding the optimal continuous model for discrete data by neural network interpolation of fractional iteration. Lecture Notes in Computer Science, 2415:1094– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151094.
                                        pdf.

Rojo-Alvarez:2002:SVR

[1746] José Luis Rojo-Álvarez, Arcadi García-Alberola, Manel Martínez-Ramón, Mariano Valdés, Aníbal R. Figueiras-Vidal, and Antonio Artés-Rodríguez. Support vector robust algorithms for non-parametric spectral analysis. Lecture Notes in Computer Science, 2415: 1100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24151100.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24151100.
pdf.

Rojo-Alvarez:2002:SVM

[1747] José Luis Rojo-Álvarez, Manel Martínez-Ramón, Aníbal R. Figueiras-Vidal, Mario de Prado-Cumplido, and Antonio Artés-Rodríguez. Support vector method for ARMA system identification: A robust cost interpretation. Lecture Notes in Computer Science, 2415:1106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151106.
pdf.

Basalyga:2002:DIH

[1748] Gleb Basalyga and Magnus Rattray. Dynamics of ICA for highdimensional data. Lecture Notes in Computer Science, 2415:1112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151112.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151112.
                                        pdf.

Boscolo:2002:BCI

[1749] Riccardo Boscolo, Hong Pan, and Vwani P. Roychowdhury. Beyond Comon’s identifiability theorem for independent component analysis. Lecture Notes in Computer Science, 2415:1119– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151119.
                                        pdf.

Hauser:2002:TPB

[1750] Alexandre Hauser, Pierre-Edouard Sottas, and José del R. Millán. Temporal processing of brain activity for the recognition of EEG patterns. Lecture Notes in Computer Science, 2415:1125– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151125.
                                        pdf.

Andreou:2002:CAO

[1751] Panayiotis Ch. Andreou, Chris Charalambous, and Spiros H. Martzoukos. Critical assessment of option pricing methods using artificial neural networks. Lecture Notes in Computer Science, 2415:1131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151131.
pdf.

Blankertz:2002:STD

[1752] Benjamin Blankertz, Christin Schäfer, Guido Dornhege, and Gabriel Curio. Single trial detection of EEG error potentials: A tool for increasing BCI transmission rates. Lecture Notes in Computer Science, 2415:1137– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151137.
pdf.

Sottas:2002:DNA

[1753] Pierre-Edouard Sottas and Wulfram Gerstner. Dynamic noise annealing for learning temporal sequences with recurrent neural networks. Lecture Notes in Computer Science, 2415:1144– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151144.
pdf.

Lopez-Risueno:2002:CNN

[1754] Gustavo López-Risueño, Jesús Grajal, Simon Haykin, and Rosa Díaz-Oliver. Convolutional neural networks for radar detection. Lecture Notes in Computer Science, 2415:1150– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151150.
                                        pdf.

Kohlmorgen:2002:SGM

[1755] Jens Kohlmorgen and Benjamin Blankertz. A simple generative model for single-trial EEG classification. Lecture Notes in Computer Science, 2415: 1156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2415/24151156.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151156.
                                        pdf.

Georgiev:2002:RBS

[1756] Pando Georgiev and Andrzej Cichocki. Robust blind source separation utilizing second and fourth order statistics. Lecture Notes in Computer Science, 2415:1162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151162.
                                        pdf.

Choi:2002:ADD

[1757] Seungjin Choi. Adaptive differential decorrelation: A natural gradient algorithm. Lecture Notes in Computer Science, 2415:1168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151168.
pdf.

Pelaez-Moreno:2002:ASL

[1758] Carmen Peláez-Moreno, Emilio Parrado-Hernández, Ascensión Gallardo-Antolín, Adrián Zambrano-Miranda, and Fernando Díaz de María. An application of SVM to lost packets reconstruction in voice-enabled services. Lecture Notes in Computer Science, 2415:1174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151174.
pdf.

Otterpohl:2002:BWL

[1759] Jens R. Otterpohl. Baum-Welch learning in discrete hidden Markov models with linear factorial constraints. Lecture Notes in Computer Science, 2415: 1180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24151180.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24151180.
pdf.

Suarez:2002:MAM

[1760] Alberto Suárez. Mixtures of autoregressive models for financial risk analysis. Lecture Notes in Computer Science, 2415:1186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151186.
                                        pdf.

Barla:2002:KBO

[1761] Annalisa Barla and Francesca Odone. Kernel-based 3D object representation. Lecture Notes in Computer Science, 2415:1195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2415/24151195.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151195.
                                        pdf.

Ghennam:2002:MSA

[1762] Souheila Ghennam and Khier Benmahammed. Multiresolution support for adaptive image restoration using neural networks. Lecture Notes in Computer Science, 2415:1201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151201.
                                        pdf.

Seguier:2002:AVS

[1763] Renaud Séguier and David Mercier. Audio-visual speech recognition one pass learning with spiking neurons. Lecture Notes in Computer Science, 2415:1207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2415/24151207.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2415/24151207.
pdf.

Turiel:2002:AIR

[1764] Antonio Turiel, Jean-Pierre Nadal, and Néstor Parga. An algorithm for image representation as independent levels of resolution. Lecture Notes in Computer Science, 2415:1213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151213.
pdf.

Gastaldo:2002:CBP

[1765] Paolo Gastaldo, Rodolfo Zunino, Ingrid Heynderickx, and Elena Vicario. Circular back-propagation networks for measuring displayed image quality. Lecture Notes in Computer Science, 2415:1219– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151219.
pdf.

Wersing:2002:ULC

[1766] Heiko Wersing and Edgar Körner.

Unsupervised learning of combination features for hierarchical recognition models. Lecture Notes in Computer Science, 2415:1225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151225.
pdf.

Aizenberg:2002:TBB

[1767] Igor Aizenberg, Taras Bregin, Constantine Butakoff, Victor Karnaukhov, Nickolay Merzlyakov, and Olga Milukova. Type of blur and blur parameters identification using neural network and its application to image restoration. Lecture Notes in Computer Science, 2415:1231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151231.
pdf.

Backer:2002:UNF

[1768] Gerriet Backer and Bärbel Mertsching. Using neural field dynamics in the context of attentional control. Lecture Notes in Computer Science, 2415:1237– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151237.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2415/24151237.
pdf.

Teichert:2002:CAA

[1769] Jens Teichert and Rainer Malaka. A component association architecture for image understanding. Lecture Notes in Computer Science, 2415:1243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151243.
pdf.

Owens:2002:NDV

[1770] Jonathan Owens, Andrew Hunter, and Eric Fletcher. Novelty detection in video surveillance using hierarchical neural networks. Lecture Notes in Computer Science, 2415:1249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151249.
pdf.

Sturzl:2002:VCD

[1771] Wolfgang Stürzl, Ulrich Hoffmann, and Hanspeter A. Mallot. Vergence control and disparity estimation with energy neurons: Theory and implementation. Lecture Notes in Computer Science, 2415:1255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2415/24151255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151255.
                                        pdf.

Ludtke:2002:PCM

[1772] Niklas Lüdtke, Richard C. Wilson, and Edwin R. Hancock. Population coding of multiple edge orientation. Lecture Notes in Computer Science, 2415:1261– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151261.
                                        pdf.

Planta:2002:NMF

[1773] Cyrill Planta, Jörg Conradt, Adrian Jencik, and Paul Verschure. A neural model of the fly visual system applied to navigational tasks. Lecture Notes in Computer Science, 2415:1268– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151268.
                                        pdf.

Shimomura:2002:NNM

[1774] Masao Shimomura, Shunji Satoh, Shogo Miyake, and Hirotomo Aso. A neural network model for pattern recognition based on hypothesis and verification with moving region of attention. Lecture Notes in Computer Science, 2415:1275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151275.
pdf.

Ceguerra:2002:AFV

[1775] Anna Ceguerra and Irena Koprinska. Automatic fingerprint verification using neural networks. Lecture Notes in Computer Science, 2415:1281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151281.
pdf.

Li:2002:FIM

[1776] Shutao Li, James T. Kwok, and Yaonan Wang. Fusing images with multiple focuses using support vector machines. Lecture Notes in Computer Science, 2415:1287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151287.
pdf.

Heittmann:2002:AVP

[1777] Arne Heittmann, Ulrich Ramacher, Daniel Matolin, Jörg Schreiter, and Rene Schüffny. An analog VLSI pulsed neural network for image segmentation using adaptive connection weights. Lecture Notes in Computer Science, 2415:1293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151293.
pdf.

Foucher:2002:KMA

[1778] Christophe Foucher, Daniel Le Guennec, and Gilles Vaucher. Kohonen maps applied to fast image vector quantization. Lecture Notes in Computer Science, 2415:1299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151299.
pdf.

Acciani:2002:UNN

[1779] Giuseppe Acciani, Ernesto Chiarantoni, Daniela Girimonte, and Cataldo Guaragnella. Unsupervised — neural network approach for efficient video description. Lecture Notes in Computer Science, 2415:1305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151305.
pdf.

Ntalianis:2002:NNR

[1780] Klimis S. Ntalianis, Nikolaos D.

Doulamis, Anastasios D. Doulamis, and Stefanos D. Kollias. Neural networks retraining for unsupervised video object segmentation of videoconference sequences. Lecture Notes in Computer Science, 2415:1312–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151312.
pdf.

Behnke:2002:LFL

[1781] Sven Behnke. Learning face localization using hierarchical recurrent networks. Lecture Notes in Computer Science, 2415:1319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151319.
pdf.

Fasel:2002:CFD

[1782] Ian R. Fasel and Javier R. Movellan. A comparison of face detection algorithms. Lecture Notes in Computer Science, 2415:1325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151325.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2415/24151325.
                                        pdf.

Boni:2002:AMS

[1783] Andrea Boni. Adaptive model selection for digital linear classifiers. Lecture Notes in Computer Science, 2415:1333– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151333.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151333.
                                        pdf.

deAngulo:2002:SLF

[1784] Vicente Ruiz de Angulo and Carme Torras. Sequential learning in feedforward networks: Proactive and retroactive interference minimization. Lecture Notes in Computer Science, 2415:1339– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151339.
                                        pdf.

Anguita:2002:AHT

[1785] Davide Anguita, Sandro Ridella, Fabio Rivieccio, and Rodolfo Zunino. Automatic hyperparameter tuning for support vector machines. Lecture Notes in Computer Science, 2415:1345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151345.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2415/24151345.
pdf.

Schraudolph:2002:CDS

[1786] Nicol N. Schraudolph and Thore Graepel. Conjugate directions for stochastic gradient descent. Lecture Notes in Computer Science, 2415:1351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151351.
pdf.

Tino:2002:ABR

[1787] Peter Tiňo and Barbara Hammer. Architectural bias in recurrent neural networks — fractal analysis. Lecture Notes in Computer Science, 2415:1359– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2415/24151359.
pdf.

Atencia:2002:CSH

[1788] Miguel A. Atencia, Gonzalo Joya, and Francisco Sandoval. Continuousstate Hopfield dynamics based on implicit numerical methods. Lecture Notes in Computer Science, 2415:1365– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2415/24151365.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2415/24151365.
                                        pdf.

Riaza:2002:TSR

[1789] Ricardo Riaza and Pedro J. Zufiria. Time-scaling in recurrent neural learning. Lecture Notes in Computer Science, 2415:1371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2415/24151371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2415/24151371.
                                        pdf.

Anonymous:2002:AIbm

[1790] Anonymous. Author index. Lecture Notes in Computer Science, 2415:1377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2415/2415auth.pdf.